A Complete Bibliography of Publications in the Journal of Statistical Computation and

Simulation

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

19 September 2025 Version 2.02

Title word cross-reference

(5, 3; 5 × 3 − 1) [Agg87]. (λ = 2/3) [Ant95]. (M, S) [REK81]. (y_{1}, y_{2}) = (1/2)(y_{1} − y_{2})^{2} [Dan80b]. −2 log(likelihoodratio) [Goo77]. 1 [Ede94]. 2 [And95c, Gha01, LS99, Mou95]. 2 × 2 [ABH82, Lat82, Mag81, Pig91, Ram89]. 2 × 2 × 2 [Hab84, O’N82]. 2 × s [GL83]. 4 [Dan80b, DL75]. 5 [DL75]. ^{15} [KW96]. ^{2} [AG81, Mur78]. ^{k} [LZ92]. n [Goo78f]. ^{(}R) [HCW07]. _{3}F_{2} [RACSGJ05]. A [Mye98, Won95]. α [PP06b]. ar(1) [FL82]. Ȳ [Jen89b]. BIBD(s^{2} + s + 1, s + 1, 1) [SK90b]. BIBD(s^{2}, s, 1) [SK90b]. C [BK96, BR03, Kös06]. χ^{2} [Ant95, Cha94, GH83a, Goo90j, KG80, Law92, Por88, Wes72, WK72, ST74]. D [Jai85, Dög01]. Dβ; v_{1}, v_{2} [Rod76]. E [Sin90b, Tsa06]. E(X) = E[E(X/Y)] [Mar98]. E(XIY = y) [Mar98]. e^{x} [Mar86b]. ED(p) [McL78]. ϵ [AS04]. F [Agg87, HH92, LKKL07, Lev78a, ST74, Too72, Wil75, WK02, Cha94, El 82, KS97, Roh88, tKWY95]. f(x) [Roh88]. F (y) [Dog92]. G [DM78, Goo81v, MY90]. H_{0} [Rod76]. I [BW01b, WS04]. ∫_{0}^{∞} exp(−x^{2})g(x) dx [KTB82]. I × J × K [Hos87]. K [BLC04, VS91, AB09, BL05, CG77, DSL06, Goo80l, KS04, Lev78c, Mad78, MM80, NS08b, Poo80, Ros95, TO04]. L [Hua91, Nar96, Dyk85, Won95]. L_{0} [WM99]. L_{1} [NSG91, WM99]. L_{16} [LAR09]. L_{2} [Pak99]. L_{∞} [WM99]. L_{p} [LK83a, SA95]. ln(z) [SB85]. M [RS86, CC01, DC00, LF82, Wu01]. M_{T} [Llo93]. m ≥ 3 [Bha01]. N [LY99, Var81, Goo80a, Goo85r, Mar81]. N(−1, 1) [Var81]. N(1, 1) [Var81]. n = 12(1)18 [Fra88]. n = 3 [Dan80b]. O(n^{4}) [SQ02]. P [Goo93c, KK99, KB87, RG93b, Wil04, Alb83, Gil07, Goo80h, Goo83b, Goo86o, Goo89l, Goo89s, Goo90i, Goo90a, Goo92b, Hol85a, Lam05, LPSL05, Mad78, Moo90, SQ07, Tho89]. P (Y < X) [CKT89, SCAH05]. par(1) [AAL02]. ϕ [MPPZ05]. π [DM05]. pr[X > Y] [IJP86]. Q [EEÖ88]. R [GMSS95, JH90b, JH90a, RS89, NS08b]. R^{1} [Nic05]. R^{k} [Jen93c]. R^{n} [Jen94, JT80]. r × s × t [Lee85]. S [Che05, HV05, Sim00, VR92, GH82b, KM06]. S^{2} [Jen89c, JH90b, JH90a, Jen89b, Jen82b]. S_{1}(S_{1} + S_{2})^{−1} [SWK73]. S_{U} [Tue01, SB82a]. \sqrt{b_{1}} [Dod83]. T [CR93, Gue78, BD09, Che78, Cic89, CVS98, CG84, DD77, El 82, Eti81, FB90, GB99, GL92, Goo86s, GS86, Goo86a, KC73, LF80, LS86, LF82, ND84, Nor84, Pos78, Pos79, Pos82, Pos94, ST74, SS75, SI72, VG05, WK75, Wei79, Zim04]. T^{2} [Jen83, Jen95c, Jen95b, KAST05]. τ [Gat09]. u [Hun02]. V [BS79]. V (n, k) [RR77]. W [LK88, Roy89, Par92]. x! [Goo85e]. X [KSK93, MW92, SCAH05, Goo85e]. X + X [RS89]. X^{2} [Goo79t, Goo79u, Goo79f, Ken79d, Goo77]. Y [KSK93, SCAH05]. z [WP81].

-1 [BG78]. -ahead [NS08b]. -chart [MW92, Sim00]. -class [MM80]. -contaminated [AS04]. -curves [Tue01]. -densities [DD77]. -Dimensional [Dög01]. -distribution [BD09]. -distributions [HV05]. -divergence [MPPZ05]. -Estimates [Nar96, RS89, RS86]. -estimation-based [Wu01]. -estimators [CC01, GMSS95]. -fold [GH82b]. -frames [Goo80l]. -linear- [NS08b]. -matrices [Hen95]. -Matrix [LK88]. -means [Ros95]. -minimax [DM78]. -optimal [Won95, REK81]. -optimum [Jai85]. -parameter [LS99]. -probabilities [GB99]. -projections [Dyk85]. -ratios [Wil75]. -records [AB09]. -rectangles [Agg87]. -sample [BL05, Kös06, MY90, Poo80, TO04, BK96, BLC04, BW01b, BR03, WS04]. -shaped [Hun02]. -smoother [DC00]. -Stage [Mad78, LY99]. -Statistics [Eti81, Hua91]. -system [Che05, VR92]. -Test [CG84, KC73, GL92, Goo86s, LS86, Pos78, Pos79, Pos82, SS75, SI72, HH92, LKKL07, Lev78a, Too72]. -tests [ND84]. -Tobit [Moo90]. -tuples [Goo85r]. -unit [KS04]. -Value [SQ07, Goo80h, Goo83b, Goo90a, KB87]. -Values [Mad78, Gil07, Goo86o, Goo89l, Goo89s, Goo90i, Goo92b, Hol85a, Lam05, WK75, Goo93c, KK99, RG93b, Wil04]. -variables [Gue78].

14 [Ano03b]. 142 [Mar90b].

2 [Sha03c]. 213 [Sar98]. 22 [Mar90b]. 228 [Goo89p]. 2374 [Mar90a]. 2SLS [Smi78].

305 [Ano90a]. 307 [Ano90a]. 310 [Ano90a]. 311 [Ano90a]. 334 [Ano90a]. 37 [Goo89p]. 3d [Goo86i].

50 [Kap87]. 51 [Sar98].

72 [Ano03b].

A-optimal [Ang03]. Aalen [CFOS02]. Abrupt [RG02, Ho93]. Absolute [KPS01, ANPV97, Bon05, CCHM08, DR94, DR02, Die05, Die06, GMR82, KN89, Pap80, SNP93]. absorbers [How85]. Abstracts [Ano73a, Ano73b, Ano73c, Ano73d, Ano74a, Ano74b, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano76a, Ano76b, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano82a, Ano82b, Ano82c, HA78, And79, Ano79m, Ano80o, Ano80p, Ano80q, Ano80r, Ano81q, Ano81r, Ano81s, Ano81t, Ano81u, Ano82n, Ano83d, Ano83e, Ano84f, Ano84g, Ger78, Hel78, HN83, Hel86, HZ88, SF77]. absurdum [Goo81b]. accelerate [NC96]. Accelerated [AHAH06, AHAH08, DP93, EM86, Mou95, Nic96, OP00a, OP00b]. accelerated-sequential [Nic96]. Accelerating [MTO08]. accelerators [RS97]. acceptable [Goo87c]. access [RK05]. accident [CC89, KC97b, RT90]. accidents [RT88]. account [KE03]. accuracies [SC75]. Accuracy [KA93, OM88, CW74, Dee76, El 82, GT90, GW94, Goo85e, Hol85a, KB87, Lev82, PP80, PP85, PQ89, WG92]. Accurate [Atk91, CL97, Mar87, Auc79]. achieve [Goo84i]. achievement [Goo84s]. actinides [ELB97]. action [TBT95]. active [LiT95]. activity [FcC05]. ad [Goo81b]. Adaptive [BR03, OB92, RC98, Sch92, TKS78, ACNT05, BS05a, BF83, BT00, CRT07, CS05, DA87, EK03, GRVV08, Jam01, O’G06, O’G09, Par80, PP06c, PS09, Shu08, SW75, SY92, SJ06, Wor89]. Addenda [Goo89p]. addendum [Goo82l, Goo89o]. addition [LL93]. additional [JG83a, MP81, San89]. additions [Edw85]. additive [Bén94, BLP09, FL82, GH82a, LL08b, O’N82, SQ07]. Adenine [Goo78m]. adequacy [Hua01]. adhockery [Goo83c]. Adjusted [FFCN07, SC09a, CNZ04, Jön07, Llo05, NS08a]. adjustment [BCH08, Dan79, Goo79d, LGPP96]. adjustments [SE90]. admissibility [Bak87]. Adsorption [Dög01]. advanced [Ano05d]. advertising [Kel94]. advice [Goo78b]. affected [EH01]. Affymetrix [HCW07]. after [GL92, Goo90d, Neu07, PN98]. again [Goo86e]. Against [BSS02, AG85, Bha01, DL92, Eve88, Goo81i, Goo86q, Kös06, KAS96, Lev78a, Mar76, Sef92, VB83]. age [Mar95, Goo86m]. age-specific [Mar95].

Aggregate [Kuk87]. Aggregated [Van84, HR07, Van87]. Aggregating [TH05]. aggregation [Dem90, Mol99, SB96]. aggregation/disaggregation [SB96]. Agreement [QD83, BT09]. ahead [NS08b]. AHP [TL96]. AI [Goo89w]. Aiding [Fea79]. aids [TJKB00, MW94]. air [RS97]. Akaike [KE03]. Alexander [Mye98]. Algorithm [Aok02, AEL03, AL93, AL94, ÁGR06, Ang03, BLP09, But99, Cai09, CD92a, Çel09, CCD96, CS94, CR73, CS05, CYB90, DM79, Dut99, EN90, Fre09, Gle89a, GSF78, Goo78l, Goo83k, GH07, Gup84, Hat86, Hig97, HH93, JG92, Kk90, KP96a, KC97c, Lai82a, LB08a, LF04, LiT95, Liu08, MOS94, Mal06, MZZ89, MW94, Mar95, MM80, NW86, PK72, PS98, PP06a, Pos94, PSW98, QFG87, RL08, SH72, TTF07, TC92, Tue01, VPB03, Vir07, Whi95, Wil89, YL85, YW09, ZR93, Zha02, ZR07]. Algorithmic [Wel82]. Algorithms [REK81, Atk92, BC08, CH91, Cha94, Che06, CY91, Dee76, FPRS92, KCS88, KKE07, KA93, Lee02b, LS88, Möh05, NP81, PHO05, QT92, RRT99, SC75, TW91, WK90b]. alignment [ÁGR06]. Allied [Sha05c]. allocation [CZ02, Coa92, Coa95, NA09, Pot81]. allocations [WNB07]. allowed [Goo80a]. allowing [Goo85r, LhKN05, NTK09]. almost [Wan99]. alpha [LAR09]. also [Goo88a]. alternating [Fri79]. Alternative [DG95, Sch86, TWLC07, AS81, CS95b, Dag95, Dow02, Eve88, HM85, JW97, KX03, KA93, Mar76, MM08, Owe81, SK08, SFSS85, SHW93, Wil01]. alternatives [AG85, BM86, Bha01, BK96, CRM06, Goo81i, GT81, KS97, KC73, Kös06, KAS96, Lev78a, MM93, PCS09, Poi92, Poo80, Rod76, Sef92, WK72]. always [Goo95d]. amalgamation [Goo86f]. ambient [DE06]. American [CCY04]. among [BCL93, BE86, HM98, LL07, Sha87b]. analogous [Goo92e]. analogue [Goo82o, GP95, Goo81i]. analogy [And89]. analyses [Bel93b, FC96, Goo81r, Hor97, Kle97, SJ07]. analysing [WmGT95]. Analysis [GLC99, Kim07, LB96, McL80, MK08, PH95, PP06a, Sch02, SHH85, Sha06b, Sha07h, SD02, ScK97, TYY02, AD01, AB03, And97, Ano05e, Ano06a, ACNT05, Ban78, BBP04, BH73, BW01b, BP78, BG99, BD09, CW72, CP76, CCHM08, CG93, CLP93, CDJ02, CM05b, DF80, DL80, Dor01, DSS06, cDJgS93, ELB97, EBS86, EH07, FH86, Fan03, FcC05, Fra74, GK90, Goo78h, Goo80f, Goo80m, Goo83-30, Goo90n, GP95, Gue89, GL90, HK00, HHI90, Hel97, Hin97a, HBC07, HK92, HJ85, Hon90, Hun87, IMP^{+}97, JCKS09, JR09, JMY96, JH72, KP09, Kan75, KW96, KMK87, KS97, KSH73, Kim92, Kle97, KHSG83, KL93, KC97b, Kru88, Kru89a, Krz83, Kun98, LSL97, Lee02a, LC92, Lev78d, LW95, LDC73, Lun06, MS01, MRR84, MC91, MB97, MO91, MGG78, Nab83, Nan98]. analysis [NC06, NJJ92, NS09, OP04, PSB03, PY99, PR84, RR03, RG06, RR06, RFGE86, Rid03, RMH88, RWL95, SM03, SSW95, Sar98, Sch75, SND89, Sha08a, SB92, SR97, SP01a, Soh00, Spe06, SC82, SS88, TP98, Tan01, TJKB00, TPG93, UP01, VGD06, VP89, Vir07, VR92, WK90a, WK90b, WB09, WDR86, Whi94, WS04, XPC03, XW07, YM96, ZR07]. analytic [IMP^{+}97, TL96]. analytical [FCN99, HP95]. Analyzing [CYC99, SJ06, LN77, NL77a, Soh94]. ’ancillarity [Goo87g, Goo84l]. and/or [LL08b, Sha06b]. Anderson [BB80, TO04]. animals [EB90]. annealing [GGSNR09, Whi95]. Annenda [Goo81-28]. annual [GB86]. Anomalies [JR09, Win75]. ANOVA [Lev78a, ASS97, Dow02, Goo86j, Gri02, Hos91, KM83, RS88a, RG93b, Rod76, SS92]. Anova-like [ASS97]. Anselin [Gri04]. antedependence [ZNAEB98]. anticipation [Goo00]. Antisurprise [Goo83h]. antithetic [BP92, Do92, RW77, RS85]. any [Hol85b, Koh81, NC72]. aperiodic [RV08b]. APL [How85]. apology [Goo89r, Goo95b]. Apparently [Kun93]. applicable [Goo82b]. Application [BK84, GLC99, Hin97a, RRCD97, CMQ03, Con95, CH88, DE06, DF80, DM78, GB99, Goo79g, Goo85o, GB86, GGSNR09, HTC07, Kap87, KMK87, Kul90, LDR92, May06, Mel80, NBB00, PK72, RS89, SZ02, Sch75, Sho86, Zac80]. Applications [BCP02, Ano06f, BS95, CG77, CGN04, CW99, FB90, Gly84, Goo84u, HS86a, IK03, KPKPB95, Kel94, KM99, Ma97, RRT99, RH76, SK08, SB88, Sha05e, TJKB00, WK90a, Wan08d, And90a]. Applied [McL80, SD02, BL09, CBPW97, DCA03, Goo84m, HG85, RFGE86, SE02, WV79, YM96]. Applying [LP00, LH79]. appraisal [HS77]. Approach [DG02, LK88, McL80, AL01, AS04, AK85, Bre93, Çel09, Cho98, EB90, FBC09, Fon90, Fon92, GS07, Goo81v, Goo86f, Goo86j, GM77, GL96, HKL08, KKE07, Kim93, LFC92, LZNL08, MZ08, MK97a, NW83, OBW05, Pan99b, PNM83, PV93, RS09, SS97a, SSM95, SB92, SFC08, SP97, Spi98, SJ06, Sy01a, TB88, WB09, WCF79, WO93, Won95]. Approaches [WDR86, BB84, Goo83-27, HHKD02, JK89, LL90, NJJ92, SG96, Sel08]. approximants [NP81]. Approximate [Ada97, AMAE97, Bal89, Cra05, FC96, OHN93, Ouy06, SM94, Wan92, ZL07, Bal92, BC94, FS75, Gol77, GSF78, Hol85a, Hou85, OM88, Pad82, Pat76, STS94, kSB90, TR75, ZW01]. Approximating [Fro04, RS90b, She83, And90b, And95c, Sol90, Tan82]. approximation [BA77, BSBS08, Cad94, Cha79, Eti81, GMR82, Goo81r, GLP72, GJ77, HP95, Jen76, Lev82, LPZ02, Mar86b, Mcd74, Mye98, OZ81, PF04, Par87, PPRW06, Por88, RG93a, SB08, Sho86, SA95, Whe75, YI01]. Approximations [BS01a, NL77b, Wes72, And95a, Arc80, BC94, BBHW95, BS94b, Che78, DL81, Dev82, DN94, Gat09, GH76, Goo86p, Jen91, Ma97, Ma99, Mag75a, Mol79, PP80, RR82, ST74]. arabinoside [Goo78m]. arbitrary [AkBA05, DM79, GL96, HC06, TO04]. arch [Kun93, Fur04, HKT04, MM08]. Archimedean [McN08]. area [GMLM^{+}08, Goo82n, Goo84d, Goo84j, Goo84k, Goo85b, GS86, MZZ89]. areas [Fie93]. ARFIMA [LOR04, TH09]. argmax [DC99]. argument [CL91, Goo79j, Goo89e, Goo89b, Goo90h, Goo94b]. ARIMA [AK85, AMYY07, Ans80, Den77]. arising [Möh05, Pat76]. arithmetic [NO75]. ARMA [AH09, De 81, DCA03, HT83, Hie81, KA82, KA85, Mah96, McG89, McK86, Pap93, SP97, TC08, Wil79]. armed [Coa95]. Armitage [TNG^{+}06]. army [NBB00]. array [LAR09]. arrival [AMYY07]. art [FF84]. aspects [CW72, Dut77, Dyk85, Sny88, ZNAEB98]. Aspin [Gol72]. assays [ME72, RB92]. assertions [Goo89t]. Assessing [CM98, Men00, RV08a, STS94, CH98, Gra86, PCS09, Sch74, SFC08].

Assessment [OS02, VG01, CM01, DBVK02, DFY08, Goo88d, LMB08, Lee01, Goo81m]. assigned [Mal06]. assisted [BSBS08]. associated [BS72, Die94, FSBN07, GA95, Lev78d, RS89, TG73, Zim04]. association [Goo81d, GL83, Goo85p, LPSL05, RS90b]. assuming [Fro89, SE90]. assumption [DFY08, Fri85]. assumptions [CS86, GJJLGS06, Too72, You93]. astronomers [Goo98a]. asymmetric [AFGP06, CM04, TB86, TB88, Wan99]. asymmetrical [Che94]. Asymptotic [Ada96, Agu99, ANPV97, Gat09, GM08, GN89, JG96, MP81, MS95, Nab83, Oga07, OP00b, Sha06a, SB00, TMW73, And90b, And95c, And80, Ant95, BS05c, CK96, Cro74, Day87, For97, Gly84, GH76, NS89, PP85, Smi84]. asymptotically [BG07, MH72]. asymptotics [FUOCN97, KW78]. atmospheric [CW74, GZP05]. atomic [Goo88d]. atoms [Goo98a]. attack [GHDB89]. attacks [Goo89j]. attempts [Hut93a]. attributable [SNC09]. attribute [GR08]. auditing [Cla96]. auditors [Joh94]. augmentation [SD01a]. augmented [CL98]. augmenting [DG95]. authorship [Goo81l]. autocorrelated [AD01, AD03, DR94, DM93, Lak81, MB86, MB90, SL09, TA94, WHF80, WD93]. Autocorrelation [KA03, AL02, And95b, BR00, DL92, Goo81t, JAK93, MW92, Sha87b, SGW94, SR00, TY86]. autocorrelations [De 81, RR01]. autocovariance [Akn07, Ans80, Pap93]. autocovariance-based [Pap93]. autocovariances [AH09, De 81, KA82, McK86]. autogressive [YL85]. autologistic [SAC06]. automated [LF04]. Automatic [CJ00, DC00, Gue89]. autonomous [Wan90]. autoregression [She83, Tse84]. Autoregressive [BSS02, LT95, OVL02, AK86a, BC08, Bar77, Cho98, Del83, Dic78, EN90, Gri04, Kim98, KP99, KP04, KL09, NP81, Nic96, OR92, Pol94, SM03, Sma05, Soh92, TA92, Tsu93, WLT08, YKB86]. available [Din08]. Average [JH01, KP87b, AG78, AK86a, Cho98, Dam82, Del83, Gan90, Gan93a, GA01b, KP04, Rig95, SL09, Shu08, TA94, VGD06, VC78, WG92, WGC04, XDZ09]. averaged [SG96]. averages [Gal02, Jen78, Ros06]. averaging [KSK93]. avoiding [Goo78j]. Axiomatic [JK89].

B [KL94, KP96a]. B-Spline [KP96a]. B-splines [KL94]. back [MW94]. back-projection [MW94]. background [RS92]. Backorder [Ber73]. backpropagation [MTO08]. backward [Ber80, Ho93]. Bahadur [DAM98, Hab92]. balanced [AM87, CD92a, CY89, Do92, FS03, LL79, Pie94b, Pol94, STH09, Sin90c, ScK97, Tay72, TC73]. balancing [Dav93]. Balaton [KT85]. ball [Goo81m, Goo88a]. bands [Far90a, HP90, LP92, LJZB05, Mee93, QH80, SW82, SW83, WM91, XDZ09]. Bandwidth [LP09, BR00, CJ00, DC00, Din08, Far90a, FH09, KP96b, KPS01, PB00, ZT01]. bar [GS78]. Barndorff [Goo83v]. Bartlett [CCNAF95, Cor95, GMR82, LH79]. Bartlett-type [Cor95]. Based [LC02, Ada96, Ada97, Agu02, AA09b, Ana09, AZ05, Bak07, Bal83, Bal89, Bal92, BL02, BL03, BRF08, BG99, Bog01, Bow85, Bow92, CC05, Cao87, CL91, CS95b, Che03, CL98, Cho08, Cra05, CNFO05, Dic78, Dog89, DCA03, Ede94, Edi83, Edw85, FP07, Fos95, Fur07, Gan93b, Goo86d, GSC87, GV81, HS73, HL05, HB06, HL92a, HKT04, JK08, JS00, KM83, Kib04, KP82, KP96a, KC97a, KM99, KP96b, L’E97, LSCN08, LB08b, Lil01, LF82, LB01, LW95, LB07, LJ00, MH78, Mai77, MMPP05, Mar81, MMP08, Mar92, Mas03, Mei08, Mei09, MPPZ05, Mou95, MR03, Mou05, Mun06, NW09, NA09, OP00a, PS99, Pap93, PP03, PX02, Pet02, PP80, PR84, QL01, RA01, RS09, SJEs93, SB08, SN83, Smi89c, SSM93, ScK97, Sy01a, TB86]. based [TB88, TA08, TW08, Tsi02, Tut90, WK90b, Wan08a, WS00, WWS04, WB94, Wu01, WL07b, WCC07, WSPC09, XM09, YX03, ZNAEB98]. bathtub [WWL09]. Bayes [AM87, AMAE97, AJAH07, Alb83, Alb84, AS04, CA89, DP93, Fra74, FMC09, Goo79j, Goo81w, Goo83-29, Goo83-28, Goo83w, Goo84p, Goo85b, Goo86g, Goo86q, Goo86d, Goo88a, Goo94e, Goo00, Hie81, IP86, Kri77, KY93, Lai82b, ML74, MK74, Mou95, OHN93, SK80, Var81, WV79, Wil04]. Bayes/non [Goo83-29, Goo86g, Goo86d]. Bayes/non-Bayes [Goo83-29, Goo86g, Goo86d]. Bayesian [AHJ92, Alb87, Alb92, AB05, AB97, BN96, Bog01, CS86, CW72, CCNA09, CLLN04, CDJ02, Con06, DG02, DV95, EH07, Eve01, FB90, Fan95, FcC05, FBC09, Fon90, Fon92, FTS09, GS07, GG78, Goo78g, Goo79i, Goo80i, Goo81r, Goo82g, Goo82e, Goo83r, Goo84l, Goo84k, Goo86n, Goo89v, HMK00, Hoe89, HKL08, JCKS09, JUP86, LL99, LZNL08, May06, Nan98, NK02, NC06, NTK09, PT89, PH95, PY99, PGV04, Qu06, RM02a, RM05, SM03, Sah02, Sen02, Sha07i, SFC08, ST87, SP01a, Spe06, TB86, TB88, Tsi03, TJKB00, TPG93, VGD06, WK75, WW95, Wec94, Wil01, Wil07, WCF79, You93, ZR93]. Bayesianism [Goo86d]. BDS [GHW99]. be [AD03, Ano78g, BH96, Goo81a, Goo81o, Goo81f, Goo84d, Goo88c, Goo93a, Goo96a, Kru88, Kru89a, Kun93, MF02, Phi91]. beer [Goo90j, Goo92b]. before [Goo81o]. behaved [DL81]. behavior [Fur96, MP81]. Behaviour [DAM98, MGN99, Rod07, Rud86, SLL00, SB96]. behavioural [CL97]. Behrens [Fun79, Gol72, Goo86j, KC96, Mol79]. being [HS86b]. bending [And89]. Bernoulli [BF83, Coa92, Goo83f, Van05]. Bessel [Goo90h, IK03]. Bessel-function [Goo90h]. Best [BP86, GV81, HSB85, LB08b, Sha07i, BF83, Fro89, Goo84h, HSW75, KB87, Ram89, SS97a, TKS78]. Beta [CNV02, Bj78, BS92, CDDCN97, Day87, DW80, Éri97, KW78, LL91, Lee92, LL99, Lou84, Ong95, SBC03, SKC75, Sef92, TAY02, VM00, WS90, YI01]. beta-binomial [LL99, WS90]. Better [HMS89, SW75]. Between [Krz82, WW03, BG78, BCH08, BS94a, Bay90, BD82, CM07a, CFOS02, Di 05, EH01, GG80, Goo80c, Goo80b, Goo81e, Goo82j, Goo83d, GL83, Goo85s, GL86, Goo90l, GK04, HB78, HK00, HH92, Hie81, HH93, Kel94, LFC92, Mei09, MP96, NS09, SS92, Var81, WLT08]. Between-group [Krz82, BD82]. Bhapkar [BS79]. bi [RW93]. bi- [RW93]. bia [GP07]. Bias [BS05b, BT09, Cad94, CSC00, CDDCN97, LS01, McG89, MCCC04, Pau92, YAAB87, CVS98, CW99, Fos95, GRVV08, GS84, Kar09, Men00, SC09b, Ray85]. Bias-corrected [CDDCN97, LS01, GRVV08]. biased [CH90, oE89, EL75, GB86, HP00, ML79, PLD88, PO05, SM96]. biased-robust [SM96]. biases [CRR99]. bibliography [Sah79]. Bidimensional [OVL02]. Bienayme [GG90]. bilinear [CS97, WS92]. billiard [Goo79j, Goo88a, Ken79e]. billiard-table [Goo79j, Ken79e]. bimodality [FT96]. bin [SJEs93]. bin-based [SJEs93]. Binary [SHH85, AL01, BS05a, Bow01, CC05, DAM98, Don97, LF97, Lee98b, MH07, MP96, Pan99b, PSB03, Rid03, RL08, SB92, SAC06, TS09, TTF07]. binding [RB92]. Bingham [MZ77, Str89]. Binning [ZC03]. Binomial [YM99, AL93, Alb84, Bea85, BR06, CC01, DM78, Fam98, For97, Gan90, Gao04, Gou93, Hab92, HSW75, Hör93, Jon86, JJH09, KCS88, KAS96, LL99, LN77, LB96, LP85, Lu99, MZ08, ML74, NL77a, Poo80, PO06, SB82b, SL89b, SW75, SP01b, TT82, TTNC09, Van93, Van05, Wil07, WS90, Zac80, vdTGL97]. Binomial-Logit [YM99]. bioassay [Cob89, CC90a, EOD86]. bioequivalence [CM01]. Bioinformatics [Sha07a, Sha09a, Sha09h]. biological [ME72]. biomathematics [Sha05e]. Biostatistical [Sha03a]. Biostatistics [Sha09g, Ano05d]. biparametric [VD08]. Birnbaum [DRYL08, LSSP08, LSCN08, PT03, RR79]. bisection [Goo85d]. Bivariate [AB08, KL94, WK02, AL94, Ahm87, ASY80, BN01a, BGR94, Bon06, Cha79, Con95, DW90, Ebr93, GA09, GHDB89, GW73, GJ77, GJJL02, GJJLGS06, HL98, HM85, HCW07, HS91, Jen76, JS02, JAK93, JJH09, KKB85, KR89, LH97, LM84, LSS93, Lou84, LP85, NK07, Ong95, Par81, PPK77, Pru93, SL98a, TW91, Tho92, WK90b, Jen93b]. Blind [BHZ08]. block [Ang03, BC82, BS78, CR73, Die93, GZP05, HC75, REK81, Sin90c, Sin90b, SB91, Tay72, TC73, Too72, WB94]. block-size [SB91]. Blockinq [Mih74]. blocks [TC73]. blue [Ada97]. blues [Ada96]. blurred [Goo95b]. board [Ano77f, Ano77g, Ano78h, Ano78i, Ano78j, Ano78k, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano82e, Ano82f, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano83a, Ano84a, Ano84b, Ano84c, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano88a, Ano88b, Ano88c, Ano88d, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano91a, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h]. board [Ano92i, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano94a, Ano94b, Ano94c, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano01m, Ano01n, Ano01o]. Bonferroni [BH95]. bonus [CI81]. Book [AC99, Ano72a, Ano75f, Ano02a, Ano02b, Ano02c, Ano02d, Ano04a, Ano04b, Ano05d, Ano05e, Ano05a, Ano05b, Ano05c, Ano06b, Ano06a, Ano06c, Ano06d, Ano06e, Ano06f, Ano08, Bor75, Che87, Goo80o, Kru73, Mye79, Mye83, RS96, Sha01a, Sha02c, Sha02a, Sha02b, Sha03b, Sha04n, Sha04a, Sha04b, Sha04c, Sha04d, Sha04e, Sha04f, Sha04g, Sha04h, Sha04i, Sha04j, Sha04k, Sha04l, Sha04m, Sha05c, Sha05d, Sha05e, Sha05f, Sha05a, Sha05b, Sha06b, Sha06a, Sha06c, Sha06d, Sha07i, Sha07j, Sha07h, Sha07c, Sha07d, Sha07e, Sha07f, Sha07g, Sha08b, Sha08c, Sha09c, Smi89a, Smi89b, War84, Sha07b, Sha08a, Sha09b]. Books [Ano06b, Sha05c, Sha05d, Sha06b, Sha06a, Sha06c, Sha06d, Sha07i, Sha07j, Sha07h, Sha09d, Sha09e, Sha09f, Ano05d, Ano05e, Ano06a, Ano06d, Ano06e, Ano06f, Sha03e, Sha04n, Sha05e, Sha05f, Sha07a, Sha09a, Sha09j, Sha09k, Sha09l]. Boolean [Goo81p, Goo79n]. Booststrapped [MS01]. Bootstrap [CM01, DV86, DR02, Far90a, GMLM^{+}08, HKT04, HMZ05, JAK93, Kar09, Kim98, Kuk87, LSCN08, McL80, Pet02, kSwXR93, SNP93, WW95, AB05, Ana09, ASS97, BG01b, BK98, BR03, CD92b, Cla96, CNZ01, CNZ04, DS89, Do92, Fan95, FH09, FCN99, HMS89, HP90, Hua91, Kuk89, Lou95, MZ08, MGG09, PT89, PP06c, Pol97, VA08, Var81, Won85]. Bootstrap-based [LSCN08]. bootstrapped [RWD95]. Bootstrapping [CKT89, DY92, HCA96, Kar90, ST88]. Borel [GJ77, GJ79]. Both [Van84, Bho84, DSL06, Goo80a, Han86]. bound [AS01, Har03, Shi07]. boundaries [Gan93a, PP73]. Boundary [Chu01, LD87]. Bounds [KB87, AB97, Bar81, BSK90, Har91, Jen93c, Jen94, Jen95c, Ma99, OM88, PT03, SDP93]. box [Har89, Goo83-29, HT93, KAST05, Spi82]. branching [Goo89m]. break [Pop08]. breaking [BG01a]. breaks [CGA09]. bridges [Kul90]. Brief [RR08]. brittle [SDP93]. Brown [Tan82]. Brownian [Car07, DC99, Har87a, Kul90]. Bruijn [AA09b]. Buckley [WZ95, YW02, YW09]. budget [Joh78]. building [BC08, De 97, KPJ09, RRCD97]. bumpy [Goo81i]. Burr [AHJ92, Mou95, WCC07]. butterfly [Goo94d].

C1 [Wat77]. C10 [Goo78a]. C100 [Goo81a]. C101 [Goo81b]. C102 [Goo81c, Hut93a]. C103 [Goo81d]. C104 [Goo81e]. C105 [Goo81f]. C106 [Goo81g]. C107 [Goo81h]. C108 [Goo81i]. C109 [Goo81j]. C11 [Goo78b]. C110 [Goo81k]. C111 [Mag81]. C112 [Bon81]. C113 [BR81]. C114 [Goo81l]. C115 [Goo81m]. C116 [Goo81n]. C117 [Goo81o]. C118 [AG81]. C12 [Ano78g]. C120 [FJT82]. C121 [Ano82d]. C122 [Goo82a]. C123 [Goo82b]. C124 [Goo82c]. C125 [Jen82a]. C126 [Goo82d]. C127 [Goo82e]. C128 [Jen82b]. C129 [Goo82f]. C13 [Jen78]. C130 [Goo82g]. C131 [GJ82]. C132 [Gue82]. C133 [Goo82h]. C134 [Goo82i]. C135 [Goo82j]. C136 [Goo82k]. C137 [Joh82]. C138 [Goo82l]. C139 [Goo82m]. C14 [Goo78c].

C140 [Goo82n]. C141 [GH82a]. C142 [GH82b]. C143 [Goo83a]. C144 [Goo83b]. C145 [Goo83c]. C146 [Goo83d]. C147 [GL83]. C148 [Goo83e]. C149 [Goo83f]. C15 [Goo78d]. C150 [Goo83g]. C151 [Goo83h]. C152 [Goo83i]. C153 [Jen83]. C154 [Goo83j]. C155 [GS83]. C156 [Del83]. C157 [Goo83k]. C158 [Goo83l]. C159 [Pap83]. C16 [Goo78e]. C160 [Goo83m]. C161 [Goo83n]. C162 [Goo83o]. C163 [Goo83p]. C164 [Goo83q]. C165 [Goo83r]. C166 [Goo83s]. C167 [Goo83t]. C168 [Goo83u, Goo89n]. C169 [Goo83v]. C17 [Goo78f]. C170 [Goo83w]. C171 [Goo83x]. C172 [Goo83y, Goo89o]. C173 [Goo83z]. C174 [Goo83-30, Goo83-28, Goo83-27]. C175 [Goo83-28]. C176 [Goo83-29]. C177 [Goo83-30]. C178 [Goo83-31]. C179 [GH83a]. C18 [Goo78g]. C180 [GH83b]. C181 [SA83]. C182 [J.83]. C183 [Edi83]. C184 [GS84]. C185 [GL84]. C186 [Goo84a]. C187 [Goo84b]. C188 [Smi84]. C189 [Goo84c]. C19 [GG78]. C190 [Goo84x]. C191 [Goo84d]. C192 [Goo84e]. C193 [Goo84f]. C194 [Goo84g, Kem84]. C195 [Goo84g]. C196 [HL84]. C197 [Goo84h]. C198 [Goo84i]. C199 [Goo84j]. C2 [Goo77]. C20 [LH78]. C200 [Goo84k]. C201 [Goo84l]. C202 [Ano90a, Goo84m]. C203 [Goo84n]. C204 [Goo84o]. C205 [Goo84p]. C206 [BD84]. C207 [Bar84]. C208 [Goo84q]. C209 [Goo84r]. C21 [Pad78]. C210 [Goo84s]. C211 [BR84]. C212 [GBSS84]. C213 [Goo84t]. C214 [Goo84u]. C215 [Goo84v]. C216 [Goo85a]. C217 [Goo85b]. C218 [Goo85c]. C219 [Bea85]. C22 [Goo78h]. C220 [Goo85d]. C221 [GT85]. C222 [Hol85a]. C223 [Goo85e]. C224 [Goo85f]. C225 [Goo85g]. C226 [Goo85h]. C227 [Goo85i]. C228 [Goo85j]. C229 [Goo85k]. C23 [Goo78i]. C230 [Ray85]. C231 [Fri85]. C232 [Goo85l]. C233 [Goo85m]. C234 [Goo85n]. C235 [Goo85o]. C236 [Goo85p]. C237 [Sha85]. C238 [Goo85q]. C239 [SB85]. C24 [Goo78j]. C240 [GS85, Mar90b]. C241 [Goo85r]. C242 [HG85]. C243 [Hol85b]. C244 [Goo85s, Goo86b]. C245 [Jon86]. C246 [Goo86a]. C247 [GS86]. C248 [Goo86b]. C249 [Mar86b]. C25 [Goo79a]. C250 [Goo86c]. C257 [Goo86s, LS86]. C258 [Goo86s]. C259 [Kru86a]. C26 [Goo79b]. C260 [Goo86d]. C261 [Goo86e]. C262 [Kru86b]. C263 [TY86]. C264 [Goo86f]. C265 [Goo86g]. C266 [Goo86h]. C267 [Goo86i]. C268 [Goo86j]. C269 [Goo86k]. C27 [Dan79, Dan80b]. C270 [Goo86l]. C271 [Goo86m]. C272 [Goo86n]. C273 [Goo86o]. C274 [Goo86p]. C275 [Goo86q]. C276 [Goo86r]. C277 [Goo87a]. C278 [Goo87b]. C279 [McN87]. C28 [Goo79c]. C280 [GS87]. C281 [Hil87]. C282 [SB87]. C283 [Goo87c]. C284 [Goo87d]. C285 [Goo87e]. C286 [Goo87f]. C287 [BF87]. C288 [Goo87g]. C289 [Goo87h]. C29 [Goo79d]. C290 [Ter87]. C291 [Har87a]. C292 [Bak87]. C293 [SL87]. C294 [Goo87i]. C295 [Agg87]. C296 [Sha87a]. C297 [Goo88a]. C298 [Roh88]. C299 [Goo88b]. C3 [Hut77]. C30 [Ken79a]. C300 [Goo88c]. C301 [RT88]. C302 [Goo88d]. C303 [Goo88e]. C304 [GL88]. C305 [Goo88f]. C306 [Tho89]. C307 [And89]. C308 [Kru89b]. C309 [MZZ89]. C31 [Ken79b]. C310 [JK89]. C311 [Goo89a]. C312 [Goo89b]. C313 [Goo89c]. C314 [Goo89d]. C315 [Goo89e]. C316 [Goo89f]. C317 [Goo89g]. C318 [Goo89h]. C319 [Goo89i]. C32 [Ken79c]. C320 [Goo89j]. C321 [Goo89k].

C322 [Goo89l]. C323 [Goo89m]. C324 [Goo89n]. C325 [Goo89o]. C326 [Cre89]. C327 [Goo89p]. C328 [GHDB89]. C329 [GF89]. C33 [Ken79d]. C330 [Goo89q]. C331 [Goo89r]. C332 [Goo89s]. C333 [Goo89t]. C334 [Goo89u]. C335 [Goo89v]. C336 [Jen89a]. C337 [Jen89b]. C338 [Jen89c]. C339 [CC89]. C34 [Ken79e]. C340 [GC90]. C341 [Joh90a]. C342 [Goo90a, Joh90b]. C342.5 [Goo90a]. C343 [JH90a, RS90a]. C344 [Goo90b]. C345 [GJH90]. C346 [GG90]. C347 [Goo90c]. C348 [JH90b]. C35 [Dav79]. C350 [Ano90a]. C351 [RT90]. C352 [Sin90a]. C353 [And90a]. C354 [SK90a]. C355 [Lac90]. C356 [Goo90d]. C357 [Goo90e]. C358 [Kru90a]. C359 [Goo90f]. C36 [Kru79]. C360 [Ter90]. C361 [Goo90g]. C362 [Kru90b]. C363 [Goo90h]. C364 [Bay90]. C365 [Sin90b]. C366 [SK90b]. C367 [Goo90i, Sin90c]. C368 [Goo90j]. C369 [Goo90k]. C37 [Arn79]. C370 [Goo90l]. C371 [Mar90b]. C372 [Goo90m]. C373 [Dem90]. C375 [And90b]. C376 [CYB90]. C377 [Goo90n]. C378 [Goo90o]. C379 [Aok91]. C38 [Goo79m, Goo79e]. C380 [CY91]. C381 [SB91]. C382 [Jen91]. C383 [Goo91]. C384 [And92]. C385 [RS92]. C386 [Goo92a]. C387 [Goo92b]. C388 [Goo92c]. C389 [Goo92d]. C39 [Ken79f]. C39.5 [MM93]. C390 [Goo92e]. C391 [ZR93]. C392 [KSK93]. C393 [Jen93a]. C394 [Goo93a]. C396 [Roy93]. C397 [Goo93b]. C398 [Goo93c]. C399 [Goo93d]. C4 [GG77]. C40 [Goo79f]. C400 [Goo93e]. C401 [SDP93]. C403 [Hut93b]. C404 [PYC93]. C405 [Jen93b]. C406 [BS94b, Jen93c]. C407 [And93, Goo94a]. C408 [Jen94]. C409 [Pie94b]. C41 [Goo79h, Goo79g]. C410 [Pie94a]. C411 [BC94]. C412 [Joh94]. C413 [Tre94]. C414 [Ede94]. C415 [Goo94b]. C416 [Goo94c]. C417 [Goo94d]. C418 [Goo94e]. C419 [Goo94f]. C42 [Goo79h]. C420 [Goo94g]. C421 [Goo94h]. C422 [Che94]. C423 [Goo94i]. C424 [Goo94j]. C425 [Wec94]. C426 [Goo94k]. C427 [And95a]. C428 [And95b]. C429 [And95c]. C43 [Goo79i]. C430 [Tre95]. C431 [SSM95]. C432 [Jen95a]. C433 [Jen95b]. C434 [Jen95c]. C435 [Goo95a]. C436 [Goo95b]. C437 [Goo95c]. C438 [Goo95d]. C439 [Goo95e]. C44 [Goo79j]. C440 [KP96b]. C441 [JR96a]. C442 [JR96b]. C443 [Goo96a]. C444 [BH96, Goo96b]. C445 [Goo96b]. C446 [Goo98a]. C447 [Lee98a]. C448 [Mar98]. C449 [Wil98]. C45 [Goo79k]. C450 [Cho98]. C451 [Lee98b]. C452 [Sar98]. C46 [Goo79l]. C462 [Ano03b]. C47 [Goo79m]. C470 [Che03]. C471 [Kru03]. C472 [Har03]. C473 [Goo03]. C474 [Ano03a]. C475 [Ano03b]. C48 [Mal79]. C49 [Goo79n, Goo80e, Goo81p]. C5 [Goo78k]. C50 [Goo79o]. C51 [Joh79]. C52 [Goo79p]. C53 [Goo79q]. C54 [Goo79r]. C55 [Auc79]. C56 [Goo79s]. C57 [Goo79t]. C58 [Goo79u]. C59 [Goo80g, Goo80a]. C6 [Goo78l]. C60 [Dan80a]. C61 [Dan80b]. C62 [GG80]. C63 [Zac80]. C64 [Goo80b]. C65 [JF80]. C66 [Jen80]. C67 [Goo80c]. C68 [Goo80d]. C69 [MG80]. C7 [Goo78m]. C70 [Goo80e]. C71 [Goo80f]. C72 [Goo80g]. C73 [Goo80h]. C74 [Goo80i, Goo83-29]. C75 [DF80]. C76 [Pap80]. C77 [Goo80j]. C78 [Goo80k]. C79 [Goo80l]. C8 [KW78]. C80 [Goo80m]. C81 [Goo80n]. C82 [Jöc81]. C83 [Fou81]. C84 [Goo81p]. C85 [Hen81]. C86 [RA81]. C87 [Goo81q]. C88 [Goo81r]. C89 [Goo81s]. C9 [Goo78n]. C90 [Goo81t]. C91 [Ano81f]. C92 [Goo81u]. C93 [GT81]. C94 [Goo81v]. C95 [Goo81w]. C96 [Goo81x]. C97 [Goo81y]. C98 [Goo81z]. C99 [Goo81-27]. calculated [KB87, LFC92]. Calculating [AH09, Bea85, HL98, HQ00, RG93b, SB05, TT82, Smi89c]. calculation [Ber73, Cla86, CCMV07, GB99, Gil07, Goo81-29, Goo82l, Goo83k, Goo87c, Jun08, Owe77, SQ07, Smi94]. calculations [Akn07, BHG01]. calculator [Gue78]. calculus [Goo80n, KW94]. calibrated [JR09]. Calibration [BCP02, DV95, Fra74, GZP05, Kru72, WCF79, WW07]. Call [Bro74]. Cambridge [Goo89p]. Can [Goo81a, Goo96a]. cancer [GSC87]. candidate [CS05, PM02]. candidates [Goo80a]. canonical [BH73, WK03]. capability [CK96, PX02, PX05, Wri95, WSPC09]. carbon [Goo98a]. carcinogenesis [KHSG83]. carcinogenicity [SW90]. Carlo [AS00, ÁGR06, AT05, BF83, BP78, CG91, CS95a, Che06, CMD74, CM76, ES88, FPRS92, FW88, Fis73b, FC89, Fou81, Fun79, GHW99, GPW80, Gat91, Goo81w, Goo81o, Goo81j, Goo86o, HM95, HT85, Hig97, HM80, HL92c, JB91, JL88, JUP86, JH72, hJ93, KPKPB95, Ken79b, KP87a, KC97c, LF80, Lee90, LF04, Lev05, LN77, Lil01, Mae87, MM05, MGG09, MT80, Mil79, ML79, MM99, NL77a, ND84, NP98, Nic96, Nor91, OPS82, OB92, Pau84, PNM83, PQ84, Poi92, Poo80, Pot81, PS85, RRP06, Rid03, Rud86, SFSS85, SND89, SCW79, SCW81, SR97, SCM90, SW90, SC82, SB93, SS88, TZ97, TZ04, VM96, WB73, WZ95, ZL09]. cars [pB97]. Case [Sha03d, AFGP06, BLC04, BL05, CCD96, CKT89, Dea86, Goo89u, GGSNR09, HS86b, MT01, Ng96, RR08, RB88]. cases [Atk91, Goo81f, GS84, HJ85]. Categorical [RK05, AB08, Goo83e, KK05, LF97, LiT95, NC06, NTK09, OHLH82, SG96, SP01a, TB85, Tut90]. categorical-scaled [OHLH82]. categorical-valued [LF97]. categories [Law01]. Categorisation [RS88a]. categorized [BGR94, MGG78]. category [NTK09]. Cauchy [Jen78, Vau94]. Causal [GT90, Goo80n, Goo94b]. Causality [Akn07]. causation [Goo80f, Goo85c, Goo93e]. cause [Kun93, PB04]. caused [TC08]. causes [Goo93a]. Cautionary [HSWF07, TKT89]. Cautions [Hay97]. cdf [WK90a]. CEC [KC97b]. CEC/USNRC [KC97b]. cell [FGSV09, GM08, HH93, PP73, RV98]. cells [BR06, Goo81-29, Goo82l, IJW03, Ken79d, Law92, Mac83]. censored [AEL03, Ada96, Ada97, ABE83, Bal89, Bal92, BL02, BL03, BRF08, Bal95, BE94, CZ07, CC97, DRYL08, DL80, EM86, GD92, Gri02, GL96, HS73, HBK92, HZ03, JB91, Kar90, KH09, KP96b, KPS01, KY93, LL06, LL09a, Lei83, LB01, Mas03, MK08, Mou95, Mou01, NCA^{+}00, NW09, OPB08, OM88, Pan98, Phi91, Pou04, PW83, QY95, RM02a, SB08, SN83, SC82, TB86, TB88, TJKB00, Wan08a, WZ95, WL07b, WCC07, WBE80, XW07, YH85, YW09, ZZCS09, Zue96]. Censoring [TYY02, AMAE97, ABE85, Atk92, BLC04, BKLW04, BL05, BH07, EH92, GSS87, HH74, HL92c, Jen03, Kru79, Ng96, Wu03, WWL09, YT96]. censorship [Gho95, Wan08c, Lia11]. center [AGNS91, LP89, LTV90]. Centering [LS88]. Central [WK02, Bic03, Che05, Fie93, Jen95a, Lee92, Ma99, MS98, Rod76]. centrally [AGNS91]. centres [NS09]. Ceres [Wet96]. certain [BR81, BR84, JG80, LH72, SNB07b, SK80, Smi94]. certainty [WS82]. chain [ÁGR06, AT05, Har91, KS87a, LF04, Lev05, PSW75, WM90, Zan08]. chains [Fan95, Gly84, Goo82k, Har93, Har89, KK85, Li88, MS80, MO91, RRT99]. chance [AE87, WECC00]. Change [RG02, Wan08d, Coo07, FcC05, HM88, Ho93, Khe08, LhKN05, MBH91, NS08b, OP04, Pet80, WW06, WO97, ZA09]. Change-in-mean [Wan08d]. change-point [NS08b, OP04, Pet80, WO97]. changes [BCH08, GS07, RR03, RR06, TC08, WGP00]. channel [BHZ08, CM05a, Goo89c]. Chaos [LC02]. Chaotics [Goo89d]. characterisitc [WS00]. characteristic [ES86, Goo81j, GJ82, JL83, KM99, Mei09, pha95]. characteristics [AASAM03, KM06, PSB03]. characterization [GC03, Pap83]. characterizations [CW99]. charged [How85]. Charlier [Dod83]. chart [GR08, Gan90, Gan91, MW92, Rig95, SL09, Shu08, Sim00, WGC04, YR06]. charts [CG94, CSAR93, DCA03, Gan89, Gan93a, Gan93b, JH01, Jen82b, JH90b, JH90a, LL03, SR00, Zha00]. Chebychev [Jen80, Jen94]. checking [GK05, KL09, You93]. checks [Goo87a]. chemical [ME72]. Chi [WK02, Ant95, BR06, Bog95, El 82, Fea79, Fie93, For97, GG80, GMR82, Goo78n, Goo81i, Goo81j, Goo81n, Goo81-29, Goo82o, Goo82l, Goo83k, GS84, GR83, GW73, pH78, Han79, HB78, Hos87, IJW03, LH72, Mag81, Mar96a, Mar96b, NL77b, OZ81, Ray85, Wei79, Ter87]. chi- [GR83]. Chi-square [WK02, Bog95, GMR82, GW73, HB78, Hos87, IJW03, LH72, Mar96a, Mar96b, NL77b, Wei79]. chi-squared [Ant95, BR06, El 82, For97, Goo78n, Goo81i, Goo81j, Goo81n, Goo81-29, Goo82o, Goo82l, Goo83k, GS84, pH78, Han79, Mag81, OZ81, Ray85, Ter87]. chi-squareds [GG80]. chi-squares [Fie93]. Choice [Sah02, Arn79, BD84, DP93, Fea79, GZP05, GRVV08, Goo79e, Goo83m, GL86, Har85, Hut93a, IC00, Krz83, SdFdGCM08, TKT89]. choices [NTK09]. Cholesky [Die93, Mai77]. choose [Adk96]. Choosing [BG99, Goo89w, WJ05, GL86, HSW75, Pes80]. Chronic [KHSG83]. circular [NJG88]. Class [NK02, BW93, CT82, HANMA98, JG80, MM80, Möh05, PS98, SW75]. classes [CL91, Goo78f, Goo84u]. classic [Ano06a]. classical [Del83, Mei09, NJJ92]. Classification [Gra88, BG78, CYC99, CG93, DM79, HL05, Hos78, KP82, KPJ09, Lee98a, Lee98b, MS95, NJJ92, NCA^{+}00, NA09, RWL95, STG^{+}01, SMP^{+}06, Var81]. classifiers [TH05]. classifying [Mah96]. Clinical [LJDK02, BS05a, BL78, CI81, Coa95, Don97, Goo81g, HT93, NS09, Sch92, SJ06]. clones [Fan95]. Clopper [PO06]. Closed [DA87, BF83]. closeness [Kru72]. closer [And93]. Closure [KK93, DSS06]. Closure-state [KK93]. cluster [Cai09, CG93, GW01, HBC07, PP06c, Ros95]. clustered [BL09, Bow01, Hin97a, SG96, YM96]. clustering [AT95, BG99, Dod00, Goo79p, Goo86a, GS01, Har75, Mur83, PH03, Sie78]. clusters [Goo82f]. Co [Coo07, WB09]. Co-integration [Coo07, WB09]. Cochran [TNG^{+}06]. Cochrane [Alt92]. code [GS78]. codes [KC97b]. codifference [Ros06]. coefficient [AASAM03, ASY80, Alb92, Auc79, BN01b, BN01a, Che01, DY92, Éri97, Eri83, Goo89f, HMS89, KC02, KKB85, LP09, Liu81, Mag75b, MP81, PL99, Sha85]. coefficients [ASY82, AL99, BCH08, BF08, BG07, CHH91, CMD74, DV86, DR02, DB84b, Gol77, Goo83l, Goo92e, HE00, Lev78c, Loh75, O’G06, OHLH82, PF04, Par92, SM03, SA04a, TG73, VA08, Wat77, WP81, Wes85, WCEC94]. coincidence [TG78]. cointegrated [BCH08, GA09]. cointegration [CV08, Khe08, MS01, SCA07, SP00]. collaboration [Ell00]. Collaborators [Ano80n, Ano93h, Ano97m, Ano73e, Ano75g, Ano77h, Ano78l, Ano80m, Ano81m, Ano82l, Ano84d, Ano86i, Ano87i, Ano89h, Ano89i, Ano90i, Ano92j, Ano94d, Ano95h, Ano96g, Ano97n]. collecting [Goo81s]. collection [Ano06a, MZZ89]. collinear [Sch86]. collinearity [CG91]. colored [J.83]. Colton [IHM78]. column [GH76, REK81]. columns [Goo81d, hL92b]. Combination [HHR02, JS02, BS01a, Goo79g, Goo89a, JG96, JG83b, LGP90, SSW95, Sar98, Sho86]. combinations [Fie93, Gue82, TBG^{+}90, Wan92]. combinatorial [Goo83x, Goo85j]. Combined [Gan89, KT94, SM96]. Combining [Gao04, LM74, MK97b, RB92, Goo84t, JK89, KL05, Lee98a]. Comment [Goo84w, Bak87, Goo79d, Goo79m, Goo79a, Goo81p, Goo83m, Goo84g, Goo85i, Goo86o, GG90, Goo90a, Goo91, Goo94a, Goo94i, Goo95a, Jöc81, Bea85, Goo79h, Goo80e, Goo80g, Goo86b, Pad78]. Comments [Ano02e, Edi04, Goo86s, Goo98b, Goo06a, Goo06b, Kru86a, Lia11, Goo81h, Goo90k, Goo92b, Goo93b, Kem84, Aok91, Goo90o, Hut93b]. common [Bha01, BG07, CM06, EOD86, GR79b, KR89, KL05, Pig91, PB04, RS85, Sha07b, Sha09b, Wil04]. commonly [Dee76, SC95]. commonsense [Goo84c]. communality [RG06]. Communication [Ell00, Dod00, Goo89c]. Comparability [Mye98]. Comparative [AW95, BCP02, DN06, Pet02, Sel08, AR00, Bak07, Bak08, BM90, Bow85, Coa92, DP06, Dog89, Fos95, Hab80, Lee02b, Neu07, PN86, Ram89, SCM90, SA04b, Soh94, SW83, TTNC09, TB85]. compared [Goo83-27, KC73]. Comparing [ÁGR06, LSS93, RR03, SJ07, BDB08, CH06b, DPS01, DC02, DSL06, Goo86n, Hab92, HM85, IHM78, Mar87, MJ93, RL08, RB92, TT86, VA08, ZA09, JR96b, LH78]. Comparison [AMYY07, BG93, BK08, BL09, CG93, CB97, DD78, DF98, FS75, Fur91, HBK92, Hie81, HSW75, Hon90, HZ03, HCW07, JG83a, KC89, Kim83, LC92, Leg00, LL07, jLP05, LD97, LPSL05, MJ08, McS06, NSG91, Par99, SC75, SNB07b, Van84, WK90b, Wil04, AG92, AT95, AL99, Ant95, BCH08, Bay90, BG01a, Cho08, CMD74, CM06, CH90, Con06, CB03, DA87, DY92, DW80, Dor88, Dut77, EOD86, FPRS92, For97, GMSS95, Goo79b, Hab84, HT93, HL97, HL05, Han78, HV93, HT85, Hin97b, HL92a, JH01, JUP86, hJ93, KP09, KT97, KK91, Ker87, KC02, KG80, Krz82, LL79, LN77, Lev78a, LW95, LOR04, LTV90, Mar86a, ML74, MRR84, MYS01, MM05, Mei09, MGG09, MM08, ML79, NL77a, ND84, NJJ92, NL08, NK08, O’G09, OHLH82, O’N82]. comparison [Pap93, PLD88, PBWU78, Pie97, Pos82, Pou04, Pru93, RHH80, RJ95, Rud86, SAD03, SC95, SSM95, Sie78, SB82c, SLL00, SB93, TB88, TA08, TW91, Tse84, Tsi02, Tsu93, VW78, VC78, WW95, War74, WK72, WDB75, WSPC09, YY07]. Comparisons [HM98, ST74, AP85, CFOS02, DDB09, Goo84c, Goo84d, Goo84f, Hal72, HH92, Jen95a, NP98, Ng96, RW96, SC95, SS99, TNG^{+}06, WK75, Kru90a]. compelling [Goo89i]. Compenis [GG77]. competing [Agu99, BK08, NM98, QC03, She09, Tsi02]. competitors [BR03, Goo80d]. complement [NO75]. complementarity [QFG87]. Complete [Fra88, HZ03, Sho95, HC06, Lot82, SFSS85, WBE80]. completely [Hen81, Tre94]. completion [HV05]. complex [DF80, Gup73, GN89, LL79]. complex-valued [DF80]. complexity [CHR03, Goo78b, Wel82, JR96a, JR96b]. complicated [ZW01]. Component [Dor01, Har91, AHADA00, BE86, CT82, CR73, Gen76, GS85, Krz83, LL07, Mar90a, SB88, SL98b, SS92, WW03]. component-of-variance [SB88]. Components [CH99, AM87, AJAH07, BR06, BB84, CH91, CD92a, EG89, FS03, HB93, Hoe89, JMY96, Krz82, Lai82a, LD97, MMR92, Mil79, NK07, OHW97, Paw01, RMH88, TBG^{+}90, WL07a]. Composite [GB86, Che94, Jen95a, KT97, MS98, WK72]. composition [Ana09, Kk90]. composition-search [Kk90]. compound [AJAH07, BG78, EBS86, Var81]. compounds [Goo98a]. Comprehensive [FRS06]. compromise [Goo83-29, Goo86g, Goo90l, jLP05]. compromises [Goo86d]. Computation [Ada91, Ans80, CY99, CZ07, CH88, Den77, HS91, Lai82a, Mar95, Ong95, PSW75, TPG93, BS79, BR84, Che01, Che05, Dha85, DW90, El 82, ES88, Far78, GB99, GO90, Goo81w, Goo81j, GGSNR09, pH78, Han79, Hua91, Kat78, KMK87, Lee92, MM80, MMWM83, TA92, WM95, pha95, Hol85b]. computationai [CW72]. Computational [Bur74, Dyk85, GR79a, KM83, Liu86, RACSGJ05, Sny88, SH72, ZNAEB98, BHLH78, Die94, Dut77, Rid80, TMW73, Wel82, Wil79, YL85, YI01]. Computationally [AC08, Che06, Die93, Zim89, Har04, NCA^{+}00]. Computations [LD87, WK75, WK02, GS73, JCM72, LT90, Mai77]. compute [BPP00]. computed [Cro74, OM88]. Computer [Fam98, KS85, LK83a, Nar90, Pet88, Tay72, CH97, FF84, Goo84i, Goo86g, Li94, LK83b, Lou84, MJ08, ME72, NSG91, NC72, PK72, Pin78, Pot81, SD01a, Sch78, Joh79, Joh82]. computer-induced [Pin78]. computer-intensive [Goo86g]. computers [NO75]. Computing [Dea86, GG75, Har87b, KA85, LGPP96, Mar96a, Mar96b, Mee93, Ruf07, SQ02, SNB07a, Str89, Wan90, AK85, CH91, DL75, FC94, Fre09, Goo87e, JL83, Mal06, Sha09f, Sie78, Zan08]. computional [GA95]. concave [Car07, Ruf07]. Concentration [QH80, Goo87b]. concept [AZ05]. Concepts [Sha05d, Ano06b, Ano06f, Ell00]. concerning [Goo79g, Goo80e, Goo80g, Goo81c, Goo83r, GH83b, Goo84w, Goo85c, Goo85i, Goo86o, Goo87g, Goo90o, Goo90g, GG90, Goo91, Goo92b, Goo93b, Goo94i, Goo95a, KHA04]. Conclusions [Ano02e, Goo06b, Edi04, Goo86s, Goo98b, Goo06a, Kru86a]. concomitant [WAP84]. condition [Del83]. Conditional [CH99, Mas03, TS09, BDG04, Har03, JB91, KK99, MYS01, Par80, SL98b, SP00, TZ97, vBBGOR06]. conditionally [BC08]. conditioned [Atk91, Lil01]. conditioning [LS88]. Conditions [BCP02, Akn07, ELB97, SK80]. conferences [TP06]. Confidence [BF08, Bon05, Bon06, BE86, BBG86, Eti81, HK00, HB93, LJDK02, PB02, SW82, SL88, TW08, TBG^{+}90, Wan08b, AH90, AB09, Alb83, AB97, Ars86, BJ99, BSK90, CT82, Che97, CB03, DV86, DF98, Ede94, Eri83, Far90a, FW88, Fra74, GMSS95, Gho95, GZP05, HMS89, HP90, HMZ05, JK08, JR83, Kar09, Kwo96, LL91, LP89, LL07, LP92, LJZB05, LD97, MMP08, Mar92, Mas03, Mee93, Mun06, OM88, OHW97, PS99, PT89, PT03, PP06c, PO06, RS89, Roy93, SB88, SB73, SM94, kSwXR93, SW83, SW84, SNP93, TG73, TTNC09, Wan92, WW95, WM91, XM09, XDZ09, Yan98, YXW07, Goo87d, Hil87]. configuration [DSL06, Hol85b]. configurations [Sin90a]. confirmation [Goo89h]. conflicting [Kib04]. conformance [PX02]. congruential [KW96, Pes80]. conjecture [Goo83j, Goo90g, Goo94i, SB85]. conjectured [Goo84w, Goo87a]. Conjectures [Ano02e, Goo06b, Edi04, Goo83q, Goo86s, Goo98b, Goo06a, Kru86a, GH83b]. connected [Goo81k, MW04]. connection [Di 05, Goo00]. Conover [AS00, BHLH78]. consequence [Goo83f, KC97b, LEG85]. conservational [Wil75]. conservative [Wil75]. considerations [FF84, GM77, RELW09, SD01a]. consistency [Goo79c, Goo87a, Sha87b, TL96, Joh90b]. Consistent [MK97a, SPK09, CNZ01, CNFO05, CNdGAL09, Fur96, GL96, Jen78, MGG09, SGR04]. Constant [LJZB05, Bén94, GA95, Mar98]. Constrained [PH03, RB00b, Hat86, Lee81, LJZB05, RV98, WECC00, Won95]. constraints [Gbu81, GG75, HAS89, LiT95, MPPZ05, VM00, WNB07]. Constructing [BJ82, CB03, Fou80, Gup84, SW83, ZR93]. Construction [TC73, Eri83, Jöc81, REK81, Tay72, Won95]. contaminated [AS04, Men00, VGD06]. contamination [FL82]. contest [WLT08]. context [Fir97, GA01b, GA02, GA09, GO03]. contigency [Mur78]. Contingency [Aok02, ABH82, AT05, Dey84, GG80, Goo79b, Goo79u, Goo79f, Goo81r, Goo81w, Goo81v, Goo83e, Goo85p, Goo86q, GF89, Hab84, Hos87, Ken79d, Lat82, Law01, Lee85, LPSL05, MH72, MPPZ05, O’N82, PD03, XYT08]. continued [DM05, Goo83u, Goo83t, Goo83y, Goo89p, Goo89n, SB85]. continuity [Mag81]. Continuous [SHH85, WO97, Bic03, BZB08, Cai09, DFY08, SB96, VR92, XDZ09]. Continuous-time [WO97, SB96]. continuum [GS85, Mar90a]. contradict [Goo81b]. contrast [BF08]. contrasts [CB03, GB99]. Contributing [Ken79d]. contributions [Die05]. Control [SS88, BW93, CM07a, DC02, DCA03, GR08, Gan90, Gan91, Gan93a, Gan93b, JH01, JW97, LL03, Mai03, Pol97, Rig95, SN05, SB86, SL09, Shu08, SS99, SR00, WS82, WR98, WR94, WGC04, YR06, Zha00]. controlled [RS77]. controls [DSL06, RRT99]. convenient [Dan80a]. Conventional [PP06c]. Convergence [RRT99, CRR99, JG92, NC96, Shi07, YKB86]. convergent [Cro74]. converging [Kat78]. conversational [War74]. convex [AGNS91, Gle89b, Gle91, Har87b, LLKJ09, MZZ89]. convolution [FB90, XPC03]. copulas [McN08]. core [Lot82]. corpuscle [Wil89]. correct [Goo88f, Ven89]. Corrected [BCV98, Cor04, SBC03, VG05, Chu01, CDDCN97, Dog89, GRVV08, LS01, Wil75, YAAB87]. Correcting [PP03, Roy89, Tho89]. Correction [Chi07, Goo83r, Tho06, Fos95, Goo85c, Goo87g, Kar09, Mag81, Mar90b, MB90, MCCC04, Pau92, Sar98]. corrections [CCNAF95, FCN99, SC09b]. Correlated [Van84, YM99, AL01, Ana09, BP01, BRL82, Bho84, BHG01, DN94, Dow02, FK04, GR83, HM98, KL93, KL05, Kul90, LB96, PS98, SCW07, SB92, Soh00, Spi98, TTF07, TWLC07, VA08, Zur93]. correlation [AASAM03, ASY80, ASY82, Alb92, And89, And90b, And95c, AJ82, Auc79, BN01b, BN01a, BGR94, BF08, Bow01, Cha75a, Che01, CMD74, CH82, Dag89, Dag95, DY92, DS89, DM91, FMOR06, Gol77, Goo78h, Goo79q, Goo92d, Goo92e, HMS89, Jen93b, Joh95, JW80, KCS88, KKB85, KT94, LL88, Leg00, Lev78c, LH94, Liu08, Mag75b, MP81, Mai77, MH72, MB97, MJ93, NC06, Owe81, PL99, PF04, PSB03, PPRW06, PB04, RS92, Rya80, SdFdGCM08, TBT95, WP81, ZQ97, Zim04, Goo83d, Goo87f]. correlational [PNM83]. correlations [Dem06, Dog93, Goo81e, Lee85, MHS04, Son97]. correspondence [Kim92]. Corrigendum [Bal92]. corroboration [Goo89g]. Corss [Lee90]. Corss-validation [Lee90]. cost [Goo81s, PQ89, Sha08a, TC75]. costs [BC82]. cotangents [Goo93d]. could [Kun93]. count [CYC99, Dre08, Joh95, Kim07, NK08, SCW07, SJ06, VG01]. counts [Fra79]. course [Goo95d]. Covariance [SD02, AC08, BD82, BDFR97, BBHW95, CP91, CCNA09, Che85, CCSG81, CYB90, CH82, CNZ01, CNFO05, Die93, DN94, Die94, FH86, Fur96, GG80, Gup73, HK92, JS02, Lee81, hL92b, LS01, NCÁHC^{+}09, PGT09, RWD95, RML90, SE90, SA79, Sim93, SN95, Tre95, Wil75, Zim89]. covariances [Ada91, Bur74, Nag75, Vau94]. covariate [ZR07]. covariates [Kar90, Kim07, LWH08, SCW07, SL93, Soh96]. covariation [Gal02]. Coverage [WM91, YM90, CS95a, CY99, Goo82m, MY90, STS94, YL85]. coverage-probabilities [STS94]. covert [Goo86d]. Cox [FGSV09, HT93, KAST05, MCCC04, Spi82]. CPIT [QD83]. Crabbe [Aok91]. craps [AE87]. credibility [Goo90l]. credible [ZR93]. credit [DS09, KPJ09]. Cressie [Rud86]. criminals [Goo78i]. criteria [BHK05, Chi79, HHI90, HT85, KPS01, Kun92, LG09, jLP05, NCA^{+}00, SD92, TJ96]. criterion [Bak87, BR00, CH90, GC03, Hol85b, KE03, MS99, Nag75, OPS82, PGT09, Sy01b, Wu01]. Critical [ASY81, BHG01, Fun88, JG83b, Auc79, Bru75, Chi07, Hua01, RR93a, SL98b]. Cross [BDFR97, Krz83, Whi94, CYC99, GOS09, Gri92, Lat82, LH94, LP09, SNC09, SJF06, Tay90, TKT89]. cross-classification [CYC99]. cross-ratio [Lat82]. cross-section [GOS09]. cross-sectional [SNC09, SJF06]. Cross-validation [BDFR97, LP09, Tay90]. Cross-validatory [Krz83, TKT89]. crossed [HB93]. crossing [LD87]. crossings [JW97, WR94]. cryptanalysis [Goo00]. cube [Lev78b]. cube-root [Lev78b]. cubic [Chu01, DB84a, WmGT95]. cumulants [Far90b, Goo79s, Goo82o, GH83b, Goo83q, Goo83x, Goo83o, SZ02]. Cumulative [GR83, Nue08, Gan89, OP00a, Par81, Par87, SB82c, Zan08]. curiosity [Goo85j]. current [Goo83-29, Mur78, RB88]. curvature [HAS89]. Curve [HM95, OS02, FPRS92, Ham77, Han86, HHKD02, LW82, LC92, Par99, Pru93, QL01, SFSS85, SND89, SA95, Soh92]. curves [Kap87, ME72, Tue01]. Customizing [Fel02]. Cuthbert [Goo79d]. cycles [TC73]. Cyclic [Har93, GA01a, RH76, Tay72]. cyclical [GA09]. cylindrical [AC00].

daily [Tsi02]. damage [OP00a]. dampings [Kah93]. Daniel [Goo79d, Goo83f]. Darling [TO04]. Data [Bog95, Bog01, Coa92, GLC99, IJW03, KT97, Pet02, Sha06b, SD02, Van84, AEL03, AL93, AS89, AC00, AB08, Ano06a, ACNT05, BF87, BP01, BRL82, BG01a, BE94, BS05b, BGR94, BR06, Bho73a, Bho84, Bic03, Bow01, CC01, CYC99, CP95, CK04, CL91, CK01, CZ07, CH97, CMQ03, CC85, CC97, DE06, DAM98, Dem06, DRYL08, Dow02, DL92, EB90, EM86, FPRS92, Gan84, Gao04, GD92, Gen76, GW01, GOS09, Goo82d, Goo83e, Goo90n, Gri02, GL96, Ham77, HT93, Han86, Har04, HV05, HK07, HBK92, HM85, Hin97a, HBC07, HSSY04, HCW07, HTC07, JB91, JR09, Joh95, JG92, Jön07, Jun08, Kar09, Kar90, KK93, Kel94, KK05, Kim07, KM99, KP96b, KPS01, Kun93, KY93, Lam05, LY99, LFC92, LGP90, LF97, LW82, LC92, LN77, LEG85]. data [LB96, LD97, Mar83, MW94, MS95, MK08, MGG78, Mou95, Mou01, MP96, Mye98, NL77a, NCA^{+}00, NK08, NBB00, O’G09, OHLH82, OM88, PL99, Pan98, Pan99b, PK72, PSB03, PH03, PPRW06, PY99, Phi91, PF93, Pou04, PB04, PV93, QC03, QY95, RM02a, RFGE86, RK05, Rod07, SB08, SG96, Sch86, STG^{+}01, SFSS85, SND89, SCW07, Sha08a, Sha09d, SB92, She09, SFC08, Sie78, SP01a, SC82, SSM93, SJ06, Sy01a, TAY02, TTF07, TB85, TWLC07, TJKB00, TS04, VP89, VG01, WmGT95, WC91, Wet96, WB94, WZ95, Wu01, YM90, YM96, YH85, YW09, ZS05, ZZCS09, vdTGL97]. data- [KP96b]. Data-dependent [Coa92, BR06]. data-oriented [Wu01]. data-sets [PL99]. date [DL80]. dealing [KJ09]. decision [BAB86, GS07, Hor97, IMP^{+}97, Kri77]. Decisions [KR02]. decomposable [MW04]. Decomposition [NS86, Kim00, Mai77, RS90a]. decompositions [Goo80l]. deconvolution [CRT07, KP96a, LT90]. defeat [Goo95c]. defense [Goo85n]. define [FS94]. definite [AS01, BD82, QFG87]. definitions [O’N82]. degradation [PY99]. Degree [WB02, FC96]. Degrees [LPZ02, Goo85c, Han79, VG05, Goo80f]. Degrees-Of-Freedom [LPZ02]. delay [DBVK02]. delayed [HSW75]. delta [Her75]. demand [Ber73]. demarcation [Goo89g]. demons [KW94]. dendrograms [FS81, Mur83]. densities [DD77, FT96, Gne97, Goo78d, Goo82c, NP09]. Density [Bow92, Dia98, Fok07, GW73, AEL03, Ahm87, Bow85, CJ95, CGN04, CRT07, CS05, Dog89, Ede94, Edi83, EB90, Fos95, Goo79r, Goo81q, Goo82d, Goo84n, Goo85k, GHDB89, Goo95e, HP00, Har85, HK07, IC00, KP96b, KPS01, LT90, MW04, NP98, Ruf07, SJEs93, Smi94, TPG93, WV79, Weg72, WL07a, Zhe88, Goo83g]. Department [Arn00]. departure [Llo93]. departures [CYB90, LL08b]. dependence [Ebr93, Fri07, GOS09, Goo85p, GJJLGS06]. dependent [BR06, Coa92, DV86, EH07, Eri83, Eti81, Gle91, JS02, Lev78c, MH07, PF04, SL87, SL93, Son97, Van05, WHF80, JH90b]. depends [Goo85q, Goo95d]. deposition [How85]. depth [Koh81]. derformination [Don97]. derivation [Cha94, McK86]. derivatives [AK85, KA85]. derive [SC09b]. derived [CW74, Goo90h]. descent [PK72]. descriptions [Goo81m]. Design [GM77, RELW09, Sha07h, AS00, BE86, BBG86, Gan93b, Goo84c, HMK00, HB93, Jai85, JR96a, JR96b, LZ92, LL03, LB07, May06, RW96, Too72, YR06]. designed [MP96]. Designing [BB99]. Designs [PT02, WGC04, Ang03, BS78, CG94, DK05, EGM^{+}04, FS03, Goo90f, Gup84, Hun02, Jen95a, JR96b, JH72, KW04, MS98, May99, Mih74, PM02, Pol94, RJ95, REK81, Sin90a, SK90b, Sin90c, SK90a, Sin90b, SB91, SC00, Tay72, TC73, VB83, Wil75, WS04, Won95]. desk [Gue78]. details [GHDB89]. Detecting [Kun92, PT02, WK06, AT95, CM06, DM91]. Detection [CK04, DE06, IM82, LH93, MH99, Sim87, WGP00, Wil08]. detectors [Goo90c]. deterioration [GR08]. determinant [Goo84w, ZL07]. determinants [Ano81f, Goo81y, GL83]. determination [BL78, Éri97, Goo81l, Goo83r, Goo89f, Jen03, KJ81, Sha85]. determinations [SJR07]. determine [EH92, Tue01]. determined [CM76, GsW01]. Determining [BAB86, MMR92, Ram89, Auc79, GJ82]. determinism [DM91]. detrending [Coo07]. developed [SW88]. Developing [Cor00, Goo89w]. Development [KK85]. developmental [AB05]. Deviance [OPB08, Cor95]. deviates [Fam98, RJ06]. deviation [AKW92, Bal89, Bal92, Bon05, CR93, FKW80, Rhi86, RS09]. deviation-based [RS09]. device [FBC09, Goo83-30]. devices [FBC09, HK84]. diagnosing [Goo89c]. diagnosis [Goo84x, GHDB89]. Diagnostic [VBS07, KL09, RL08, SCK01]. Diagnostics [CHM09, XW07, XW09, CRR99, LL09b, RK05, SP01b, XYT08, You93]. dichotomous [EP92, PW86]. dichotomy [KSH73]. Dickey [CL98, DG95, Goo84r]. dictionaries [Ano06c]. Difference [LFC92, PB02, Auc86, Bak07, BCL93, Bho84, CM07a, Gue78, HB78, Pap80, SND89, SM94, TA08]. differenced [Rod07]. differences [BS78, MF02, RR77, SB88, WA72]. differencing [TH09]. different [CNFO05, Gan84, GK01, Lac92, PSB03, TNG^{+}06]. differential [Goo92e, MS99, Nic05, SFC08]. diffuse [Sny88]. Diffusion [Mol99, Dem90, GGSNR09]. Diffusion-limited [Mol99, Dem90]. digamma [Dev92]. Diggle [HC06]. digital [GLP72, Sch78]. digits [Goo88f]. digraph [AA09b]. dimension [AT95, CY92, DM91, PS09, SMO03]. Dimensional [Dög01, Goo89p, JSS09, LZNL08, Spi06]. dimensionality [BVP90, Goo90d]. dimensions [DL75, Goo83u, Goo86q, Goo89n]. dimethylsulfide [CG97]. diminishing [Goo80h, Goo83b]. Diophantine [Goo84m]. direct [AG78, Dor88, SS97a, Tut90]. directed [LP00]. Dirichlet [Goo81v, Goo83-28, Goo84a, Goo86p, HV93, IMLG09, Nar90, Ron89, Wil98]. dirty [Goo84n]. Disaggregated [Van84]. disaggregation [SB96]. discharges [Dog93]. discovered [Goo85g, Goo86k]. discrepancy [BDFR97, GL85, PX97]. Discrete [WV79, Ait83, BHLH78, BP92, Fro95, GAM09, Goo83b, JG80, LK83b, Mel80, NK08, NC72, PX05, RACSGJ05, Roh88, SHW93]. discretisation [pha95]. discretized [PF93]. Discriminant [GLC99, McL80, SHH85, Sch75, Amo85, BBP04, Bel93a, Bel93b, BG99, BVP90, Dor01, GM79, GK90, Goo82c, GL90, Hon90, JMY96, KP09, KL93, LSL97, Lee98b, LW95, MS95, MYS01, MR03, Mou05, NJJ92, OHLH82, PLD88, RFGE86, SLL00, Whi94]. discriminate [LDC73]. Discriminating [GK04]. Discrimination [MP96, KP79, RHH80]. discriminiation [Goo79c]. disparity [BS94a, JS00]. dispersed [CYC99]. Dispersion [PT02, BCX93, BRL82, Dor08, FPRS92, FS94, Lem77, LL08b, MS98, May99, May06, Pie97, Son97, TAY02, WA72]. disproportionate [LN77]. Distance [BCP02, BDFR97, CM98, EW91, Goo90k, GM06, Pak99]. distances [Goo85b]. distinct [Goo80j]. distinguishability [WK90a]. Distorted [SD02]. distortion [CM04, CM05b]. distributed [Aza73, CJ73, HK84, Mak00, SS86, YT96]. Distribution [AB09, BS02, CNV02, Dod83, Fam99, Fra88, HPY79, KPKPB95, LC02, PB02, RG02, Sho95, WB02, WK02, AA09a, AL93, AL94, Ada91, Ada96, Ada97, ABE83, ABE85, AG92, ASY80, ASY82, AS90, AS01, AFGP06, AB97, Arc80, AR00, Bal85, BP86, Bal89, Bal92, BS95, BN01b, BN01a, BL03, BKLW04, BH07, BRF08, BG94, Bal95, Bj78, BB74, BR06, BRT07, Bho73a, BS01a, BD09, BS92, Cha75b, CH76, CL91, CS94, CJ95, Cha79, Che78, Che82, CK96, Che97, Che05, CC09, CB97, Chi08, Cho08, CKM01, CH88, CDDCN97, CAT78, Dag78, Dea86, DW80, Dog92, Dup96b, Dup96a, DW01, Dut99, DC99, EH92, El 82, Éri97, ESRV98, Eve88, Fle95, FS75, Fro04, Gan90, GG77, Gat09, GAM09, Goo78k, Goo79g, Goo80b]. distribution [Goo81z, Goo81-29, Goo82l, GH83b, Goo83q, GH83a, Goo83x, Goo83k, Goo84a, Goo85r, Goo85s, GS86, Goo86a, Goo90j, GC90, Goo96a, GR83, GV81, GW73, Gup73, GJ77, GN89, GK01, GJJL02, GJJLGS06, HB78, Har87b, HZ96, HSB85, HK07, Hol85b, How85, IK03, J.83, Jen82a, Joh79, Joh82, Jön07, KC73, KM06, KC96, KC02, KH09, KKB85, KH04, KN89, KC97a, KJ81, KR89, KL05, Kul90, KS04, Kwo95, LH97, LM84, Lat82, LL91, LH72, LFC92, Law92, LB08a, Lee99, Lee01, LSSP08, Lem77, LSCN08, LP89, LZNL08, LB01, LP92, LP85, Luc08, LJ00, LTV90, Mal06, Mar96a, Mar96b, MZ77, MK74, Mas03, Mcd74, MK08, Mol79, NK07, Nag75, NS89, NL77b, NW09, Nor84, NC72, Nue08, OZ81, Ong95, Pad82, PSW75, PF04, PPK77, PRS87, PGT09]. distribution [Pie94a, PT99, PB00, Pos94, PN98, PW83, QQ82, RG93a, RA01, RM05, RR93b, RR77, SQ02, SB08, SBC03, SY08, SF93, SWK73, SKC75, SA92, SL98a, kSB90, Sho86, SCM90, SA04b, SS92, SDP93, Sny88, Soh96, SB82c, SW82, SW83, SW84, SI72, SVW88, Tad81, TAY02, TC94, Tsa06, Tsu93, TL96, Tza09, UP01, Van93, Van05, VM00, Vau94, Wan08b, Wan08a, Wes72, Wil75, Win75, WA72, WL90, Woz94, Wu03, WL07b, WCC07, WWL09, WBE80, WS90, XDZ09, Yan99, YL01, YXW07, Zan79, ZK86, Zan08, ZW01, ZY04]. Distribution-Free [PB02, AB09, Lem77, Pie94a, RR77, WA72]. Distributions [RB00b, Ahm88, AMAE97, Ahm92, Ait83, Amo85, BJM92, BG93, BL02, BHLH78, Bar79, Bar81, Beh72, Beh73, Bha01, Bho73b, Bon06, BP92, BJ82, BT00, Cha94, Che94, CM76, Cro74, CM98, CM01, Dem06, Dev92, DH77, DF98, EG89, FB90, Fea79, Fou80, FJT82, FS81, Fro89, Fro95, Fro04, GK00, GH76, Gol77, Goo78g, Goo80j, Goo81e, Goo83z, Goo87e, Gue78, Gue82, GT78, GJ79, GC03, GK04, pH78, Han79, HV05, HANMA98, HM80, HD77, Hos78, HL09, Hut77, IJP86, IMLG09, Jah05, JG80, JCKS09, Jen76, Jöc81, Jon86, JAK93, KJ09, KP79, KW94, Lak81, LS99, LK83b, Ma97, Ma99, MF02, Mat79, MB97, MY90, MHS04, Möh05, MM99, MR03, Nie06, NK08, Ouy06, PT03, PS98, PP06a, Pat76, RP01, RR82, RACSGJ05, Ron89, RW93, ST74, SB88]. distributions [SB73, Sha06c, SHW93, SN83, SC09b, Str89, SH72, Tad81, Tan82, Tse84, Tue01, VW78, VR92, WK03, Wei79, Whe75, Zac80, ZW01, lS96]. disturbance [BH85]. Disturbances [Van84, AAR93, Bar77, DM93, OR92, Poi92, SP00]. divariate [Pie97]. Divergence [MMP08, Ant95, FP07, Goo84p, MPPZ05, PD03, PB03]. Divergence-based [MMP08]. divergent [Goo90b]. diversities [LH78]. diversity [GS83, HL84]. divisible [PS98, Sin90b, WK03]. DNA [ÁGR06]. do [FMC09, Koh81, WC91, WG92]. does [Dam82, Goo81b, Dav79, Goo77, Kru03]. dogs [Goo98a]. domain [AGNS91, But99, Hal82]. dominance [NP98, SP86, TZ04]. done [Cor00]. Donoho [BD01]. dose [ANPV97, HHKD02, Hun02, JCS07, Kap87, LZNL08, RS97]. dose-response [ANPV97, HHKD02, Hun02, Kap87]. doses [HHKD02]. Double [BC08, CH76, Kuk89, GV81, KL09]. Doubling [AE87]. Doubling-up [AE87]. doubly [Ada96, Ada97, LB01, RM02a, ZZCS09]. down [JG83a, Ker87]. Downton [BN01b, SL98a]. dragged [Ano78g]. draw [Li94]. drift [Car07, DC99, Gan91]. driven [Bog95, Bog01, IJW03, KT97]. dropped [Goo81f]. drug [SSW95, Sar98]. dual [Gne97, RB00a, Goo95e]. Duality [CCHM08]. duplicate [SK90a]. duplicated [Goo83i]. duration [CP76, TBT95, Van87]. Durbin [AAL02, DS89, GL92, SA92, Smi78]. dynamic [Goo87i, Kim93, MB86, MB90, OR92, Owe81, RRCD97, Van87, Goo81b]. dynamics [KT85].

each [JCS07, MTO08]. earlier [Goo90i]. earliest [Goo86m]. early [Bar84, GS83, Goo84o, Goo84q, Goo88d]. earnings [GB86]. Easy [KH04]. EBLUP [GMLM^{+}08]. Eccentricities [Mag75a]. econometric [BH85, Mae87, Owe77]. econometrics [Sha09e]. economic [Goo87b, LL03]. economical [YR06]. ED [Kap87]. ED100 [PP06b]. ED50 [Ker87]. Eddington [Goo98a]. EDF [Fam99]. edge [Dog89]. edge-corrected [Dog89]. Edgeworth [KKB85, PP80]. edited [Sha03e]. editor [Hog73, Ano06g, Goo88g]. Editorial [Ano73e, Ano75g, Ano77f, Ano77g, Ano77h, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano80m, Ano80n, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano82e, Ano82f, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano82l, Ano83a, Ano84a, Ano84b, Ano84c, Ano84d, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano89i, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f]. Editorial [Ano90g, Ano90h, Ano90i, Ano91a, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano92j, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94a, Ano94b, Ano94c, Ano94d, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano97m, Ano97n, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano01a, Ano01b]. Editorial [Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano01m, Ano01n, Ano01o, Goo89w, SS97b, Goo90a, Goo84g, Goo86s, Goo90g]. Eeaviside [HG85]. Effect [Cha75b, CW74, Edw95, AkBA05, AL02, BS94a, CR93, EA78, Fur08, Gbu81, Goo81d, Goo84s, Goo86i, Goo94d, GSS87, HBC07, KSH73, MJ08, MW92, MH99, Rhi86, SS75, SP00, Too72]. Effective [LL91, Gue89, HHKD02, JCS07, LZNL08, MTO08, WL07a]. effectiveness [Cla96, Sha08a]. Effects [Bow01, GLP72, KAST05, KKB85, PT02, SD02, YM99, BD96, CH88, HK00, HZ96, Hud83, JO85, JSS09, KE03, Kru88, Kru89a, Kun93, LZ92, LC92, Lev78b, LL08b, MS98, May99, May06, Mun06, PW86, RR01, RVZRZP08, Sha87b, SC82, SCB07, Van87, WW03, WB94]. efficacy [Ano03b, DM91]. efficiencies [BE94, CH02, Hab92, Mar81, PR84, SSSB00, SNB07b, TMW73]. Efficiency [AD01, Van84, BS94a, BT09, BK96, GJJL02, Hal82, JR96a, KS87a, KM83, KA93, Liu86, MS95, McL80, Mou05, Won95, Jen95a]. Efficient [BS79, Cla86, El 82, Gil07, GJJLGS06, Hua91, RJ06, YX03, ZA09, BLP09, Bic03, BG07, Che06, Die93, Dio81, GO90, GSF78, Liu08, MJ08, MH72, Nar96, TTF07, Wil79]. Efron [McL80]. eigen [Cha75a]. eigen-structure [Cha75a]. electro [FBC09]. electro-explosive [FBC09]. electron [Goo89v]. elementary [GJH90]. elements [KP99]. Elimination [Pin78]. ellipsoid [AH90, CH02]. ellipsoidal [Mag75a]. elliptical [JCKS09, LL88]. embedded [KS91]. Empirical [AM87, Alb84, AS04, Bal83, BRL82, BHK05, CFOS02, KP09, KY93, LC02, Lai82b, Mou95, Tsu93, Adi98, AL99, CP76, CZ07, CM98, DH81b, ES86, FJ83, Goo00, HL05, HKT04, Hos87, JL83, KC97a, KM99, Lev78a, Lev78b, Lev78d, ML74, MK74, Mei08, MY90, MMW91, NP98, PSB03, PRS87, Pie97, Sco02, SK80, SI72, Tse84, WV79, WT06, WS00, pha95]. EMS [MW94]. encounter [BM92]. encountered [SC95]. Encyclopedia [Sha09g, Sha09h, Sha09i, Sha03c]. End [DBVK02]. End-to-end [DBVK02]. endpoints [Don97, TAM^{+}03]. energy [AZ05, How85, RRCD97]. enforce [AK86a]. enhanced [CS94]. Enigma [Goo00]. enlightened [HS86b]. ensemble [HL05]. entropy [CH06a, Cho08, LJ00, PP03]. entropy-based [LJ00]. Entry [Kel94]. enumeration [Hen95]. environmental [DE06]. Environmetrics [Sha03c]. enzyme [GHDB89]. epidemiology [Sha03a]. Equal [WNB07, ASY82, BBG86, LH97, LH79, Ng96]. Equal-precision [WNB07]. equality [Bai89, BM90, Bha01, CB07, Hal72, Hay97, HH92, IP86, JS02, KAS96, LL09a, Lev78c, Loh75, Mar98, Mar76, Mat79, Nag75, Pan99a, Pie97, RML90, RA81, SHW93, SN83]. equation [BC94, BD84, CG91, DB84a, EP92, Lye91, MB86, MB90, Rig95, SB92, TL07a, XPC03]. equations [Fir97, Goo84m, HS77, HAS89, HL92c, MS99, MH07, Nic05, OR92, PYC93, Sha03e, SE02, VBS07]. equiprobable [GG77, GH83a, GS84]. equivalence [Don97, TA08, WS82, Wil07]. Equivalent [LL08a]. Ergodicity [Cre89]. Erlang [FS75]. Errata [Ano72b, Ano73f, Ano75h, Ano81n, Ano81o, Ano81p, Ano82m, Ano83b, Ano83c, Ano84e, Ano85i, Ano85j, Ano87j, Ano87k, Ano87l, Ano88f, Ano89j, Goo81-28, Ano90a]. Erratum [Ano79l, Ano88g, Ano03b, Goo89o, Goo89n]. Error [GK90, KMK87, McL80, Sch02, XPC03, AR01, AKW92, Bel93a, CHH91, Edw95, FMC09, GMR82, GMLM^{+}08, Goo79n, Goo81p, Goo81c, Goo82m, Goo84r, Goo87c, Han86, HT99, Hon90, Jam01, JO85, KP09, Ken79a, KAST05, KL93, Kru88, Kru89a, KPS01, LMB08, LSL97, LKKL07, LAR09, MYS01, MM05, Oht98, PLD88, PB93, SGR04, SLL00, SOU04, SP86, Tho06, TL07b, TC08, VC78, ZL96, pha95]. Error-free [KMK87]. Error-rate [GK90, Bel93a, MYS01]. Errors [HMS02, VB81, AAL02, AD01, AD03, Ana09, ANPV97, CH97, CVS98, DR94, DB84b, EH01, EH07, FMC09, Gol72, Goo80c, Goo83j, GT85, Goo90m, Jen93a, JL83, Kim07, KJ09, Kul90, Loh75, Mak00, Mar87, McG89, MB86, MB90, MK97b, Pin78, SS86, SPK09, Sha07b, Sha09b, Sha87b, Sma05, Soh92, SNP93, VB83, WHF80, WW07, XW09, Ano03a]. errors-in-variables [WW07]. establishing [Don97]. estimate [GR79a, Goo84b, Lai82b, RWD95, RR93b, SL87, Tay94]. estimated [BN95, BD82, CY92, FW88, GM79, Goo82m, KK85, Kel94, LS99, Owe81, WL90, Woz94]. Estimates [Mag75b, Nar96, YW02, Atk92, BB96, CRM06, Cad94, CH91, Cao87, CVS98, CM76, DP06, Dic78, Die06, Dup96b, Dup96a, EP92, EN90, Far78, Goo78a, Goo79b, HN80, HM80, JAK93, KK05, Lai82b, LSS93, LL79, LM74, Lil01, LK83a, LT90, MW84, MP81, MZ77, Mar95, Men00, MM99, OHLH82, PP03, PB00, RS89, RS86, Tut90, VG05, Wan90, WLK06]. Estimating [ANAA97, ASS04, AkBA05, BJM92, BGR94, BG07, Dan87, Dup96b, Dup96a, Ebr93, Fra79, Ham77, KL93, Kun98, LLKJ09, LL79, Llo05, MS99, MKDM94, PB01, RRP06, Rid80, RB88, RB00b, RM02b, SS86, SCAH05, SD01b, Van93, Van87, Ven89, BLP09, BH82, EP92, EB90, FS03, Fro95, Gen76, Goo78g, Goo78d, Goo81z, GA95, HT93, HV93, Kah93, LH97, LL91, MOS94, MRR84, McL78, MK97a, MH07, Pet80, Rhi86, Sha87a, SB92, SL98b, VBS07, WW06, ZL09, Har87a]. Estimation [AS90, Amo85, Bar77, BS02, CA89, CH99, Cob89, CS97, CNV02, Dam82, DRYL08, DR94, DB84b, FS94, GAM09, GV02, GR79b, HR07, HV05, HAS89, IJP86, Kap81, KS87a, KH09, KP87b, KC97a, KR89, LL88, Liu81, LP85, McL80, OBW05, Pet88, Pol94, Pou04, QC03, RP01, RA01, RS83, Sah02, SN05, SMO03, Sha01b, SKJ86, SP86, SCB07, Tah90, TAY02, VS91, WHF80, Wu03, Abd89, AA09a, AEL03, AL93, AL94, AM87, Ahm88, AMAE97, Ahm92, Ahm99, Ait83, AHJ92, AC08, Alb84, Alb85, Alb87, Alb92, AHH07, AD01, AGNS91, ABH82, AKW92, Atk91, Bal89, Bal92, BN01b, BN01a, BL09, BK82, Bj78, BDG04, Bel93a, BW01a, BF08, BS78, Bic03, BS92, BB84, But99, CDF05, CC05, CD92a, CZ02, CS95a, CLLN04, CM07a]. estimation [CMQ03, Cho98, CR73, CC85, CDDCN97, CNZ01, CMH04, DM78, Dey84, DP93, Dia98, Die94, Din08, DW80, Dog92, DV95, Ebb73, Edi83, EBA00, EL75, ESRV98, EW91, FPRS92, FKW80, FL82, FC89, Fon90, Fon92, Fri07, Fri79, FTS09, Fur08, GK90, GHCJ09, GPW80, GD92, GO03, Goo79i, Goo79g, Goo79o, Goo81q, Goo82d, Goo82e, Goo83e, Goo84n, Goo85k, GL85, GL86, Goo86o, Goo87h, GL88, GHDB89, Gri04, GJJLGS06, HT83, Han94, Har85, Hie81, Hoe89, Hon90, HH96, HZ03, How88, HM98, HHKD02, HMZ05, HL92c, HL09, IC00, Jam01, Jon75, Jos01, KKE07, Kap83, Kap87, KX03, Kar09, Kim93, KPH05, KS87b, KH04, KP87a, KJ81, Kru90b, KW00, KP96b, KPS01, Kuk89, KT94, LY96, LY99, LSL97, LB08a, Lee81, Lee85, LL99, LL90, LOR04, Luc08, MZ08]. estimation [Mar87, Mas03, MC91, McN87, Mil79, MR09, NCÁHC^{+}09, NP98, NL08, NW09, Nic96, Nie06, PS99, PT89, Pak99, PL99, PHO05, Paw01, Pep93, PP06c, PNW06, PW83, QL01, Qu06, RG06, RWD95, RM96, Rol01, Ron89, RV98, SJEs93, San89, SY08, Sco02, Sen02, SGR04, SPK09, STH09, She09, SAC06, Shi85, SA04a, Sim87, Sma05, Soh96, SB82c, SW90, Spi98, SOU04, SVW88, TWLC07, Tza09, VD08, VC78, VS93, VB81, WV79, WK03, Weg72, WO97, WS82, WW03, Win75, Wu01, WS90, YL01, YX03, YT96, Zac80, ZK86, ZT01, ZNAEB98]. estimations [GK01]. Estimator [OVL02, AASAM03, Agu99, AN03, AS81, And80, AG78, BD01, BS72, CLP93, CNZ01, Del83, EOD86, Fur96, GP07, Kru86a, Lem80, jLP05, MGN99, MK74, MPPZ05, MB90, PD03, Sha87b, SMH97, SS92, Smi89d, TA92, Tsu93, Wan99, WZ95, You08, YW09]. Estimators [ASY80, KB96, Kuk87, SS95, Van84, Abd95, AG92, AFGP06, AAR93, ANPV97, BP86, BR92, BBP04, BE94, Bow85, BS05c, Brä92, BS01b, BG07, CC01, CHáS03, CHH91, CR03, Cla86, CC90a, CFOS02, CH90, CKT89, CNFO05, CNdGAL09, CH02, Dag89, Dag95, DN06, Dio81, Dog89, DM93, EOD86, FH86, FP07, Fos95, FRB06, Fur91, GMSS95, GRVV08, GV81, GL96, GJJL02, Hal82, HL97, HP00, Han78, HSB85, HBK92, HCA96, Hua91, IC80, JUP86, hJ93, KP09, KK91, Ker87, Kib04, KC02, Kim83, KL05, KTM05, LSCN08, LM74, LB08b, LS01, LF82, jLP05, LTV90, Lye91, Mae87, MH72, ML74, MM80, McG89, MYS01, ML79, MB86, Mou95, NS89, Oga07, OHN93, OR92, Oht98, Owe77, Par99, PP06b, Par80, PSB03, PLD88, Pig91, PR84, PN98]. estimators [Pru93, Ruf07, Sch86, STS94, SNB07a, SK80, She83, ST88, SCM90, SA04b, Soh94, TMW73, Tay90, Tse84, VS94, YAAB87, Zan79, ZL96]. etc [Goo83-28]. Ethical [Goo78e]. Euclidean [SBO81]. evaluate [Con95, MJ08]. evaluating [Cla96, SL88, SGH75]. Evaluation [GA00, Gue78, Kwo95, Par81, Par87, RW93, Zhe88, CSC00, CAT78, CNFO05, DD77, DN06, Dha85, DCA03, FW88, Fis73a, GK00, Gle89a, Gol77, Goo89v, Goo94a, GS78, HN80, JG96, KY93, MTO08, Nor84, SCW79, SCW81, SCM90, SA04b, TC92, TAM^{+}03, MM93]. evaluations [Goo89q]. evanescent [SB96]. even [Goo98a, Han79]. event [Goo81x, SCW79, SCW81, Soh94]. events [BD96, BR81, DM05]. Evidence [Tsi02, Goo80c, Goo81c, Goo81x, Goo83o, Goo83s, Goo84h, Goo84x, Goo85l, Goo89a, Goo89i, Goo89b, JK89, Joh94, MB90, Poi92, RR82]. evolution [Bar72]. Evolutionary [Nie06, Ang03]. Ewma [Gan91, AL02, CG94, Gan93b]. ’exact [Goo84o, Aok02, BL02, BLC04, BL05, BRF08, BD82, Che97, Edw85, Eve01, FB90, Fra88, Gat91, GH76, IP86, Jen03, LB01, SKC75, SMV76, TNG^{+}06, Ant95, AT05, Ars86, CLLN04, Den77, For97, Fou80, FS75, Gol77, Goo81-29, Goo82l, GH83a, Goo83k, Goo85r, Goo90j, HN80, HANMA98, Hut77, Jöc81, KK99, KB96, Lam05, Lat82, Lee99, LPSL05, Mal06, Mat79, MY90, Nag75, PGT09, Roy93, Sha87a, SGW94, TA08, TC92, TA92, Vau94, WWL09]. examination [LM74]. examinations [Arn79, Goo79e, Goo83m]. examining [Wes85]. example [And92, And93, Edw89, Goo87i, Goo89c, Har87a, HJ85]. example-multiple [HJ85]. examples [Ell00, PL99]. exceed [Goo77]. Excellent [Ano06c, Sha06c]. Excess [PNW06, Goo98a]. exchange [Tsi02]. exchangeability [Alb87]. exchangeable [Alb85, DAM98]. excluded [Goo85n]. excluding [SNB07a, SNB07b]. exemplary [AL96]. exhaustive [KW96]. existence [Goo94g]. existing [BD96]. exit [Kel94]. exogeneous [Mae87]. expansion [Day87, Liu08, Oga07, SB82a]. expansions [Bar81, Kat78, Nab83, OP00b, PP80, PP85]. expectations [BG93, Law92]. expected [Goo83n, Hab92, Har03, PLD88, Vau94, Wan92]. expenditure [Joh78]. Experiment [SB82b, Goo86m, LN77, NL77a, PN86]. Experimental [RR82, VW78, CCD96, Hun02, JR96a, JR96b, KKE07, MTO08, May06, Mih74, VB83]. experiments [BB99, CL97, CH97, FC96, Ker87, Khe08, KT94, Mac83, MP96, Pas03, Phi91, RELW09, SD01a, SJ07]. expert [Edw89]. explanatory [KS91, LGP90, Goo85q]. explication [Goo94b]. explicatum [Goo84h, Goo89b]. exploitable [Zim89]. Exploratory [WK90a, Goo90n]. exploring [DSS06]. explosive [FBC09]. Exponential [OS02, RG02, TYY02, AA09a, Ahm99, ASSY79, ASY80, ASY82, AR00, BN01b, BN01a, BL02, BLC04, BL05, BRF08, BG94, Bal95, Beh73, BLP09, BS01b, Chu80, CCNAF95, CC97, EG89, EBS86, FUOCN97, GS07, Goo83p, GK01, GK04, HM88, Jah05, JS00, Joh79, Kun92, LB01, Mar76, Mei08, MK08, Ng96, PS99, Pau92, PO05, RP01, RA01, RM05, RJ06, SB88, SB08, SY08, SL98a, SCAH05, SC09a, SN83, Son97, VD08, Wan08a, YST90]. exponentiality [Bal83, BL03, DH81b]. Exponentially [Gan93a, Gan90, HK84, Rig95, Shu08, TA94]. exposition [Goo87f, PV93]. exposures [PW86]. expression [SFC08]. expressions [GJ82]. Extended [Hou85, BG78, Goo79j, GF89, Gri92, Haw79, KM06, LL06, STH09, Var81]. extending [FRS06]. extension [Gri04, TO04, YR06, Dan80b]. Extensions [CMD74, BGS07]. extra [Gou93, Lu99]. extra-binomial [Gou93, Lu99]. extraction [Goo83l, Lee02a]. extreme [ANM09, AsY86, BG93, BKLW04, BDG04, BSK90, CKW73, Dup96a, DF98, LS99, Nad99, PW83, SWK73, Tay77]. extreme-value [Dup96a, DF98, LS99, PW83]. extremes [BG93]. eye [Goo81m]. eye-ball [Goo81m]. Eyraud [Joh82].

F [FC96]. F-tests [FC96]. face [DG95, LS88]. facets [SM91]. factor [AS00, BS05b, CA89, CB03, cDJgS93, Goo84p, Goo86q, Gup84, Hab84, HB93, Hos87, LN77, NL77a, Ouy06, Pas03, RG06, Str89, Vir07, VB81, VB83]. Factorial [PT02, AB03, Gup84, LZ92, Mac83, PN86, SB91, SC00, Goo90f]. factorization [Die93, Goo93d]. factors [Goo81w, Goo85b, MJ08, Mee93, Wil04]. fail [Dam82]. Failure [Bal95, PH95, AHJ92, BK08, EBS86, HKL08, LL03, OP00b, PT89, RB88, TC75, WWL09, YL01, YH85]. Failure-censored [Bal95]. failures [PP06a]. fallacy [Goo98a]. false [MMP08]. false-positive [MMP08]. familial [JSS09]. families [Che80, GM06, Jen95c, PS99]. family [Ahm88, BF08, CCNAF95, FP07, FUOCN97, GS07, JS00, Pau92, SC09a, SC00, Tad81, VD08, VS93, WK90a, ZW01]. Farima [BG01b]. Fast [Deá80, GHCJ09, KCS88, pB97, BLP09, Goo81-29, Goo82l, Goo83k, Goo81-29, Goo82l, Goo87e, Hol85b]. Faster [Gri04]. Fatal [Mae87]. fatigue [Pas03]. fatigue-limit [Pas03]. fault [Goo89c]. faulty [SKJ86]. favorable [DSL06]. favourites [WC91]. FD [AHH07]. feasible [AN03, AC08, Har04]. feature [Goo89e]. features [EH07]. Federalist [HL84]. fees [Goo79e, Goo79m]. feneralized [VG01]. Feynman [Goo80m]. Fibonacci [Goo93d, Goo94c]. Field [SK80]. fields [But99, Cai09, DN94, QT92]. Fieller [Her75]. fifty [Arn00, Goo90m]. file [Lau79]. filling [Nor91]. filter [Hig97]. filtering [Sel08, Sny88]. finance [Kun93]. financial [KM99, SJF06]. Find [YW02, DSL06, NW86]. finding [DM79, YK04]. Finetti [Goo88e, Goo89r]. fingerprints [Goo83i]. Finite [Brä92, CR03, CRT07, DK05, HMS02, Jön07, Lye91, Ouy06, Rid03, Rod06, Rod07, SAD03, SMH97, Van84, WB09, AHAH08, Agu99, AHO97, AHADA00, AHAH06, AJAH07, CGA03, Cro74, CH02, CM98, Dor88, Fan95, GA00, GM77, HP80, KX03, Kri77, MS80, NS08b, OBW05, PO05, SD01b, Spi98, Tsu93, Zha02, Zha06, And90b, And95c]. finite-mean [Cro74]. Finite-sample [Jön07, Lye91, Agu99, CH02]. First [LT95, Sho95, Arn00, Bar77, Dag95, Di 05, DL80, Gol72, Goo80c, Goo83-29, Goo83j, Goo83-30, Goo83-28, Nic96, Pol94, She83, Tse84, TA92, YL85, ZNAEB98]. First-Order [LT95, Dag95, DL80, Nic96, She83, TA92, ZNAEB98]. Fisher [GF89, Aok02, Bar84, BVP90, BS72, Fun79, Gol72, Goo84o, Goo84q, Goo85m, Goo86j, IP86, KC96, Mol79, PF93, Sco02, SL98a, TC92]. fisherman [Goo98a]. fit [Agu02, ABE83, ABE85, Aly90, Ant95, BRT07, Bow92, CW01, Cho08, EA78, Fel02, For97, FJ83, FJ84, Goo78n, Hab80, JS00, JL88, LL06, LS99, LG84, Luc08, LJ00, Mei09, PP03, PP73, SB08, SY89, Spi06, Wan08a, Wei79, WS00, WL90, Woz94, ZC03]. fitted [JG92, QQ82]. Fitting [BD96, CK01, Han86, Vir07, Cel81, FH09, HK07, KB87, Lee02b, Pat76, SA95, TTF07]. Five [SHH85]. fixed [ABH82, DM78, Don97, Goo83b, HZ96, HH74, Kru88, Kru89a, Kru90b, KSK93, Wor89]. flat [Kem84]. Flexible [Sch02, WmGT95, Goo86n, Goo81z]. flights [Mol99]. flock [OBW05]. floods [ESRV98]. Floor [SD02]. flow [CBPW97]. Fluctuation [KK05]. flux [CG97]. flybutter [Goo94d]. Fmax [Yan99]. focus [DS09]. fold [BE86, BBG86, GH82a, GH82b]. folded [Sum83]. Followed [RG02]. Following [LJDK02, CH99, HPY79]. footrule [SQ02]. forcing [WSPC09]. forecast [DCA03, TA94, VG01]. Forecasting [FK04, LGP90, SK87, XYR09, AMYY07, BHK05, GT90, GB86, Gue89, RMH88]. forecasts [VG01, WG92]. forensic [CA89]. forgotten [Fri85]. form [AS01, Arn79, Cha79, GHCJ09, Goo86g, QQ82]. forms [Chi79, FJT82, Goo86r, WK90a]. formula [Bar84, BR84, Dan80a, Goo83f, Goo84o, Goo84q, Goo85e, Smi89c]. Formulae [WM99, FC94]. Formulas [HP80, Gly84, JK89, KTB82]. Formulating [CZ02]. formulation [War74]. Forsythe [Tan82]. forthcoming [Ano73a, Ano73b, Ano73c, Ano73d, Ano74a, Ano74b, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano76a, Ano76b, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano82a, Ano82b, Ano82c]. Fortran [NO75]. Forward [Ber80, Ho93]. foundations [Ano06b, Goo86l]. four [DW01, OPS82, Sha03e, Soh94]. four-parameter [DW01]. Fourier [Goo81-29, Goo82l, Goo87e, Hol85b, Har85, Smi89c]. fourth [Ano03a, Goo94b, Par81, RR93a]. Foutz [Edi83, Goo81h, JL88]. fractal [CY92]. Fractiles [Bar79, Haw79]. fraction [DM05, HL98, SB85]. Fractional [AR01, CGA03, Cor97, GA01a, Goo89p, Goo89n, SC00, TH09, Goo79s, Goo83u]. fractionally [GA01b, GA02, GA09]. fractions [Goo83u, Goo83t, Goo83y, Goo89p, Goo89n]. fracture [SDP93]. frailty [CK04, GSH04]. frames [Goo80l]. Free [PB02, AB09, CH76, Goo83n, HPY79, KMK87, Lem77, Pie94a, RR77, WA72]. Freedom [LPZ02, FC96, Han79, VG05]. frequencies [BLP09, Kah93, MOS94, MK97a]. frequency [But99, Hal82, KX03, LFC92, SMP^{+}06]. frequently [Dan80a]. Friedman [AS00]. Frobenius [BDFR97]. Full [UP01]. Fuller [CL98, DG95]. Fully [vBBGOR06]. Function [LC02, RB00b, SW83, Ahm87, AN03, AS01, Amo85, And95b, AB97, Ans80, BA77, BN95, BVP90, BDFR97, Che78, Chi08, CFOS02, Day87, Den77, Ebr93, Ede94, oE89, ES86, Fis73a, GM79, GK90, GD92, Goo81j, GJ82, Goo83v, Goo84n, Goo87c, Goo88c, Goo90h, GR83, HK07, HBK92, HG85, Jon75, JL83, Jos01, Kat78, KPH05, KC97a, KM99, LH72, Mar86b, MS99, MS95, Mcd74, MYS01, Mei09, MW04, MR03, Mou05, Nue08, OHLH82, PRS87, PLD88, PB00, Pos94, RR93b, RB88, RACSGJ05, Ros06, Ruf07, SdFdGCM08, STH09, Sho86, Tah90, TC92, TC75, WM91, WS00, WWL09, XDZ09, YK04, pha95]. function-based [Mei09]. functional [Abd89, Baí09, HBC07, LWH08, Pat76]. functionals [CSS83, Goo85j, OP04]. functions [Ahm94, Ahm99, AHJ92, AFGP06, Arc80, CJ95, Dey84, El 82, Goo78h, Goo81t, Goo82c, Goo83d, Goo95e, Gue82, GW73, KW94, L’E97, LLKJ09, LP92, Mar87, Ong95, Pap80, PP85, RM96, Smi94, SVW88, Win75, ZW01, Goo80j]. functors [JG83a]. fundamental [LDR92]. further [Goo83m, Goo84q, GL88, Goo92c, HK07, LM74, Poi92, WCF79, Goo79t, Goo83x, Goo92b]. future [AR00, Lee99, Pad82]. fuzzy [Goo78g, Goo94e].

Galerkin [DB84a]. Galton [GG90]. game [Wec94]. games [AE87, Goo90f, Rod06]. Gamma [BS02, BG93, Bha01, BB75, Che05, CKT89, Dag78, Dod00, Far90b, Fis73a, GW01, Goo79r, GK04, IJP86, KW78, Kim83, KC97a, LM84, Mar86b, NK07, Ong95, OPB08, Pep93, PB72, SY89, kSB90, Sim93, TR75, Tad81, Tza09, Whe75, WWS04, YI01, YAAB87, ZY04]. gaps [Goo80b, Goo85s]. GARCH [HR07, Men00, TC08]. gauges [CG06]. Gaussian [AG92, Amo85, AG78, But99, CP76, CSS83, Cha79, DFY08, DN94, Gal02, Gup73, GN89, JG87, Jen76, Jen93c, Jen95b, KTB82, KM06, KC02, KA85, LY99, Lee99, Lee01, NP09, OP00a, Pad78, Pad82, PY99, Ros06, SY08, SdFdGCM08, Sim00, Sma05, TC94, YL01, ZL07]. Gaussian-quadrature [KTB82]. GEE1 [BL09]. GEE2 [BL09]. Gehan [HH74]. gene [HTC07]. General [Van84, BL02, BLC04, CS97, De 81, FS94, GS87, GS78, LC92, LGPP96, Pes80, PR84, TC75, ZR93]. Generalised [Ars86, PO06, JL96, KL94, OPB08]. Generalization [Yan99, BB80, BTL93, GL83, GC90, Hsu02]. Generalizations [Fel02, Goo85p, Goo86q, Goo79u]. Generalized [Agu02, DH77, Fam99, GK01, Kuk99, NK02, PT02, San89, WO93, AA09a, AN03, ANAA97, AC08, Ano81f, BS95, BH07, CYC99, Che06, DM79, Dup96b, Dup96a, DF98, EP92, FC96, Fam98, Fro04, Gil95, Goo79r, Goo81y, GL83, Goo86j, GK04, HH74, Hin97b, Joh82, KHA04, KAST05, LSSP08, Li01, LL08b, Mas03, Mei08, MK08, MH07, NC96, NP09, NL08, PB03, Paw01, PW83, RA01, RM05, RVZRZP08, SA04b, SJ06, TTF07, TH09, VBS07, Wan99, WAP84, XW07, XW09, XM09, ZS05, ZL96, Zim89, Ano81f, Goo81y]. generally [BHG01, SL09]. generallzations [Goo79t]. generate [WD93]. generated [Gan90, LK83a, RACSGJ05]. Generating [Ahm87, AL96, Bar78, DN94, HS86a, Lak81, Son97, AL01, Chi08, CYB90, Deá80, Pap80, PS98, PB72, Rya80, TR75, YST90, MG80]. Generation [Cha75a, Che85, hL92b, Pul79, Sch78, Sim93, Arc80, Dem06, Dev82, Dev84, Dev92, Fam98, FS75, Hör93, Joh79, JW80, Joh82, JL96, KS85, Kk90, KC97a, Lou84, Nar90, NC72, Pad78, RT78, RJ06, Sah79, SL93, WC91, Wil98, GL84]. Generator [BD76, LL73, AA09b, Bay90, Pes80, Zur93]. generators [BH85, DM05, Fle95, KW96, LSSP08, NO75, Pau84, RH76, Tan01]. Genetic [KKE07, BC08, Çel09, Goo81k, Hig97, PHO05, Sha03a]. genetics [Ano05d, Möh05, Sha03a, Sha09h]. genomics [Sha09h]. Genstat [Ban78]. geodesics [Dod00]. geometric [Ada91, Nue08, SB87, Smi94, TL07a]. geometrical [Goo86f, Sin90a]. Geometry [Goo83z, Sha04n]. germination [CMQ03]. germination-growth [CMQ03]. get [AG81]. Gibbs [Har04, NC96, Pan99b, Shi07, UP01]. Gini [LH78]. Given [Van84, Bon81, Cha75a, Che85, Dem06, Goo78g, Goo84a, Hol85b, Jen91, hL92b, MG80, MB97, Pul79, Sim93, Son97]. gives [Goo83s]. glance [HL84]. GLMM [YM96]. global [LZ92, YK04]. glossary [Mih72]. GLS [Cho98]. GM [Abd95]. GM-estimators [Abd95]. GMM [WLT08]. God [Goo85g]. Godambe [Goo80k]. Gompertz [AHADA00, AJAH07, MR03]. good [Mag75a, Mar90b]. Goodness [ABE83, ABE85, BRT07, Fel02, Mei09, Wan08a, WS00, WL90, Agu02, Aly90, Ant95, Bow92, CW01, Cho08, EA78, For97, FJ83, FJ84, Goo78n, Hab80, JS00, JL88, Lac90, LS99, Luc08, LJ00, PP03, PP73, SB08, SY89, Spi06, Woz94, ZC03]. Goodness-of-fit [ABE83, ABE85, Fel02, Mei09, Wan08a, WS00, WL90, Agu02, Aly90, Ant95, Bow92, Cho08, For97, FJ83, FJ84, Goo78n, JS00, JL88, LS99, PP73, SB08, Spi06, ZC03]. goodness-of-link [Lac90]. Govern [Mye98]. GPX [Liu08]. grading [Goo80d]. Gram [Dod83]. Granger [CM06]. graph [Fra79]. graphic [Gra86]. Graphical [Wet96, MS01, VP89, Won95]. graphics [Har75]. Grenander [And80]. groundwater [ELB97]. Group [LJDK02, BD82, BT09, BE86, CB07, DDB09, GR08, GK90, HK00, Krz82, LF83, LL07, MYS01, MS80, RRCD97, Sin90b, WW03]. group-testing [BT09]. grouped [AL93, Ham77, JG92, MO91, QC03, WB73]. grouping [Chu80, SC82]. groups [JCS07, TG90]. Growth [HM95, CMQ03, FPRS92, GM08, LW82, LC92, SFSS85, SND89, Soh92]. Grundy [Li94]. guaranteed [CS95a, CY99]. guaranteed-coverage [CS95a, CY99]. guided [SR97]. Gumbel [ESRV98].

hadron [Goo90o]. Haenszel [BP01, BA77, Lev82]. Hakvunic [Goo84f].

Hakvunic-mean [Goo84f]. half [Bal85, BP86, BS95, BS94b, Ter90]. half-normal [BS94b]. half-space [Ter90]. handbooks [Ano06c]. Haphazardness [Goo95c]. hard [CM07b, Lot82, Wel82]. harder [Dav79]. hardware [Sha09k]. harmonic [Goo84u, Goo84e, Goo84t, Goo92a, HHI90]. harmonic-mean [Goo84u, Goo84e, Goo84t, Goo92a]. Hartley [BCL93, Yan99]. Hastings [CS05]. having [Goo80j]. hazard [Agu02, GD92, Gri02]. hazard-based [Agu02]. hazards [Atk91, CSC00, Gho95, Lou95, QyT92, SB05]. health [Sha05e]. heart [GHDB89]. heavy [ABE85, lS96]. heavy-tailed [lS96]. Hellinger [EW91]. Hermite [GT78]. herring [Goo88a]. heterogeneity [CM05b, Lev78a, NS09, TS00, Wil75]. heterogeneous [GOS09, LKKL07]. Heteroscedastic [Mak00, BC08, Mye98, WW03]. Heteroscedasticity [LL09b, LT95, Dor88, Goo86j, HS04, LS01, TT86, TZ97, WK06]. heteroskedastic [CNZ04, Fur91]. Heteroskedasticity [CNZ01, CNdGAL09, CNFO05, Fur96, Fur08, SP00]. Heteroskedasticity-consistent [CNZ01, CNdGAL09, CNFO05]. Heuristic [hJ93, GRVV08]. hidden [RRT99, Spe06]. Hidiroglou [GH07]. Hierarchic [RML90, Mur83]. hierarchical [CDJ02, Eve01, Goo94e, LiT95, Nan98, SN95, TAM^{+}03, Goo83-28]. hierarchy [TL96]. high [Ken79d, SMP^{+}06, Spi06, Wil79]. Higher [GV02, BHZ08, Goo86q, SZ02, Sie78, SH72]. Higher-Order [GV02, BHZ08]. highest [TPG93]. highly [FBC09]. Hill [GRVV08, MGN99]. Hiroshima [HJ85]. histograms [HS91, Goo78f]. historical [Goo85i, Goo85f, Goo86o, TP06]. history [Bar84, Goo83j, GS83, Goo84o, Goo84q]. Hitler [Goo95c]. HIV [Mar95, ZZCS09]. HNBUE [HANMA98]. Hodges [Cla86]. homogeneity [BH95, Goo81n, Goo86h, GN89, JG87, Kös06, Lev78b, Llo93, yL87, MPPZ05]. Homogeneous [SD02, Spe06]. homomorphisms [AA09b]. homoscedasticity [Pie94a, Pie94b]. honest [Ano82d]. Horvitz [SNB07a]. hot [Goo98a]. Hotelling [Jen83, Jen95b, KAST05]. Houston [DE06]. huge [WLT08]. hull [Gle89b, Gle91]. Hybrid [DG02, AMYY07, Atk92, BRF08, Dia98]. hypergeometric [KS85, Ma99]. hypotheses [Alb87, Bho73b, BH95, Che80, FC96, GsW01, Goo83s, Goo94g, Goo95b, HB93, HTC07, JF80, KT97, LW82, Mal79]. Hypothesis [LW04, Spi82, BCL93, DR02, Die06, Goo83r, Goo85b, Goo90o, Hab84, Hos87, JO85, KK85, LL93, SFSS85, SL87, Val07, WWL09]. hypothesis-dependent [SL87]. hypothyroidism [Goo94i].

ICM [Gle89a]. idea [Goo96a]. ideas [Ano05e, Goo89q]. identical [Goo80j]. identifiability [Mcd74]. Identification [De 97, RG93a, Zue96, BHZ08, CY92, DK05, oE89, Gal02, KP99, Ros06, SZ02, SP97, WS92]. identifications [JCS07]. identify [GR08]. Identifying [Sho95, SCB07]. identities [Goo83x]. IF [Goo83-30]. ignoring [KSH73]. II [Bal89, Bal92, BL02, BL03, BLC04, BKLW04, BL05, BH07, BE94, Con95, EM86, Goo85s, Goo89w, Goo92d, Jen89b, JH90a, JR96b, KH09, KY93, LB01, MM05, SN83, SS88, TB88, TYY02, Wan08a]. III [Jen89c]. ill [Atk91, LS88]. ill-conditioned [Atk91]. ill-conditioning [LS88]. image [Gle89a, Tay94]. imaginary [GT81, Goo83-30]. imaging [Sha09l]. Impact [AHH07, Pau84, CR93, Coo07]. Implementation [AG78, Fan03, NO75, PGV04, Li94, PK72, ZL07]. Importance [CD92b, Con95, Kuk99, Lee02a, Goo79c, OB92, RC98, SD01b, YK04]. important [De 97, Goo94j]. Imposing [CH06a]. impracticable [Goo81o]. imprecisions [Han94]. improper [Kem84]. improve [HTC07, Hun02, Lee98a, WG92]. Improved [BN01a, BR06, BH95, CHH91, CC97, FFCN08, HHKD02, KP87a, LT95, RG06, AH90, Arn79, CLLN04, FC94, Goo83k, Goo86p, LSCN08, Loh73, MW94]. Improvement [Cho08, Goo78n, Goo82i]. Improvements [CLP93, NP98]. Improving [Aok02, BD76, Dor08, GH07, JG92, CS08, Goo87a, PLD88, Por88, SSW95, Sar98, Bay90, Cho98]. Imputation [Li88, ZZCS09, Bel93b, LPZ02, vBBGOR06]. imputed [BS05b]. imputing [DFY08]. in-control [Rig95]. Incidence [Mar95, MW94]. including [Goo83q]. inclusion [SNB07a]. income [Goo81s]. Incomplete [SP01a, Ang03, AT05, Bho73a, Bho84, CR73, Day87, EH92, FPRS92, Fis73a, KK93, LW82, LC92, Mar86b, SFSS85, Shi85, TP98, Tay72, TC73, YI01, ZR07]. incomplete-data [LC92]. Incorporating [HTC07, NC06]. incorrect [CW74]. increase [GW94]. increases [Goo80h, Goo83b]. increasing [WWL09]. increments [Kun93]. indefinite [SA79]. Independence [Pet02, ASSY79, ASY81, ASY82, Alb85, Alb87, Del83, FP07, Fle95, Goo86q, KK99, Mal79, TG90, WM90, BR81, Jen89b]. independent [Bak07, BS01a, CS05, Fre07, Goo94f, HB78, JG83b, Mat79, MBH91, Mel80, MK97b, SB73, Sum83, Vir07]. index [CK96, GO03, GP07, Goo79k, Goo82f, GS83, Goo85i, KHA04, LH78, MR09, PX02, PX05, Smi84, TL96, Wri95, WSPC09]. indexed [Lau79]. indexes [Goo82j, Goo89s]. Indices [Goo87b]. Indirect [Cor97]. individual [Soh92, WDR86]. individuals [SR00]. induced [CRR99, Pak99, Pin78]. induction [Goo81a, Goo83a, Goo83-27, Goo85a, KCS88]. inequalities [Dev84, FJT82, GsW01, Jen80, RS92]. inequality [CH06a, Goo88b, Goo90h, QFG87]. inevitable [Goo85a]. Inexact [Pol97]. inexactification [Goo93b]. inexpensive [NCA^{+}00]. Inference [Bak07, Bak08, BKLW04, BH07, KYF08, LMB08, Mou01, Sch83, SCA07, AJAH07, BS95, BL02, BLC04, BL05, BRF08, BS94a, BN96, CS86, CL97, CC09, CM05a, Cor97, EA78, Edw85, Eve01, FB90, FFCN08, KS91, LB01, Mar92, MH99, NTK09, PB03, RM05, RJ95, kSB90, ST87, TS09, Tsi03, WWS04, WCC07]. inferences [LWH08, Rod06]. infinite [Bar81]. infinitely [PS98, WK03]. Inflated [Zim04]. Influence [Gil95, Soh96, XYT08, AFGP06, Gra86, Gra88, Kim92, LMB08, MG06, ZT01]. Influential [Hos91]. influentiality [Goo79k]. informal [Goo86m]. Informatics [Sha09j]. Information [Beh72, Beh73, Dod00, HHI90, HD77, Sy01a, Adk96, ASS04, BHK05, CL98, Dha85, KE03, Kun92, LG09, Lye91, MB86, MB90, NCÁHC^{+}09, PF93, RW96, Sco02, SL98a, Sy01b, ZR07]. information-based [CL98]. Information-statistical [Sy01a]. information-theoretic [Sy01b]. informational [GL83, Goo81e]. informative [MB97]. initial [HV93, HB06, Mou05]. initiating [TJKB00]. innovation [SP97]. innovational [Pop08]. input [CH97, PCS09]. input-output [PCS09]. insect [TBT95]. insight [TB86]. inspection [SKJ86]. Instead [YW02]. instrumental [FAS82, Owe81]. instruments [Han78, Han94]. integer [PB72]. integral [DW90, Hut93b, Rig95, TL07a]. integrals [DD77, Goo80m, LM84, Möh05, RS90b, Smi89c, XPC03]. Integrated [TYY02, GA01b, GA02]. Integration [PS85, BD84, CGA03, Coo07, GA01a, OB92, PQ89, She91, WB09, XPC03]. intensity [Ho93, RB88, Sen02, Soh96]. intensive [Goo86g]. Interacting [GS01]. interaction [Gen76, Hab84, HB93, HSSY04, Hos87, O’N82, PR84]. Interactions [Goo82j]. intercept [LMB08, SA92]. interchanging [HH93]. intercomponent [Tho92]. interesting [And95b]. interests [Goo81f]. interface [VP89]. interior [Win75]. intermediate [Goo83s, SK90b]. internal [Din08]. internally [CY89]. International [Goo81g]. Internet [Sha06d]. interneuron [TBT95]. Interpolation [Goo78k]. interpretation [Goo83r, Goo84l, Goo84j, Goo84k, Goo89f]. Interpreting [CC90b]. interquantile [Goo03]. interrupted [RR03, RR06]. intersection [CP95]. intersubject [Zim04]. Interval [BS78, CH99, EBA00, MZ08, SY08, TYY02, WS90, BJ99, BF08, Bon05, Bon06, Che97, CC85, CNdGAL09, Ede94, Gho95, HHKD02, HMZ05, Kar09, KJ81, Mas03, NW09, OM88, PT89, Pad82, Pan98, PP06c, Pot81, QY95, Roy93, SB88, STS94, SL88, TPG93, tKWY95, WK90b, WW03, YXW07, YW09]. interval-censored [YW09]. Intervals [LJDK02, PB02, AB09, AR00, BG01b, BE86, BBG86, CT82, CS95a, CY99, CSAR93, CB03, DV86, DF98, FW88, Fra74, GMSS95, GZP05, HMS89, HK00, HB93, Her75, JK08, JR83, Kwo96, LL91, LP89, LL07, LJZB05, LD97, MMP08, Mar92, Mun06, OHW97, PS99, PO06, RS77, SB73, SM94, Sha01b, kSwXR93, SW84, SNP93, TW08, TTNC09, TBG^{+}90, Wan92, WW95, Wan08b, WL07b, XM09, YM90]. intervention [Gue89]. intraclass [BF08, CCSG81, CH82, NC06, PSB03, WP81]. invariant [Goo78a, Goo79i, Goo81q, Pan98, Goo90k]. invention [RR01]. Inverse [OP00a, AG92, AS81, Amo85, BGS07, BG01a, CJ95, Che78, Fea79, Fra74, Hen81, KMK87, KC02, KP82, Lee99, Lee01, Mas03, Pad78, Pad82, PY99, SY08, Sho86, TC94, Tre94, YL01, Zac80, Jon86]. inverse-sampling [BG01a]. inverses [DM79, San89]. inversion [Dee76, MS80]. inversions [She91]. inverted [AA09a, Ong95]. Inverting [SN95, Dan80a]. investigated [FS90]. Investigating [SdFdGCM08, Won85]. Investigation [FT96, ZS05, AR01, BHLH78, De 81, FJ83, GHW99, Hos87, Nat82, PSB03, PQ84, RR08, TZ04]. investigations [BH73, Tho74]. involved [JL83]. involving [DD77, Fre07]. ion [CM05a]. IP [DBVK02]. irregular [Goo82k, JH72]. irregularly [PH03, TY90]. irrelevance [Goo88e]. Irreversible [Nor91]. Irwin [Bar84, Goo84o, Goo84q]. Ising [Shi07]. Isolikelihoods [Goo90e]. Issues [PH95, DC02, Die94]. Item [Sah02]. items [Hut93a, TP98]. iterated [Alt92, Mai03, Pol97, SA04a]. iterative [LY99, MK97a, SB96].

Jackknife [Ahm88, QyT92, Hua91, LL79]. Jackknifing [ST88]. Jacobians [Goo81y]. James [CLP93, KC89, KTM05, Mye98, WZ95, YW02, YW09]. Jeffreys [Goo83-27]. Jensen [NK07]. Jessen [Gha01]. Johansen [MS01]. Johnson [CK01, Nie06, SB82a, SVW88, Tue01]. Joint [LH72, DL81, FJT82, GA01a, KC97b, Ma97, Par87, SWK73, Too72]. JSCS [Ano03b, Mar90b, Sar98]. judgement [Ano82d, Goo89t]. judgements [Goo79g]. judgments [Goo87a]. judicial [GT85]. jump [ÁGR06, Gle97]. jumps [Mol99]. Jupiter [Goo84b, Goo85f]. Just [YW02, Goo84s].

K-philosophy [Kru89b]. Kalman [Sny88]. Kaplan [CFOS02]. kappa [DW01, KB96, VA08]. kappa-exact [KB96]. Kempthorne [Goo84g]. Kendall [Gat09, LH79, Mag75b, MP81]. Kernel [Ait83, Woo85, Bel93b, Bow85, DP06, Dog89, Fos95, GD92, Jos01, Li01, NP98, PB00, RHH80, Tut90, Wor89, ZT01]. kernel- [Fos95]. kernel-based [Bow85, Dog89]. key [Goo94f]. Khintchine [BJ82]. kind [Ano03a]. kinds [Gol72, Goo80c, Goo80i, Goo83j]. knotting [WK90a]. knowledge [CR93, SGR04]. known [BG07, KC02, Kim83, NS86, NK08, PT99]. Koen [Chi07]. Koheleth [Goo85g]. Kolmogorov [BHLH78, Fel02, GSF78, Goo78c, Loh73, WLK06]. Kotz [GJJLGS06]. Kullback [PB03]. kurtosis [CR93, HL92a].

L. [Ken79e]. Lack [DL92, LL06]. lack-of-fit [LL06]. lady [Goo90j, Goo92b]. lag [Aza73, CL98]. lagged [LH94, WHF80]. Lagrange [Fur08, Sef92, SB82a]. Laguerre [SB08]. Lake [KT85]. language [JCM72]. Laplace [Goo84b, Kuk99, SW82]. lapses [Goo87f]. Large [EGM^{+}04, Gho95, And97, And80, AG81, Goo83e, Goo89f, Gri04, Har04, Hin97a, Kru03, Lau79, Moo90, PN98, WZ95]. large-scale [WZ95]. larger [MKDM94, SMO03]. largest [CT82, Fon92]. latent [Dre08, LL88, TAY02, ZL09]. lattices [Nor91]. Lavallée [GH07]. law [Goo81-27, Goo85n, Goo85g, Goo93e, Kru03, Sen02, ST87]. laws [Mei08]. layout [DD78, Pie94a, Pie94b, PR84, XM09]. Least [Cel81, Die05, KW00, Mai77, QFG87, RR01, SVW88, Car07, CCHM08, DSL06, Dee76, DR94, DR02, Die06, DM79, DM93, EN90, FC96, Far78, Goo79c, Goo83v, Goo84v, GM77, Hie81, HH96, KC89, KMK87, Lee02b, MM80, PV93, RR03, RR82, San89, SMH97, SA04a, SB82c, TA92, Tsu93, VM96, Wei79, Wat77]. Least-squares [RR01, SVW88, GM77, PV93, RR03, RR82]. left [MK08, Pan98, Ter87]. left-tail [Ter87]. Legendre [RR93b]. Lehmann [Cla86, HPY79]. Leibler [PB03]. lemma [Goo80l]. Length [CW99, PO05, Gan93b, JH01, Mac92, Rig95, SB82b, Zan08]. Length-bias [CW99]. Length-biased [PO05]. lengthed [And95c]. lengths [CI81]. Lepage [BT00]. Letter [Hog73]. level [BF87, CM04, HCW07, JW97, RR03, RR06, Rod07, TC08, WR94]. levels [Eti81, JR83, SW90, VB81, VB83]. Levene [yL87, Neu07, Pan99a]. Leverage [CNZ04]. Leverage-adjusted [CNZ04]. Liang [LSS93]. lie [Goo90c]. lie-detectors [Goo90c]. Life [HJ85, Wan08c, AHAH08, Ahm88, AHAH06, BJ73, DP93, EH92, EM86, HANMA98, Mou95, PL99, Lia11]. lifetime [AHADA00, CK04, KS04, PY99]. Lifetimes [TYY02, HK84, YT96]. lifetme [Ahm99]. like [ASS97, WA72]. Likelihood [BS02, CKM01, CNV02, NK02, Sch02, Sha06a, SN83, Tsi02, WWS04, Adi98, AH09, AGNS91, AG78, AK85, Ant95, Bal89, Bal92, BRF08, BCV98, BBP04, Bj78, BB74, BS72, BS05c, BB84, But99, Cad94, CH91, CC05, CD92a, CHáS03, CG93, CZ07, CMQ03, CR73, CDDCN97, CC97, Cor04, DAM98, Den77, Dic78, Die94, cDJgS93, DMF83, FH86, FFCN08, GG77, GR79a, GO03, Goo78a, Goo79f, Goo82g, Goo82e, Goo83-28, Goo83e, Goo86j, Goo86p, Goo87h, Gri04, GN89, HN80, Hie81, Hoe89, Hut77, KK99, KA85, KG80, Kru86a, Lai82b, Lam05, LY96, LY99, Lee85, LSS93, Lil01, Luc08, LP00, MZ77, MH72, McN87, Men00, Mil79, Nag75, OHLH82, PGT09, PT99, PW83, RML90, Ron89, RV98, Rud86, Ruf07, SBC03, Sco02, Shi85, ST88, Smi89d, SB82c]. likelihood [SB93, SJ06, Tza09, WK72, WO93, YX03, YK04, ZK86, ZY04, ZNAEB98, Joh90a]. Likelihood-based [Tsi02, CC05, ZNAEB98]. likelihood-ratio [Goo86j, Lam05, ZY04]. likelihood/minimum [GL85]. likelihoods [CLLN04, FFCN07, FT96, JK89]. Lilliefors [GSF78]. limit [Bar78, HC06, Mac92, Pas03]. Limited [MB90, Dem90, Lye91, MB86, Mol99]. limited-information [Lye91]. limiting [Har91, WK90a]. limits [AL02, Eri83, HS73, Har87b, Har89, Hil87, KA03, KS04, Lee99, Lee01, RS89, RVZRZP08, TC94, TG73]. LIML [Mcd74]. Lindley [GAM09]. line [EL75, EB90, KYF08, MT01]. Linear [HMS02, How88, Kuk99, MGG78, NK02, PT02, RG02, Van84, Abd89, AH90, AC08, AS81, AD01, AD03, And95b, AL99, Arc80, Baí09, BP86, BL02, Bel93b, BF08, BVP90, CYC99, CG91, CHM09, Che06, CBS06, CCMV07, CS97, CJ73, Dor01, Ebb73, Edw95, EH01, FH86, Fie93, Fra74, Fri79, Gan91, GM79, GK90, GsW01, GO90, GG75, GA01a, Gil95, Gil07, GS73, GV81, GLP72, GM77, Gri92, GJJL02, GJJLGS06, HZ96, HSB85, HV05, Hin97b, Hou85, HM98, Hsu02, HH93, Hua01, Hud83, HMZ05, Joh78, JMY96, KHA04, KMK87, KL94, LY99, Li01, LB08b, LP09, LL09b, LK83a, LB01, LJZB05, LWH08, LP00, MS95, MYS01, Mee93, MMWM83, MMW91, NC96, NW83, NSG91, NJJ92, NL08, Nic05, NS08b, NFD00, Oht98, PK72, Pat76, Paw01]. linear [PYC93, Pes80, PR84, Qu06, QFG87, RHH80, RS88b, Rol01, SK08, SCA07, SA95, Soh92, SLL00, SOU04, SMV76, SJF06, ScK97, SCB07, SP01b, TS09, Tah90, TT88, TTF07, TAM^{+}03, TBG^{+}90, VG01, Wan92, Wei79, WCEC94, WECC00, WW07, WO93, YM90, ZL96, Ken79a]. linear-bilinear [CS97]. linear-statistic [Gil07]. linearity [CG06, Fur07, Goo87f, Li01, Zue96]. lines [CM76, RR03, TT86]. LINEX [AN03]. link [Lac90, MH07]. linkage [BS72, FS81]. LISREL [BF87]. little [Goo94h]. LLRANDOM [LL73]. LM [Kib04]. load [Dav93]. loadings [BS05b]. Local [NBB00, ZT01, CCHM08, Coo07, EK03, FH09, Hoe89, LMB08, QL01, Jen95c]. local-to-unity [Coo07]. Locally [SY92, JG83b, XYR09]. location [BP86, BW01a, BK96, BT00, BR03, CC09, Cla86, Fri07, GV81, Jen89a, Jen89b, Jen89c, Kap81, LF83, Lem87, Nat82, Ng96, PQ84, RA01, RR77, TMW73, TAY02, Tho92, Wan90, WGP00, WA72]. location-scale [BT00]. log [ABH82, BB74, Che97, Hud83, KG80, NL08, NFD00, OPB08, Ruf07, SJF06, SP01b]. log-concave [Ruf07]. log-gamma [OPB08]. log-likelihood-ratio [KG80]. log-linear [Hud83, NL08, NFD00, SP01b]. log-logistic [Che97]. logarithm [Far90b, Hut77, ZL07]. Logarithmic [GA02, Goo83f]. logic [Goo85n, Joh90a]. logistic [Ada96, AsY86, Bal85, BP86, BS95, BH07, BL09, BB89, CDF05, Che97, CB97, DPS01, DC02, FTS09, Gao04, Ham77, HSB85, Hsu02, HHKD02, Jai85, Jon75, KP09, KW04, Lam05, LSS93, LB96, Nor84, NBB00, OHLH82, RV08a, RELW09, Sch86, STG^{+}01, SB05, SW84, Sum83, Tho06, TL07b, VW78, WL07b, vdTGL97]. logistic-normal [LB96]. Logit [YM99, McL78]. loglinear [FP07]. lognormal [Che82, Ebb73, GH83b, Goo83q, Goo83x, GGSNR09, J.83, KH04, OM88, Pou04, RFGE86, SB87, UP01, Win75]. Long [TBT95, CGA09, CR03, Cra05, DN06, FRB06, NM98, RRP06, SB96, WW06, Wan08d, WGP00]. Long-duration [TBT95]. long-memory [CGA09, DN06, FRB06, RRP06]. long-term [NM98, SB96]. longitudinal [BS05a, Kar09, SJ06, WmGT95]. look [And93, IHM78, LDC73, WCF79]. look-up [LDC73]. loops [Koh81]. Lorenz [OS02]. loss [AN03, Goo88c, STH09, Wan99, WSPC09]. loss-based [WSPC09]. lot [MMR92]. low [Kk90, PX97]. low-parameter [Kk90]. Lower [PT03, BC94, PT89, Jen93c]. LR [BE94, Kib04]. LR-estimators [BE94]. LS [Sma05]. LTB [LAR09]. Lucas [Goo93d, Goo94c]. luck [Goo93c].

M [AFGP06]. M-estimators [AFGP06]. MA [CHáS03, SZ02]. machine [KPJ09]. macro [KK91, Kel94]. main [MJ08]. maintenance [Sha03d]. majorant [Car07]. make [Dav79]. Making [Har04, BAB86]. Mallows [CM98]. management [OBW05]. Mann [Lem87, SL89a]. MANOVA [OPS82, Hos91]. Mantel [AB82, BP01, BA77, Leg00, Lev82, Sie78]. many [Mac83, Mal79, RT88]. Maple [BS05c, CS08, SC09b]. Marginal [NK02, Dem06, Goo81f, Joh78, MPPZ05, WS90]. marginals [DL81, Ebr93, Sim93]. margins [ABH82, GG80, Son97]. marked [CM07b]. Market [WT06, LGP90, RMH88, Sch75]. Marketing [GLC99, GW94, Sha09e, Soh00]. Markov [ÁGR06, AT05, Cai09, DL80, Fan95, Gly84, Goo89m, Har93, Har89, Har91, HP80, KK85, KS87a, KK91, KS91, KK93, Kel94, Kim93, LF04, Lev05, Li88, MS80, MO91, PSW75, QT92, RRT99, SB86, Spe06, TH09, WM90, Zan08]. Markov-chain [Zan08]. Markov-process [KK91, KK93, Kel94]. Markov-switching [TH09]. Markovian [CM05b]. Markovity [Goo83u, Goo89n]. martin [Aza73]. mass [Goo84b, Goo85f, Goo89v]. masses [Goo89u, GJH90, Mal06]. matched [Goo85r, Phi91]. matching [Tue01]. Mathematica [SC09b, Tza09]. mathematical [BPP00, Goo79c]. Mathematically [Goo85p]. Matrices [KP87b, BD82, BBHW95, CP91, Cha75a, CYB90, GH76, Goo86r, Hen95, JS02, JW80, KP99, hL92b, MJ93, PCS09, RML90, Rya80, SN95, WM95, Zim89]. Matrix [FC94, LK88, Beh72, Beh73, BDFR97, CCNA09, Che85, CKW73, CH82, CNZ01, CNFO05, Dan80a, Dee76, Dha85, Die93, DN94, Dut99, FPRS92, FS94, Fur96, Goo88b, Gup73, Hen81, HK92, LS01, Liu08, Mai77, MS80, PGT09, SWK73, SKC75, SA79, Sco02, Sim93, Tre94, Tre95, ZL07, Jen91]. maxima [CSS83, Hoe89, YK04]. maximizes [Goo86p]. Maximum [Bj78, BS02, CMQ03, CNV02, GMR82, GO03, Lee85, Luc08, MH72, Mil79, PW83, Ron89, RV98, SS95, Sco02, Shi85, Tza09, WB02, AGNS91, AG78, Bal89, Bal92, BS72, BS05c, BB84, But99, Cad94, CH91, CD92a, CH06a, CHáS03, CG93, CR73, CH88, CDDCN97, Dic78, Die94, cDJgS93, FH86, GR79a, Goo78a, Goo82e, Goo83e, GL85, Goo86q, Gri04, HN80, Hie81, Hoe89, Kul90, Lai82b, LY96, LY99, LSS93, Lil01, MZ77, Men00, OHLH82, Ruf07, Sch78, Smi89d, SB82c, SC00, ZK86, Goo87h, McN87]. maximum-likelihood [Goo78a, GL85]. maximum-likelihood/minimum-discrepancy [GL85]. MaxMin [AL02]. Maxwell [KW94]. may [Kru88, Kru89a, MF02, Phi91]. MCEM [TTF07]. MCMC [CRR99, RRT99, Vir07]. mean [AG92, AS90, Bak07, Bak08, Bal89, Bal92, Bho73a, Bho73b, BLP09, Bon81, Bon05, BTL93, CCNA09, CT82, Che94, CGN04, Che85, CM07a, CKM01, CCSG81, CV07, Cro74, EBA00, FB90, Fon92, GsW01, GMLM^{+}08, Goo84u, Goo84e, GBSS84, Goo84f, Goo84t, Goo92a, GR79b, HP80, HT99, HMZ05, Jön07, KC02, KJ81, KR89, KL05, Kru86a, LH97, LFC92, hL92b, Mac83, MF02, MG80, MK74, Nic96, Oht98, OBW05, Pad82, PBWU78, PB93, Pul79, RS09, SL09, SB87, kSB90, VC78, VS91, WV79, Wan92, Wan08b, Wan08d, Zha00]. mean-adjusted [Jön07]. mean-square [PB93]. meaning [Goo79i]. meaningfully [Goo81a]. means [Alb83, Alb85, Auc86, BCX93, BW01b, BB96, Bho84, BR84, Bre93, BHG01, CSAR93, CH06b, CH82, Fon90, GHDB89, HH92, HS86a, HH93, JG83a, JUP86, hJ93, KG80, KTM05, Mar76, MK97b, Nag75, Pep93, Ros95, Roy93, SB88, SMO03, SSM95, TB88, WS04, WWS04]. measure [Bic03, Goo83c, GL83, Goo85h, Goo93e, SOU04, Tan01, WK03]. Measurement [Sch02, CHH91, CG06, Edw95, Goo82h, Hol85a, Kim07, LMB08, RS88b, Tho06, TL07b, ZL96]. measurements [CW74, DE06, GHDB89, Mar86a, WDR86]. Measures [Kim92, AS00, ASS97, BZB08, BBG86, CBPW97, FS94, FP07, Goo81d, GL83, Goo84p, Goo85p, GS87, Gra88, HT93, HL92a, IM82, Jen93c, MC91, PY99, Pol94, PB04, SE90, Wan90, Wil75, WM95]. meat [DV95]. mechanics [Goo82b]. mechanism [LL03]. media [Kel94]. Medialized [Goo89l, Goo90i]. median [AHO97, Gan93b, Hil87, KJ81, PB01, SAD03, SMH97, SB87]. Medians [PB02]. medical [Ell00, Goo84x, IHM78]. medoids [VPB03]. Mee [RVZRZP08]. Meier [CFOS02]. Mellin [Mei08]. memory [CGA09, CCY04, CR03, Cra05, DN06, FRB06, Goo86m, RRP06, WW06, Wan08d, WGP00]. Mersenne [Goo80b, Goo85s]. Mes [Oht98]. Meta [Soh00, CDJ02, JCKS09]. meta-analysis [JCKS09]. metamodel [KT94]. metamodels [SN05]. metaprinciple [Goo89i]. Method [Nic05, PT02, Wu02, Bel93a, CCY04, CCMV07, Dem06, FS03, Fou80, Fra74, Gha01, Goo78g, Goo78j, Goo81o, Goo82d, Goo84n, GL85, Goo87a, GK01, HT93, HK07, Her75, JL88, Jun08, Kah93, KX03, Lee81, LAR09, Mar81, May06, Mcd74, MT80, MK97a, PGV04, Rya80, SL98b, Smi89d, Sol90, Som85, TR75, VA08, WL07a, WD93, YM96, YST90, Zan08, ZL09]. methodology [GS78, ZC03]. Methods [SHH85, Ait83, AB05, AL99, And97, Bay90, BG01b, CC05, CH98, CM06, Cla96, CB03, CNZ04, CM01, DSL06, DW80, Do92, DSS06, DH81a, oE89, EK03, FC89, FS75, GHCJ09, GG75, GAM09, Goo80i, Goo82b, GS85, GL86, GSC87, HT93, HL05, HK92, HSW75, HZ03, HM98, Hun87, JL96, Kar09, Ken79a, KT94, LY99, Lee98b, Leg00, LN77, LW95, LOR04, Lou84, MJ08, MRR84, Mei09, MH99, Mur78, NL77a, NL08, Pap93, PPRW06, RV08a, Roy93, STG^{+}01, Sel08, SW83, Tay94, VM96, WmGT95, Weg72, Wet96, Wor89, WSPC09, ZK86, ZZCS09]. Metropolis [CS05]. MEWMA [LL03]. micro [KK91, KK93]. microarray [HTC07, SFC08]. mid [LPSL05]. mid- [LPSL05]. middle [Goo85n]. midrank [HG85]. Miller [Goo89e]. min [SSW95, Sar98]. mineral [Goo79g]. minima [BC82]. minimal [Goo78l, TC75]. minimal-vocabulary [Goo78l]. Minimally [MB97]. minimax [CCHM08, DM78, Goo94k, PM02, STS94]. Minimizing [VC78]. Minimum [EW91, PD03, PB93, WLK06, ANPV97, AZ05, BS94a, BW01a, CH02, JCS07, MPPZ05, Oht98, Pak99, Sch78, SNP93]. minimum-discrepancy [GL85]. mining [Sha09d, Sy01a]. minus [Hut77]. mis [tKWY95]. mis-specified [tKWY95]. miscellaneous [Goo90k]. misclassification [FMC09, MG06, MMP08]. Mises [BCV98, LJ00]. missing [AL94, Bho73b, Mac83, STG^{+}01]. Misspecification [Fur08, DB84b, HKT04, MH07, PNW06, RB00a, WB94, Yan98]. misspecified [BJM92, Lye91, Jen89a, Jen89b, Jen89c]. misstatement [Goo83w]. Mixed [Kuk99, AC08, BS05a, Che06, CB03, FH86, Goo81v, Goo83-28, HZ96, HCW07, JSS09, Kri77, LP00, NL08, Paw01, Pes80, RV08a, Rol01, Spi98, TTF07, WO93]. mixed-effects [HZ96]. mixture [AHAH08, AHO97, AHADA00, AHAH06, AJAH07, Amo85, AL96, Atk92, Beh72, Beh73, BZB08, CG93, CCD96, Cor00, CMH04, EG89, Eve88, For97, GM79, HMK00, Jah05, JG92, LB08a, MR03, Mou05, NM98, OZ81, Pap83, PT99, RC98, SJF06, SCK01, Zha02, Zha06, vdTGL97]. Mixtures [SHH85, AMAE97, Éri97, Gne97, Hat86, HD77, Hos78, HL09, IMLG09, Jen95b, KX03, WL07a]. Ml [AL93, AL94, Atk92, CR03, McG89]. MLE [Dut99, MR03, WLT08]. modal [PP06a]. mode [And80, Bic03, NS89, War74]. Model [AW95, BCP02, CY89, HMS02, LT95, OS02, PT02, PH95, Pet02, RB00a, Sah02, Sch02, Van84, YM99, AG85, AH90, AHJ92, AHADA00, AJAH07, AR01, AC00, AL99, ANPV97, AK86a, Atk92, Aza73, BC08, BS94a, BD96, BH85, BG99, BHK05, BB89, CDF05, CGA09, CHR03, Cel81, CY92, CSC00, Con06, CB03, CY91, Dag95, DAM98, DRYL08, DK05, DL80, DB84b, EBA00, EBS86, FH86, FP07, Fok07, Fri85, FTS09, Gao04, Gat00, GW01, Gho95, Goo83b, Goo86n, Goo86c, Goo89m, GL90, GGSNR09, HK00, HZ96, HPY79, HCA96, HP95, Hsu02, Hua01, HL92c, IHM78, Jai85, JO85, Joh95, JG92, KHA04, KM06, KW04, KA85, KT85, KL09, LMB08, LB08a, LF97, LL99, Lee02a, LL06, LKKL07, LW82, LL07, LB08b, Llo93, MJ08, MMP08, MBE00]. model [McL78, MM08, MCCC04, Mou95, NC96, Nan98, NCA^{+}00, NL08, NM98, Oga07, Oht98, Pap93, Pas03, PNW06, PN98, QyT92, Rod07, SB86, SS97a, SGR04, SPK09, SKJ86, Soh92, SB00, ScK97, SD92, Sy01b, TAY02, Tho92, TA92, Tsu93, Vir07, VB81, WK90a, tKWY95, WB09, WK06, WW03, WW07, Wu01, XW07, Xio99, YKB86, YR06, YH85, Zha02, ZS05, Zha06, vdTGL97]. Model-based [Pet02, BG99]. model-robust [MBE00]. Modeling [BZB08, Law01, Sch02, Bow01, Gue89, Hun02, NK08, RB00a, Sha03e, Wet96]. Modelling [Pan99b, SCW07, CCNA09, CM05a, SZ02, SB96, TS04]. Models [BSS02, HM95, Kuk99, NK02, OVL02, OS02, Sah02, AHAH08, AM87, Ahm99, Akn07, AH09, AHAH06, AsY86, AC08, Alb85, AD01, AD03, AAR93, AL96, AMYY07, BS05a, BCV98, BB96, BG99, BT09, Bon05, Brä92, BS01b, BG07, CC05, CD92a, CYC99, CG91, CBPW97, CCNA09, CHH91, CR03, CW01, Che06, CBS06, Cho98, CDJ02, CR73, CCNAF95, CVS98, Cor04, Cor97, CS97, Cor00, CMH04, Dag89, Dam82, De 97, Dha85, Dic78, DM79, Dod00, DS09, Ebb73, oE89, Edw95, EW91, Eve01, FK04, FUOCN97, FCN99, Fon90, Fur91, GHCJ09, GS07, GG75, GW94, GA01b, Gil95, GS73, GS01, GLP72, GA95, Gri04, GSH04, HR07, HE00, HN80, HHI90, HV05, HMK00, HT85, Hie81, Hin97b, Hos91, Hou85, HCW07, HM98, HH93, Hud83, JG87]. models [Jen93b, JS00, JSS09, JJH09, KKE07, KS87a, KK91, KS91, KK93, Kel94, Ken79c, Kim93, Kim98, KL94, KP04, Kun92, Law01, Lee81, Lee90, Lee98a, LG09, LC92, Li01, LP09, LL09b, LK83a, LB96, LL08b, Liu81, LW04, LWH08, Lot82, Lou95, Lu97, Lye91, LP00, Mah96, Mak00, McG89, McK86, Men00, MGG78, MP96, NJG88, NS08a, NFD00, OR92, OP00a, OP00b, OPB08, Ouy06, Owe81, PK72, PM02, Pau92, Paw01, PR84, PT99, Qu06, QQ82, RRCD97, Rid03, RV08a, Rol01, RS90b, RVZRZP08, RV98, RWL95, SM03, SK08, SK87, SB05, She09, SAC06, Shi07, SBO81, SC09a, Sma05, Son97, SW90, Spe06, Spi98, Spi82, SN95, SJF06, SCK01, SCB07, SP01b, SS88, TS09, TT88, TTF07, TNG^{+}06, TL07b, TAM^{+}03, TC08, Tsi02, VGD06, VG05, VS93, VS94, VG01]. models [WLT08, Wan08d, Wil79, WO93, WS92, WB94, WM95, XW09, ZL96, ZL09, Zim89, ZNAEB98, Zue96]. moderate [El 82, MY90]. Modern [MH99]. Modification [Gha01, KE03, SL89a, Cor95, Edi83, LLKJ09]. modifications [LhKN05, yL87, NM98]. Modified [BVP90, GJ79, SY89, TT88, WECC00, BB80, BP01, DC00, DS89, EEÖ88, Eti81, Gan90, HH93, Law92, Lem87, LP00, MOS94, YX03, And92]. modulus [KW96]. Moment [Sho95, AASAM03, And89, JG80, KX03, KC97a, RG93a, RR93a, Tue01, WK90a]. Moments [AB82, Mel80, And95a, BS95, BD01, BR84, CJ95, De 81, Dod83, Éri97, Goo79s, GH82b, Goo83p, Goo83s, Goo83y, Goo84p, Hut77, IMLG09, Jon86, KPKPB95, Mar96a, Mar96b, Mcd74, Nor84, Owe77, PP03, Par81, Roh88, SKJ86, Sie78, SDP93, SH72, SS88, TA92, Zac80, Goo78c, Goo79r, Goo83o]. monitor [RS09]. Monitoring [CG06, Gan90, GS07, SL09, TA94, CG94, DE06, EK03, Shu08, Zha00]. monocular [Goo86i]. monotone [DP06, Die83, HP00, LLKJ09]. Monotonic [Goo84p]. monotonicity [GK05]. Monozygotic [Goo78i]. Monro [McL78]. Monte [LF04, AS00, ÁGR06, AT05, BF83, BP78, CG91, CS95a, Che06, CMD74, CM76, ES88, FPRS92, FW88, Fis73b, FC89, Fou81, Fun79, GHW99, GPW80, Gat91, Goo81w, Goo81o, Goo81j, Goo86o, HM95, HT85, Hig97, HM80, HL92c, JB91, JL88, JUP86, JH72, hJ93, KPKPB95, Ken79b, KP87a, KC97c, LF80, Lee90, Lev05, LN77, Lil01, Mae87, MM05, MGG09, MT80, Mil79, ML79, MM99, NL77a, ND84, NP98, Nic96, Nor91, OPS82, OB92, Pau84, PNM83, PQ84, Poi92, Poo80, Pot81, PS85, RRP06, Rid03, Rud86, SFSS85, SND89, SCW79, SCW81, SR97, SCM90, SW90, SC82, SB93, SS88, TZ97, TZ04, VM96, WB73, WZ95, ZL09]. Monte-Carlo [CMD74, HT85]. monthly [Mar83]. Moore [KMK87, Tre94]. Moran [BN01b, CT88, SM91]. Most [Phi91]. motion [Car07, DC99, Har87a]. motions [CM07b]. move [EB90]. Moving [KP87b, AG78, AK86a, BTL93, Cho98, Dam82, Del83, Gal02, Gan90, Gan93a, GA01b, KP04, Rig95, Ros06, SL09, Shu08, TA94, VGD06, WGC04]. moving-average [Cho98, Del83]. much [BCL93, WG92]. Multi [GR08, MH78, AB03, Bal83, Coa95, DBVK02, DH81b, GPW80, IMLG09, JSS09, LF83, LZNL08, PP06a, Pas03, SQ07, TB82]. multi-armed [Coa95]. Multi-attribute [GR08]. multi-dimensional [JSS09, LZNL08]. multi-factor [Pas03]. multi-factorial [AB03]. multi-modal [PP06a]. multi-sample [Bal83, DH81b, LF83, TB82]. multi-service [DBVK02]. multi-Smirnov [GPW80]. Multi-stage [MH78, SQ07]. multi-state [IMLG09]. multicategorical [TS04]. multicentre [NS09]. multicollinearity [IC80, SM96]. multicomponent [AP06]. Multicovariate [NS08a]. Multicovariate-adjusted [NS08a]. multidimensional [CC90b, FTS09, Goo81q, SBO81]. multifactor [HK92]. multiflicative [GH82a]. multilayer [BB99]. Multilevel [SNC09, CC05]. multinomial [BN96, Fre09, GG77, Goo78g, Goo81-29, Goo82l, GH83a, Goo84a, GL85, Goo86p, Hut77, JS00, Kwo96, Law92, Ma97, Nan98, Pap83, RV98, SH72, ZS05, Goo86p]. multinomials [Goo79j, GL85, GL88]. multinominal [Jon75]. multinormal [Bho73a, Dea86]. multinormality [GS87]. Multiple [Ano06d, CSAR93, LPZ02, AASAM03, AA09b, Arn79, CB07, Che01, CB97, CH88, DC02, DS89, FC96, GB99, Goo79e, Goo83m, Goo84c, GL90, Haw79, HJ85, HTC07, Hut93a, jLP05, LJZB05, MMW91, NP81, San89, SC95, SP00, SGH85, Som84, SOU04, Tan01, Tho06, WK75, WDR86, WGP00]. multiple-choice [Arn79, Goo79e, Goo83m, Hut93a]. multiple-group [CB07]. multiplications [Goo85d]. multiplicative [BLP09, KW96, O’N82]. Multiplicity [cDJgS93]. multiplier [Fur08, Lee81, Sef92]. multiply [DFY08, KH09]. multiplying [Goo85d]. multipopulation [KT94]. Multiscale [WGP00]. multiset [RS89]. Multistage [PB00, PNW06]. Multivariate [Ban78, CP91, ESRV98, Gat00, HHR02, LJDK02, YR06, AS89, Ahm92, ASY82, AS90, Ano06a, Ano06d, AZ05, BN01b, BPP00, BGS07, BW01a, BP78, Bho73b, BD01, BJ82, CD92a, CGA09, CP95, CT82, Che85, CH82, Dem06, DFY08, Die83, DCA03, EP92, ES88, Eve88, FSBN07, FJ84, Fun88, GB99, GK00, Goo82d, GH83b, GR79b, GN89, HR07, HL97, HS04, HL92a, Hut93b, J.83, JH01, KL05, Kwo95, LC92, hL92b, LZNL08, Loh86, LK83b, Ma99, MK74, MM05, MHS04, Nad99, Nat82, Par87, PRS87, RFGE86, Rig95, RS90b, SMO03, SKC75, She91, SGH85, Sol90, SW90, Wil08, Wor89, ZL09, vBBGOR06, Goo90d]. muscle [DV95]. Must [BH96]. Mutually [Agg87]. MV [Sin90b]. MV-optimal [Sin90b]. MVUE [HL98]. my [Goo83r, Goo83-27]. myeloma [HJ85].

N [BG78, LY96]. N-stage [LY96]. Nagasaki [HJ85]. Nair [Yan99]. naïveté [HS86b]. natural [Goo85p]. Naval [Goo00, LL73]. Near [May99, May06]. Near-saturated [May99, May06]. Nearest [GD92, Dog89, Dog92, LW95, Whi94]. Nearly [CNV02, VD08, Kem84, Sin90c]. nearness [CH90, KC89]. necessary [Goo93a]. need [Goo89k]. Negative [JS00, AJ82, CC01, DSL06, Fam98, Jon86, JJH09, LP85, Ma99, MW04, Pap83, SB82b, SL89b, SP01b, TTNC09, Van93, Zac80]. negative- [Jon86]. negative-binomial [Zac80]. negatively [EG89]. neighboring [SNB07b]. neighbour [Dog89, Dog92, GD92, GS01, LW95, Whi94]. neighbouring [SNB07a]. Nelson [CFOS02]. Neologisms [Goo81u]. nested [BJM92, BE86, BBG86, EBA00, McN08, Whi95]. nesting [Koh81]. nests [Gup84]. nets [Goo81k]. Network [Aok02, DSS06, DE06, HP95]. networks [AMYY07, DBVK02, DSS06, STG^{+}01, SL98b]. neural [AMYY07, HP95, STG^{+}01]. neuronal [Dog93]. neurotransmitters [CMQ03]. neutron [Goo89u, RS97]. Newcomb [Goo86l]. newly [Goo86k]. News [And79, Ano79m, Ano80o, Ano80p, Ano80q, Ano80r, Ano81q, Ano81r, Ano81s, Ano81t, Ano81u, Ano82n, Ano83d, Ano83e, Ano84f, Ano84g, Ger78, Hel78, HA78, HN83, Hel86, HZ88, SF77]. Newton [BR84, KC97c]. Neyman [Bog01, Goo82a, Goo83r, Goo84r, QD83]. Nielsen [Goo83v]. no [Hab84, Hos87]. Node [WB02]. noise [BLP09, CR03]. nominal [LL08b]. Non [DG02, Fur07, MBH91, Mun06, SD02, SJF06, Wei79, AP85, AAR93, And95b, BJM92, Bak07, Bar77, Bar78, BRL82, BD82, CHM09, CS95a, Che05, Chi79, DDB09, DFY08, DB84b, Fie93, Fri79, GK05, GH83a, GS84, Gri09, KP79, Lei83, Lev78b, LEG85, LW95, LOR04, Ma99, Mak00, MF02, MW04, MK97a, Mye98, Nic05, PYC93, Pet80, PB72, Rhi86, Rod76, SK08, SPK09, SA92, SP00, Sim00, SS92, Sma05, Spe06, SR00, Tan82, TT86, Tsu93, Wes72, ZA09, Zue96]. non- [Bar78]. non-Bayes [Goo83-29, Goo86g, Goo86d]. non-central [Che05, Fie93, Ma99, Rod76]. non-equiprobable [GH83a, GS84]. non-Gaussian [DFY08, Sim00, Sma05]. Non-Homogeneous [SD02, Spe06]. non-integer [PB72]. non-iterative [MK97a]. Non-linear [SJF06, Wei79, And95b, CHM09, Fri79, Nic05, PYC93, SK08]. Non-linearity [Fur07, Zue96]. non-monotonicity [GK05]. non-negative [MW04]. non-nested [BJM92]. non-normal [AAR93, Bak07, Bar77, CS95a, KP79, MF02, Mye98, SPK09, SP00, SS92, Tan82]. non-normality [DB84b, Lev78b, Rhi86, SA92, SR00, TT86]. non-normally [Mak00]. non-null [Wes72]. non-orthogonal [LEG85]. Non-Parametric [DG02, MBH91, Mun06, AP85, BRL82, Chi79, DDB09, Gri09, Lei83, LW95, Pet80, ZA09]. non-positive [BD82]. non-stationary [Gri09, LOR04, Tsu93]. noncausality [CM06]. noncentral [Cha94, CS94, CJ95, Éri97, GJ82, Gue78, pH78, Han79, Lee92, Mar96a, Mar96b, Pos94]. nonconforming [HL97, HL98]. nonfeasible [Dup96b, Dup96a]. nonhomogeneous [Har89, KW00, SL93]. nonlinear [CCNA09, CVS98, Cor04, DH77, DH81a, FUOCN97, HS77, HL92c, KKE07, Kar09, LL08b, ME72, MR03, Mou05, NCÁHC^{+}09, Pau92, SN05, Sel08, SC09a, SS88, TS09]. nonlinearity [TY90]. Nonmonotonic [CV07]. nonnegative [Wat77]. nonnormal [Bon06, CY99, DM93, Hon90, Sch83]. nonnormality [DS89, KKB85]. nonnull [GN89]. nonorthogonal [EL75]. Nonparametric [ACNT05, BDG04, CHáS03, CC90a, CMH04, Kap87, KS97, KP79, KP82, RS77, WK03, Weg72, Woo85, ASY81, Alt92, AB97, BP78, Bow85, BR00, CJ00, CM01, DP06, Die83, Dio81, DMF83, Far90a, Goo84s, Goo84n, HMS89, HT93, HB06, JG96, KC73, KS87b, Lai82b, LT90, Mar92, Neu07, NA09, PP06b, WM91, WDB75, ZC03]. nonrandom [MS95]. Nonregular [BCP02]. Nonstationarity [JH90a]. nonstationary [KS91, Rid03, SP97]. nonsymmetric [Bar79, Cro74]. nonzero [KP99]. Norm [Wu02, OP04, SA95]. Normal [Gne97, Ahm92, ASS04, AS90, AAR93, Atk92, Bak07, Bal89, Bal92, BPP00, Bar77, Bar78, BCX93, BCL93, Beh72, Ber73, BGR94, BB96, Bho73b, Bon81, BD09, BG07, Bre93, Bru75, BS94b, CP95, CCNA09, CT82, CS95a, CGN04, Che85, Cho08, CKM01, CH06b, CAT78, Deá80, DM93, DW90, Dut99, DL75, EW91, Eti81, ES88, Eve01, FB90, Fis73b, Fon92, GM79, GsW01, Gao04, Gat00, GK00, Goo78k, Goo81z, Goo87c, GT78, GC03, HS73, HL97, HL98, Hat86, HH92, HD77, Hos78, Hut93b, KP79, KP82, KN89, KR89, KL05, KTM05, Kwo95, LMB08, LH97, Lev82, hL92b, LB96, jLP05, MF02, MG80, MK74, Mat79, MK97b, Mye98, Oga07, PP06c, PT99, Pul79, RA81, RS90b, RJ06, RW93, SM03, SMO03, SB73, SSM95, SPK09, SP00]. normal [SS92, Sol90, Som85, SW83, SLL00, Tan82, TW91, TKS78, VS93, VS94, WK90b, Wan08b, vdTGL97]. normal-theory [Oga07]. Normality [HHR02, SD02, And80, BM86, BN95, BDB08, CS95b, DFY08, DB84b, DL92, EA78, EEÖ88, Eve88, FRS06, FSBN07, Goo96a, HL92a, Lev78b, Loh86, MM05, Mei09, Par92, PRS87, Poi92, Rhi86, SA92, SR00, TT86, Val07, YY07, You93, Che03]. normalization [Str89]. Normalized [Aly90, CS95b]. normalizing [GA95]. normally [Goo87c, Mak00]. normals [JS02]. normative [Goo89k]. Norms [WM99]. northeastern [FcC05]. notation [And92]. Note [Jam01, MM80, WM90, AN03, Ano00j, Ano00k, Ano00l, Ano06g, AK86a, BF87, Baí09, BC94, BR92, BD84, BP92, Che94, CY91, CH02, Dev82, Die94, Ebb73, Fis73a, Fre07, Fri79, Goo83l, Goo84q, Goo85f, Goo86s, GL88, Goo88g, Goo92c, HT83, Han94, Hoe89, HSSY04, HSWF07, KA82, KL09, Kwo96, Lac90, Lam05, LL99, Lee99, LS86, Lei83, Lev05, Mad77, MG06, Mar98, McK86, Mur78, PT99, PS09, Rid80, SAJ08, SL87, SA83, Smi78, Tho89, TKT89, Wan99, Wec94, Wil98, ZR93]. Notes [YKB86]. noticed [Goo84r]. NP [Wel82]. NP-hard [Wel82]. nuclear [CC89, RT88, RT90]. nuisance [Lil01, Pep93]. Null [Fra88, Coo07, Fou80, Goo85b, Goo94g, HTC07, Jöc81, LMB08, MY90, PGT09, SA92, Wes72, ZY04]. Number [BD76, LL73, Bay90, DM05, Fle95, Goo80a, Goo81-29, Goo82l, GH82b, Goo83i, Goo85j, Goo85d, Goo87i, Goo88f, Goo89d, HM88, HTC07, KW96, Kun92, Kun98, LSSP08, MBH91, NO75, Pad78, Pau84, Pes80, RH76, Sah79, SSSB00, Tan01, WL07a, Won85]. numbers [AAR93, Goo79b, Goo93d, Goo94c, Goo98a, HJ85, Kru03, L’E97, Lak81, PX97, SC95, Son97, WD93, Zur93]. Numerical [Bar72, BH82, CAT78, CNFO05, DL81, DD77, Dut77, DH81a, GB99, GK00, GS73, KS87b, LDR92, PS85, She91, TL07a, TA92, BD84, CBPW97, DSL06, Dee76, Goo85d, Mal06, Nor84, PX97, PP80, PQ89, PV93, Rod76, TA08, Dan79]. numerological [Goo89t, Goo89u]. numerology [Goo88f, Goo89v, Goo92c, Goo94a, Goo95a, GJH90]. nutrition [Goo78b].

O.K. [Goo82e]. object [Dog92]. objectivity [Goo84s]. observation [SND89, WL07b]. Observations [LJDK02, AAR93, Bho73b, EH92, EH07, Gan90, Goo83n, Goo89h, GL90, HSW75, Hos91, HM98, JG83a, KL93, Lac92, Lee99, NCÁHC^{+}09, OPB08, Pad82, San89, Sch83, SL09, Van87]. observe [FMC09]. observer [EB90]. obtain [SNB07a]. obtained [Tad81]. obtaining [BD82, Dup96b, Dup96a, KA82]. obverse [Goo94d]. occurrence [She09]. occurring [Dan80a, Goo84w]. Ocean [CG97]. odd [Goo83t]. odds [Alb87, ABH82, Pig91]. oedemology [GG78]. off [BS94a]. offs [GA95]. old [LM74]. OLS [Sha87b]. omissions [Goo80a]. omitted [Oht98, SGW94]. omnibus [ES86]. Omnitab [JCM72]. On-line [EL75]. One [Abd95, HS73, Kru88, RR77, RVZRZP08, SS99, YW02, AT95, ABH82, BE86, Cao87, CR93, CG94, Cic89, CCNAF95, Don97, Dow02, Edi83, EEÖ88, FBC09, Goo78c, GL83, Goo86j, HK00, HK84, JO85, Kwo95, LKKL07, LS86, Lem77, Loh73, Mar92, NO75, Nat82, PQ84, Pie94b, Pos79, SL09, Tan01, WB73, WW03, XM09, Goo84e]. one- [Nat82]. one-fold [BE86]. one-parameter [CCNAF95]. One-sample [RR77, CR93, Cic89, Edi83, EEÖ88, Goo78c, LS86, Lem77, Loh73, Mar92, PQ84, Pos79, WB73]. one-shot [FBC09, HK84]. One-sided [HS73, RVZRZP08, SS99, CG94, Don97, Kwo95]. One-step [Abd95]. One-way [Kru88, Dow02, Goo86j, HK00, JO85, LKKL07, Pie94b, RVZRZP08, WW03, XM09]. Online [EK03]. only [Goo84s]. only-just [Goo84s]. operating [PSB03, Sho86]. operation [War74]. operations [Sha05f]. operator [Gri92]. operators [Hig97]. ophthalmologic [AB08]. opposition [Dav79]. optical [GS78]. Optimal [CG94, IC00, KW04, Pot81, RW77, Wor89, Ang03, BAB86, Gan93b, HH92, HMK00, HP95, HS91, Jen82b, LiT95, LGPP96, MRR84, REK81, Sin90b, Won95]. optimality [Jen83]. optimization [Hua85, PK72, SCW79, SCW81, Sy01b]. Optimizing [RS88b]. Optimum [RM02b, CZ02, Jai85, JG83b, MM08]. optional [Goo91]. options [CCY04]. Orcutt [Alt92]. Order [Bal85, Gal02, GV02, LB07, LT95, OS02, Ros06, Sch72, Ada91, Agg87, AC00, And92, BS95, Bar77, BHZ08, Che03, Chi00, CVS98, Dag95, DL80, EN90, EK98, FUOCN97, Fre07, Fur07, GP07, GV81, HSB85, HS86a, IMLG09, JR96b, Kim98, KP99, Mas03, Mye98, Nic96, O’N82, Par81, Pol94, RT78, RA01, RS88b, SZ02, SMO03, She83, SH72, Tse84, TA92, Vau94, Wil79, YL85, ZNAEB98]. ordered [BB96, BK96, HM85, KAS96, KTM05, Law01, LB08b, LiT95, MW84, Mar76, Poo80, RM96, SHW93, TB85, WL07b]. Ordering [RB00b, BN96, KG80]. orderings [Goo80a]. orders [Tan01]. Ordinal [Pet02, AB08, BD96, BZB08, Dem06, MGG78, RWL95, TAY02, TH05, XYT08, BF87]. ordinary [Lee02b]. organic [Goo98a]. oriented [Wu01]. original [YM90]. orthant [Di 05, DL75, RW93, Sol90]. orthodox [Joh90a]. Orthogonal [Tay90, Abd95, Agg87, DN94, Gup84, LEG85, SB91]. orthonormal [hL92b]. other [And80, AE87, CSS83, Dow02, Goo83e, Goo93b, MH72, ML79, TR75, VC78, WNB07]. outcomes [DFY08, EP92]. Outlier [DE06, IM82, Sim87, CP91, CK04, Haw79, MH99, Pop08, SM96, SCB07, Wil08]. outlier-multicollinearity [SM96]. Outliers [LH93, Ano78g, BG94, CP95, CB97, CBS06, Fun88, Hou85, ZY04]. output [BB99, CH98, Con06, PCS09, Phi97]. outputs [kSwXR93]. Outsourcing [Sha06d]. Outstanding [Sha07j]. over-dispersed [CYC99]. over-relaxation [Hun87]. overall [BH95]. overcome [Win75]. overdispersed [vdTGL97]. overdispersion [Hin97b, Joh95, JJH09, Lu97]. overflows [Goo78j]. overparameterized [HH93]. overshoot [CP76]. overview [MBE00]. Owen [RVZRZP08]. ozone [DE06].

P [BBHW95]. P-values [BBHW95]. Package [LL73, Sha03e]. packing [JT80, Lot82, Zhe88]. Page [Hol85b]. pair [Phi91]. paired [DA87, Jun08, Mun06, O’G09, SCAH05, Goo84f]. paired-comparison [DA87]. pairs [GL84, GL88, Shi85]. pairwise [BH95, HM98, MG06, Ng96]. Palm [Chi08]. panel [GOS09, Jön07, Kim07, PPRW06]. paper [Bru75]. papers [Ano73a, Ano73b, Ano73c, Ano73d, Ano74a, Ano74b, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano76a, Ano76b, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano82a, Ano82b, Ano82c, Bro74, HL84]. paradox [Goo80k, Goo82a, Goo86f, Goo89h]. paradoxical [Goo94h]. paragraph [Goo83-29, Goo83-30, Goo83-28]. Parallel [PK72, Con06, KS04, PSW98, Goo84t]. parallelism [SFSS85]. Parallelizing [MMW91]. Parameter [BS92, BS02, Kap83, Liu08, Sho95, Ada97, AHH07, And97, BLC04, BG94, Bha01, BG07, CH06a, Che97, CC09, Cob89, CCNAF95, CVS98, Dag89, Dup96b, Dup96a, DW01, FFCN07, FFCN08, FL82, FRB06, FTS09, Goo86p, GGSNR09, JK08, Jam01, KKE07, Kk90, KPH05, KH09, KH04, KC97a, LSCN08, LS99, Luc08, ML74, Ng96, Pep93, PB72, PW83, RP01, Rod76, Rol01, SN83, SA04b, SC09b, TH09, Tza09, UP01, Van93, WK90a, WW06, Wan08b, WJ05, WLK06, Win75, WL90, Woz94, Wu03, WWL09, WBE80, Xio99, YX03, YXW07, YAAB87, Zan79, ZK86]. parameterization [KH04]. Parameters [YT96, Ada96, ABE83, ABE85, AHJ92, ANAA97, BP86, Bha01, BS01b, Dav93, De 97, DRYL08, Die94, Fro95, GZP05, HV93, HH96, HZ03, HS91, LL91, LB08b, LS99, Lil01, Lou95, MW84, MTO08, MS99, MKDM94, Ng96, NW09, Nie06, Pes80, Pol94, Pou04, PW83, QC03, RA01, SS86, SM94, SPK09, SB82a, TAY02, Tay94, Tue01, VM00, VB81, Wor89, WL90, Woz94]. Parametric [DG02, HL09, PP06b, PN86, AP85, BRL82, BG01a, CG97, Chi79, CMH04, DDB09, Dey84, FRB06, GG75, GO03, Goo87h, Gri09, GSH04, Lei83, LP09, LP92, LW95, MYS01, MBH91, Mun06, Pet80, QQ82, RB00a, ZA09]. parent [Fro89, VW78]. Paretian [LL93, ST87, Tsi02]. Pareto [BS01b, Dup96b, Mou01, SA04b, Wu03]. part [And95c]. Partial [Bén94, Sho95, AL99, Goo83t, Goo83y, Goo92e, JAK93, Leg00, Mag75b, MP81, MS95, RW96, TC73, Goo79q, Goo92d, Goo92e, Kru79]. partially [AHAH08, DL80, Goo89v, KHA04, LP09, PK72, Qu06]. partially-censored [DL80]. particle [CM07b, RS97, Sel08]. particles [How85]. Partioning [GW94]. partition [TO04]. partitioning [GT90, Mur78, VPB03, Goo79f]. passage [Di 05]. passing [How85]. Path [Aok02, Goo80m, DF80, Goo78h]. patients [RL08]. pattern [Sy01a]. patterned [WM95]. patterns [DM79]. PBIB [Sin90a]. peak [Kös06]. Pearson [Ant95, Bog95, DH77, Goo81j, Goo81-29, Goo82o, Goo82l, Goo82a, Goo83r, Goo83k, GS84, Goo90j, GJJL02, KG80, Pos78, Pos79, Pos82, PO06, Ray85, Rud86, SC09a, Wes72, YAAB87]. Pearson- [KG80]. Peirce [Goo81c, Goo83r]. Peizer [YI01]. penalization [CRT07]. penalized [GR79a, Goo83-28, Goo83e, Goo87h, LY96, LY99, Goo84v]. penalties [Goo81q]. penalty [Win75, Wu01]. penetration [NBB00]. Penrose [AG85, KMK87, Ken79e, Tre94]. Percentage [BCX93, Bho73b, SWK73, CT88, Dan87, HB78, KC96, Kwo95, Mat79, Par92, SKC75]. Percentiles [CL91, BJM92, GPW80, KW78, PT03, ST74, Kru86b, Kru90b]. perceptrons [BB99]. Perfect [Dög01, CS05]. Performance [BP01, CC05, Cor95, HMS02, Kib04, LG09, OVL02, SHH85, Var81, WLT08, ÁGR06, Atk92, BG78, BF83, Bel93a, BS01b, CM98, Dag89, Dag95, DK05, HCW07, Hua91, Khe08, Kun98, Lac92, Mar83, MB86, Moo90, Oht98, RV08a, STG^{+}01, SHW93, SL88, WB09]. performances [Don97, Fro89]. Periodic [TC75, AH09, Sma05]. periodicity [AHH07]. periodogram [Gle97, McS06]. Permanent [RMH88, SD01b]. permanents [GL83]. permutability [Goo86h]. Permutation [AB03, AL99, CRM06, Dow02, Gil07, Goo82f, GP95, JR83, Leg00, Too72]. permutations [Goo81d, O’G06]. perspective [SP01a, TP06]. Petersburg [Rod06]. pH [DV95]. Phase [Con95, LL09b, Sch83]. Philos [Goo89p]. philosophy [Ano06b, Goo83-29, Goo93e, Goo94f, Kru89b]. physical [Goo88f, Goo92c, Goo95a]. phytoplankton [KT85]. piece [Goo88f]. pieces [Goo89a]. piecewise [Arc80, HV05, SA95, WmGT95]. Pitfalls [BB75]. Pitman [CH90, KC89]. pivotal [XM09]. Pivots [GV02]. plan [Jen03]. plane [CM76, HQ00, VGD06]. Planning [EM86]. plans [Bal95, RW77, SSSB00, SNB07a, SNB07b]. Plausibility [DFY08, Goo83v, Joh90b]. play [CI81, WC91]. play-the-winner [CI81]. pling [Goo86e]. plot [EEÖ88, EH07, FC96, RJ95, SJ07]. plots [MT01, QH80, Wet96, Won95]. Plotting [Por88, LG84]. plug [ACNT05, PB00]. plug-in [ACNT05, PB00]. poem [Goo86k]. Point [GV02, PS99, BK98, BHG01, CH98, Chi08, CM07b, Dog89, Dog92, Goo82g, Goo83g, Goo86c, GS01, Har85, HH74, KK91, Mal06, NS08b, OP04, Pet80, WO97]. point-object [Dog92]. Points [RG02, Ano06d, BCX93, BS79, Bho73b, Che82, CT88, Dan87, FcC05, HM88, HB78, KPSW83, KC96, Kwo95, Mal06, Mat79, MBH91, Par92, SWK73, SKC75]. Poisson [GF89, AL93, AL94, ANAA97, Alb83, Alb85, BB74, BRT07, Brä92, CC01, CJ73, Fam99, FGSV09, Fro04, GW01, GAM09, GBSS84, HQ00, HM80, Ho93, JSS09, hJ93, KCS88, KX03, KC73, Kk90, KW00, Lu97, Mel80, MM99, Nue08, Pap83, PPK77, SM94, SL93, Sim93, Smi89c, Soh94, WV79, WO97, XW07, XW09]. Poisson-distributed [CJ73]. Poisson-gamma [GW01]. poles [YKB86]. Policy [Ano88e]. polychoric [Alb92, Lee85, LL88]. polygons [CM76, HM80]. Polyhedral [SS95]. Polynomial [KP87b, AR01, DN94, FH09, Gle89b, KP99, LWH08, SB08]. polynomials [BR84, Goo85d, GT78, WmGT95]. polypeptide [Goo82k]. Pooling [AS89, Ahm94]. Poor [BD76]. Popper [Goo89e, Goo89g]. Popper/Miller [Goo89e]. Population [SG96, AASAM03, ASS04, AkBA05, Bak08, BM90, BF83, CR93, CT82, CM01, Fro89, Goo84a, GR79b, HS73, HSW75, Hon90, Kri77, Möh05, OBW05, PO05, Rhi86, SD01b, TW08, Val07, VS91]. Population-averaged [SG96]. populations [Bak07, BF08, BN96, BD09, CG77, Chu80, Cla96, GM79, GM77, GH07, GN89, HPY79, HM85, Mat79, Mou05, RA81, SW75, Som85, TKS78, VS91]. Portmanteau [FMOR06]. positions [LG84]. positive [AS01, BD82, DSL06, Ebr93, FS94, FL82, Goo81v, Goo83n, GR79b, MMP08, QFG87, Smi89d]. positive-definite [QFG87]. Possible [CRR99, GS84, Goo84j, Kim00, Liu86, NTK09]. possibly [Goo86k]. Post [KC97b, Pet88, SLL00, SOU04]. Post-processing [KC97b]. posterior [FT96, HZ96, Lat82, TPG93, UP01]. posteriors [DL81]. Postgraduate [LL73]. postmortem [DV95]. postulate [Ano82d]. potential [Goo79g, Goo83w]. potentially [Soh00]. potentials [TBT95]. Power [CL98, HH92, Pet88, QM77, SJR07, SD02, Woz94, AAL02, Ano05e, Ant95, Bal83, BK82, BK84, BB74, BA77, BL78, BDB08, BHG01, Cho08, CV07, CJ73, DDB09, Die83, Dod83, DH81b, FMOR06, FP07, GB99, GA01b, GA02, Goo78h, Goo79l, Goo83d, Goo83l, Goo85q, GS86, HK07, HTC07, Jen95c, JG83a, Joh79, Kan75, Kap93, Kat78, KC73, KW04, KAST05, Lev82, Lev78b, Lev78d, Llo05, Loh73, Mar86a, Mar87, NP98, NS08b, Owe81, PD03, PP85, Poo80, Pos82, Rod76, RR79, SAD03, SB05, Sen02, SA92, Smi78, SSM93, Tay77, TC92, Wes85, YH85, YST90, Jen93a, Jen95b]. power-divergence [Ant95, FP07, PD03]. power-shift [BK84]. power-summing [Goo79l]. powerful [Goo81i, LhKN05, Phi91, SAD03]. powers [Goo93d, Hab80, KT97, Mar76, SGW94, SCB07, VW78]. pp [Ano03b, Mar90b, Sar98]. Practical [GV02, Day87, SC95]. pragmatic [Goo88e, NW83]. Pratt [YI01]. Pre [CH99, GK05, VS94, DG95, HBC07]. Pre-checking [GK05]. pre-smoothing [HBC07]. Pre-test [CH99, VS94]. pre-testing [DG95]. precision [And95a, Bur74, DM78, Han78, MMR92, Sch74, WNB07]. Prediction [AP85, AR00, BG01b, HM95, WL07b, BL02, BLC04, BL05, Goo86p, GM77, LL99, Lee99, Lia11, Mee93, Mou01, Pad82, RS77, RM02a, Sen02, Sha01b, Wan08c, YM90]. Predictive [STG^{+}01, LZNL08, WG92, Goo79o, Goo83a]. predictor [LJZB05]. predictors [EH07]. preference [Goo89j]. preliminary [Ahm92, EA78, GL92, HS77, Kib04, VC78]. preposterior [Goo83-30]. preprocessors [FL82]. presence [AL94, BK08, CM04, Dre08, FL82, GOS09, Hel97, NM98, Pep93, Pou04, PB04, WGP00, Wil01]. prespecified [KB87]. prevalent [She09]. prevision [LDR92]. pricing [CCY04]. primary [LL79]. prime [KW96]. primes [Goo80b, Goo85s]. principal [GS85, JMY96, Krz82, Krz83, Mar90a]. principle [CH06a, Goo83v]. Printer [Har75]. Prior [Adk96, Alb87, CS86, DP93, FB90, Fea79, Kem84, Lai82b, SdFdGCM08, Sny88, Soh96]. priori [Dha85]. prioritization [Hor97]. priors [AS04, Goo81z, Goo83-28, Kem84]. Probabilistic [CBPW97, BH96, Goo85n, Goo93e, Goo85a]. Probabilities [WK02, Alb84, Ant95, BK08, BPP00, Bea85, BD82, BS01a, BH82, CG77, Con95, Dea86, Di 05, DL75, ES88, For97, Fre09, GB99, Goo78g, Goo79o, Goo82n, Goo84d, Goo84j, Goo85q, Goo85b, GL85, Gue78, Hab92, KAS96, Lee92, LD87, Ma99, Mag75a, NA09, Par81, Par87, QD83, RV98, RW93, STS94, SNB07a, Sol90, SP86, SMV76, Ter87, TT82, WM91, Goo83y]. Probability [Jon75, Kuk87, LM84, Mur83, Sy01b, Ano03b, Arc80, BB80, Bru75, CC89, Che82, Dev84, Dup96b, Dup96a, Edi83, FS81, Fre07, Gne97, Goo78d, Goo80b, Goo81b, Goo81-27, Goo81q, Goo83t, GL84, Goo84i, Goo84k, Goo85s, Goo86l, GS86, Goo87i, Goo87a, GL88, GHDB89, Goo90l, Goo94e, Goo95d, Gue82, HK07, KN89, KG80, KW94, Kuk89, LG84, Pap80, RT90, RS90b, Sha06c, Smi89d, Tan82, Tue01, Ven89, VR92, WK90b, Weg72, WS90, YL85, YM90, Goo82c, Goo83e, Goo84m, Goo85k]. probe [HCW07]. probit [Dag95, Jos01, Spi98, Tah90]. problem [AZ05, BC82, BS72, BK96, BR03, CZ02, CA89, DM78, Di 05, DH81a, ES86, Fun79, Gol72, Goo79c, GBSS84, Goo86j, Goo86l, Goo86e, GJJLGS06, HC75, Hos78, Ken79e, Kru72, Lem87, LT90, Lot82, LGPP96, Nat82, NJJ92, QFG87, RS86, SM96, Ter90, Wan08d, Wil89, WS92]. Problems [BS02, CYC99, CSS83, Dut77, EL75, Goo79s, Goo90b, Hua85, Kem84, LF83, LK83a, SGH75, TA08, Wel82]. Proc [Goo89p]. procedure [Alt92, AB82, BF83, CH76, Fro95, GR79a, GSH04, KK05, KHSG83, Lev82, Lev78c, Li94, Mad78, McL78, MA74, RVZRZP08, Ros95, SC95, SA95, Som84, SB96, TT88, Tho74, TWLC07, Won85, YI01]. Procedures [PB72, AD01, AR01, BK82, BH95, CB07, CH06b, DA87, Deá80, DR02, Don97, Dow02, Fro89, Goo84s, HG85, Hon90, Hua85, JG96, KP79, LL91, LF83, MMWM83, MGG09, PN86, Pot81, RJ95, SFSS85, Sha01b, kSB90, Spi06, TB86, TT86, TB85, TKS78, VC78, Wil04, Wil79, XPC03, ZA09]. process [ANAA97, AG78, Ans80, AK85, CP76, CG94, CK96, Chi08, Del83, Den77, Dre08, EN90, Fro04, GR08, Gle91, Goo82i, Goo89m, GG90, HM80, Ho93, HKT04, JW97, KK91, KK93, Kel94, KA82, LD87, LL03, Nic96, Nor84, PX02, PX05, Pol94, RS09, RB88, SB86, Sen02, SL93, Shu08, Soh94, TL07a, TWLC07, TH09, WO97, WR98, WR94, Wri95, YL85, Zha00, ZL07, TL96]. Processes [KP87b, BG01b, BK98, CSS83, CMQ03, CM07b, Cra05, Cre89, De 81, Dog89, Dog92, DCA03, GA02, GS01, HT83, HP80, Jen83, KP99, KW00, LOR04, OHN93, PX05, RRP06, SdFdGCM08, SL89b, SP97, SB96, TA94, XYR09]. Processing [Phi97, KC97b, RS88b]. product [AASAM03, BS95, CAT78, PF04]. production [MMR92]. products [DD77, Goo88b, Sha87a]. professional [Goo88h]. profile [BP78, FFCN07, YX03]. profiles [CW74]. prognostic [NCA^{+}00]. program [How85, NC72]. programming [BPP00, DH77, Goo90f, Kim93, NJJ92, WCEC94, WECC00]. programs [Dee76, KM83, NSG91]. progressive [BKLW04, BH07, Wu03, WWL09, YT96]. progressively [BL02, BL03, Mou01, NW09, SB08, Wan08a, WCC07]. Projection [SS95, Goo82d, MW94]. projection-tomographic [Goo82d]. projections [Dyk85]. Prony [MOS94]. Proper [Goo79e, Goo79m]. Properties [CC01, NL08, Agu99, And90b, And95c, ANPV97, And80, Bar78, BR92, BS72, Brä92, BB89, CGA03, CR03, CMD74, Dio81, EP92, FJ83, GL92, GH82a, Goo84p, GM08, KB96, Loh73, Lye91, PQ84, Rid03, SMH97, SE02, TY86, Wil08, JH90b, JH90a]. property [GJJL02, Jen82b, Par80, Sha87b]. proportion [HL97, HQ00, MZ08, PX02, PO06, TW08, TTNC09]. Proportional [Kuk87, Atk91, CSC00, Gho95, Goo87c, Kuk89, Lou95, QyT92, SB05, Smi84]. proportionally [EH92]. proportions [Hab92, IP86, KP87a, Kwo96, PT99, RS77, TS00, TA08]. proposal [Goo79l, Goo82i]. proposed [PQ84, Wil08]. propositional [KW94]. protein [SA83]. proteomics [Sha09h]. Protom [Goo82d]. proton [Goo89u, Goo89v]. provided [Goo81x]. Pseudo [CS05, Alb83, AL96, BH85, Fle95, JW80, Lak81, PX97, WO93]. pseudo- [JW80]. pseudo-Bayes [Alb83]. pseudo-disturbance [BH85]. pseudo-likelihood [WO93]. Pseudo-perfect [CS05]. pseudo-random [AL96, Lak81, PX97]. pseudorandom [Mac92, Pes80]. psi [Kat78]. Psychometric [Sha07j]. publication [Goo88a]. pupillary [SMP^{+}06]. PVAR [Akn07].

QH [Goo96a]. quadrant [Ebr93]. Quadratic [Chi79, AS01, Bel93b, Car07, Cha79, DC99, Fan03, FJT82, GJ82, GS83, GV81, Lee98b, PLD88, RHH80, SLL00]. quadrature [KTB82, PB93]. qualitative [PK72]. quantal [CC85, CC90a, Goo90o, HSSY04]. Quantile [RB00b, CRM06, Fur04, Fur07, HSB85, Kar09, MP81, Tsi03, VM00, WK06]. Quantiles [TB82, Adi98, AB09, BDG04, BSK90, CSS83, Gho95, Goo79g, HT93, Har87a, Mag75b, MMPP05, NW86]. quantitative [AD01, AD03]. quantities [XM09, YXW07]. quantization [Sel08]. quarter [VGD06]. quasi [FP07, MPPZ05, Rod07, ST88, SJ06]. quasi-differenced [Rod07]. quasi-independence [FP07]. quasi-likelihood [ST88, SJ06]. quasi-symmetry [MPPZ05]. quasimodel [Goo86c]. question [Goo83y]. questioned [Goo81a]. questionnaire [BS05b]. questionnaires [Goo90c]. queueing [WAP84]. queuing [DBVK02, RH76]. quick [Goo84n, Auc79]. quotients [Goo83t, Goo83y].

r [AG81]. RA [OVL02]. race [SKC75]. radiation [SZ02]. Radii [Dög01, Por88]. Ramanujan [GBSS84, Mar86b]. Random [BD76, BH85, Dev92, DM05, Dög01, FS03, Hud83, JT80, LL73, LSSP08, RH76, YM99, AM87, Ahm87, AL96, Bay90, BC82, BG01a, BD96, Cai09, CM76, DF80, Dev82, Dev84, DN94, Fle95, Gho95, Gle89b, GH76, Goo79p, Goo80j, GR83, HM88, HK00, HZ96, HM80, Hör93, JO85, Joh78, JW80, JL96, JSS09, KS85, KPKPB95, KW96, L’E97, Lak81, Lau79, LKKL07, LC92, hL92b, Lia11, LL08b, Liu81, Lou84, Mal06, MBH91, MW04, Mur83, NO75, Nar90, Nor91, NC72, Oga07, Pad78, PX97, Pap80, PS98, Pas03, Pau84, QT92, RS90a, RVZRZP08, RS85, SM03, Sah79, SWK73, SND89, Sho86, Sim93, Smi94, Soh94, Son97, SB96, Tan01, TL96, Wan08c, WD93, Wil98, Wil07, WW03, WNB07, WCC07, WWL09, YT96, Zhe88, Zur93, Goo90c]. random-effects [LC92]. Random-response [Goo90c]. random-sampling [BG01a]. randomization [Bai89, BW01b, CB07, Goo92b, Hay97, Pie94b, WS04]. randomized [LPSL05, Too72]. randomly [AEL03, JB91, PP73, Goo81k]. randomness [Chi07, Goo95c, HC06, RV08b, SA83]. range [BCX93, BHG01, CH88, Goo85k, Goo03, HH92, Rhi86, Som84]. Rank [CRM06, Pan98, SA04a, Bel93a, BJ73, CMD74, FS90, GO90, Goo85o, Goo94i, Jen91, KP82, Kös06, Mad77, Mag75b, MP81, MB97, MMWM83, Nat82, PN86, SJR07, SMV76, Tho92, WB73, WA72, ZQ97]. rank-like [WA72]. ranked [ANM09, HMZ05, HSWF07, LB08b, Mel80, OBW05, SAD03, TW08]. ranked-set [HSWF07]. Ranking [Fon90, RW96, TP98]. rankings [CG77, Mar81]. ranks [HB06, MH78, Mun06, PR84]. Rao [SB93]. Raphson [KC97c]. rapid [TC92]. rapidly [Kat78]. rare [DM05]. Rate [McL80, AHJ92, AR01, Bel93a, GK90, GS83, Goo84a, Goo89c, JG92, KP09, KAST05, LSL97, MYS01, PT89, PLD88, She09, Soh94, SLL00, WWL09, YKB86, YL01]. rate-estimators [KP09]. rates [Hon90, KL93, KP87a, Lee98a, MM05, Tsi02, ZA09]. Ratio [CH99, PB02, ABH82, Ant95, BCV98, BB80, BB74, Bon06, CZ07, CKM01, CC97, Cor04, DAM98, DMF83, Fok07, GG77, Goo79f, Goo86j, Goo89v, GN89, Hut77, JL96, KK99, KN89, KG80, Lam05, Lat82, LL79, MKDM94, Nag75, PNM83, Pig91, PT99, RML90, Roy93, Rud86, SBC03, SF93, Shi85, SN83, SB93, tKWY95, WK72, WWS04, YST90, ZY04]. ratio-of-uniforms [JL96, YST90]. rational [Mar86b]. rationality [Goo89k, Goo94k]. ratios [Alb87, Her75, JG80, LH72, NK07, RG93a, SB73, WK90a, Wil75, YI01]. raw [CL91, LFC92, SS88]. Rayleigh [Ada97, Jen76, KH09, RM02a]. re [Goo79o]. re-use [Goo79o]. reactor [CC89, Goo88d, RT88, RT90]. Read [Rud86]. ready [Ken79a]. real [GGSNR09, PL99]. reality [Gup73]. realization [Cao87]. really [AG81, Goo94h, Kru03]. recapture [CL97, Llo93]. recently [Wil08]. receptor [RB92]. Recognizability [Goo86h]. Reconstructed [Mac83]. reconstruction [Gle89a, Tay94]. Record [BG01a, AR00, Jah05]. Record-breaking [BG01a]. records [AB09]. Recovering [Gou93, Jen82a]. rectangle [WK90b]. rectangles [Agg87, Dea86]. rectangular [Fre09, Hen95, SSSB00]. Recurrence [BS95, CJ95]. recurrent [BD96]. Recursive [PV93, Rol01, WM95, Dic78, EN90, Goo85d, NS08b, San89, Tan01]. recursively [Jön07]. red [Goo88a]. Reduced [Chi08, GP07]. Reducing [Mai03, SS95, KPKPB95, Kru86b, SSM93]. reductio [Goo81b]. reduction [AJ82, BVP90, CCY04, CY89, CY91, Dor01, Fis72, Fis73b, Goo90d, NS86, PS09, RS85, SGH75, WAP84, XDZ09]. redundancy [MMW91]. redundant [HKL08]. refereeing [Goo82i]. reference [AB08, Goo90i, KJ09]. reflecting [Gan93a]. Refutation [Goo93b]. regarding [Goo85f, Goo86s]. region [Alb83, SSSB00, Yan98]. regions [Ars86, ZR93]. Regression [Ano06e, DG02, LG84, Sch02, SS95, SB92, SD02, Woo85, Wu02, Abd95, AH90, Adk96, AN03, AsY86, AS81, Alt92, AL99, AS04, AKW92, Baí09, BL09, BGS07, BCV98, BT09, Bon05, Brä92, BR00, CDF05, CRM06, CC05, CZ02, CHM09, CHH91, CR03, CW01, CJ00, CCMV07, CH90, CVS98, Cor04, CJ73, DPS01, DC02, DS89, DV86, DRYL08, DR94, DR02, Die05, Die06, DB84b, Dut77, DH81a, oE89, EL75, Eri83, FK04, Far90a, FFCN08, Fir97, Fra74, Fri79, Gbu81, GMSS95, Goo80f, Goo92a, Gra86, Gri92, GB86, HE00, HP90, HMZ05, IC80, Jam01, Joh95, KKE07, Kar09, KC89, Ken79c, Kib04, Kim93, KYF08, KJ09, KSK93, Kul90, Lam05, LSS93, Lee02b, LL06, Lei83, LS88, LB08b, LL09b, Liu81, LJZB05, LWH08, Lou95, Lu97, Mak00]. regression [MRR84, MBE00, McG89, MA74, Mee93, MT80, MMW91, MH07, MCCC04, NW83, NW86, NSG91, NS08a, NBB00, O’G06, Oga07, Oht98, OPB08, Poi92, RK05, RV08a, RS88b, RS86, Roy93, RELW09, Sch86, Sch83, SS86, STG^{+}01, SC75, SAJ08, SPK09, SY89, SA04a, SM96, SGH85, SNP93, SB00, SOU04, SD92, TT88, TJ96, Tho06, TJKB00, Tut90, VG05, VG01, VB81, Wan99, WHF80, Wat77, WDR86, Wes85, WCEC94, WK06, WB94, XW07, XW09, YM90, YH85, YAAB87, ZL07, ZR07, Zue96, Ken79f]. regressions [CCHM08, CLLN04, FCN99, Fur04, Fur07, HCA96, Kim00, LL08a, Liu86, Mar90a, Rid80, SA92, TG78, War74]. regressive [BB89]. regressor [KB87]. regressors [Oht98, SGW94]. regret [Goo82h]. regular [GM08, SC00]. Regularized [Lee98b]. Rejection [CH99, Goo93b]. relate [GH76]. Related [Sha05d, ASS04, Ano05e, Ano06e, Ano06f, EH92, Fon90, Goo78g, Goo83x, GL84, GL85, GL88, Goo95b, Kle97, NJG88, Sah79, YXW07]. Relation [JR96a]. relations [BS95, CJ95, Ken79a, LH94]. relationship [Abd89, EH01, GA09, Goo80c, GL83, Goo84p, Pat76]. relationships [LWH08, MS01]. Relative [BE94, SSSB00, Ano03b, BA77, Hal82, HM85, Joh90b, Loh75, MS95, SNB07b]. relaxation [Hun87]. relaxations [QT92]. relevance [GT81, GL83]. relevant [Goo93e, Oht98, SGW94]. Reliability [AA09a, Sha05d, Agu99, Ahm94, AHJ92, AP06, Bal95, BH96, LL93, LB07, OP00b, PT89, Sha03d, Sha09k, SGH75, YXW07, Goo95d]. reliability-related [YXW07]. Reliable [Sha05d, FBC09]. Reml [MC91, CR03, Har04, McG89]. remote [RK05]. removals [Wu03, WCC07, WWL09, YT96]. renewal [BC94, BD84, DB84a, Fro04, HBK92, XPC03]. repair [BC82, TC75]. repairable [Agu02, HKL08]. reparameterization [KM83]. Reparameterizing [NC96, AK86a]. repeat [GS83, Goo84a]. repeated [AS00, AK86b, BZB08, FS94, HT93, Mar86a, MC91, Pol94, SE90, Wil75, WM95]. repeats [Goo81-29, Goo82l, GH82a, GH82b]. replacement [BC82, HC75, McN87, PO05, TC75]. replicated [SPK09, WW07]. replicates [LZ92]. replication [LL79, LN77, NL77a]. reply [Goo96b]. representation [AG78, CY89]. Representations [Ma99]. representing [IMLG09]. required [Joh94, Rod76]. resampling [CD92b, Do92, FS03]. research [And90a, DV95, Hor97, Sha05f, TP06]. research-estimation [DV95]. reset [SE02]. residual [Ana09, Bon05, LH94, Lou95]. residual-based [Ana09]. residuals [AsY86, DV86, Eri83, Gil95, HKT04, NS08b, O’G06, OPB08, QQ82, SC09a]. residues [Goo83y]. resistance [ZQ97]. Resistant [BW01a]. resolution [SC00]. resolvability [SK90a]. resolving [JF80]. resources [Goo79g]. respect [Goo84a, HM80, Lev78b, STS94]. Response [Sah02, ANPV97, CL97, EB90, Fan03, Gbu81, GW94, Goo90c, GSC87, HSSY04, HHKD02, Hun02, Kap87, LAR09, MGG78, Par99, PW86, Rid03, SCW79, SCW81, TAY02, TH05, VM96]. responses [Bho84, LiT95, SMP^{+}06, Spi98]. restricted [AG85, Bha01, BB84, CH91, CD92a, Die94, Gil95, Goo92b, Hoe89, Wan08b]. restriction [SMO03, tKWY95]. restrictions [CH06a, QFG87, SK08, STS94, SCA07]. result [Gri04]. Results [KHA04, Ars86, GM79, Goo83-30, JB91, Joh78, JH72, Krz82, Krz83, Lee90, Lev78d, MGN99, PX97, Pet80, PN98, PV93, Rol01, RB92, SB92, Soh00, Spi98, SB00, TZ97, WK90b, WBE80]. retrieval [Lau79]. retrospective [Hin97a, MW92]. returns [SJF06]. reuse [Kap83]. Reverse [Dem90]. reversible [ÁGR06]. Review [Ano02a, Ano02b, Ano02c, Ano02d, IC00, Sha01a, Sha02a, Sha02b, Sha03b, Ano72a, Ano75f, Ano04a, Ano04b, Ano05d, Ano05e, Ano05a, Ano05b, Ano05c, Ano06b, Ano06a, Ano06c, Ano06d, Ano06e, Ano06f, Ano08, Che87, Goo80o, Kle97, Mye83, Pie97, Sha04n, Sha04a, Sha04b, Sha04c, Sha04d, Sha04e, Sha04f, Sha04g, Sha04h, Sha04i, Sha04j, Sha04k, Sha04l, Sha04m, Sha05c, Sha05d, Sha05e, Sha05f, Sha05a, Sha05b, Sha06b, Sha06a, Sha06c, Sha06d, Sha07i, Sha07j, Sha07h, Sha07c, Sha07d, Sha07e, Sha07f, Sha07g, War84, ZK86]. reviewed [Ano06e, Sha07a, Sha07b, Sha08a, Sha09a, Sha09b, Sha09d, Sha09e, Sha09f, Sha09g, Sha09h, Sha09i, Sha09j, Sha09k, Sha09l]. Reviews [Mye79, Sha02c, Smi89a, AC99, Bor75, Kru73, RS96, Sha08b, Sha08c, Sha09c, Smi89b]. revised [RWL95]. revisited [GO03]. reweighted [SA04a]. Rho [Fra88]. Ridge [FW88, Fir97, AN03, DM93, EL75, Fan03, HE00, Kib04, ML79, Wan99, Goo92a]. ridging [EK03]. Riemann [Phi97]. right [BL03]. Risk [SS95, SR97, BG78, BA77, CLP93, DS09, Goo88d, KY93, PNW06, SNC09, She09]. risks [Agu99, BK08, NM98, QC03]. Rissanen [GC90]. riverflow [Lun06]. Robbins [McL78]. Robinson [CGA03, GA00, GA01b]. Robust [Abd89, ANM09, BJ99, BBP04, BSS02, BW01b, BB96, Bic03, BSK90, CM05a, FKW80, Fri07, Gri09, LF83, LP89, LH94, LP92, MW84, OVL02, Pak99, RR06, BD01, BS01b, CC01, CZ02, CJ00, Dut77, DH81a, Fur96, HM85, IM82, Jam01, Kap93, Kim83, Lev78a, Lev78d, LS01, LL90, MBE00, MJ93, PB03, PQ84, PGV04, RS86, Sim87, SM96, SB00, TG78, TB86, TT86, TB88, VM96, WS82, WS04]. Robustness [Ana09, BK82, BK96, CJ73, HM95, Kan75, KC73, PH95, PB04, SA92, SR00, SD02, TG73, VB83, AASAM03, BS94a, BJ73, Cic89, DC02, DG95, DL92, FCN99, IHM78, LF80, OPS82, Oga07, Pos78, Pos79, SCW81, SBO81, WJ05, Yan98]. ROC [QL01]. root [CM04, CM05b, DN94, FRB06, Jön07, Lev78b, LhKN05, Pop08, You08]. Roots [KR02, CKW73, SWK73]. rotation [Goo81q]. rough [Ken79a]. Roughness [Goo81q]. routines [BB75]. routing [BW93]. row [REK81]. row-column [REK81]. rows [Goo81d]. RPMH [Liu08]. Rukhin [MMPP05]. rule [Auc79, Bog01, Goo84u, Goo84e, Goo84f, Goo84t, Goo89g, Goo92a, JW97, Kri77, STH09]. rules [Ano03b, BG78, CI81, Coa92, Coa95, She91, SLL00, Var81]. run [CI81, Gan93b, JH01, Rig95, WR98, Zan08]. run-length [Zan08]. runs [GR08, Goo79a, JW97].

S [Ken79e, Dan80b, GH82a, Ars86]. s-fold [GH82a]. Saddlepoint [BBHW95, Gat00, Ma97, OHW97, PPRW06, BSBS08]. sales [GT90]. sam [Goo86e]. same [Gan84]. Sample [BL78, BK96, CCMV07, GSC87, HMS02, Jun08, KJ81, SD01a, Sho95, Van84, Agu99, ASS04, Amo85, And90b, And95c, And95a, And80, ABH82, Ant95, AZ05, Bal83, BLC04, BL05, BRF08, BAB86, BB74, BW01b, Bon81, BS72, Brä92, BS01b, BT00, BR03, CGA03, CR93, CH76, CR03, Che80, Che82, CY99, Che01, CGN04, Che85, Chi79, Cic89, CC85, Cob89, CRT07, CG84, CH02, CM98, Dag89, DPS01, De 81, DK05, Dha85, Die83, Dio81, DW80, Don97, Dor88, DL92, DH81b, Ede94, Edi83, EEÖ88, EP92, EOD86, ES86, FS90, For97, Fur96, GM79, Gat09, Gho95, Goo78c, Goo79o, Goo80h, Goo82m, Goo83b, Goo84d, GSS87, GC03, Hab80, HP80, HP95, HK84, Jen03, Jön07, JAK93, Kap83, KH09, Kös06, KB96, LF80, LF83, LL08a, LS86]. sample [LL79, Lem77, Lem87, LN77, hL92b, Loh73, Loh75, Lye91, MMPP05, Mar81, MG80, Mar92, MY90, MB86, Moo90, MK97b, MR03, NL77a, Nat82, Neu07, NW09, Ouy06, PLD88, PQ84, Poo80, Pos78, Pos79, Pos82, PN98, PB01, Pru93, Pul79, RR08, RR77, Rid03, Rod07, Rud86, RW96, SAD03, SE90, SB05, SB92, SMH97, SJR07, SB82b, SE02, SW88, Spi98, SI72, TY86, TB82, TW08, TO04, Tho92, Tre95, Tse84, Tsu93, VW78, Wan08a, WB09, WB73, Wil08, WS04, ZY04]. sampled [PH03]. sampler [Har04, NC96, Shi07, UP01]. samples [Ada96, Ada97, Bak07, Bal89, Bal92, BL02, BE94, BP78, BS79, Che85, CMD74, DC02, DM91, GA00, HS73, HL97, HP00, HZ03, Kim83, KG80, KP96b, LL09a, LB08b, LB01, MF02, MGG78, Mou05, Pie97, PW83, SB87, SN83, SS92, TB86, TB88, WL07b, WCC07]. Sampling [And97, Dag78, HS77, Kuk87, Kuk99, McN08, SS92, TL96, ABE83, ANM09, AL96, Bal95, BG01a, BM92, CSAR93, CB97, Con95, Dod83, GW01, Goo83-31, Goo86m, GL96, Hin97a, HSW75, HMZ05, HSWF07, KC73, Krz82, Krz83, Kuk89, Lee02a, LL79, Liu08, LK83b, LB07, McN87, Möh05, OB92, OBW05, Pan99b, PP06c, PO05, RC98, RW77, Rod76, RA81, Sch92, SB73, SSSB00, SNB07a, SNB07b, SB82b, SW75, SD01b, Tho74, WDR86, Wil75, WNB07, WBE80]. sandwich [LS01]. SAS [KHSG83, PYC93, Sha06d, Sha09f, Whi94]. satellite [CW74]. Saturated [Goo86c, May99, May06]. Saunders [DRYL08, LSSP08, LSCN08, PT03]. say [Kru03]. scale [Ada97, BP86, Bha01, BT00, Gne97, GM06, How88, IM82, Jen89a, Jen89b, Jen89c, Jen93b, Jen95c, KH09, Kim83, MKDM94, Ng96, RA01, RS86, Sim87, WDB75, WZ95]. scaled [LB01, OHLH82]. scaling [CC90b, LS88, LiT95, SBO81]. scatter [EH07]. scattered [RS97]. scatterplots [SL88]. Scheffé [SC95]. Scheffé-type [SC95]. Scheme [TYY02, Bar77, BD84, CBS06, NA09, Zan08]. schemes [BG01a, LGPP96, SW75]. School [LL73]. science [CA89, DV95, Goo94f, Goo94j]. Sciences [Sha09i]. scientific [Goo81a, Goo81-27, Goo83a, Goo83-27]. scientist [Goo79k]. Score [JJH09, Lu97, MMWM83, BN95, CRM06, FC94, FUOCN97, FCN99, SBC03, Zha06]. scorecard [KPJ09]. scores [BS05b, KP82]. scoring [Ano03b, DS09, FPRS92]. screening [RRCD97, Tan01]. script [CS08]. search [Gan84, Kk90, LiT95, NFD00]. Seasonal [KR02, KE03, Lun06, RRP06, SCA07, SAJ08]. seasonality [Mar83]. Second [FUOCN97, ASS04, AC00, CVS98, EK98, Gol72, GP07, Goo80c, Goo83j, IMLG09, JR96b, Mye98, O’N82, Hut93a]. second-order [CVS98, GP07, IMLG09, JR96b, Mye98, O’N82]. section [GOS09]. sectional [SNC09, SJF06]. sections [Jen82a]. See [Dan80b, Goo83-29, Goo83-30, Goo83-28]. seemingly [Fir97, HCA96, WHF80]. segmented [KYF08]. segments [MT01]. seismic [FcC05]. Selected [CT88, PP73, VS91]. selecting [BF83, MM08, SSSB00, SW75, TKS78, WL07a]. Selection [Dav93, GLC99, NW83, Woo85, AR01, Ber80, BHK05, BR00, Bre93, CHR03, CL98, Chi00, CJ00, CDJ02, DA87, DC00, Far90a, Fok07, Fro89, Gri92, HP00, HT85, HP95, Hua85, Kim98, KP99, KP96b, KPS01, LSL97, LP09, MA74, NCA^{+}00, NFD00, Pap93, PB00, PSW98, PN98, RS89, SS97a, SGH85, Soh92, Som84, Som85, SLL00, SOU04, SD92, Sy01b, Ven89, VS93, VS94, WB09, Wu01]. Selective [FAS82, Sha06d]. selector [FH09, KPH05]. Self [WK02, GL96, JK08, MW04, WK90b, Zan08]. self-consistent [GL96]. self-decomposable [MW04]. self-similarity [JK08]. self-starting [Zan08]. Self-validated [WK02, WK90b]. semi [CMH04, DL80, FRB06, GO03, LP09]. semi-Markov [DL80]. semi-parametric [CMH04, FRB06, GO03, LP09]. Semiparametric [CDJ02, She09, TS04, DN06, Lee90, Zha02, Zha06]. sen [FS90]. Sensational [Ano06f]. sense [KC89]. sensitive [CK96, RR77, Wri95]. Sensitivity [ASS97, BD09, CS86, Cao87, CH97, Hor97, IMP^{+}97, Kle97, RS97, And97, BG07, CG97, CCHM08, CBPW97, FH86, Hel97, SdFdGCM08]. separate [Che80, Edw95]. separated [Lam05]. separateliess [Goo82j]. separateness [Goo82f, Goo85i]. separation [PBWU78]. sequence [BG78, HM88, Mac92, MBH91, TBT95]. Sequences [PX97, ÁGR06, NS08b, Rod06, SA83]. Sequential [Coa95, CH06b, Dög01, JH72, Lau79, LJDK02, LF82, Nic96, QY95, WB73, BB80, BF83, Cha75b, Con95, CG84, DA87, DDB09, DK05, EL75, FS90, JT80, KSH73, LF80, LLKJ09, Mad77, MH78, MT80, Nab83, Nor91, SCW79, SCW81, SW88, SI72, Tho74, TKS78, Zhe88, ZA09]. sequentially [MP96]. serial [And90b, And95c, Dag89, Dag95, DS89, FMOR06, Gol77, Joh95, Owe81, PPRW06, And89]. Serially [Van84, Ana09, FK04]. Series [AW95, Sha06b, And90a, And90b, And92, And93, And95c, AP06, AMYY07, BHK05, CHáS03, CY92, CR03, CW01, CLLN04, Cor97, CY91, CV07, Cre89, Cro74, DF80, Dha85, Dre08, GHCJ09, GW94, GZP05, Goo83l, GS86, Goo90b, Gri09, Gue89, Har85, HHI90, Kat78, KKB85, LGP90, LH93, LF97, Lee02a, LL08a, LH94, LL90, Nab83, Pau84, PPRW06, RR01, RR03, RR06, RWD95, RR93b, RMH88, SAJ08, Smi94, TS09, Tay90, TY90, VG01, Wan08d, WS92]. serious [CC89, RT90]. servers [RK05]. service [DBVK02]. set [ANM09, Alb84, ABH82, CL91, Har93, Har91, Hin97a, HMZ05, HSWF07, LB08b, OBW05, Rod76, SAD03, TW08, Var81]. set-chain [Har91]. set-chains [Har93]. sets [And97, Chu80, Har87b, Har91, Har04, MZZ89, PL99, PB72]. setting [OBW05]. several [BF08, BP78, CH06b, DPS01, Gat91, Hab80, HH92, JCS07, KAS96, LN77, LTV90, Mar83, ML74, NL77a, O’G09, Owe77, PSB03, PLD88, Pig91, PQ89, SMO03, SE90, SCW79, SCW81, TT86, TG90, TA08, TKS78]. Sewall [Goo80m]. Shakespeare [Goo86k]. shape [Bha01, Che97, Cob89, FFCN07, FFCN08, Goo88c, Kim83, LLKJ09, Van93, WWL09, YX03, YAAB87]. Shaped [RB00b, Hun02]. shapes [Goo82k]. Shapiro [Roy89]. share [LGP90]. shared [Rol01]. shares [Joh78]. sharp [Goo94g]. Sharpened [Jen80]. sharpening [Goo84t]. Sharper [FJT82]. Sherman [Edi83, Goo81h]. Shewhart [CSAR93, Gan89]. shift [BK84, CC90a, CV07, Dio81, Mun06, Rod07]. shifts [CM04, Kel94, TC08]. shock [YR06]. short [DL75]. shortcomings [Mae87]. shot [FBC09, HK84]. should [Goo81f, Goo84d, Goo88c, Ano78g]. Shrinkage [Ahm92, CM07a, PL99, RWD95, hJ93, Kib04]. shrinking [Goo86p]. shuffling [Bay90]. sic [GJH90]. sided [Car07, CG94, Don97, DC99, GS07, HS73, Kwo95, RVZRZP08, SS99]. sigma [Ano06f, PB04]. sign [Goo94i, SAD03, TBG^{+}90, Wil01]. sign-rank [Goo94i]. Signal [NCÁHC^{+}09, Lee02a]. signals [BLP09, Kun92]. signed [Goo85o, Mad77, SJR07, WB73]. signed-rank [Goo85o, Mad77, SJR07]. Significance [Chi00, JR83, KPSW83, AD03, BF87, BS79, CH88, Gan84, Gil07, Goo80h, Goo82f, Goo83b, Goo86a, Goo87g, Hab92, Her75, Joh90a, KP87a, Mah96, SW90, Zim04, Goo94j]. significant [Goo88f]. signs [Goo78d, Goo79a]. similarity [CM98, CM01, GS87, JK08, Smi84]. Simple [Lou84, Mol79, kSB90, AL01, AD03, CH88, Day87, Dor88, EK98, FH09, GP07, Goo83f, GG90, Gra86, Hie81, JCM72, KSH73, LB08b, NSG91, RELW09, SSM95, SCB07, SD92, Tho74, VS94, WK06, YM90, Zan79, ZW01]. simplicity [Goo81f]. simplification [Goo80n]. simplifications [GA95]. Simplified [HL92c, TAM^{+}03, CG84]. simular [Mih74]. Simulated [PR84, WP81, Mar87, MY90, Whi95]. Simulating [EG89, MHS04, BJ82, CSS83, GGSNR09, KW94, Nic05]. Simulation [BSBS08, Che82, Dög01, EH01, FGSV09, IK03, Khe08, LC02, Lot82, Möh05, Nad99, NP09, PS85, RM02b, RR79, SHH85, SL89b, Tay94, WC91, YM96, YM99, AAR93, And90a, AJ82, Aza73, BH73, Bel93b, BH85, BB75, BM92, BB89, CH97, CH98, Cla96, Cra05, Dan87, De 97, DDB09, DFY08, Do92, DM93, DCA03, Edw85, FSBN07, FS90, Fis72, FF84, Gly84, GS78, Hab80, Har87a, IC00, Joh78, JMY96, Jun08, KYF08, KL93, Kru86b, Kru90a, KP96b, KAS96, KT94, LAR09, Loh86, Mar92, Mih72, Nor91, OR92, Pap93, Pau84, PCS09, Phi97, Pig91, RRCD97, Rod06, Ros95, SN05, Sch83, Sch78, Sch72, SNB07b, SL98b, kSwXR93, SBO81, SCM90, SP97, SW88, Sma05, SGH75, Tho74, Tho06, TL07b, Whe75, WAP84, WA72, Zan79, ZL96, Zhe88]. Simulation-assisted [BSBS08]. simulation-based [Edw85]. Simulations [Hun02, LT90, KPKPB95, Lil01, Sch74, WT06]. simulative [RHH80]. Simultaneous [Alb85, HP90, JCS07, Pep93, SW84, VS93, XM09, CG91, CKM01, Dey84, HS77, HL92c, Kwo96, LJZB05, Lye91, Mee93, OR92, PW86, SSM95, SS99, WM91]. single [AS00, BS95, BB99, CM05a, FS81, KHA04]. single-factor [AS00]. single-index [KHA04]. single-linkage [FS81]. singular [BPP00, GK00, Goo80l, Goo88b, Kwo95, RW93]. singularity [Tre95]. sinusoidal [MOS94, MK97a]. sinusoids [Kun98]. situations [LC92, Wil04]. six [Ano06f, Pru93]. Size [CM04, Kuk87, AkBA05, Amo85, BAB86, BL78, Che80, CY99, CCMV07, CM05b, Dha85, DG95, Don97, Ede94, GZP05, Goo80h, Goo83b, Goo84d, HP95, Jen03, Jun08, KJ81, Kuk89, LKKL07, Llo05, MMR92, MR03, SD01a, SB05, SS75, SJR07, SB82b, SB91, SI72, Too72]. size-robustness [DG95]. sizes [BF08, GSC87, LL08a, MY90, Sin90b, TO04, YK04]. skew [CGN04, LMB08, Val07]. skew-normal [LMB08]. skew-normality [Val07]. skewed [Bak08, Cla96, GH07, JCKS09, Lee98b]. skewness [BS05c, CR93, CK96, Dod83, HL92a, SSM93, Wri95]. Skovgaard [LP00]. sliced [BGS07]. slippage [HPY79]. slope [AD03]. Slopes [SD02, DPS01, DC02, WK06]. Small [Ant95, BS72, BS01b, CC85, Dag89, DPS01, Die83, Dio81, DW80, EOD86, For97, Fur96, GM79, HK84, Mar81, Moo90, Pru93, SE90, Amo85, ABH82, BB74, BE94, BS79, Che80, Che82, CMD74, Cob89, DC02, DM91, EP92, GMLM^{+}08, Goo84i, Goo85e, Goo86a, Hab80, HJ85, KG80, KB96, Law92, LN77, MY90, MB86, MGG78, MR03, NL77a, PN98, Rud86, RW96, RELW09, SSSB00, SB92, SE02, Tse84, Var81, Wil08, Zhe88, Zim04, TY86]. small-area [GMLM^{+}08]. Small-sample [Ant95, CC85, DPS01, Dio81, For97, SE90, Cob89, RW96, Wil08, TY86]. smaller [FMC09, KW96, MKDM94]. smallest [Goo88b]. smallsample [LC92]. Smirnov [BHLH78, Fel02, GPW80, GSF78, Goo78c, Gri04, Loh73, TB82, WLK06]. Smirnov-type [Loh73]. Smith [Mar90b]. Smooth [BSS02, MR09, KT97, MK74, QD83]. Smoothed [Tut90, CD92b, Wil89]. smoother [DC00, EK98]. smoothers [FL82]. Smoothing [Jos01, SA79, AK86b, ACNT05, BS94a, Chu01, DY92, EK03, FPRS92, Gri09, HBC07, Hua01, KPH05, KA93, Li01, LW04, QL01, Smi89d, SY92, Tay94, TB85, WV79, WW95, Wor89, ZT01]. Soc [Goo89p]. society [Goo82b]. Socransky [Smi84]. soft [CM07b]. Software [Ano88h, Sha06d, Sha09k, BP01, BD96, Goo88h, Goo89w]. solar [SZ02]. Solution [DB84a, BC94, Dut77, Gol72, TL07a, Wil89]. solutions [cDJgS93, WCEC94]. solve [PYC93]. Solving [VM00, BD84]. Some [BGS07, BH73, CH91, CW72, CSS83, DC02, GL92, Gly84, GS01, Hal72, Hua85, Joh78, Kös06, Lev78d, yL87, MGN99, PH95, PRS87, Pet80, SD92, TT86, Wil08, Wil79, WBE80, AP85, ASY81, Ars86, BJM92, BRL82, BE94, BP78, BBHW95, BB75, Bow85, BK96, Cai09, CMD74, CM06, Coa92, CW99, Dee76, ELB97, EG89, EP92, EOD86, ES88, Fos95, Fro89, Fro95, Gol77, Goo79b, Goo79s, Goo83q, Goo84u, Goo85p, Goo87f, Goo90k, HB78, JG83a, JH72, KPSW83, KC73, Kem84, Kim83, Kle97, Krz82, Krz83, Law92, Lee90, LM74, Lot82, MJ08, Mar86a, Mar76, MM93, MH72, ML79, MB86, ND84, NS86, NK08, Pap93, Poo80, QM77, QD83, RS92, Rol01, SC75, SB92, SHW93, SGW94, Spi98, SW83, TZ97, TZ04, VS93, Wes85, WDB75, Wil04]. some [XPC03, Zac80, Zan79, Goo81h, Goo86d, Goo87a, GS87, Goo89v, Hut77, Sin90b]. somewhat [Goo87h]. Southern [CG97]. Sowey [Sah79]. space [AK85, CY89, GHCJ09, Kah93, Sie78, Ter90, TO04, Wan08b]. space-time [Sie78]. spaced [TY90]. spacings [Aly90, Bal83, BL03, CS95b, L’E97]. Spanning [WB02, BW01a]. Sparse [HK92]. Spatial [Sha06b, BR92, Chi07, Chi08, Cre89, GM08, GA95, Gri04, HC06, LG09, Lot82, PH03, SAC06, TY86, WLT08, Zim89]. spatially [Dow02]. SPC [TWLC07]. Spearman [Fra88, SQ02, Tho89]. special [Goo86r]. species [Goo82m, Goo86e, ZS05]. species-sam [Goo86e]. Specific [KK85, Lak81, Mar95, MHS04]. Specification [TC08, KK93, KP04, vBBGOR06]. specifications [SdFdGCM08]. specified [CYB90, Goo89m, tKWY95]. spectral [DF80, HT83, IC00, LF97]. speculative [Goo89q]. speed [MMW91]. sphere [KPSW83]. Spheres [Dög01]. spherical [Goo79q, Goo92d, Jen83, Jen93a, Mag75a]. sphericity [CYB90]. Spiegelhalter [FRS06]. spike [Dog93]. Spline [AK86b, KP96a, AKW92, Jos01, KA93, KPJ09, Lee02b, LW04, Tah90]. Splines [DG02, Chu01, DB84a, Dia98, KL94, SY92, WW95]. split [FC96, RJ95, SJ07]. split-plot [FC96, RJ95, SJ07]. Splitting [TJ96, LY96, LY99]. Spread [AW95, Lem80, Tay77]. SPRT [KKB85]. SPSS [Sha06d, Sha09f]. spurious [SAJ08]. spuriousness [PNM83]. square [Bog95, Cel81, DN94, DM79, GMR82, GW73, HB78, Hos87, IJW03, LH72, Law01, Mar96a, Mar96b, NL77b, PB93, Wei79, WK02, Zhe88]. squared [Ant95, BR06, El 82, For97, GMLM^{+}08, Goo78n, Goo81i, Goo81j, Goo81n, Goo81-29, Goo82o, Goo82l, Goo83k, GS84, GR83, pH78, Han79, HT99, HKT04, Mag81, Oht98, OZ81, Por88, Ray85, Ter87, VC78]. squareds [GG80]. squares [CCHM08, Dan79, Dee76, DM93, EN90, FC96, Far78, Fie93, Goo79d, GM77, Hie81, HH96, KC89, KMK87, KW00, Lee02b, Mac83, Mai77, MM80, PV93, QFG87, RR01, RR03, RR82, San89, SMH97, SA04a, SB82c, SVW88, TA92, Tsu93, VM96, Wan92, Wat77, Wei79]. SRCOS [TP06]. St. [Rod06]. Stability [WCEC94, KM99, LL93, SMH97, Wes85]. stability-under-addition [LL93]. Stabilizing [AW95]. stable [Bar77, Bar78, Bar79, Bar81, BM86, Cro74, Kun93, LL93, Poi92, Tsi02]. stable-Paretian [LL93]. Stage [LY99, Mad78, ACNT05, EBA00, Far78, Fro95, GW01, HMK00, LY96, MH78, MS98, May99, May06, Nan98, Nic96, RR82, Rod76, Sch92, SQ07]. staged [Paw01]. Stahel [BD01]. standard [AKW92, Atk92, Bal89, Bal92, CR93, DPS01, FKW80, Goo82m, Jam01, LPSL05, Mar87, ME72, MK97b, Rhi86]. Standardized [BW93, Goo84d, RR93a, Goo82n, Goo84j]. Star [RB00b]. Star-Shaped [RB00b]. starting [Zan08]. state [AK85, CY89, Fan95, FF84, GHCJ09, HP80, IMLG09, Kah93, KK93, WM90]. state-space [GHCJ09]. states [Nor84]. stationarity [AK86a, GOS09, NP81]. stationary [CP76, Deá80, Gri09, KM06, LY99, LOR04, Mae87, NJG88, Nic96, PSW75, TA92, Tsu93, XYR09, Zha00]. Statistic [OS02, Adi98, Ant95, BCL93, BP01, BS79, Bho73a, CT88, CH88, DAM98, Dod83, Fun79, GHW99, GG77, Gil07, GK05, Goo78c, Goo85r, Goo89w, GN89, KS97, LP00, Mcd74, Mur78, Mye98, PT99, RR77, Sie78, SM91, Tan82, WLK06, Wes72]. Statistical [Ano88h, CC09, Din08, Edw89, EBS86, FF84, KS91, LWH08, NW09, Sha09i, TYY02, WR98, WR94, WCC07, Zha00, Ano05d, Ano05e, Ano06b, Che01, Ell00, Fun88, Goo79b, Goo81l, Goo82f, Goo82b, GHDB89, Goo89w, Goo90h, Goo90b, GA95, GGSNR09, HT99, HCW07, IMP^{+}97, JF80, JW97, JCM72, Kle97, KHSG83, KP87a, Mar83, MO91, PHO05, Sha05f, Sha06c, Sha08a, ST87, SS88, Sy01a, Tan01, TL07b, Zim04]. statistician [Goo84s, Goo88h]. statisticians [Ano06c]. Statistics [Eti81, Goo88h, Sha05c, Sha05e, Sha06a, Sha07i, Ada91, ASY81, AT95, Ant95, Bal85, BS95, BP78, Bho73b, BR84, BHZ08, Bur74, Che03, CS08, CW74, DDB09, For97, Fou80, Fre07, Fur07, Gat91, GO90, Goo84w, Goo90k, GV81, HSB85, HS86a, Hua91, Jah05, JH90b, Jöc81, JG83b, KG80, Kru00, LB07, Mas03, MY90, Pak99, Pet80, RT78, RA01, Rud86, Sch72, Sha04n, Sha07b, Sha09b, SCB07, TB82, Vau94, Wel82, Yan99, Arn00]. steepest [PK72]. Stein [CLP93, KC89, KTM05, Rod76, STH09]. Stein-rule [STH09]. Step [AHAH08, Xio99, Abd95, JG83a, MTO08]. step-down [JG83a]. stepping [Ber80]. Stieltjes [SB85]. Stirling [Goo85e]. Stmultaneous [Ahm99]. Stochastic [Aza73, CCD96, Har89, NFD00, QT92, RM02b, ASY81, BH85, BN96, Cao87, CM07b, FS90, GM08, GGSNR09, Hel97, KT85, MS99, NP98, Nic05, Nor84, OHN93, SW88, SP86, TZ04, WA72]. stochastically [RM96]. stock [Sch75]. stooge [Goo84s]. stoogian [Goo89h]. stopping [Goo91]. storage [Kru86b]. Straddling [Goo03]. straight [MT01, TT86]. straightforward [Day87]. strategies [DY92, DG95, Lev05]. stratification [GH07]. stratified [BL78, GM77, NTK09, PW86, WNB07, XYT08]. stratum [LL79]. stream [AA09b]. strength [PP06a]. stress [CC90b, Xio99]. stronger [Dav79]. Structural [Sch02, CGA09, Coo07, Khe08, LhKN05, MB86, MB90, OR92, PS09, Sha03e]. structure [AC08, And90b, And95c, Cha75a, CCSG81, CH82, Die93, FH86, Gup84, Lee81, SB91, TH05, Zim89]. structured [AK85, CP91, NCA^{+}00]. structures [BDFR97, GA09, SE90]. Student [CVS98, El 82, GS86, Jen89a, Jen93a, VW78, VG05, Zim04]. Studentised [HH92]. Studentized [BCX93, BHG01, CH88, Fun79]. Student’s [Che78]. Student’s- [Che78]. Studies [SBO81, AB05, BH73, FSBN07, Fis72, Fis73b, GSC87, HS77, KHSG83, MT80, Pot81, QM77, RS97, RR79, Sha03d, SNC09, WC91, WB73]. Study [AW95, HM95, LC02, Pet88, Pet02, SHH85, YM99, AS00, AAR93, AB08, AJ82, AR00, Aza73, Bak07, Bak08, Bal83, BM90, BF83, BRL82, Bel93b, BP78, Bow85, BG07, BM92, BB89, CG91, CCD96, Cla96, Coa92, DDB09, DP06, Do92, Dog89, DM93, DH81b, DCA03, EH01, FRS06, Fos95, Fun79, Goo89u, Hab80, HL92c, JMY96, KYF08, KL93, KAS96, LF80, Lee02b, Lev78b, Mar92, ME72, Mil79, MMW91, NP98, Neu07, Nic96, NS08b, OPS82, OR92, Pap93, PN86, Pig91, Poo80, RFGE86, RRP06, RACSGJ05, Rod06, Ros95, Sch83, SFSS85, SND89, SNB07b, SP97, Sma05, Soh94, Soh00, SW83, SC82, SI72, TMW73, Tho06, TL07b, TTNC09, TB85, VM96, WA72, WZ95, YM96, Zan79, ZL96]. Subdivision [LH79]. subintervals [Goo85k]. subject [GG75, Han86, HAS89, Ken79a]. subjective [Goo90l, Hel97]. subjectivity [Goo85m]. subjectspecific [SG96]. subpopulations [Won85]. Subsampling [GZP05, BBG86]. subsequent [EA78, MH99]. subset [Bre93, Kim00, MA74, O’G06, PSW98, SS97a, Som84, Som85]. subsets [HS86a, RT78]. success [WS90]. Successive [Hun87, SND89]. sufficient [Goo93a, Kru00]. Sufftctmcy [Goo85l]. suggested [YI01]. suited [War74]. sum [ANPV97, BG93, Gan89, GBSS84, GR83, KPKPB95, L’E97, Nab83, OZ81, SNP93, Zan08]. sum-functions [L’E97]. summands [KM06]. summation [Smi89c]. summer [TP06]. summing [Goo79l]. sums [BG93, Dan79, Goo79d, Goo93d, Phi97, Sum83, WR98]. sup [OP04]. sup-norm [OP04]. superimposed [BLP09]. superiority [Fou81]. superpopulation [Fri85]. supersaturated [EGM^{+}04]. supplement [Goo81n, Sah79]. surface [GSC87, VM96]. surfaces [Fan03]. surprise [Goo83v, Goo84v, Goo85h, Goo93c, Goo89s]. Surprising [BS94b]. surrogate [TAM^{+}03]. survey [Edw85, Goo87e, ML74, Sel08, TP98, WB94]. surveys [LL79]. survival [CFOS02, Ebr93, GSC87, GP95, HV05, JB91, Jun08, NCA^{+}00, NM98, Pru93, RM96, She09, SC82, WM91, YM96]. survivor [Ahm99]. suspicious [Goo89e]. Sutton [Che94]. sweep [Gri92, HH93, Lev05]. swindled [HS86b]. switching [HKL08, Kim93, Rid80, TH09]. symmetric [AGNS91, BT00, CV08, Cor04, Hen81, LP89, LTV90, QFG87, RS90a, Tre94, lS96]. symmetrization [Dor08]. Symmetry [HMS02, AS90, FRS06, Kap93, Law01, Mad77, MPPZ05, WB73]. synergistic [PW86]. synthesis [Goo83-29]. System [RM02b, Agu99, Agu02, AP06, Cao87, CK01, Che05, Fir97, HKL08, Joh78, KS04, PYC93, Pos78, Pos79, Pos82, SZ02, SGH75, SVW88, Tue01, VR92]. Systematic [BM92, Kuk89]. systematically [BDB08]. systems [BCH08, Edw89, Hal82, HAS89, IMLG09, SE02].

T [CS94]. T-distribution [CS94]. Table [KC96, MZ77, NS89, Dey84, DL75, GG80, Goo79j, Goo86q, HB78, HJ85, Ken79e, Lat82, LDC73, NC06]. Tables [Aok02, Bru75, CKW73, Cro74, FJ84, HH74, Rod76, SB73, ABH82, AT05, CS97, Goo78k, Goo79b, Goo79u, Goo79f, Goo81r, Goo81w, Goo81v, Goo83e, GL83, Goo85p, GF89, Hab84, Hos87, Hou85, Ken79d, Law01, Lee85, LPSL05, Mag81, MG06, MH72, MPPZ05, O’N82, PD03, Par92, Pig91, Smi89d, XYT08]. Tabulating [OP04]. Taguchi [LAR09]. Tail [Fie93, Ant95, BS01b, For97, Fro89, GO03, GP07, Goo82n, Goo84d, Goo84e, Goo84j, Goo84k, GS86, Ma99, MR09, RS77, Ter87, Goo83-31, Goo85b]. tail-area [Goo82n, Goo84d, Goo84j, Goo84k, GS86, Goo85b]. tail-index [MR09]. tailed [lS96]. tails [Goo84e, Goo86a, Ken79b]. Taiwan [FcC05]. tangents [PK72]. Tanner [GJ77]. Target [CDF05]. Tarone [GP95]. tasting [Goo90j, Goo92b]. tax [Goo81s]. tea [Goo92b]. Tech [Arn00]. Technique [GLC99, DM91, DB84a, Lev78d, Mac83, SR97, SM96, Smi89c, SGH75]. techniques [AP85, ASS97, Bel93b, BH82, DPS01, DM78, Fos95, GL86, GB86, Kle97, KS87b, KC97b, LH79, LL79, MS80, MM08, NP09, NS86, PBWU78, RFGE86, RACSGJ05, SCW79, SCW81, VBS07, WC91, Wil08]. technology [Goo94j]. temperature [CW74]. Temporally [Van84, HR07, Van87]. temporary [RMH88, TC08]. ten [AD01]. Tendencies [Goo93a]. tendency [Bic03, Goo94b]. Tentative [Goo93e]. term [NM98, SB96]. terminating [TJKB00]. termination [Tho74]. terminology [Goo82g, Goo83g, Goo86c, Goo86r, Goo93b, Mih72, Goo80i]. terms [AB09]. Terrell [Goo90g]. tessellation [GM08, HQ00, HM80, MM99]. Test [Aok02, Chi00, CG84, Eve88, Fel02, HHR02, JO85, KC73, MH07, YK04, Ahm92, ASY82, AAL02, AZ05, Auc86, Bal83, BL03, BHLH78, BB80, BCX93, BCL93, BA77, Bog95, Bog01, BHG01, Bru75, BT00, CR93, Cha75b, Che94, CS95b, CL98, Chi79, CH99, Cho08, CKM01, Cic89, CV08, CS08, DSL06, DS89, DAM98, Dee76, DG95, Dow02, DMF83, Edi83, EEÖ88, ES86, FRS06, FS90, Fou80, FJ83, FJ84, GPW80, Gat00, GA01a, GL92, Goo78n, Goo78d, Goo81h, Goo82f, Goo84o, GS84, Goo85o, Goo86s, Goo89m, Goo94i, Goo96a, GP95, Gup73, Hab80, HH74, Haw79, HH92, HANMA98, HM85, HC06, HB06, HT99, Hou85, Hut93a, IJW03, Jen03, Jen89a, JL88, Jöc81, JG83b, Jön07, KK99, Kib04, LL93, Law92, LH79, LZ92, LF80, LKKL07, LS86, Lem77, Lem87, Lev78a, Lev78b]. test [LS99, LK83a, Llo05, yL87, Loh75, LJ00, LPSL05, Mad77, Mad78, Mah96, MMPP05, MM93, Mcd74, MY90, MJ93, Mye98, NS08b, O’G06, OP00a, Pak99, Pan99a, Par92, PQ84, Pie94a, Pie94b, PP73, Pop08, Pos78, Pos79, Pos82, Ram89, RM02a, Ray85, RR77, RR79, SK08, SAD03, SB08, SSW95, Sar98, SMO03, SS75, Sef92, SA92, SJR07, Shi85, SL89a, SE02, SN83, SW88, Smi78, SI72, SCB07, TG78, TO04, TC92, Too72, Val07, VC78, VS94, Wan08a, WLK06, Wil01, WA72, WS04, WS00, WWL09, ZC03, ZY04, Zha06, Zim04]. tested [Rod76]. testes [TY90]. testimony [Goo81x]. Testing [Bai89, Bha01, CB07, CH82, Dre08, EH92, GOS09, Gle97, GM06, HM88, HC06, HS04, Hua01, KM99, LC02, LW82, Li01, LL08b, Llo93, Loh86, Lu99, MF02, PGT09, PW86, Poi92, RV08b, TG90, TZ97, WW06, WB94, lS96, AG85, AkBA05, AD03, Bar72, BR06, BJ73, Bho73a, Bho73b, BT09, BH95, CRM06, CP95, Che82, CBS06, Chi07, CKM01, Coo07, CH88, DP93, DR02, DG95, Edw95, FBC09, Fun88, Goo81n, Goo96a, GN89, Hay97, HTC07, HK84, JG87, LL09a, Lev78c, LW04, Mat79, MBH91, Mei09, Mou95, NJG88, Nag75, PPRW06, QY95, RR03, RR06, SFSS85, Spi82, Spi06, WM90, Wil04, Wil07, Won85, RA81]. Tests [ASSY79, BG94, BSS02, BM86, BN95, BCP02, CCSG81, Fam99, GsW01, GA00, HE00, Joh95, L’E97, LJDK02, Lil01, LT95, Mei08, Pet02, Pie97, AHAH08, Agu02, ABE83, ABE85, AHAH06, ASY81, AS00, Aly90, AL99, AB03, AT05, Bai89, BM90, BCH08, BCV98, BRL82, BB74, BW01b, BJ73, BBHW95, Bow92, Brä92, BDB08, BK96, BR03, CGA03, CP91, CK04, CH76, Che80, Che03, CB97, CM04, CM05b, CC97, Cor04, CV07, CNFO05, DD78, DDB09, Die06, Die83, Don97, Dor88, Dor08, DL92, DH81b, EA78, EM86, FC96, FMOR06, FRS06, FSBN07, FC94, FUOCN97, FCN99, Fle95, Fou80, Fou81, FRB06, Fur04, Fur07, Fur08, Gan84, Gat91, GA01b, Gil07, GSF78, Goo84t, Goo86a, Goo87g, Goo90d, Goo94j, GSS87, Hab84, Hab92, Hal72, HPY79, Hay97]. tests [HB93, HMS02, Hin97b, Ho93, HSSY04, HL92a, HKT04, Hos87, IP86, JB91, Jen93a, JS00, JS02, Jöc81, JG83a, JR83, Joh90a, JJH09, KT97, KPSW83, KK85, Khe08, Kib04, KP87a, Kös06, KAS96, Lac90, Lac92, LL06, Leg00, LhKN05, LH94, LF82, Loh73, LG84, Lu97, LPSL05, MH78, MS01, Mar86a, Mar76, Mar81, Mar83, McS06, MM05, MPPZ05, MY90, Moo90, ND84, Neu07, OPS82, O’G09, OP00b, Owe81, PX97, Pan98, PP03, PRS87, Pes80, Phi91, PR84, Poo80, QM77, QD83, RR01, RR08, RML90, RL08, SBC03, SE90, SY89, SQ07, SHW93, SP00, SGW94, SW90, SMV76, SB93, TNG^{+}06, TA08, Tho92, TL07b, TZ04, VW78, WB73, WK72, Wes85, WDB75, Woz94, YY07, Ken79c]. Texas [DE06]. theatres [Sho86]. Their [LL91, BR03, Cro74, GH76, GH82a, GS87, MH99, MK97b, SdFdGCM08, WW95, WP81, Zac80]. theorem [Bar78, BJ82, Goo83w, Goo88e, Goo89g, Goo94h, Her75, LDR92]. theoretic [Kun92, Sy01b]. Theoretical [Pes80, And95b, Ans80, KA82, McK86, PX97]. theories [Bar72, Sha06c]. theory [And90b, And92, And93, And95c, DB84a, Goo82a, Goo87i, Goo89p, Goo89d, Goo89k, IMP^{+}97, Mur83, NP81, Oga07, Wec94, WT06]. therapeutic [Don97]. therapy [Goo78m, GSC87]. there [GL85, Jen03, Mal79, NS09, PM02]. thermal [ACNT05, De 97]. thin [How85]. Thompson [EOD86, SNB07a]. those [Goo87a]. thoughts [Sha05f]. Three [Ano05e, GGSNR09, Loh73, BK82, CC09, CB03, Far78, FTS09, Hab84, HM85, Hos87, HM98, KH04, KC97a, Luc08, MGG09, Nan98, Nic96, PD03, PW83, SA04b, Tza09, UP01, WK90a, WM90, Wel82, WBE80, Zan79, ZK86]. three-factor [CB03, Hab84, Hos87]. Three-parameter [GGSNR09, CC09, FTS09, KH04, KC97a, Luc08, PW83, SA04b, Tza09, WBE80, Zan79, ZK86]. three-stage [Nan98, Nic96]. three-way [PD03]. threshold [BC08, GRVV08, TY90, Xio99]. threshold-type [TY90]. Thresholding [Wu02]. thumb [Auc79, Goo84u, Goo84e, Goo84f, Goo84t, Goo92a]. tied [SMV76]. ties [Goo80a, Goo85r, Hol85b, Roy89, Tho89, Wil01]. Tiku [TB86]. Time [AW95, Sha06b, And90a, And90b, And92, And93, And95c, AMYY07, BHK05, CHáS03, CY92, CR03, CW01, CLLN04, Cor97, CY91, CV07, Cre89, DF80, Dha85, Di 05, EH92, GHCJ09, GW94, GA01a, GZP05, Gri09, Gue89, Hal82, HHI90, Jen03, Kru86b, LGP90, LH93, LF97, Lee02a, LL08a, LH94, LL90, Nor84, OP00b, Pau84, PPRW06, Pop08, Pot81, RR01, RR03, RR06, RM02a, RWD95, RB88, RMH88, SQ02, SCW79, SCW81, SAJ08, SL93, Shi07, Sie78, SB96, TS09, TY90, VG01, Wan08d, WO97, WS92, YH85]. time-series [And90b, And95c, CLLN04, GHCJ09, RR01, RR03, RR06]. times [CP76, SND89, TJKB00]. Tobit [Dag89, Moo90, ZL09]. Toeplitz [Die93]. Tolerance [KS04, TC94, AB09, AL02, CS95a, CY99, HS73, KA03, Lee01, RVZRZP08, TG73]. tolerant [Goo84k]. Tomographic [Sha09l, Goo82d]. too [Goo84i]. tool [BDB08, LGPP96, PHO05, SCK01]. tool-wear [LGPP96]. topic [Goo94f]. topic-independent [Goo94f]. Topics [Sha05c, Sah79]. total [AASAM03, ASY81, Goo81s, GM06, RM02a]. totally [Smi89d]. totals [GH76]. tourist [AMYY07]. toxicity [AB05]. Trade [GA95, BS94a]. trade-off [BS94a]. Trade-offs [GA95]. traditional [DR02]. traffic [Soh96]. Trails [LJDK02]. training [KL93, MS95]. transect [EB90]. transfer [oE89]. transferability [PV93]. transform [Bel93a, Mei08, Nat82, PN86, Goo81-29, Goo82l, Goo87e, Hol85b]. transformation [And95b, BK82, BK84, Cel81, HT93, HK07, Kap93, KAST05, Neu07, Spi82, SSM93, Tad81, Yan98, YH85]. Transformations [AW95, JL96, GA02, KP82]. transformed [Gri02]. transforming [Mei09]. transforms [Sie78, WP81]. Transition [BSS02, LB08a]. transitive [DSS06]. transitivity [Goo89j]. translation [Jen93b, SVW88]. translation-scale [Jen93b]. transport [CBPW97]. transvariation [NA09]. trap [Goo90a]. treatment [BS78, CM07a, SB88]. treatments [DSL06, Goo78e, Goo86n, IHM78, ZA09]. Tree [NCA^{+}00, LW95]. tree-based [LW95]. Tree-structured [NCA^{+}00]. Trees [WB02, BW01a, HL05, J.83, STG^{+}01, TJ96]. Trend [RG02, AHH07, Die83, GA01a, GP95, HB06, KP87a, TNG^{+}06]. trends [Lun06]. triad [Fra79]. trial [Bea85, Ram89, TT82]. trials [BS05a, BL78, CI81, Coa95, Con95, Goo81g, NS09, SSW95, Sch92, SJ06, Van05, Sar98]. triangles [HQ00]. triangular [Kim00]. trick [GF89]. trigamma [Dev92]. Trigonometric [Gue82]. Trigonometrical [Goo93d]. trigonometry [Goo79q, Goo92d, Goo94c]. trimmed [HS86a, JUP86]. Trimming [Aok02]. trimmings [HS86a]. trinomial [CL91]. trivariate [ESRV98, GT78, LK83b, RW93]. trouble [Goo83-31]. true [HTC07]. truncated [BRT07, BD09, Dag78, Gri02, GT78, JG92, Mal06, Mar96a, Mar96b, Pan98, PPK77, RB88, SI72, vdTGL97]. truncating [Cha75b]. Truncation [Bar81, Har85, Mal06, SS75]. truths [WT06]. trying [AG81]. tudies [Mae87]. Tukey [FL82]. tuning [WJ05]. tuples [Goo85r]. Turing [Goo94h, Goo00]. Tutorial [Ano80s, FF84, VW85, VR92]. twice [Hut77]. Two [Aok02, Car07, CH98, Dic78, Don97, Fle95, Kru89a, LL90, MS98, MT01, Paw01, Pos82, Sha04n, Sha05f, Sho95, SM91, AMAE97, AHADA00, AJAH07, Amo85, AT95, AT05, ACNT05, AZ05, Bak07, BM90, BLC04, BG94, Bay90, Beh72, Beh73, BT00, BBG86, CYC99, Cha94, Chi79, CMD74, CS97, CAT78, CG84, DD78, DL92, DC99, EBA00, EH01, ES86, FS90, Fon90, Fro95, GK90, GS07, GW01, Goo78g, Goo81e, Goo84e, Goo86n, GHDB89, GSS87, Gue78, Hab92, Han78, Han94, HMK00, HB93, HT85, HD77, Hos78, HL09, IHM78, IP86, Jah05, JS02, JG83b, KN89, LF80, LL09a, LL79, Lem87, LSCN08, LN77, LL09b, Loh75, MG06, MH72, Mar92, MRR84, May99, May06, MYS01, MPPZ05, MKDM94, MR03, Mou05, MP96, NC06, NL77a]. two [Neu07, PL99, Pan99a, PF04, Pap80, Par99, PP03, PM02, PN86, Pie94a, PR84, Pos78, RP01, RR08, RR82, Rod76, RA81, RL08, SK08, SB88, SMO03, Sch83, Sch92, SF93, SB73, SM94, SHW93, SW75, SN83, SW88, Smi89d, SC09b, SCB07, TB88, Tan01, Tho92, Too72, VW78, VM00, WL90, Woz94, Wu03, WWL09, YXW07, ZY04, ZA09]. two- [LN77]. two-component [AHADA00]. two-compound [AJAH07]. two-factor [HB93, NL77a]. two-fold [BBG86]. two-group [GK90, MYS01]. Two-Parameter [Sho95, BLC04, BG94, LSCN08, RP01, SN83, WL90, Woz94, Wu03, WWL09, YXW07]. two-phase [LL09b, Sch83]. Two-point [CH98]. Two-sample [Pos82, AZ05, BT00, Chi79, ES86, FS90, GSS87, LF80, Lem87, Mar92, Neu07, Pos78, RR08, SW88, Tho92, VW78]. Two-sided [Car07, DC99, GS07]. Two-stage [MS98, ACNT05, EBA00, Fro95, GW01, HMK00, May99, May06, RR82, Rod76, Sch92]. Two-staged [Paw01]. two-tails [Goo84e]. Two-way [Aok02, Kru89a, MT01, AT05, CYC99, CS97, DD78, Fon90, MG06, MPPZ05, NC06, PN86, Pie94a, PR84, Smi89d]. Type [AHJ92, Mou95, AMAE97, BW01b, BT00, Cor95, KM06, KTM05, Loh73, Pan99a, Pet80, SK08, SK87, SC95, SH72, TB82, TY90, VS94, WS04, AR01, Bal89, Bal92, BL02, BL03, BLC04, BKLW04, BL05, BH07, BE94, EM86, GJJL02, KH09, KAST05, KY93, LS99, LB01, MM05, Ng96, NW09, SN83, TB88, TYY02, Wan08a, WCC07]. type- [Mou95]. type-2 [AMAE97]. Type-I [LS99, NW09]. Type-II [BL02, BL03, BLC04, BKLW04, BL05, KH09, LB01, TB88, TYY02, Wan08a]. typical [FGSV09].

ultrastructural [SGR04, SPK09]. umbrella [Kös06]. unbalanced [BE86, BBG86, FC96, FS03, Gen76, HK00, HB93, HK92, HSWF07, LL07, LD97, RJ95, RVZRZP08, SE90, SJ07, SN95, ScK97, WW03, WS04]. Unbiased [CNV02, RS83, BP86, GV81, HSB85, LM74, LB08b, SB73, VD08, Wan99]. uncertain [Goo81x, NCÁHC^{+}09]. Uncertainty [ELB97, Hel97, Hor97, IMLG09, KC97b, MB97, VG01, Wan90, WECC00]. unclassified [Mou05]. unconditional [TA08]. undamped [Kun92]. underflows [Goo78j]. underlying [EH92]. undermining [Goo89g]. unequal [AAR93, Bea85, BF08, IJW03, JO85, Kru88, Kru89a, LH97, LF83, MF02, RG93b, Sin90b, TT82, SB91]. unexpected [RS90a]. unextended [Var81]. Unified [CBS06, BRF08, YXW07]. uniform [Fle95, HL09, L’E97, Mee93, MKDM94, WD93, Wu03, Jen94]. uniformity [Bog95, Bog01, IJW03, KPSW83, MMPP05, QM77, QH80, Sef92]. uniformly [SS86]. uniforms [JL96, KPKPB95, YST90]. unimodal [Ede94]. Union [CP95, MZZ89]. Union-intersection [CP95]. Unit [FRB06, KR02, CM04, CM05b, GR08, HG85, Jön07, KS04, LhKN05, Pop08, You08]. units [LL79, SSSB00, SNB07a, SNB07b, WDR86]. unity [Coo07]. Univariate [WW07, AB08, CMH04, GS73, HL97, Hat86, HCW07, NS89, SW90, VR92]. universal [GSH04, Goo83a]. universally [Dio81]. universe [SS92]. universes [Tan82]. unknown [ABE83, ABE85, Bha01, Ede94, Fou80, Jöc81, KJ09, Kös06, Pop08, VG05, XDZ09]. Unobserved [CM05b]. unordered [LPSL05]. unpaired [Goo84d, Goo84f]. unranked [TP98]. unrelated [Fir97, HCA96, WHF80]. Unreplicated [PT02, SGR04]. unrestricted [TBG^{+}90]. unsolved [Goo79s, Kem84]. Unweighted [HH96]. up-and-down [Ker87]. Updated [RR93a]. Updating [KK99, MS80, PCS09, GJH90]. upon [KP82]. upper [AS01, Ant95, BG94, BC94, BSK90, For97, Har03, HC06, Jen93c, Shi07, ZY04]. usable [Mac92]. Use [CY92, KJ09, Spi06, ASS97, BS79, Dev84, FAS82, Goo79o, Goo87e, Goo90b, Har04, HK07, Hos78, Jen89c, JMY96, MB97, Phi91, Rhi86, RS85, SK80, Tza09, tKWY95, Win75]. used [ASY81, Bho73a, Chi79, Dee76, Goo96a, Nab83, OP04]. useful [Ano06c, PHO05, RWL95]. usefulness [SGR04]. users [Whi94]. Using [BPP00, DSL06, Pet02, RW96, SC09b, AP85, AS04, Bar78, BC94, BBP04, BCL93, BW01a, BN95, BD96, BR06, BDFR97, BHZ08, BS05c, BJ82, CYC99, CH06a, CK04, CSAR93, CC85, CCMV07, CH90, Coo07, DS89, DB84a, DSS06, EH92, ES86, Eti81, FBC09, FRB06, Fro04, Gal02, Gan90, GK90, Gao04, Gat91, Goo81-29, Goo82l, Goo84s, Goo84i, GM77, GM06, HSB85, Hig97, IM82, JCKS09, Jon75, Kim00, Kru90a, Kun92, Lai82a, Lai82b, LGP90, Li88, Li01, LAR09, LPSL05, Mad78, Mai03, MTO08, Mar95, MM80, MS99, Mcd74, MS80, MH07, Mur83, NCÁHC^{+}09, O’G06, PP06a, Par81, Pet80, PP73, RM02a, RWD95, Rod07, RM02b, RS90b, Ros06, Roy93, SZ02, SAD03, SN05, Sco02, SL87, SA04a, SW90, Spi82, Sy01b, Tah90, UP01, VG01, Vir07, WK90a]. using [WmGT95, Wet96, WCEC94, WR98, WR94, WAP84, Won95, XYR09, PYC93]. USNRC [KC97b]. usual [Goo88c, Yan98]. utility [Goo83f]. Utilization [CI81]. utilizing [ASS04, Wil04].

v [IP86, AH09]. Valand [AB82]. valid [AD03]. validated [WK90b, WK02]. Validating [DS09]. validation [BDFR97, CCMV07, Gri92, Lee90, LP09, NCA^{+}00, Tay90, Whi94]. validatory [Krz83, TKT89]. validity [Ken79c, Pat76]. Value [SQ07, Auc79, BKLW04, DR94, DR02, Die05, Die06, Dup96a, DF98, Goo80h, Goo81r, Goo83b, Goo83n, Goo84w, Goo88b, Goo90a, Har03, KB87, LS99, PW83, Sch78, TP98]. valued [DF80, Gle91, LF97]. Values [Mad78, AL94, ASY81, AK86b, AR00, BBHW95, Bru75, CCHM08, Chi07, Fun88, Gil07, Goo86o, GS87, Goo89l, Goo89s, Goo90i, Goo92b, Goo93c, HV93, Hol85a, JG83b, KK99, Kat78, KN89, Lam05, Nad99, RR93a, RG93b, Vau94, WK75, Wil04]. variability [BBG86, CH98, Goo89q, RR08, RS09]. Variable [GLC99, LSL97, SGH85, Soh92, Ber80, Cel81, CSAR93, Far90b, FAS82, Gbu81, Gri92, Ken79f, KB87, LAR09, Mal06, MW04, NTK09, NFD00, RS90a, SOU04, ZL09]. Variables [SHH85, ASSY79, AL96, BW93, CAT78, CJ73, DF80, EH01, Eti81, Goo79p, GR83, Gue78, HM88, Han86, HB78, KPKPB95, KS91, Ken79a, KN89, LGP90, LL88, LJZB05, Lou84, Mae87, MA74, MBH91, Mel80, NW83, NC72, OZ81, Owe81, Pap80, PS98, PNM83, PQ89, RS85, SF93, Sho86, SCAH05, Sim00, TG90, WHF80, Wil07, WAP84, WW07, XW09]. Variance [AJ82, CH99, Die06, Fis72, Fis73b, Kuk87, RS85, WAP84, AM87, ASS04, AB03, Ano81f, Bai89, BJ99, BB74, Bon81, BS05c, BB84, BE86, CH91, CD92a, CG94, CB07, CKM01, CR73, Del83, FS03, Gan89, Gen76, HK00, Har03, Hay97, HP80, HB93, HK92, Hoe89, Hua91, Hun87, JH72, Kan75, KS97, Kru88, Kru89a, Kuk89, Lai82a, Lev78a, Lev78b, Lev78d, LL07, jLP05, yL87, LD97, Mai03, MF02, MG80, Mil79, NS86, OHW97, OBW05, Paw01, PB01, Pul79, SB88, SA79, Sha87a, SL87, SS92, Smi84, SGH75, TBG^{+}90, WGP00, WW03, WM95, lS96]. variance-covariance [SA79]. variances [Ada91, BM90, BCL93, BW01b, Bon06, Bur74, Hal72, JG87, JO85, Ken79f, Kru88, Kru89a, Lac92, LH97, LH79, LF83, LKKL07, LL79, MF02, Mat79, Nag75, Pan99a, RG93b, RA81, SB73, SGR04, Shu08, SS88, Vau94, WS04]. variants [CS05, Sum83]. Variate [SL93, BB75, Dev82, Dev84, Dev92, FS75, JL96]. variates [AL01, Bar78, BW93, Bho84, BP92, Dag78, Fis73b, Hör93, KS85, LH72, hL92b, Mai03, Mel80, PB72, Pol97, Pul79, SN05, SS88, TR75, Wil98, YST90]. variation [BG07, Gou93, GM06, KC02, Loh75]. variational [RV08a]. Variations [McL78, Lu99]. various [AG92, BM90, Goo80i, Sie78, YY07]. Varma [HN80]. varying [Fro89, LP09, LL08b]. varying-coefficient [LP09]. Vector [KP87b, AS90, Ans80, AK86a, Bho73a, CCSG81, GsW01, Gle91, KP99, KP04, SP97]. vector-valued [Gle91]. vectors [Ahm87, Bho73b, Deá80, GH76, Goo80j, GL84, Nar90, Sim93]. vehicles [Wan90]. vendor [HP95]. verified [RL08]. version [RR93a]. versions [Bog95, Bog01, CCD96, LPSL05]. versus [And90b, And95c, AB08, DR02, Goo82c, Goo84e, Goo84f, Goo93b, Goo94k, KK91, Moo90, PN86, SB87, TS09, WT06]. verticles [GJH90]. very [Gri04, Hin97a, Goo86a]. via [AL93, AL94, Con06, CY89, Dia98, GJ82, HMZ05, IMLG09, Kap83, LLKJ09, RV08b, WK06, WW07, Hua01]. viable [Bic03]. view [Ano06d, IMP^{+}97, KPH05]. view-points [Ano06d]. VII [GJJL02]. Vinograde [Goo80l]. violation [Too72]. Virginia [Arn00]. visual [TBT95]. vocabulary [Goo78l, Goo82m]. Volterra [BC94, DB84a, SK87]. Volume [FC89, Sha03c, CH02]. Voronoi [FGSV09, FGSV09, HQ00, HM80, MM99]. voting [Goo79l]. vs. [Her75, Lau79, Luc08].

W [Kib04]. Wald [GK05, SK08, SB93]. walk [Ahm87]. Walker [BR92]. warning [Whi94]. Watson [GG90, AAL02, DS89, GL92, SA92, Smi78]. Wavelet [JK08, SMP^{+}06, Wu02, AS04, Cra05, Qu06, XYR09]. Wavelet-based [JK08, Cra05]. way [Aok02, AT05, CYC99, CS97, DD78, Dow02, Fon90, Goo86j, HK00, JO85, Kru88, Kru89a, LKKL07, MG06, MH72, MPPZ05, MT01, NC06, PD03, PN86, Pie94a, Pie94b, PR84, RVZRZP08, Smi89d, WW03, XM09]. Weak [Goo89j]. weakly [Zur93]. wear [LGPP96]. Weibull [AMAE97, ABE83, ABE85, CK04, Che82, CC09, Cob89, FFCN07, FFCN08, Goo96a, HH96, HZ03, Jen03, Kap81, LS99, LL03, Luc08, Mas03, Mou05, NW09, OM88, PT03, PH95, QC03, RB88, SCM90, WK90a, WL90, Woz94, WBE80, YX03, YR06, YXW07, YT96, Zan79, ZK86]. weight [Fro89, Goo80c, Goo81c, Goo81x, Goo83s, Goo84h, Goo85l, Goo89b, Goo89i]. Weighted [Gri02, IC80, RS89, BBP04, BS01a, Che03, CV08, EG89, Gan90, Gan93a, JG96, KPKPB95, OZ81, OP04, Rig95, Ros95, SL09, Shu08, SJ06, TA94]. weighting [MK97b]. weights [Goo83o, MRR84, Goo84x]. Welch [Gol72, Lev78d, MM93]. well [DL81, NS86, NK08]. well-behaved [DL81]. well-known [NS86, NK08]. where [EB90, Mar98]. which [Lev78a, Mag75a, WA72]. White [CNZ01, Dor88]. Whitney [Lem87, SL89a]. Who [Goo85g, HS86b]. Wicksell [Wil89]. width [IC00, Kru90b, LJZB05, XDZ09]. Widths [Woo85]. Wiener [LD87]. Wilcoxon [Fun79, Goo85o, Goo94i, HH74, Lem87, Pos82, SJR07, VW78]. Wilk [Roy89]. will [Goo81o]. win [Dav79]. Window [Woo85, IC00]. winner [CI81]. Winsorized [TT88]. Wishart [CKW73, GJ82, Jen82a]. with-replacement [PO05]. withdrawal [GSS87]. withheld [Goo88a]. within [Goo81-29, Goo82l, Soh92]. within-individual [Soh92]. without [AG81, HC06, LZ92, LN77, McN87, NL77a, RR01, SA92]. words [RV08b]. wrapped [Gat00]. Wright [CK96, Goo80m].

x [Mur78]. XII [AHJ92, Mou95, WCC07].

Yates [Bar84, Goo84o, Goo84q, Li94]. year [Goo81o]. years [Arn00, Goo90m]. Young [Aok91]. Yule [BR92, GL83].

Zeger [LSS93]. zero [BRT07, BLP09, KX03]. zero-mean [BLP09]. zerotruncated [BRT07]. zeros [Gle89b, VS93].

References

Abouammoh:2009:REG

[AA09a] A. M. Abouammoh and Arwa M. Alshingiti. Reliability estimation of generalized inverted exponential distribution. Journal of Statistical Computation and Simulation, 79(11):1301–1315, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alhakim:2009:MSG

[AA09b] Abbas Alhakim and Mufutau Akinwande. A multiple stream generator based on de Bruijn digraph homomorphisms. Journal of Statistical Computation and Simulation, 79(11):1371–1380, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Albertson:2002:PDW

[AAL02] K. Albertson, J. Aylen, and K. B. Lim. The power of the Durbin–Watson test when the errors are par(1). Journal of Statistical Computation and Simulation, 72(6):507–516, 2002. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amirkhalkhali:1993:SSE

[AAR93] Saleh Amirkhalkhali, Samad Amirkhalkhali, and Gouranga U. L. Rao. A simulation study of estimators of SUR models with unequal numbers of observations and with non-normal disturbances. Journal of Statistical Computation and Simulation, 48 (3–4):181–194, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Agarwal:2003:EPT

[AASAM03] Satish Agarwal, Jamal Al-Saleh, and D. K. Al-Mutairi. An estimator of population total in multiple characteristics and its robustness to product moment correlation coefficient. Journal of Statistical Computation and Simulation, 73(5):377–386, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ascher:1982:MMV

[AB82] Steven Ascher and John Bailar. Moments of the Mantel–Valand procedure. Journal of Statistical Computation and Simulation, 14(2):101–111, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Andrew:1997:BNC

[AB97] A. Neath Andrew and Kevin Bodden. Bayesian nonparametric confidence bounds for a distribution function. Journal of Statistical Computation and Simulation, 59(2):147–160, 1997. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anderson:2003:PTM

[AB03] Marti Anderson and Cajo Ter Braak. Permutation tests for multi-factorial analysis of variance. Journal of Statistical Computation and Simulation, 73(2):85–113, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aldridge:2005:BBM

[AB05] Gwen Aldridge and Dale Bowman. Bayesian bootstrap methods for developmental toxicity studies. Journal of Statistical Computation and Simulation, 75(2):81–91, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Angers:2008:BVU

[AB08] Jean-François Angers and Atanu Biswas. Bivariate versus univariate ordinal categorical data with reference to an ophthalmologic study. Journal of Statistical Computation and Simulation, 78(6):489–502, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmadi:2009:DFC

[AB09] J. Ahmadi and N. Balakrishnan. Distribution-free confidence intervals for quantiles and tolerance intervals in terms of k-records. Journal of Statistical Computation and Simulation, 79(10):1219– 1233, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abdullah:1989:REL

[Abd89] Mokhtar B. Abdullah. Robust estimation of a linear functional relationship. Journal of Statistical Computation and Simulation, 33(2):101–123, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abdullah:1995:OSG

[Abd95] Mokhtar B. Abdullah. One-step GM-estimators for orthogonal regression. Journal of Statistical Computation and Simulation, 52(4):301–309, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aho:1983:GFT

[ABE83] Michael Aho, Lee J. Bain, and Max Engelhardt. Goodness-of-fit tests for the Weibull distribution with unknown parameters and censored sampling. Journal of Statistical Computation and Simulation, 18(1):59–69, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aho:1985:GFT

[ABE85] Michael Aho, Lee J. Bain, and Max Engelhardt. Goodness-of-fit tests for the Weibull distribution with unknown parameters and heavy censoring. Journal of Statistical Computation and Simulation, 21(3–4):213–225, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ansell:1982:SSE

[ABH82] J. Ansell, A. Bendell, and S. Humble. On the small sample estimation of the log odds ratio in 2 × 2 contingency tables with one set of fixed margins. Journal of Statistical Computation and Simulation, 14(3–4):305–320, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anderson-Cook:1999:BR

[AC99] C. M. Anderson-Cook. Book reviews. Journal of Statistical Computation and Simulation, 64(4):383–386, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anderson-Cook:2000:SOM

[AC00] Christine M. Anderson-Cook. A second order model for cylindrical data. Journal of Statistical Computation and Simulation, 66(1):51–65, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alam:2008:CFE

[AC08] Md. Moudud Alam and Kenneth Carling. Computationally feasible estimation of the covariance structure in generalized linear mixed models. Journal of Statistical Computation and Simulation, 78(12):1229–1239, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Artiaga:2005:NTS

[ACNT05] Ramón Artiaga, Ricardo Cao, Salvador Naya, and Javier Trillo. Nonparametric two-stage plug-in adaptive smoothing for thermal analysis data. Journal of Statistical Computation and Simulation, 75(1):39–54, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alpargu:2001:EAT

[AD01] Gulhan Alpargu and Pierre Dutilleul. Efficiency analysis of ten estimation procedures for quantitative linear models with autocorrelated errors. Journal of Statistical Computation and Simulation, 69(3):257–275, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alpargu:2003:VTS

[AD03] Gülhan Alpargu and Pierre Dutilleul. To be or not to be valid in testing the significance of the slope in simple quantitative linear models with autocorrelated errors. Journal of Statistical Computation and Simulation, 73(3):165–180, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Adatia:1991:CVC

[Ada91] A. Adatia. Computation of variances and covariances of order statistics from the geometric distribution. Journal of Statistical Computation and Simulation, 39(1–2):91–94, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Adatia:1996:ABP

[Ada96] A. Adatia. Asymptotic blues of the parameters of the logistic distribution based on doubly censored samples. Journal of Statistical Computation and Simulation, 55(3):201–211, 1996. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Adatia:1997:ABS

[Ada97] A. Adatia. Approximate blue of the scale parameter of the Rayleigh distribution based on doubly censored samples. Journal of Statistical Computation and Simulation, 56(3):251–271, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Adimari:1998:ELS

[Adi98] Gianfranco Adimari. An empirical likelihood statistic for quantiles. Journal of Statistical Computation and Simulation, 60(1): 85–95, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Adkins:1996:PIR

[Adk96] Lee C. Adkins. Prior information in regression: to choose or not to choose? Journal of Statistical Computation and Simulation, 55(1–2):31–48, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aucamp:1987:DCO

[AE87] Donald C. Aucamp and Walter J. Eckardt, Jr. Doubling-up in craps and other games of chance. Journal of Statistical Computation and Simulation, 27(1):35–43, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Acusta:2003:EAD

[AEL03] Andre Acusta, Paul Eggermont, and Vincent Lariccia. An em algorithm for density estimation with randomly censored data. Journal of Statistical Computation and Simulation, 73(3):223– 232, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Allende:2006:MEA

[AFGP06] Héctor Allende, Alejandro C. Frery, Jorge Galbiati, and Luis Pizarro. M-estimators with asymmetric influence functions: the distribution case. Journal of Statistical Computation and Simulation, 76(11):941–956, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Angell:1978:IDR

[AG78] I. O. Angell and E. J. Godolphin. Implementation of the direct representation for the maximum likelihood estimator of a Gaussian moving average process. Journal of Statistical Computation and Simulation, 8(2):145–160, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Arnold:1981:CHG

[AG81] J. C. Arnold and I. J. Good. C118. How to get a large r^{2} without really trying. Journal of Statistical Computation and Simulation, 14(1):69–71, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Adegboye:1985:TAR

[AG85] O. Sola Adegboye and A. K. Gupta. On testing against restricted alternatives for Penrose model. Journal of Statistical Computation and Simulation, 22(2):147–160, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Akman:1992:CVE

[AG92] Olcay Akman and Ramesh C. Gupta. A comparison of various estimators of the mean of an inverse Gaussian distribution. Journal of Statistical Computation and Simulation, 40(1–2):71– 81, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aggarwal:1987:CMO

[Agg87] Lalit K. Aggarwal. C295. Mutually orthogonal F -rectangles of order (5, 3; 5 × 3 − 1). Journal of Statistical Computation and Simulation, 28(4):350–353, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amey:1991:MLE

[AGNS91] A. Amey, A. K. Gupta, V. Norton, and G. J. Székely. On maximum likelihood estimation of the center of a centrally symmetric convex domain. Journal of Statistical Computation and Simulation, 39(1–2):1–7, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alvarez:2006:CPR

[ÁGR06] Luis J. Álvarez, Nancy L. Garcia, and Eliane R. Rodrigues. Comparing the performance of a reversible jump Markov chain Monte Carlo algorithm for DNA sequences alignment. Journal of Statistical Computation and Simulation, 76(7):567–584, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Agustin:1999:AFS

[Agu99] Marcus A. Agustin. Asymptotic and finite-sample properties of a reliability estimator for a competing risks system. Journal of Statistical Computation and Simulation, 64(3):221–234, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Agustin:2002:GHB

[Agu02] Ma. Zenia Agustin. Generalized hazard-based goodness-of-fit tests for a repairable system. Journal of Statistical Computation and Simulation, 72(7):519–531, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Adkins:1990:ICE

[AH90] Lee C. Adkins and R. Carter Hill. An improved confidence ellipsoid for the linear regression model. Journal of Statistical Computation and Simulation, 36(1):9–18, 1990. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aknouche:2009:CAL

[AH09] Abdelhakim Aknouche and Fayçal Hamdi. Calculating the autocovariances and the likelihood for periodic V ARMA models. Journal of Statistical Computation and Simulation, 79(3):227– 239, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Hussaini:2000:FMT

[AHADA00] Essam K. Al-Hussaini, Gannat R. Al-Dayian, and Samia A. Adham. On finite mixture of two-component Gompertz lifetime model. Journal of Statistical Computation and Simulation, 67 (1):20–67, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Hussaini:2006:ALT

[AHAH06] Essam K. Al-Hussaini and Alaa H. Abdel-Hamid. Accelerated life tests under finite mixture models. Journal of Statistical Computation and Simulation, 76(8):673–690, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abdel-Hamid:2008:SPA

[AHAH08] Alaa H. Abdel-Hamid and Essam K. Al-Hussaini. Step partially accelerated life tests under finite mixture models. Journal of Statistical Computation and Simulation, 78(10):911–924, 2008. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Almasri:2007:IPT

[AHH07] Abdullah Almasri, Björn Holmquist, and Shakir Hussain. Impact of the periodicity and trend on the FD parameter estimation. Journal of Statistical Computation and Simulation, 77(1): 79–87, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Hussaini:1992:BEP

[AHJ92] Essam K. Al-Hussaini and Zeinhum F. Jaheen. Bayesian estimation of the parameters, reliability and failure rate functions of the Burr Type XII failure model. Journal of Statistical Computation and Simulation, 41(1–2):31–40, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmad:1987:GRV

[Ahm87] Khalaf E. Ahmad. Generating random vectors from a bivariate random walk density function. Journal of Statistical Computation and Simulation, 27(3):243–250, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmad:1988:JEF

[Ahm88] Ibrahim A. Ahmad. Jackknife estimation for a family of life distributions. Journal of Statistical Computation and Simulation, 29(3):614–211, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmed:1992:SPT

[Ahm92] S. E. Ahmed. Shrinkage preliminary test estimation in multivariate normal distributions. Journal of Statistical Computation and Simulation, 43(3–4):177–195, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmed:1994:PRF

[Ahm94] S. E. Ahmed. Pooling reliability functions. Journal of Statistical Computation and Simulation, 49(1–2):85–101, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmed:1999:SES

[Ahm99] S. E. Ahmed. Stmultaneous estimation of survivor functions in exponential lifetme models. Journal of Statistical Computation and Simulation, 63(3):235–261, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Hussaini:1997:MFM

[AHO97] Essam K. Al-Hussaini and Magued I. Osman. On the median of a finite mixture. Journal of Statistical Computation and Simulation, 58(2):121–144, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aitken:1983:KME

[Ait83] C. G. G. Aitken. Kernel methods for the estimation of discrete distributions. Journal of Statistical Computation and Simulation, 16(3–4):189–200, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Arvidsen:1982:VRT

[AJ82] Nils I. Arvidsen and Tommy Johnson. Variance reduction through negative correlation, a simulation study. Journal of Statistical Computation and Simulation, 15(2-3):119–127, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Jarallah:2007:BIU

[AJAH07] R. A. Al-Jarallah and E. K. Al-Hussaini. Bayes inference under a finite mixture of two-compound Gompertz components model. Journal of Statistical Computation and Simulation, 77(11):915– 927, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ansley:1985:SSS

[AK85] C. F. Ansley and R. Kohn. A structured state space approach to computing the likelihood of an ARIMA process and its derivatives. Journal of Statistical Computation and Simulation, 21 (2):135–169, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ansley:1986:NRV

[AK86a] Craig F. Ansley and Robert Kohn. A note on reparameterizing a vector autoregressive moving average model to enforce stationarity. Journal of Statistical Computation and Simulation, 24(2):99–106, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ansley:1986:SSR

[AK86b] Craig F. Ansley and Robert Kohn. Spline smoothing with repeated values. Journal of Statistical Computation and Simulation, 25(3–4):251–258, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-kandari:2005:ETE

[AkBA05] Noriah M. Al-kandari, Sana S. BuHamra, and S. E. Ahmed. Estimating and testing effect size from an arbitrary population. Journal of Statistical Computation and Simulation, 75(12):987– 1001, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aknouche:2007:CCA

[Akn07] Abdelhakim Aknouche. Causality conditions and autocovariance calculations in PVAR models. Journal of Statistical Computation and Simulation, 77(9):769–780, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ansley:1992:EES

[AKW92] Craig F. Ansley, Robert Kohn, and Chi-Ming Wong. The estimation of error standard deviation in spline regression. Journal of Statistical Computation and Simulation, 44(1–2):1–15, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Adamidis:1993:MEP

[AL93] K. Adamidis and S. Loukas. Ml estimation in the Poisson binomial distribution with grouped data via the em algorithm. Journal of Statistical Computation and Simulation, 45(1–2):33– 39, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Adamids:1994:MEB

[AL94] K. Adamids and S. Loukas. Ml estimation in the bivariate Poisson distribution in the presence of missing values via the em algorithm. Journal of Statistical Computation and Simulation, 50 (3–4):163–172, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anderson:1996:GPR

[AL96] Jon E. Anderson and Thomas A. Louis. Generating pseudorandom variables from mixture models by exemplary sampling. Journal of Statistical Computation and Simulation, 54(1-3):45– 53, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anderson:1999:ECP

[AL99] Marti J. Anderson and Pierre Legendre. An empirical comparison of permutation methods for tests of partial regression coefficients in a linear model. Journal of Statistical Computation and Simulation, 62(3):271–303, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

AlOsh:2001:SAG

[AL01] M. A. Al Osh and S. J. Lee. A simple approach for generating correlated binary variates. Journal of Statistical Computation and Simulation, 70(3):231–255, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amin:2002:EAE

[AL02] Raid Amin and Kuiyuan Li. The effect of autocorrelation on the EWMA MaxMin tolerance limits. Journal of Statistical Computation and Simulation, 72(9):719–735, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Albert:1983:PBC

[Alb83] James H. Albert. A pseudo-Bayes confidence region for p Poisson means. Journal of Statistical Computation and Simulation, 17 (1):11–29, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Albert:1984:EBE

[Alb84] James H. Albert. Empirical Bayes estimation of a set of binomial probabilities. Journal of Statistical Computation and Simulation, 20(2):129–144, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Albert:1985:SEP

[Alb85] James H. Albert. Simultaneous estimation of Poisson means under exchangeable and independence models. Journal of Statistical Computation and Simulation, 23(1–2):1–14, 1985. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Albert:1987:BEO

[Alb87] James H. Albert. Bayesian estimation of odds ratios under prior hypotheses of independence and exchangeability. Journal of Statistical Computation and Simulation, 27(3):251–268, 1987. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Albert:1992:BEP

[Alb92] James H. Albert. Bayesian estimation of the polychoric correlation coefficient. Journal of Statistical Computation and Simulation, 44(1–2):47–61, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Altman:1992:ICO

[Alt92] N. S. Altman. An iterated Cochrane–Orcutt procedure for nonparametric regression. Journal of Statistical Computation and Simulation, 40(1–2):93–108, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aly:1990:NSG

[Aly90] Emad-Eldin A. A. Aly. Normalized spacings and goodness-of-fit tests. Journal of Statistical Computation and Simulation, 36 (2-3):117–127, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmad:1987:EBE

[AM87] R. Ahmad and S. M. Mostafa. Empirical Bayes estimation of variance components in balanced random models. Journal of Statistical Computation and Simulation, 27(2):143–153, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmad:1997:ABE

[AMAE97] K. E. Ahmad, H. M. Moustafa, and A. M. Abd-Elrahman. Approximate Bayes estimation for mixtures of two Weibull distributions under type-2 censoring. Journal of Statistical Computation and Simulation, 58(3):269–285, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amoh:1985:EDF

[Amo85] R. K. Amoh. Estimation of a discriminant function from a mixture of two inverse Gaussian distributions when sample size is small. Journal of Statistical Computation and Simulation, 20 (4):275–286, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aslanargun:2007:CAN

[AMYY07] Atilla Aslanargun, Mammadagha Mammadov, Berna Yazici, and Senay Yolacan. Comparison of ARIMA, neural networks and hybrid models in time series: tourist arrival forecasting. Journal of Statistical Computation and Simulation, 77(1):29– 53, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Akdeniz:2003:NNF

[AN03] Fikri Akdeniz and Akio Namba. A note on new feasible generalized ridge regression estimator under the LINEX loss function. Journal of Statistical Computation and Simulation, 73(4):303– 310, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anatolyev:2009:RRB

[Ana09] Stanislav Anatolyev. Robustness of residual-based bootstrap to the composition of serially correlated errors. Journal of Statistical Computation and Simulation, 79(3):315–320, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Nachawati:1997:EPG

[ANAA97] H. Al-Nachawati, I. Alwasel, and A. A. Alzaid. Estimating the parameters of the generalized Poisson AR(1) process. Journal of Statistical Computation and Simulation, 56(4):337–352, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anderson:1979:NA

[And79] O. D. Anderson. News and abstracts. Journal of Statistical Computation and Simulation, 8(3–4):317–318, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Andriano:1980:ANO

[And80] K. N. Andriano. On the asymptotic normality and other large sample properties of Grenander’s mode estimator. Journal of Statistical Computation and Simulation, 10(3–4):271–288, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anderson:1989:CSC

[And89] Oliver D. Anderson. C307. Serial correlation: the bending moment analogy. Journal of Statistical Computation and Simulation, 31(1):40–44, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See errata [Ano90a].

Anderson:1990:CAS

[And90a] Oliver D. Anderson. C353. Applications of simulation in time series research. Journal of Statistical Computation and Simulation, 36(1):31–64, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anderson:1990:CFT

[And90b] Oliver D. Anderson. C375. Finite time-series properties versus asymptotic theory in approximating sample serial correlation structure. Journal of Statistical Computation and Simulation, 37 (3–4):236–240, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anderson:1992:CMO

[And92] Oliver D. Anderson. C384. Modified order notation with an example from time series theory. Journal of Statistical Computation and Simulation, 40(3–4):257–259, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anderson:1993:CCL

[And93] Oliver D. Anderson. C407. A closer look at an example from time series theory. Journal of Statistical Computation and Simulation, 47(1–2):124–126, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anderson:1995:CPA

[And95a] O. D. Anderson. C427. The precision of approximations to moments of sample moments. Journal of Statistical Computation and Simulation, 52(2):163–165, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anderson:1995:CIN

[And95b] Oliver D. Anderson. C428. An interesting non-linear transformation of the theoretical autocorrelation function. Journal of Statistical Computation and Simulation, 52(2):165–167, 1995. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anderson:1995:CFL

[And95c] Oliver D. Anderson. C429. Finite lengthed time-series properties versus asymptotic theory in approximating sample serial correlation structure: part 2. Journal of Statistical Computation and Simulation, 52(2):167–171, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Andres:1997:SMS

[And97] T. H. Andres. Sampling methods and sensitivity analysis for large parameter sets. Journal of Statistical Computation and Simulation, 57(1-4):77–110, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Angelis:2003:EAO

[Ang03] Lefteris Angelis. An evolutionary algorithm for a-optimal incomplete block designs. Journal of Statistical Computation and Simulation, 73(10):753–771, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Nasser:2009:RER

[ANM09] Amjad D. Al-Nasser and Ahmed Bani Mustafa. Robust extreme ranked set sampling. Journal of Statistical Computation and Simulation, 79(7):859–867, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1972:BR

[Ano72a] Anonymous. Book review. Journal of Statistical Computation and Simulation, 1(2):195–196, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1972:E

[Ano72b] Anonymous. Errata. Journal of Statistical Computation and Simulation, 1(4):395, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1973:AFPa

[Ano73a] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 2(1):83–84, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Anonymous:1973:AFPb

[Ano73b] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 2(2):193–195, 1973. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1973:AFPc

[Ano73c] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 2(3):291–292, 1973. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1973:AFPd

[Ano73d] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 2(4):393–394, 1973. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1973:EC

[Ano73e] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 2(2):189–190, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1973:E

[Ano73f] Anonymous. Errata. Journal of Statistical Computation and Simulation, 2(1):149, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1974:AFPa

[Ano74a] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 3(1):96–97, 1974. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1974:AFPb

[Ano74b] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 3(2):199–200, 1974. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1975:AFPa

[Ano75a] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 3(3):96–97, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1975:AFPb

[Ano75b] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 3(4):395–396, 1975. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1975:AFPc

[Ano75c] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 4(1):83–84, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1975:AFPd

[Ano75d] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 4(2):167–168, 1975. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1975:AFPe

[Ano75e] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 4(3):245–246, 1975. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1975:BR

[Ano75f] Anonymous. Book review. Journal of Statistical Computation and Simulation, 3(4):392, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1975:EC

[Ano75g] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 3(4):394, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1975:E

[Ano75h] Anonymous. Errata. Journal of Statistical Computation and Simulation, 4(1):81–82, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1976:AFPa

[Ano76a] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 4(4):327–328, 1976. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1976:AFPb

[Ano76b] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 5(1):83–84, 1976. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1977:AFPa

[Ano77a] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 5(2):167–168, 1977. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1977:AFPb

[Ano77b] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 5(3):245–246, 1977. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1977:AFPc

[Ano77c] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 5(4):321–322, 1977. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1977:AFPd

[Ano77d] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 6(1):73–74, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1977:AFPe

[Ano77e] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 6(2):155–156, 1977. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1977:EBa

[Ano77f] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 6(1):ebi, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1977:EBb

[Ano77g] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 6(2):ebi, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1977:EC

[Ano77h] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 6(2):153–154, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1978:AFPa

[Ano78a] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 6(3–4):319–320, 1978. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1978:AFPb

[Ano78b] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 7(1):83, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1978:AFPc

[Ano78c] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 7(2):165, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1978:AFPd

[Ano78d] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 7(3–4):303, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1978:AFPe

[Ano78e] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 8(1):83–84, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1978:AFPf

[Ano78f] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 8(2):171–173, 1978. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1978:CSO

[Ano78g] Anonymous. C12. Should outliers be dragged in? Journal of Statistical Computation and Simulation, 7(2):163–164, 1978. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1978:EBa

[Ano78h] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 6(3–4):ebi, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1978:EBb

[Ano78i] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 7(1):ebi, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1978:EBc

[Ano78j] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 7(2):ebi, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1978:EBd

[Ano78k] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 7(3–4):ebi, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1978:EC

[Ano78l] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 8(2):169–170, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1979:AFPa

[Ano79a] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 8(3–4):319–320, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1979:AFPb

[Ano79b] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 9(1):81, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1979:AFPc

[Ano79c] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 9(2):167, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1979:AFPd

[Ano79d] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 9(3):251–252, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1979:AFPe

[Ano79e] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 9(4):320, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1979:AFPf

[Ano79f] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 10(1):85–86, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1979:EBa

[Ano79g] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 9(1):ebi, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1979:EBb

[Ano79h] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 9(2):ebi, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1979:EBc

[Ano79i] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 9(3):ebi, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1979:EBd

[Ano79j] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 9(4):ebi, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1979:EBe

[Ano79k] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 10(1):ebi, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1979:E

[Ano79l] Anonymous. Erratum. Journal of Statistical Computation and Simulation, 9(3):249, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1979:NA

[Ano79m] Anonymous. News and abstracts. Journal of Statistical Computation and Simulation, 9(3):247–249, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:AFPa

[Ano80a] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 10(2):175–176, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:AFPb

[Ano80b] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 10(3–4):321, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:AFPc

[Ano80c] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 11(1):77–78, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:AFPd

[Ano80d] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 11(2):161–162, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:AFPe

[Ano80e] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 11(3–4):319–320, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:AFPf

[Ano80f] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 12(1):79–80, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:EBa

[Ano80g] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 10(2):ebi, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:EBb

[Ano80h] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 10(3–4):ebi, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:EBc

[Ano80i] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 11(1):ebi, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:EBd

[Ano80j] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 11(2):ebi, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:EBe

[Ano80k] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 11(3–4):ebi, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:EBf

[Ano80l] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 12(1):ebi, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:ECa

[Ano80m] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 10(2):171–173, 1980. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:ECb

[Ano80n] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 11(3–4):315–316, 1980. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:NAa

[Ano80o] Anonymous. News and abstracts. Journal of Statistical Computation and Simulation, 10(2):169–170, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:NAb

[Ano80p] Anonymous. News and abstracts. Journal of Statistical Computation and Simulation, 10(3–4):319–320, 1980. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:NAc

[Ano80q] Anonymous. News and abstracts. Journal of Statistical Computation and Simulation, 11(1):75, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:NAd

[Ano80r] Anonymous. News and abstracts. Journal of Statistical Computation and Simulation, 11(3–4):317–318, 1980. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1980:T

[Ano80s] Anonymous. Tutorial. Journal of Statistical Computation and Simulation, 10(2):147–158, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:AFPa

[Ano81a] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 12(2):155–156, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:AFPb

[Ano81b] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 12(3–4):319–320, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:AFPc

[Ano81c] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 13(1):77–78, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:AFPd

[Ano81d] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 13(2):167–168, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:AFPe

[Ano81e] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 13(3–4):331–332, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:CGD

[Ano81f] Anonymous. C91. Generalized determinants and generalized variance. Journal of Statistical Computation and Simulation, 12 (3–4):311–313, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:EBa

[Ano81g] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 12(2):ebi, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:EBb

[Ano81h] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 12(3–4):ebi, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:EBc

[Ano81i] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 13(1):ebi, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:EBd

[Ano81j] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 13(2):ebi, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:EBe

[Ano81k] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 13(3–4):ebi, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:EBf

[Ano81l] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 14(1):ebi, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:ECa

[Ano81m] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 14(1):83–85, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:Ea

[Ano81n] Anonymous. Errata. Journal of Statistical Computation and Simulation, 13(1):63, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:Ec

[Ano81o] Anonymous. Errata. Journal of Statistical Computation and Simulation, 13(2):149, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:Ed

[Ano81p] Anonymous. Errata. Journal of Statistical Computation and Simulation, 13(3–4):329, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:NAa

[Ano81q] Anonymous. News and abstracts. Journal of Statistical Computation and Simulation, 12(2):153, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:NAb

[Ano81r] Anonymous. News and abstracts. Journal of Statistical Computation and Simulation, 12(3–4):317–318, 1981. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:NAc

[Ano81s] Anonymous. News and abstracts. Journal of Statistical Computation and Simulation, 13(1):64, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:NAd

[Ano81t] Anonymous. News and abstracts. Journal of Statistical Computation and Simulation, 13(2):64, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1981:NAe

[Ano81u] Anonymous. News and abstracts. Journal of Statistical Computation and Simulation, 14(1):81–82, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1982:AFPa

[Ano82a] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 14(2):167–168, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1982:AFPb

[Ano82b] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 14(3–4):325–326, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1982:AFPc

[Ano82c] Anonymous. Abstracts to forthcoming papers. Journal of Statistical Computation and Simulation, 15(2-3):249–250, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1982:CPH

[Ano82d] Anonymous. C121. The postulate of honest judgement. Journal of Statistical Computation and Simulation, 14(3–4):321, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1982:EBa

[Ano82e] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 14(2):ebi, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1982:EBb

[Ano82f] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 14(3–4):ebi, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1982:EBc

[Ano82g] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 15(1):ebi, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1982:EBd

[Ano82h] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 15(2-3):ebi, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1982:EBe

[Ano82i] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 15(4):ebi, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1982:EBf

[Ano82j] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 16(1):ebi, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1982:EBg

[Ano82k] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 16(2):ebi, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1982:EC

[Ano82l] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 16(2):157–159, 1982. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1982:E

[Ano82m] Anonymous. Errata. Journal of Statistical Computation and Simulation, 14(2):165, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1982:NA

[Ano82n] Anonymous. News and abstracts. Journal of Statistical Computation and Simulation, 16(1):77–79, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1983:EB

[Ano83a] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 16(3–4):ebi, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1983:Ea

[Ano83b] Anonymous. Errata. Journal of Statistical Computation and Simulation, 18(1):85, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1983:Ec

[Ano83c] Anonymous. Errata. Journal of Statistical Computation and Simulation, 18(2-3):246, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1983:NAa

[Ano83d] Anonymous. News and abstracts. Journal of Statistical Computation and Simulation, 16(3–4):317, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1983:NAb

[Ano83e] Anonymous. News and abstracts. Journal of Statistical Computation and Simulation, 17(2):157, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1984:EBa

[Ano84a] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 20(1):ebi, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1984:EBb

[Ano84b] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 20(2):ebi, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1984:EBc

[Ano84c] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 20(3):ebi, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1984:EC

[Ano84d] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 20(3):271–273, 1984. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1984:E

[Ano84e] Anonymous. Errata. Journal of Statistical Computation and Simulation, 20(1):89, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1984:NAa

[Ano84f] Anonymous. News and abstracts. Journal of Statistical Computation and Simulation, 20(1):87–88, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1984:NAb

[Ano84g] Anonymous. News and abstracts. Journal of Statistical Computation and Simulation, 20(2):181, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1985:EBa

[Ano85a] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 20(4):ebi, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1985:EBb

[Ano85b] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 21(1):ebi, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1985:EBc

[Ano85c] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 21(2):ebi, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1985:EBd

[Ano85d] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 21(3–4):ebi, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1985:EBe

[Ano85e] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 22(1):ebi, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1985:EBf

[Ano85f] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 22(2):ebi, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1985:EBg

[Ano85g] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 22(3–4):ebi, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1985:EBh

[Ano85h] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 23(1–2):ebi, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1985:Ea

[Ano85i] Anonymous. Errata. Journal of Statistical Computation and Simulation, 20(4):341, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1985:Ec

[Ano85j] Anonymous. Errata. Journal of Statistical Computation and Simulation, 22(2):187, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1986:EBa

[Ano86a] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 23(3):ebi, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1986:EBb

[Ano86b] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 23(4):ebi, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1986:EBc

[Ano86c] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 24(1):ebi, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1986:EBd

[Ano86d] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 24(2):ebi, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1986:EBe

[Ano86e] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 25(1–2):ebi, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1986:EBf

[Ano86f] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 25(3–4):ebi, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1986:EBg

[Ano86g] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 26(1–2):ebi, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1986:EBh

[Ano86h] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 26(3–4):ebi, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1986:EC

[Ano86i] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 23(4):329–331, 1986. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1987:EBa

[Ano87a] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 27(1):ebi, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1987:EBb

[Ano87b] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 27(2):ebi, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1987:EBc

[Ano87c] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 27(3):ebi, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1987:EBd

[Ano87d] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 27(4):ebi, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1987:EBe

[Ano87e] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 28(1):ebi, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1987:EBf

[Ano87f] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 28(2):ebi, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1987:EBg

[Ano87g] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 28(3):ebi, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1987:EBh

[Ano87h] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 28(4):ebi, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1987:ECa

[Ano87i] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 28(2):187–189, 1987. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1987:Ea

[Ano87j] Anonymous. Errata. Journal of Statistical Computation and Simulation, 27(1):93, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1987:Ec

[Ano87k] Anonymous. Errata. Journal of Statistical Computation and Simulation, 28(1):95–96, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1987:Ed

[Ano87l] Anonymous. Errata. Journal of Statistical Computation and Simulation, 28(2):185, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1988:EBa

[Ano88a] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 29(1):ebi, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1988:EBb

[Ano88b] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 29(2):ebi, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1988:EBc

[Ano88c] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 29(3):ebi, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1988:EBd

[Ano88d] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 29(4):ebi, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1988:EP

[Ano88e] Anonymous. Editorial policy. Journal of Statistical Computation and Simulation, 30(2):145, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1988:Ea

[Ano88f] Anonymous. Errata. Journal of Statistical Computation and Simulation, 30(2):157–158, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1988:Ec

[Ano88g] Anonymous. Erratum. Journal of Statistical Computation and Simulation, 30(4):315, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1988:SS

[Ano88h] Anonymous. Statistical software. Journal of Statistical Computation and Simulation, 30(2):147, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1989:EBa

[Ano89a] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 31(1):ebi, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1989:EBb

[Ano89b] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 31(2):ebi, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1989:EBc

[Ano89c] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 31(3):ebi, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1989:EBd

[Ano89d] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 31(4):ebi, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1989:EBe

[Ano89e] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 32(1–2):ebi, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1989:EBf

[Ano89f] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 32(3):ebi, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1989:EBg

[Ano89g] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 32(4):ebi, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1989:ECa

[Ano89h] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 31(1):65–66, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1989:ECb

[Ano89i] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 33(3):197–198, 1989. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1989:E

[Ano89j] Anonymous. Errata. Journal of Statistical Computation and Simulation, 31(1):67, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1990:CEC

[Ano90a] Anonymous. C350. Errata to C202, 305, 307, 310, 311, and 334. Journal of Statistical Computation and Simulation, 34(2-3):176, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo84m, Goo88f, And89, JK89, Goo89a, Goo89u].

Anonymous:1990:EBa

[Ano90b] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 35(1–2):ebi, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1990:EBb

[Ano90c] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 35(3–4):ebi, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1990:EBc

[Ano90d] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 36(1):ebi, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1990:EBd

[Ano90e] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 36(2-3):ebi, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1990:EBe

[Ano90f] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 36(4):ebi, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1990:EBf

[Ano90g] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 37(1–2):ebi, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1990:EBh

[Ano90h] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 37(3–4):ebi, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1990:EC

[Ano90i] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 37(3–4):249–251, 1990. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1991:EBa

[Ano91a] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 38(1-4):ebi, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1991:EBb

[Ano91b] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 39(1–2):ebi, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1991:EBc

[Ano91c] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 39(3):ebi, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1991:EBd

[Ano91d] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 39(4):ebi, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1992:EBa

[Ano92a] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 40(1–2):ebi, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1992:EBb

[Ano92b] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 40(3–4):ebi, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1992:EBc

[Ano92c] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 41(1–2):ebi, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1992:EBd

[Ano92d] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 41(3–4):ebi, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1992:EBe

[Ano92e] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 42(1–2):ebi, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1992:EBf

[Ano92f] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 42(3–4):ebi, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1992:EBg

[Ano92g] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 43(1–2):ebi, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1992:EBh

[Ano92h] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 43(3–4):ebi, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1992:EBi

[Ano92i] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 44(1–2):ebi, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1992:EC

[Ano92j] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 41(1–2):125–127, 1992. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1993:EBa

[Ano93a] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 44(3–4):ebi, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1993:EBb

[Ano93b] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 45(1–2):ebi, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1993:EBc

[Ano93c] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 45(3–4):ebi, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1993:EBd

[Ano93d] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 46(1–2):ebi, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1993:EBe

[Ano93e] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 46(3–4):ebi, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1993:EBf

[Ano93f] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 48(1–2):ebi, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1993:EBg

[Ano93g] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 48(3–4):ebi, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1993:EC

[Ano93h] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 45(3–4):257–261, 1993. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1994:EBa

[Ano94a] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 50(1–2):ebi, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1994:EBb

[Ano94b] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 50(3–4):ebi, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1994:EBc

[Ano94c] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 51(1):ebi, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1994:EC

[Ano94d] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 49(1–2):125–128, 1994. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1995:EBa

[Ano95a] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 51(2-4):ebi, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1995:EBb

[Ano95b] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 52(1):ebi, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1995:EBc

[Ano95c] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 52(2):ebi, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1995:EBd

[Ano95d] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 52(3):ebi, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1995:EBe

[Ano95e] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 52(4):ebi, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1995:EBf

[Ano95f] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 53(1–2):ebi, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1995:EBg

[Ano95g] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 53(3–4):ebi, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1995:EC

[Ano95h] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 51(2-4):405–407, 1995. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1996:EBa

[Ano96a] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 54(1-3):ebi, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1996:EBb

[Ano96b] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 54(4):ebi, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1996:EBc

[Ano96c] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 55(1–2):ebi, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1996:EBd

[Ano96d] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 55(3):ebi, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1996:EBe

[Ano96e] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 55(4):ebi, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1996:EBf

[Ano96f] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 56(1):ebi, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1996:EC

[Ano96g] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 54(4):393–395, 1996. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1997:EBa

[Ano97a] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 56(2):ebi, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1997:EBb

[Ano97b] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 56(3):ebi, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1997:EBc

[Ano97c] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 56(4):ebi, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1997:EBd

[Ano97d] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 57(1-4):ebi, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1997:EBe

[Ano97e] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 58(1):ebi, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1997:EBf

[Ano97f] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 58(2):ebi, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1997:EBg

[Ano97g] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 58(3):ebi, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1997:EBh

[Ano97h] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 58(4):ebi, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1997:EBi

[Ano97i] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 59(1):ebi, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1997:EBj

[Ano97j] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 59(2):ebi, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1997:EBk

[Ano97k] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 59(3):ebi, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1997:EBl

[Ano97l] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 59(4):ebi, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1997:ECa

[Ano97m] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 56(2):189–191, 1997. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1997:ECb

[Ano97n] Anonymous. Editorial collaborators. Journal of Statistical Computation and Simulation, 59(3):291–294, 1997. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1998:EBa

[Ano98a] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 60(1):ebi, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1998:EBb

[Ano98b] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 60(2):ebi, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1998:EBc

[Ano98c] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 60(3):ebi, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1998:EBd

[Ano98d] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 60(4):ebi, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1998:EBe

[Ano98e] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 61(1–2):ebi, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1998:EBf

[Ano98f] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 61(3):ebi, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1998:EBg

[Ano98g] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 61(4):ebi, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1999:EBa

[Ano99a] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 62(3):ebi, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1999:EBb

[Ano99b] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 62(4):ebi, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1999:EBc

[Ano99c] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 63(1):1, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1999:EBd

[Ano99d] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 63(3):ebi, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1999:EBe

[Ano99e] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 63(4):ebi, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1999:EBf

[Ano99f] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 64(1):ebi, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1999:EBg

[Ano99g] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 64(2):ebi, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1999:EBh

[Ano99h] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 64(3):ebi, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:1999:EBi

[Ano99i] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 64(4):ebi, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2000:EBa

[Ano00a] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 66(1):ebi, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2000:EBb

[Ano00b] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 66(2):ebi, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2000:EBc

[Ano00c] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 66(3):ebi, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2000:EBd

[Ano00d] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 66(4):ebi, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2000:EBe

[Ano00e] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 67(1):ebi, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2000:EBf

[Ano00f] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 67(2):ebi, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2000:EBg

[Ano00g] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 67(3):ebi, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2000:EBh

[Ano00h] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 67(4):ebi, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2000:EBi

[Ano00i] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 68(1):ebi, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2000:ENa

[Ano00j] Anonymous. Editorial note. Journal of Statistical Computation and Simulation, 66(2):iii, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2000:ENb

[Ano00k] Anonymous. Editorial note. Journal of Statistical Computation and Simulation, 66(3):iii, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2000:ENc

[Ano00l] Anonymous. Editorial note. Journal of Statistical Computation and Simulation, 66(4):iii, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2001:EBa

[Ano01a] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 68(2):ebi, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2001:EBb

[Ano01b] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 68(3):ebi, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2001:EBc

[Ano01c] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 68(4):ebi, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2001:EBd

[Ano01d] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 69(1):ebi, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2001:EBe

[Ano01e] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 69(2):ebi, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2001:EBf

[Ano01f] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 69(3):ebi, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2001:EBg

[Ano01g] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 69(4):ebi, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2001:EBh

[Ano01h] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 70(1):ebi, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2001:EBi

[Ano01i] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 70(2):ebi, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2001:EBj

[Ano01j] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 70(3):ebi, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2001:EBk

[Ano01k] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 70(4):ebi, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2001:EBl

[Ano01l] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 71(1):ebi, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2001:EBm

[Ano01m] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 71(2):ebi, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2001:EBn

[Ano01n] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 71(3):ebi, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2001:EBo

[Ano01o] Anonymous. Editorial board. Journal of Statistical Computation and Simulation, 71(4):ebi, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2002:BRa

[Ano02a] Anonymous. Book review. Journal of Statistical Computation and Simulation, 72(1):99–100, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2002:BRb

[Ano02b] Anonymous. Book review. Journal of Statistical Computation and Simulation, 72(3):261, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2002:BRc

[Ano02c] Anonymous. Book review. Journal of Statistical Computation and Simulation, 72(4):355, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2002:BRd

[Ano02d] Anonymous. Book review. Journal of Statistical Computation and Simulation, 72(7):597–598, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2002:CCC

[Ano02e] Anonymous. Comments, conjectures and conclusions. Journal of Statistical Computation and Simulation, 72(1):1–31, 2002. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2003:CEF

[Ano03a] Anonymous. C474. Errors of the fourth kind. Journal of Statistical Computation and Simulation, 73(8):617, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2003:CEC

[Ano03b] Anonymous. C475. Erratum to C462 “More on the relative efficacy of probability scoring rules”, JSCS 72, 2002, pp. 11–14. Journal of Statistical Computation and Simulation, 73(8):616, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2004:BRa

[Ano04a] Anonymous. Book review. Journal of Statistical Computation and Simulation, 74(5):387–390, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2004:BRb

[Ano04b] Anonymous. Book review. Journal of Statistical Computation and Simulation, 74(10):777, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2005:BRa

[Ano05a] Anonymous. Book review. Journal of Statistical Computation and Simulation, 75(7):589–593, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2005:BRb

[Ano05b] Anonymous. Book review. Journal of Statistical Computation and Simulation, 75(10):853–855, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2005:BRc

[Ano05c] Anonymous. Book review. Journal of Statistical Computation and Simulation, 75(11):935–939, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2005:BRR

[Ano05d] Anonymous. Book review a review of advanced biostatistics and statistical genetics books. Journal of Statistical Computation and Simulation, 75(9):761–769, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2005:BRT

[Ano05e] Anonymous. Book review three books on statistical power analysis and related ideas. Journal of Statistical Computation and Simulation, 75(12):1019–1020, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2006:BRC

[Ano06a] Anonymous. Book review a collection of classic books for multivariate data analysis. Journal of Statistical Computation and Simulation, 76(3):267–278, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2006:BRB

[Ano06b] Anonymous. Book review books on philosophy or foundations of statistical concepts. Journal of Statistical Computation and Simulation, 76(9):839–845, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2006:BRE

[Ano06c] Anonymous. Book review excellent dictionaries and handbooks useful to statisticians. Journal of Statistical Computation and Simulation, 76(2):181–184, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2006:BRM

[Ano06d] Anonymous. Book review multiple view-points of multivariate books. Journal of Statistical Computation and Simulation, 76 (6):561–565, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2006:BRR

[Ano06e] Anonymous. Book review regression and related books are reviewed. Journal of Statistical Computation and Simulation, 76 (1):83–91, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2006:BRS

[Ano06f] Anonymous. Book review sensational books about six sigma related concepts and applications. Journal of Statistical Computation and Simulation, 76(7):655–659, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2006:EN

[Ano06g] Anonymous. Editor’s note. Journal of Statistical Computation and Simulation, 76(5):371, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2008:BR

[Ano08] Anonymous. Book review. Journal of Statistical Computation and Simulation, 78(10):967–968, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Andre:1997:APM

[ANPV97] Carmen D. S. André, Subash C. Narula, Clovis A. Peres, and Gilberto A. Ventura. Asymptotic properties of the minimum sum of absolute errors estimators in a dose-response model. Journal of Statistical Computation and Simulation, 58(4):361– 370, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ansley:1980:CTA

[Ans80] Craig F. Ansley. Computation of the theoretical autocovariance function for a vector Arima process. Journal of Statistical Computation and Simulation, 12(1):15–24, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anton:1995:SSC

[Ant95] Formann K. Anton. Small-sample comparison of the exact and asymptotic upper tail probabilities of chi-squared goodness-of-fit statistics: Pearson’s χ^{2}, likelihood ratio, and power-divergence statistic (λ = 2/3). Journal of Statistical Computation and Simulation, 51(2-4):369–384, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aoki:1991:CCC

[Aok91] Masanao Aoki. C379. Comments on Crabbe and Young. Journal of Statistical Computation and Simulation, 39(3):175–178, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [CY89].

Aoki:2002:IPT

[Aok02] Satoshi Aoki. Improving path trimming in a network algorithm for Fisher’s exact test in two-way contingency tables. Journal of Statistical Computation and Simulation, 72(3):205–216, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Acar:1985:PUN

[AP85] Serpil B. Acar and A. N. Pettitt. Prediction using nonparametric techniques: some comparisons. Journal of Statistical Computation and Simulation, 21(2):117–134, 1985. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Arazyan:2006:RMS

[AP06] Alvard Y. Arazyan and Nabendu Pal. On the reliability of a multicomponent series system. Journal of Statistical Computation and Simulation, 76(1):1–20, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Awad:2000:PIF

[AR00] Adnan M. Awad and Mohammad Z. Raqab. Prediction intervals for the future record values from exponential distribution: comparative study. Journal of Statistical Computation and Simulation, 65(1-4):325–340, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ambler:2001:FPM

[AR01] Gareth Ambler and Patrick Royston. Fractional polynomial model selection procedures: investigation of Type I error rate. Journal of Statistical Computation and Simulation, 69(1):89– 108, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Archer:1980:GPL

[Arc80] N. P. Archer. The generation of piecewise linear approximations of probability distribution functions. Journal of Statistical Computation and Simulation, 11(1):21–40, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Arnold:1979:CIF

[Arn79] J. C. Arnold. C37. An improved form of multiple-choice examinations. Journal of Statistical Computation and Simulation, 9 (1):73, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Arnold:2000:VTD

[Arn00] Jesse C. Arnold. Virginia Tech Department of Statistics: the first fifty years. Journal of Statistical Computation and Simulation, 66(1):1–17, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Arsham:1986:GKC

[Ars86] Hossein Arsham. Generalised K–S confidence regions: some exact results. Journal of Statistical Computation and Simulation, 25(1–2):9–23, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ali:1981:AEI

[AS81] M. A. Ali and N. Singh. An alternative estimator in inverse linear regression. Journal of Statistical Computation and Simulation, 14(1):1–15, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmed:1989:PMD

[AS89] S. E. Ahmed and A. K. Md. E. Saleh. Pooling multivariate data. Journal of Statistical Computation and Simulation, 31(3):149– 167, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ali:1990:EMV

[AS90] Abunnabi M. Ali and Md A. K. Ehanses Saleh. Estimation of the mean vector of a multivariate normal distribution under symmetry. Journal of Statistical Computation and Simulation, 35 (3–4):209–226, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Subaihi:2000:MCS

[AS00] Ali A. Al-Subaihi. A Monte Carlo study of the Friedman and Conover tests in the single-factor repeated measures design.

Journal of Statistical Computation and Simulation, 65(1-4):203– 223, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alkarni:2001:UBD

[AS01] S. H. Alkarni and M. M. Siddiqui. An upper bound for the distribution function of a positive definite quadratic form. Journal of Statistical Computation and Simulation, 69(1):41–56, 2001. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Angelini:2004:EBA

[AS04] Claudia Angelini and Theofanis Sapatinas. Empirical Bayes approach to wavelet regression using ϵ-contaminated priors. Journal of Statistical Computation and Simulation, 74(10):741–764, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Archer:1997:SMA

[ASS97] G. E. B. Archer, A. Saltelli, and I. M. Sobol. Sensitivity measures, Anova-like techniques and the use of bootstrap. Journal of Statistical Computation and Simulation, 58(2):99–120, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Saleh:2004:EVN

[ASS04] Mohammad Fraiwan Al-Saleh and Hani M. Samawi. Estimating the variance of a normal population by utilizing the information in a sample from a second related normal population. Journal of Statistical Computation and Simulation, 74(2):79–90, 2004. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Saadi:1979:TIE

[ASSY79] S. D. Al-Saadi, D. F. Scrimshaw, and D. H. Young. Tests for independence of exponential variables. Journal of Statistical Computation and Simulation, 9(3):217–233, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Saadi:1980:ECC

[ASY80] S. D. Al-Saadi and D. H. Young. Estimators for the correlation coefficient in a bivariate exponential distribution. Journal of Statistical Computation and Simulation, 11(1):13–20, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Saadi:1981:CVS

[ASY81] S. D. Al-Saadi and D. H. Young. Critical values of some nonparametric statistics used for tests of total stochastic independence. Journal of Statistical Computation and Simulation, 12 (3–4):217–224, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Saadi:1982:TIM

[ASY82] S. D. Al-Saadi and D. H. Young. A test for independence in a multivariate exponential distribution with equal correlation coefficients. Journal of Statistical Computation and Simulation, 14 (3–4):219–227, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-sarraf:1986:ERL

[AsY86] Z. Al-sarraf and D. H. Young. The extreme residuals in logistic regression models. Journal of Statistical Computation and Simulation, 25(1–2):115–125, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anderson:1995:CTS

[AT95] N. H. Anderson and D. M. Titterington. A comparison of two statistics for detecting clustering in one dimension. Journal of Statistical Computation and Simulation, 53(1–2):103–125, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aoki:2005:MCM

[AT05] Satoshi Aoki and Akimichi Takemura. Markov chain Monte Carlo exact tests for incomplete two-way contingency tables. Journal of Statistical Computation and Simulation, 75(10):787– 812, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Atkinson:1991:APH

[Atk91] Neely E. Atkinson. Accurate proportional hazards estimation in ill-conditioned cases. Journal of Statistical Computation and Simulation, 39(4):193–203, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Atkinson:1992:PSH

[Atk92] Scott E. Atkinson. The performance of standard and hybrid em algorithms for ML estimates of the normal mixture model with censoring. Journal of Statistical Computation and Simulation, 44(1–2):105–115, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aucamp:1979:CQA

[Auc79] Donald C. Aucamp. C55. Quick and accurate rule of thumb for determining the critical value of the coefficient of correlation. Journal of Statistical Computation and Simulation, 9(4):311– 313, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aucamp:1986:TDM

[Auc86] Donald C. Aucamp. A test for the difference of means. Journal of Statistical Computation and Simulation, 24(1):33–46, 1986. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Attanasio:1995:TSS

[AW95] Massimo Attanasio and Weng Kee Wong. Transformations for stabilizing spread in a time series model: a comparative study. Journal of Statistical Computation and Simulation, 52(1):55– 70, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aslan:2005:NTM

[AZ05] B. Aslan and G. Zech. New test for the multivariate two-sample problem based on the concept of minimum energy. Journal of Statistical Computation and Simulation, 75(2):109–119, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Azalost:1973:SSS

[Aza73] Barbara L. Azalost. Stochastic simulation study of martin’ distributed lag model. Journal of Statistical Computation and Simulation, 2(1):55–64, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bennett:1977:APF

[BA77] B. M. Bennett and M. P. Allshouse. On an approximation to the power function of the Mantel Haenszel test for relative risk.

Journal of Statistical Computation and Simulation, 5(2):91–97, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Barenghi:1986:DOS

[BAB86] Carlo F. Barenghi, Amir D. Aczel, and Roger J. Best. Determining the optimal sample size for decision making. Journal of Statistical Computation and Simulation, 24(2):135–145, 1986. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bailer:1989:TVE

[Bai89] A. John Bailer. Testing variance equality with randomization tests. Journal of Statistical Computation and Simulation, 31(1): 1–8, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Baillo:2009:NFL

[Baí09] Amparo Baíllo. A note on functional linear regression. Journal of Statistical Computation and Simulation, 79(5):657–669, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Baksalary:1987:CCA

[Bak87] Jerzy K. Baksalary. C292. A comment on an admissibility criterion. Journal of Statistical Computation and Simulation, 28 (4):345–347, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Baklizi:2007:IAM

[Bak07] Ayman Baklizi. Inference about the mean difference of two nonnormal populations based on independent samples: a comparative study. Journal of Statistical Computation and Simulation, 77(7):613–624, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Baklizi:2008:IAM

[Bak08] Ayman Baklizi. Inference about the mean of a skewed population: a comparative study. Journal of Statistical Computation and Simulation, 78(3):421–435, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishanan:1983:EPS

[Bal83] N. Balakrishanan. Empirical power study of a multi-sample test of exponentiality based on spacings. Journal of Statistical Computation and Simulation, 18(4):265–271, 1983. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:1985:OSH

[Bal85] N. Balakrishnan. Order statistics from the half logistic distribution. Journal of Statistical Computation and Simulation, 20 (4):287–309, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:1989:AML

[Bal89] N. Balakrishnan. Approximate maximum likelihood estimation of the mean and standard deviation of the normal distribution based on Type II censored samples. Journal of Statistical Computation and Simulation, 32(3):137–148, 1989. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See corrigendum [Bal92].

Balakrishnan:1992:CAM

[Bal92] N. Balakrishnan. Corrigendum to approximate maximum likelihood estimation of the mean and standard deviation of the normal distribution based on Type II censored samples. Journal of Statistical Computation and Simulation, 43(1–2):123, 1992. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Bal89].

Balasooriya:1995:FCR

[Bal95] Uditha Balasooriya. Failure-censored reliability sampling plans for the exponential distribution. Journal of Statistical Computation and Simulation, 52(4):337–349, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Banfield:1978:MAG

[Ban78] C. F. Banfield. Multivariate analysis in Genstat. Journal of Statistical Computation and Simulation, 6(3–4):211–222, 1978. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Barricell:1972:NTE

[Bar72] Nils Aall Barricell. Numerical testing of evolution theories. Journal of Statistical Computation and Simulation, 1(2):97–127, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bartels:1977:EFO

[Bar77] R. Bartels. Estimation in a first order autoregressive scheme with non-normal stable disturbances. Journal of Statistical Computation and Simulation, 6(1):35–48, 1977. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bartels:1978:GNN

[Bar78] R. Bartels. Generating non- normal stable variates using limit theorem properties. Journal of Statistical Computation and Simulation, 7(3–4):199–212, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bartels:1979:FNS

[Bar79] Robert Bartels. Fractiles for the nonsymmetric stable distributions. Journal of Statistical Computation and Simulation, 9 (2):127–132, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bartels:1981:TBI

[Bar81] Robert Bartels. Truncation bounds for infinite expansions for the stable distributions. Journal of Statistical Computation and Simulation, 12(3–4):293–302, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Barnard:1984:CEH

[Bar84] G. A. Barnard. C207. The early history of the Fisher–Yates– Irwin formula. Journal of Statistical Computation and Simulation, 20(2):153–155, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bays:1990:CIR

[Bay90] Carter Bays. C364. Improving a random number generator: a comparison between two shuffling methods. Journal of Statistical Computation and Simulation, 36(1):57–59, May 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic),

1563-5163. URL http://www.tandfonline.com/doi/abs/10.

1080/00949659008811264. See [LL73, BD76] for the two nearly identical shuffling algorithms. This paper explains why the first does not lengthen the generator period, or much reduce the lattice structure of linear congruential generators, but the second improves both dramatically.

Bennet:1974:SSD

[BB74] B. M. Bennet and J. B. Birch. On the small sample distribution and power of the log likelihood ratio and variance tests for the Poisson. Journal of Statistical Computation and Simulation, 3 (1):33–40, 1974. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bowman:1975:PSG

[BB75] K. O. Bowman and J. J. Beauchamp. Pitfalls with some gamma variate simulation routines. Journal of Statistical Computation and Simulation, 4(2):141–154, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Baruah:1980:GAM

[BB80] Hemanta K. Baruah and G. P. Bhattacharjee. A generalization of Anderson’s modified sequential probability ratio test. Journal of Statistical Computation and Simulation, 11(3–4):197–208, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Brownt:1984:MLR

[BB84] Kenneth G. Brownt and Michael A. Burgess. On maximum likelihood and restricted maximum likelihood approaches to estimation of variance components. Journal of Statistical Computation and Simulation, 19(1):59–77, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Brooks:1989:SSP

[BB89] Camilla A. Brooks and George E. Bonney. A simulation study of properties of a regressive logistic model. Journal of Statistical Computation and Simulation, 32(1–2):31–43, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bhattacharya:1996:REO

[BB96] Bhaskar Bhattacharya and Ayanendranath Basu. Robust estimates of ordered means in normal models. Journal of Statistical Computation and Simulation, 54(1-3):165–175, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Belue:1999:DES

[BB99] Lisa M. Belue and Kenneth W. Bauer, Jr. Designing experiments for single output multilayer perceptrons. Journal of Statistical Computation and Simulation, 64(4):301–325, 1999. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Burdick:1986:CIM

[BBG86] Richard K. Burdick, Nancy J. Birch, and Franklin A. Graybill. Confidence intervals on measures of variability in an unbalanced two-fold nested design with equal subsampling. Journal of Statistical Computation and Simulation, 25(3–4):259–272, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Booth:1995:SAP

[BBHW95] James G. Booth, Ronald W. Butler, S. Huzurbazar, and Andrew T. A. Wood. Saddlepoint approximations for p-values of some tests of covariance matrices. Journal of Statistical Computation and Simulation, 53(3–4):165–180, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Basu:2004:RDA

[BBP04] Ayanendranath Basu, Smarajit Bose, and Sumitra Purkayastha. Robust discriminant analysis using weighted likelihood estimators. Journal of Statistical Computation and Simulation, 74(6): 445–460, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Berg:1982:BRP

[BC82] Menachem Berg and Robert Cléroux. The block replacement problem with minima! repair and random repair costs. Journal of Statistical Computation and Simulation, 15(1):1–7, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Barzanti:1994:CNA

[BC94] L. Barzanti and C. Corradi. C411. A note on the approximate solution of the Volterra renewal equation using upper and lower approximations. Journal of Statistical Computation and Simulation, 49(3–4):225–228, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Baragona:2008:DTA

[BC08] Roberto Baragona and Domenico Cucina. Double threshold autoregressive conditionally heteroscedastic model building by genetic algorithms. Journal of Statistical Computation and Simulation, 78(6):541–558, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Barassi:2008:CBT

[BCH08] Marco R. Barassi, Guglielmo Maria Caporale, and Stephen G. Hall. A comparison between tests for changes in the adjustment coefficients in cointegrated systems. Journal of Statistical Computation and Simulation, 78(1):1–17, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bau:1993:UHS

[BCL93] J. J. Bau, J. Hubert Chen, and Kin Lam. On using Hartley’s statistic to test the hypothesis of not much difference among normal variances. Journal of Statistical Computation and Simulation, 45(1–2):41–60, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bolfarine:2002:DTU

[BCP02] Heleno Bolfarine, Celso R. B. Cabral, and Gilberto A. Paula. Distance tests under nonregular conditions: Applications to the comparative calibration model. Journal of Statistical Computation and Simulation, 72(2):125–140, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Barroso:1998:CLR

[BCV98] Lùcia P. Barroso, Gauss M. Cordeiro, and Kalus L. P. Vasconcellos. Corrected likelihood ratio tests for von Mises regression models. Journal of Statistical Computation and Simulation, 61 (3):237–258, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bau:1993:PPS

[BCX93] Jinn-Jonp Bau, Hubert J. Chen, and Momiao Xiong. Percentage points of the Studentized range test for dispersion of normal means. Journal of Statistical Computation and Simulation, 44 (3–4):149–163, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bays:1976:IPR

[BD76] Carter Bays and S. D. Durham. Improving a poor random number generator. ACM Transactions on Mathematical Software, 2 (1):59–64, March 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://doi.acm.org/10. 1145/355666.355670. See also [LL73] for a slightly different, but inferior, shuffling algorithm, and [Bay90] for a comparison, both mathematical, and graphical, of the two algorithms. Reference [3] for IBM Report GC20-8011-0 is incorrectly given year 1969; the correct year is 1959.

Bhargava:1982:EPO

[BD82] Anil K. Bhargava and Damon Disch. Exact probabilities of obtaining estimated non-positive definite between-group covariance matrices. Journal of Statistical Computation and Simulation, 15(1):27–32, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bilgen:1984:CNC

[BD84] Semih Bilgen and Z. Seyda Deligönul. C206. A note on the choice of numerical integration scheme in solving the renewal equation. Journal of Statistical Computation and Simulation, 20(1):81–84, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Berridge:1996:FRE

[BD96] Damon M. Berridge and Dirley M. Dos Santos. Fitting a random effects model to ordinal recurrent events using existing software. Journal of Statistical Computation and Simulation, 55(1–2):73– 86, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bjorkestol:2001:MSD

[BD01] Kirsten Björkestöl and Michael B. Dollinger. On the moments of a Stahel Donoho robust multivariate estimator. Journal of Statistical Computation and Simulation, 70(1):89–106, 2001. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Borowiak:2009:SAD

[BD09] Dale Borowiak and Ashish Das. Sensitivity analysis of the tdistribution under truncated normal populations. Journal of Statistical Computation and Simulation, 79(5):723–729, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Breton:2008:TSC

[BDB08] Marc D. Breton, Michael D. Devore, and Donald E. Brown. A tool for systematically comparing the power of tests for normality. Journal of Statistical Computation and Simulation, 78 (7):623–638, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Biscay:1997:CVC

[BDFR97] Rolando Biscay, Eloísa Díaz-Francés, and Luis M. Rodríguez. Cross-validation of covariance structures using the Frobenius matrix distance as a discrepancy function. Journal of Statistical Computation and Simulation, 58(3):195–215, 1997. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Beirlant:2004:NEE

[BDG04] Jan Beirlant, Tertius De Wet, and Yuri Goegebeur. Nonparametric estimation of extreme conditional quantiles. Journal of Statistical Computation and Simulation, 74(8):567–580, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Burdick:1986:CIA

[BE86] Richard K. Burdick and John Eickman. Confidence intervals on the among group variance component in the unbalanced one-fold nested design. Journal of Statistical Computation and Simulation, 26(3–4):205–219, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bergeron:1994:RES

[BE94] Lisette R. Bergeron and Luis A. Escobar. Relative efficiencies of some LR-estimators with Type II censored data and small samples. Journal of Statistical Computation and Simulation, 49 (3–4):151–160, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Beaulieu:1985:CCC

[Bea85] Norman C. Beaulieu. C219. Comment on “Calculating binomial probabilities when the trial probabilities are unequal”. Journal of Statistical Computation and Simulation, 20(4):327–328, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [TT82].

Behboodian:1972:IMM

[Beh72] Javad Behboodian. Information matrix for a mixture of two normal distributions. Journal of Statistical Computation and Simulation, 1(4):295–314, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Behboodian:1973:IMM

[Beh73] Javad Behboodian. Information matrix for a mixture of two exponential distributions. Journal of Statistical Computation and Simulation, 2(1):1–16, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bello:1993:PRT

[Bel93a] A. L. Bello. On the performance of rank transform discriminant method in error-rate estimation. Journal of Statistical Computation and Simulation, 48(3–4):153–165, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bello:1993:SSI

[Bel93b] Abdul L. Bello. A simulation study of imputation techniques in linear quadratic and kernel discriminant analyses. Journal of Statistical Computation and Simulation, 48(3–4):167–180, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Benasseni:1994:PAC

[Bén94] Jacques Bénasséni. Partial additive constant. Journal of Statistical Computation and Simulation, 49(3–4):179–193, 1994. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bernstein:1973:BCN

[Ber73] N. Bernstein. Backorder calculation for normal demand. Journal of Statistical Computation and Simulation, 2(3):233–235, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Berk:1980:FBS

[Ber80] Kenneth N. Berk. Forward and backward stepping in variable selection. Journal of Statistical Computation and Simulation, 10 (3–4):177–185, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bechhofer:1983:MCS

[BF83] Robert E. Bechhofer and Thomas Frisardi. A Monte Carlo study of the performance of a closed adaptive sequential procedure for selecting the best Bernoulli population. Journal of Statistical Computation and Simulation, 18(2-3):179–213, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Babakus:1987:COD

[BF87] Emin Babakus and Carl E. Ferguson. C287. Ordinal data and LISREL: a note on significance level. Journal of Statistical Computation and Simulation, 28(2):179–182, 1987. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bhandary:2008:CIE

[BF08] Madhusudan Bhandary and Koji Fujiwara. Confidence interval estimation for a linear contrast in intraclass correlation coefficients under unequal family sizes for several populations. Journal of Statistical Computation and Simulation, 78(7):609–621, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ballard:1978:RPE

[BG78] Robert John Ballard and Dennis C. Gilliland. On the risk performance of extended sequence compound rules for classification between n(-1, 1) and n(1, 1). Journal of Statistical Computation and Simulation, 6(3–4):265–280, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bain:1993:CEE

[BG93] Lee J. Bain and Gaoxiong Gan. Comparison of expectations of the extreme of sums and the sum of extremes from gamma distributions. Journal of Statistical Computation and Simulation, 47(3–4):219–225, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balasooriya:1994:TUO

[BG94] Uditha Balasooriya and Veeresh Gadag. Tests for upper outliers in the two-parameter exponential distribution. Journal of Statistical Computation and Simulation, 50(3–4):249–259, 1994. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Biernacki:1999:CMM

[BG99] Christophe Biernacki and Gérard Govaert. Choosing models in model-based clustering and discriminant analysis. Journal of Statistical Computation and Simulation, 64(1):49–71, 1999. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Berger:2001:RBD

[BG01a] Michael Berger and Sneh Gulati. Record-breaking data: a parametric comparison of the inverse-sampling and the randomsampling schemes. Journal of Statistical Computation and Simulation, 69(3):225–238, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bisaglia:2001:PIF

[BG01b] Luisa Bisaglia and Matteo Grigoletto. Prediction intervals for Farima processes by bootstrap methods. Journal of Statistical Computation and Simulation, 68(2):185–201, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Brazauskas:2007:ECP

[BG07] Vytaras Brazauskas and Jugal Ghorai. Estimating the common parameter of normal models with known coefficients of variation: a sensitivity study of asymptotically efficient estimators. Journal of Statistical Computation and Simulation, 77(8):663– 681, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Best:1994:ECC

[BGR94] D. J. Best, D. Graham, and J. C. W. Rayner. Estimating correlation from categorized bivariate normal data. Journal of Statistical Computation and Simulation, 49(3–4):141–149, 1994. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Barreda:2007:SEM

[BGS07] Laure Barreda, Ali Gannoun, and Jérôme Saracco. Some extensions of multivariate sliced inverse regression. Journal of Statistical Computation and Simulation, 77(1):1–17, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Beauchamp:1973:SIS

[BH73] John J. Beauchamp and David G. Hoel. Some investigations and simulation studies of canonical analysis. Journal of Statistical Computation and Simulation, 2(3):197–209, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bohning:1982:NTE

[BH82] Dankmar Böhning and Karl-Heinz Hoffmann. Numerical techniques for estimating probabilities. Journal of Statistical Computation and Simulation, 14(3–4):283–293, 1982. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Berry:1985:RPD

[BH85] Keith Berry and Cliff J. Huang. Random pseudo-disturbance generators in a stochastic simulation of an econometric model. Journal of Statistical Computation and Simulation, 22(3–4):285– 302, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Blair:1995:IBP

[BH95] Clifford R. Blair and Yosef Hochberg. Improved Bonferroni procedures for testing overall and pairwise homogeneity hypotheses. Journal of Statistical Computation and Simulation, 51(2-4):281– 289, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ben-Haim:1996:CMR

[BH96] Yakov Ben-Haim. C444. Must reliability be probabilistic? Journal of Statistical Computation and Simulation, 55(3):263–265, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See reply [Goo96b].

Balakrishnan:2007:ITI

[BH07] N. Balakrishnan and Ahmed Hossain. Inference for the Type II generalized logistic distribution under progressive Type II censoring. Journal of Statistical Computation and Simulation, 77(12):1013–1031, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bhattacharya:2001:TES

[Bha01] Bhaskar Bhattacharya. Testing equality of scale parameters against restricted alternatives for m ≥ 3 gamma distributions with unknown common shape parameter. Journal of Statistical Computation and Simulation, 69(4):353–368, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bretz:2001:CPP

[BHG01] Frank Bretz, Anthony J. Hayter, and Alan Genz. Critical point and power calculations for the Studentized range test for generally correlated means. Journal of Statistical Computation and Simulation, 71(2):85–97, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Billah:2005:EIC

[BHK05] Baki Billah, Rob J. Hyndman, and Anne B. Koehler. Empirical information criteria for time series forecasting model selection. Journal of Statistical Computation and Simulation, 75(10):831– 840, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bartels:1978:CIC

[BHLH78] Richard H. Bartels, Susan D. Horn, Albert M. Liebetrau, and William L. Harris. A computational investigation of Conover’s Kolmogorov–Smirnov test for discrete distributions. Journal of Statistical Computation and Simulation, 7(2):151–166, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bhoj:1973:DSU

[Bho73a] Dinesh S. Bhoj. On the distribution of a statistic used for testing a multinormal mean vector with incomplete data. Journal of Statistical Computation and Simulation, 2(4):309–316, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bhoj:1973:PPS

[Bho73b] Dinesh S. Bhoj. Percentage points of the statistics for testing hypotheses on mean vectors of multivariate normal distributions with missing observations. Journal of Statistical Computation and Simulation, 2(3):211–224, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bhoj:1984:DMC

[Bho84] Dinesh S. Bhoj. On difference of means of correlated variates with incomplete data on both responses. Journal of Statistical Computation and Simulation, 19(4):275–289, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Boulouird:2008:BCI

[BHZ08] M. Boulouird, M. M. Hassani, and A. Zeroual. Blind channel identification using higher-order statistics. Journal of Statistical Computation and Simulation, 78(4):325–338, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bickel:2003:REE

[Bic03] David R. Bickel. Robust and efficient estimation of the mode of continuous data: the mode as a viable measure of central tendency. Journal of Statistical Computation and Simulation, 73 (12):899–912, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bhattacharyya:1973:RRT

[BJ73] G. K. Bhattacharyya and Richard Johnson. On robustness of rank tests in life testing. Journal of Statistical Computation and Simulation, 2(2):159–169, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Beckman:1978:MLE

[Bj78] R. J. Beckman and G. L. Tiet jen. Maximum likelihood estimation for the beta distribution. Journal of Statistical Computation and Simulation, 7(3–4):253–258, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bryson:1982:CSM

[BJ82] Maurice C. Bryson and Mark E. Johnson. Constructing and simulating multivariate distributions using Khintchine’s theorem. Journal of Statistical Computation and Simulation, 16(2): 129–137, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Barham:1999:RCI

[BJ99] Abdelrahim M. Barham and S. Jeyaratnam. Robust confidence interval for the variance. Journal of Statistical Computation and Simulation, 62(3):189–205, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bai:1992:EPS

[BJM92] Jun Bai, Anthony J. Jakeman, and Michael McAleer. Estimating the percentiles of some misspecified non-nested distributions. Journal of Statistical Computation and Simulation, 42(3–4):151– 159, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Beauchamp:1982:RTP

[BK82] J. J. Beauchamp and V. E. Kane. Robustness of three power transformation estimation procedures. Journal of Statistical Computation and Simulation, 15(4):251–272, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Beauchamp:1984:APS

[BK84] J. J. Beauchamp and V. E. Kane. Application of the powershift transformation. Journal of Statistical Computation and Simulation, 19(1):35–58, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Buning:1996:RES

[BK96] Herbert Büning and Wolfgang Kössler. Robustness and efficiency of some tests for ordered alternatives in the C-sample location problem. Journal of Statistical Computation and Simulation, 55(4):337–352, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Branu:1998:BPP

[BK98] W. J. Branu and R. J. Kulperger. A bootstrap for point processes. Journal of Statistical Computation and Simulation, 60 (2):129–155, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bajorunaite:2008:CFP

[BK08] Ruta Bajorunaite and John P. Klein. Comparison of failure probabilities in the presence of competing risks. Journal of Statistical Computation and Simulation, 78(10):951–966, 2008. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:2004:IEV

[BKLW04] N. Balakrishnan, N. Kannan, C. T. Lin, and S. J. S. Wu. Inference for the extreme value distribution under progressive Type-II censoring. Journal of Statistical Computation and Simulation, 74(1):25–45, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bernstein:1978:SSP

[BL78] D. Bernstein and S. W. Lagakos. Sample size and power determination for stratified clinical trials. Journal of Statistical Computation and Simulation, 8(1):65–73, 1978. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:2002:ELI

[BL02] N. Balakrishnan and Chien-Tai Lin. Exact linear inference and prediction for exponential distributions based on general progressively Type-II censored samples. Journal of Statistical Computation and Simulation, 72(8):677–686, 2002. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:2003:DTE

[BL03] N. Balakrishnan and Chien-Tai Lin. On the distribution of a test for exponentiality based on progressively Type-II right censored spacings. Journal of Statistical Computation and Simulation, 73 (4):277–283, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:2005:EIP

[BL05] N. Balakrishnan and Chien-Tai Lin. Exact inference and prediction for k-sample exponential case under Type-II censoring. Journal of Statistical Computation and Simulation, 75(5):315– 331, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balemi:2009:CGG

[BL09] Andrew Balemi and Alan Lee. Comparison of GEE1 and GEE2 estimation applied to clustered logistic regression. Journal of Statistical Computation and Simulation, 79(4):361–378, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:2004:EIP

[BLC04] N. Balakrishnan, Chien-Tai Lin, and Ping-Shing Chan. Exact inference and prediction for K-sample two-parameter exponential case under general Type-II censoring. Journal of Statistical Computation and Simulation, 74(12):867–878, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bian:2009:EFA

[BLP09] J. W. Bian, H. W. Li, and H. M. Peng. An efficient and fast algorithm for estimating the frequencies of superimposed exponential signals in zero-mean multiplicative and additive noise. Journal of Statistical Computation and Simulation, 79(12):1407– 1423, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bera:1986:TNS

[BM86] Anil K. Bera and Colin R. Mckenzie. Tests for normality with stable alternatives. Journal of Statistical Computation and Simulation, 25(1–2):37–52, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:1990:CSV

[BM90] N. Balakrishnan and C. W. Ma. A comparative study of various tests for the equality of two population variances. Journal of Statistical Computation and Simulation, 35(1–2):41–89, 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bromaghin:1992:SES

[BM92] Jeff Bromaghin and Lyman McDonald. Systematic encounter sampling: a simulation study. Journal of Statistical Computation and Simulation, 40(3–4):177–191, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bera:1995:TNU

[BN95] Anil K. Bera and Pin T. Ng. Tests for normality using estimated score function. Journal of Statistical Computation and Simulation, 52(3):273–287, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bhattacharya:1996:BIM

[BN96] Bhaskar Bhattacharya and Balgobin Nandram. Bayesian inference for multinomial populations under stochastic ordering. Journal of Statistical Computation and Simulation, 54(1-3):145– 163, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:2001:IEC

[BN01a] N. Balakrishnan and Hon Keung Tony Ng. Improved estimation of the correlation coefficient in a bivariate exponential distribution. Journal of Statistical Computation and Simulation, 68(2): 173–184, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:2001:ECC

[BN01b] N. Balakrishnan and Hon Keung Tony Ng. On estimation of the correlation coefficient in Moran–Downton multivariate exponential distribution. Journal of Statistical Computation and Simulation, 71(1):41–58, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bogdan:1995:DDV

[Bog95] Małgorzata Bogdan. Data driven versions of Pearson’s chisquare test for uniformity. Journal of Statistical Computation and Simulation, 52(3):217–237, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bogdan:2001:DDV

[Bog01] Małgorzata Bogdan. Data driven versions of Neyman’s test for uniformity based on Bayesian rule. Journal of Statistical Computation and Simulation, 68(3):203–222, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bondesson:1981:CNS

[Bon81] Lennart Bondesson. C112. A normal sample with given sample mean and variance. Journal of Statistical Computation and Simulation, 14(1):51–52, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bonett:2005:CIR

[Bon05] Douglas G. Bonett. Confidence interval for residual mean absolute deviation in regression models. Journal of Statistical Computation and Simulation, 75(8):673–678, 2005. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bonett:2006:CIR

[Bon06] Douglas G. Bonett. Confidence interval for a ratio of variances in bivariate nonnormal distributions. Journal of Statistical Computation and Simulation, 76(7):637–644, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Borowsky:1975:BR

[Bor75] Michael S. Borowsky. Book reviews. Journal of Statistical Computation and Simulation, 4(2):165–166, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bowman:1985:CSS

[Bow85] Adrian W. Bowman. A comparative study of some kernel-based nonparametric density estimators. Journal of Statistical Computation and Simulation, 21(3–4):313–327, 1985. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bowman:1992:DBT

[Bow92] A. W. Bowman. Density based tests for goodness-of-fit. Journal of Statistical Computation and Simulation, 40(1–2):1–13, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bowman:2001:ECM

[Bow01] Dale Bowman. Effects of correlation in modeling clustered binary data. Journal of Statistical Computation and Simulation, 69(4):369–389, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bhapkar:1978:MCS

[BP78] V. P. Bhapkar and Kenneth W. Patterson. A Monte Carlo study of some multivariate nonparametric statistics for profile analysis of several samples. Journal of Statistical Computation and Simulation, 6(3–4):223–237, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:1986:BLU

[BP86] N. Balakrishnan and S. Puthenpura. Best linear unbiased estimators of location and scale parameters of the half logistic distribution. Journal of Statistical Computation and Simulation, 25(3–4):193–204, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Brown:1992:NAV

[BP92] Timothy C. Brown and Diana L. Pallant. A note on antithetic variates from discrete distributions. Journal of Statistical Computation and Simulation, 40(3–4):193–199, 1992. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Begg:2001:PSM

[BP01] Melissa Dowd Begg and Andrea B. Paykin. Performance of and software for a modified Mantel–Haenszel statistic for correlated data. Journal of Statistical Computation and Simulation, 70 (2):175–195, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bansal:2000:UMP

[BPP00] Vikrant Bansal, John D. Perkins, and Efstratios N. Pistikopoulos. Using mathematical programming to compute singular multivariate normal probabilities. Journal of Statistical Computation and Simulation, 67(3):219–253, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Blass:1981:CIC

[BR81] J. Blass and V. K. Rohatgi. C113. Independence of certain events. Journal of Statistical Computation and Simulation, 14 (1):52–55, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Blass:1984:CCM

[BR84] Josef Blass and Vijay K. Rohatgi. C211. On the computation of the moments of certain statistics by means of Newton’s formula for polynomials. Journal of Statistical Computation and Simulation, 20(2):164–167, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Basu:1992:NPS

[BR92] Sabyasachi Basu and Gregory C. Reinsel. A note on properties of spatial Yule–Walker estimators. Journal of Statistical Computation and Simulation, 41(3–4):243–255, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Braun:2000:ACB

[BR00] W. John Braun and Valentin Rousson. An autocorrelation criterion for bandwidth selection in nonparametric regression. Journal of Statistical Computation and Simulation, 68(1):89–101, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Buning:2003:ABT

[BR03] Herbert Büning and Michael Rietz. Adaptive bootstrap tests and their competitors in the C-sample location problem. Journal of Statistical Computation and Simulation, 73(5):361–375, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Best:2006:ITB

[BR06] D. J. Best and J. C. W. Rayner. Improved testing for the binomial distribution using chi-squared components with datadependent cells. Journal of Statistical Computation and Simulation, 76(1):75–81, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Brannas:1992:FSP

[Brä92] Kurt Brännäs. Finite sample properties of estimators and tests in Poisson regression models. Journal of Statistical Computation and Simulation, 41(3–4):229–241, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bremner:1993:NAS

[Bre93] J. M. Bremner. A new approach to subset selection for normal means. Journal of Statistical Computation and Simulation, 44 (3–4):187–208, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:2008:ELI

[BRF08] N. Balakrishnan, Abbas Rasouli, and N. Sanjari Farsipour. Exact likelihood inference based on an unified hybrid censored sample from the exponential distribution. Journal of Statistical Computation and Simulation, 78(5):475–488, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bell:1982:ESS

[BRL82] William D. Bell, Stuart M. Rothstein, and Wai-Kee Li. Empirical study of some non-parametric tests for dispersion of correlated data. Journal of Statistical Computation and Simulation, 15(1):9–15, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Brock:1974:CP

[Bro74] Paul Brock. Call for papers. Journal of Statistical Computation and Simulation, 3(1):95, 1974. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Best:2007:GFZ

[BRT07] D. J. Best, J. C. W. Rayner, and O. Thas. Goodness of fit for the zero-truncated Poisson distribution. Journal of Statistical Computation and Simulation, 77(7):585–591, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Brunner:1975:TCV

[Bru75] Edgar Brunner. Tables of critical values of the normal probability paper test. Journal of Statistical Computation and Simulation, 3(4):387–391, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bowman:1972:SSP

[BS72] K. O. Bowman and L. R. Shenton. Small sample properties of the maximum likelihood estimator associated with Fisher’s linkage problem. Journal of Statistical Computation and Simulation, 1(1):71–85, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bhattacharya:1978:IET

[BS78] C. G. Bhattacharya and K. R. Shah. Interval estimation of treatment differences in block designs. Journal of Statistical Computation and Simulation, 6(3–4):243–255, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Bhapkar:1979:ECB

[BS79] V. P. Bhapkar and Jeffrey H. Schwartz. Efficient computation of Bhapkar’s V statistic and significance points for its use in small samples. Journal of Statistical Computation and Simulation, 10(1):1–14, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bowman:1992:PEB

[BS92] K. O. Bowman and L. R. Shenton. Parameter estimation for the beta distribution. Journal of Statistical Computation and Simulation, 43(3–4):217–228, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Basu:1994:TBR

[BS94a] Ayanendranath Basu and Sahadeb Sarkar. The trade-off between robustness and efficiency and the effect of model smoothing in minimum disparity inference. Journal of Statistical Computation and Simulation, 50(3–4):173–185, 1994. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Byers:1994:CSA

[BS94b] R. H. Byers, Jr. and L. R. Shenton. C406. Surprising approximations to the half-normal. Journal of Statistical Computation and Simulation, 49(3–4):215–216, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:1995:RRS

[BS95] N. Balakrishnan and R. A. Sandhu. Recurrence relations for single and product moments of order statistics from a generalized half logistic distribution with applications to inference. Journal of Statistical Computation and Simulation, 52(4):385–398, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bhoj:2001:ADW

[BS01a] Dinesh S. Bhoj and Klaus Schiefermayr. Approximations to the distribution of weighted combination of independent probabilities. Journal of Statistical Computation and Simulation, 68(2): 153–159, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Brazauskas:2001:SSP

[BS01b] Vytaras Brazauskas and Robert Serfling. Small sample performance of robust estimators of tail parameters for Pareto and exponential models. Journal of Statistical Computation and Simulation, 70(1):1–19, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bowman:2002:PML

[BS02] K. O. Bowman and L. R. Shenton. Problems with maximum likelihood estimation and the 3 parameter gamma distribution. Journal of Statistical Computation and Simulation, 72(5):391– 401, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bari:2005:BLM

[BS05a] Wasimul Bari and Brajendra C. Sutradhar. On binary longitudinal mixed models in adaptive clinical trials. Journal of Statistical Computation and Simulation, 75(9):687–707, 2005. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bernaards:2005:BFL

[BS05b] Coen A. Bernaards and Klaas Sijtsma. Bias of factor loadings from questionnaire data with imputed scores. Journal of Statistical Computation and Simulation, 75(1):13–23, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bowman:2005:AVS

[BS05c] K. O. Bowman and L. R. Shenton. The asymptotic variance and skewness of maximum likelihood estimators using Maple. Journal of Statistical Computation and Simulation, 75(12):975– 986, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Butler:2008:SAS

[BSBS08] R. W. Butler, R. K. Sutton, J. G. Booth, and P. Ohman Strickland. Simulation-assisted saddlepoint approximation. Journal of Statistical Computation and Simulation, 78(8):731–745, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Breiman:1990:RCB

[BSK90] Leo Breiman, Charles J. Stone, and Charles Kooperberg. Robust confidence bounds for extreme upper quantiles. Journal of Statistical Computation and Simulation, 37(3–4):127–149, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Beg:2002:RTA

[BSS02] A. B. M. Rabiul Alam Beg, Mervyn Joseph Silvapulle, and Paramsothy Silvapulle. Robust tests against smooth transition autoregressive models. Journal of Statistical Computation and Simulation, 72(2):167–178, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Buning:2000:ATS

[BT00] Herbert Büning and Thorsten Thadewald. An adaptive twosample location-scale test of Lepage type for symmetric distributions. Journal of Statistical Computation and Simulation, 65 (1-4):287–310, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bilder:2009:BEA

[BT09] Christopher R. Bilder and Joshua M. Tebbs. Bias, efficiency, and agreement for group-testing regression models. Journal of Statistical Computation and Simulation, 79(1):67–80, 2009. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bucciarelli:1993:GMM

[BTL93] B. Bucciarelli, L. G. Taff, and M. G. Lattanzi. A generalization of the moving mean. Journal of Statistical Computation and Simulation, 48(1–2):29–46, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Burrill:1974:CPS

[Bur74] Donald F. Burrill. Computational precision in statistics: variances and covariances. Journal of Statistical Computation and Simulation, 3(2):105–115, 1974. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Butlet:1999:FDA

[But99] Neil A. Butlet. A frequency domain algorithm for maximum likelihood estimation of Gaussian fields. Journal of Statistical Computation and Simulation, 64(2):151–165, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Biscay:1990:MFL

[BVP90] R. Biscay, P. Valdes, and R. Pascual. Modified Fisher’s linear discriminant function with reduction of dimensionality. Journal of Statistical Computation and Simulation, 36(1):1–8, 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bauer:1993:SRV

[BW93] Kenneth W. Bauer, Jr. and James R. Wilson. Standardized routing variables: a new class of control variates. Journal of Statistical Computation and Simulation, 46(1–2):69–78, 1993. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bennett:2001:REM

[BW01a] Mihoko V. Bennett and Thomas R. Willemain. Resistant estimation of multivariate location using minimum spanning trees. Journal of Statistical Computation and Simulation, 69(1):19– 40, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bernard:2001:RSA

[BW01b] Anthony J. Bernard and Peter S. Wludyka. Robust I-sample analysis of means type randomization tests for variances. Journal of Statistical Computation and Simulation, 69(1):57–88, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Boscardin:2008:MMO

[BZB08] W. John Boscardin, Xiao Zhang, and Thomas R. Belin. Modeling a mixture of ordinal and continuous repeated measures. Journal of Statistical Computation and Simulation, 78(10):873– 886, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chan:1989:EBF

[CA89] K. P. S. Chan and C. G. G. Aitken. Estimation of the Bayes’ factor in a forensic science problem. Journal of Statistical Computation and Simulation, 33(4):249–264, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cadigan:1994:BAM

[Cad94] Neol G. Cadigan. Bias approximation for maximum likelihood estimates. Journal of Statistical Computation and Simulation, 51(1):89–95, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cai:2009:RCA

[Cai09] Haiyan Cai. A random cluster algorithm for some continuous Markov random fields. Journal of Statistical Computation and Simulation, 79(11):1287–1299, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cao:1987:SEB

[Cao87] Xi-Ren Cao. Sensitivity estimates based on one realization of a stochastic system. Journal of Statistical Computation and Simulation, 27(3):211–232, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Carolan:2007:TSB

[Car07] C. Carolan. Two-sided Brownian motion with quadratic drift and its least concave majorant. Journal of Statistical Computation and Simulation, 77(5):379–387, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cornwell:1978:NED

[CAT78] L. W. Cornwell, L. A. Aroian, and V. S. Taneja. Numerical evaluation of the distribution of the product of two normal variables. Journal of Statistical Computation and Simulation, 7(2): 123–131, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Childs:1997:CTM

[CB97] Aaron Childs and N. Balakrishnan. Comparison of tests for multiple outliers when sampling from the logistic distribution. Journal of Statistical Computation and Simulation, 58(1):1–20, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Coombs:2003:CMC

[CB03] Toni Coombs and Richard Burdick. A comparison of methods for constructing confidence intervals on contrasts in a three-factor mixed model. Journal of Statistical Computation and Simulation, 73(8):597–602, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Charway:2007:TMG

[CB07] Harry Charway and A. John Bailer. Testing multiple-group variance equality with randomization procedures. Journal of Statistical Computation and Simulation, 77(9):797–803, 2007. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cawlfield:1997:PSM

[CBPW97] Jeffrey D. Cawlfield, Samuel Boateng, John Piggott, and Ming-Chee Wu. Probabilistic sensitivity measures applied to numerical models of flow and transport. Journal of Statistical Computation and Simulation, 57(1-4):353–364, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Childs:2006:UST

[CBS06] A. Childs, N. Balakrishnan, and M. R. Srinivasan. Unified scheme for testing for outliers in linear models. Journal of Statistical Computation and Simulation, 76(1):21–39, 2006. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cobb:1985:SSI

[CC85] E. Benton Cobb and J. D. Church. Small-sample interval estimation using quantal data. Journal of Statistical Computation and Simulation, 22(3–4):189–202, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cameron:1989:CPS

[CC89] R. F. Cameron and C. F. Clement. C339. The probability of a serious nuclear reactor accident. Journal of Statistical Computation and Simulation, 32(1–2):124–126, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cobb:1990:NES

[CC90a] Benton E. Cobb and J. D. Church. Nonparametric estimators for shift in quantal bioassay. Journal of Statistical Computation and Simulation, 37(3–4):159–170, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cox:1990:ISM

[CC90b] Mitchell A. A. Cox and Trevor F. Cox. Interpreting stress in multidimensional scaling. Journal of Statistical Computation and Simulation, 37(3–4):211–223, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:1997:ILR

[CC97] Gauss M. Cordeiro and Enrico A. Colosimo. Improved likelihood ratio tests for exponential censored data. Journal of Statistical Computation and Simulation, 56(4):303–315, 1997. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cadigan:2001:PRE

[CC01] N. G. Cadigan and J. Chen. Properties of robust m-estimators for Poisson and negative binomial data. Journal of Statistical Computation and Simulation, 70(3):273–288, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Callens:2005:PLB

[CC05] Marc Callens and Christophe Croux. Performance of likelihoodbased estimation methods for multilevel binary regression models. Journal of Statistical Computation and Simulation, 75(12): 1003–1017, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2009:SIA

[CC09] Dongming Chen and Zhenmin Chen. Statistical inference about the location parameter of the three-parameter Weibull distribution. Journal of Statistical Computation and Simulation, 79 (3):215–225, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Celeux:1996:SVE

[CCD96] Gilles Celeux, Didier Chauveau, and Jean Diebolt. Stochastic versions of the em algorithm: an experimental study in the mixture case. Journal of Statistical Computation and Simulation, 55(4):287–314, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Castillo:2008:DLS

[CCHM08] Enrique Castillo, Carmen Castillo, Ali S. Hadi, and Roberto Mìnguez. Duality and local sensitivity analysis in least squares, minimax, and least absolute values regressions. Journal of Statistical Computation and Simulation, 78(10):887–909, 2008. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Colosimo:2007:SSC

[CCMV07] E. A. Colosimo, F. R. B. Cruz, J. L. O. Miranda, and T. Van Woensel. Sample size calculation for method validation using linear regression. Journal of Statistical Computation and Simulation, 77(6):505–516, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cepeda-Cuervo:2009:BMM

[CCNA09] Edilberto Cepeda-Cuervo and Vicente Núñez-Antón. Bayesian modelling of the mean and covariance matrix in normal nonlinear models. Journal of Statistical Computation and Simulation, 79(6):837–853, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:1995:BCO

[CCNAF95] Gauss M. Cordeiro, Francisco Cribari-Neto, Elisete C. Q. Aubin, and Silvia L. P. Ferrari. Bartlett corrections for one-parameter exponential family models. Journal of Statistical Computation and Simulation, 53(3–4):211–231, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Clement:1981:TMV

[CCSG81] Bernard Clement, Sukharanyan Chakraborty, Bimal K. Sinha, and Narayan C. Giri. Tests for the mean vector under intraclass covariance structure. Journal of Statistical Computation and Simulation, 12(3–4):237–245, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chan:2004:MRM

[CCY04] Raymond H. Chan, Yong Chen, and K. M. Yeung. A memory reduction method in pricing American options. Journal of Statistical Computation and Simulation, 74(7):501–511, 2004. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Calvin:1992:ARM

[CD92a] James A. Calvin and Richard L. Dykstra. An algorithm for restricted maximum likelihood estimation in balanced multivariate variance components models. Journal of Statistical Computation and Simulation, 40(3–4):233–246, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:1992:IRS

[CD92b] Zehua Chen and Kim-Anh Do. Importance resampling for the smoothed bootstrap. Journal of Statistical Computation and Simulation, 42(1–2):107–124, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:1997:BCM

[CDDCN97] Gauss M. Cordeiro, Enivaldo C. Da Rocha, Jacira Guiro C. Da Rocha, and Francisco Cribari-Neto. Bias-corrected maximum likelihood estimation for the beta distribution. Journal of Statistical Computation and Simulation, 58(1):21–35, 1997. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cabrera:2005:TEL

[CDF05] Javier Cabrera, Vipul Devas, and Luisa T. Fernholz. Target estimation for the logistic regression model. Journal of Statistical Computation and Simulation, 75(2):121–140, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chung:2002:SHS

[CDJ02] Younshik Chung, Dipak Dey, and Junghoon Jang. Semiparametric hierarchical selection models for Bayesian meta analysis. Journal of Statistical Computation and Simulation, 72(10):825– 839, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Duan:1993:MSM

[cDJgS93] Jin chuan Duan, Jean-guy, and Simonato. Multiplicity of solutions in maximum likelihood factor analysis. Journal of Statistical Computation and Simulation, 47(1–2):37–47, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Celmins:1981:LSM

[Cel81] Aivars Celmins. Least square model fitting with transformation of variable. Journal of Statistical Computation and Simulation, 14(1):17–39, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Celebi:2009:NAG

[Çel09] Mehmed Çelebi. A new approach for the genetic algorithm. Journal of Statistical Computation and Simulation, 79(3):275– 297, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Colosimo:2002:ECB

[CFOS02] Enrico Colosimo, Flávio Ferreira, Maristela Oliveira, and Cleide Sousa. Empirical comparisons between Kaplan–Meier and Nelson–Aalen survival function estimators. Journal of Statistical Computation and Simulation, 72(4):299–308, 2002. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Carroll:1977:PRP

[CG77] Raymond J. Carroll and Shanti S. Gupta. On the probabilities of rankings of k populations with applications. Journal of Statistical Computation and Simulation, 5(2):145–157, 1977. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Crockett:1984:STS

[CG84] N. G. Crockett and V. Gebski. A simplified two sample sequential t-test. Journal of Statistical Computation and Simulation, 20(3):217–234, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Capps:1991:MCS

[CG91] Oral Capps, Jr. and William D. Grubbs. A Monte Carlo study of collinearity in linear simultaneous equation models. Journal of Statistical Computation and Simulation, 39(3):139–162, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Celeux:1993:CMC

[CG93] Gilles Celeux and Gérard Govaert. Comparison of the mixture and the classification maximum likelihood in cluster analysis.

Journal of Statistical Computation and Simulation, 47(3–4):127– 146, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chang:1994:ODO

[CG94] T. C. Chang and F. F. Gan. Optimal designs of one-sided EWMA charts for monitoring a process variance. Journal of Statistical Computation and Simulation, 49(1–2):33–48, 1994. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Campolongo:1997:PSD

[CG97] Francesca Campolongo and Albert Gabric. The parametric sensitivity of dimethylsulfide flux in the Southern Ocean. Journal of Statistical Computation and Simulation, 57(1-4):337–352, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chang:2006:MLM

[CG06] Tee-Chin Chang and Fah-Fatt Gan. Monitoring linearity of measurement gauges. Journal of Statistical Computation and Simulation, 76(10):889–911, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Candelon:2003:FSP

[CGA03] Bertrand Candelon and Luis Gil-Alana. On finite sample properties of the tests of Robinson (1994) for fractional integration. Journal of Statistical Computation and Simulation, 73(6):445– 464, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Caporale:2009:MLM

[CGA09] Guglielmo Maria Caporale and Luis A. Gil-Alana. A multivariate long-memory model with structural breaks. Journal of Statistical Computation and Simulation, 79(8):1001–1013, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2004:DSN

[CGN04] John T. Chen, Arjun K. Gupta, and Truc T. Nguyen. The density of the skew normal sample mean and its applications.

Journal of Statistical Computation and Simulation, 74(7):487– 494, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chase:1976:DST

[CH76] Gerald R. Chase and John E. Hewett. Double sample tests — a distribution free procedure. Journal of Statistical Computation and Simulation, 4(4):247–257, 1976. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cox:1982:TMM

[CH82] C. Philip Cox and Chien-Pai Han. Testing multivariate means when the covariance matrix has intraclass correlation structure. Journal of Statistical Computation and Simulation, 16(2):97– 107, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Copenhaver:1988:CDM

[CH88] Margaret Diponzio Copenhaver and Burt Holland. Computation of the distribution of the maximum Studentized range statistic with application to multiple significance testing of simple effects. Journal of Statistical Computation and Simulation, 30(1):1–15, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Conerly:1990:CBR

[CH90] Michael D. Conerly and J. Michael Hardin. A comparison of biased regression estimators using a Pitman nearness criterion. Journal of Statistical Computation and Simulation, 34(2-3):75– 95, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Callanan:1991:SNA

[CH91] Terrance P. Callanan and David A. Harville. Some new algorithms for computing restricted maximum likelihood estimates of variance components. Journal of Statistical Computation and Simulation, 38(1-4):239–259, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cheng:1997:SCS

[CH97] Russell C. H. Cheng and Wayne Holloand. Sensitivity of computer simulation experiments to errors in input data. Journal of Statistical Computation and Simulation, 57(1-4):219–241, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cheng:1998:TPM

[CH98] Russell C. H. Cheng and Wayne Holland. Two-point methods for assessing variability in simulation output. Journal of Statistical Computation and Simulation, 60(3):183–205, 1998. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chiou:1999:CIE

[CH99] Paul Chiou and Chien-Pai Han. Conditional interval estimation of the ratio of variance components following rejection of a pretest. Journal of Statistical Computation and Simulation, 63(2): 105–119, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Croux:2002:NFS

[CH02] Christophe Croux and Gentiane Haesbroeck. A note on finitesample efficiencies of estimators for the minimum volume ellipsoid. Journal of Statistical Computation and Simulation, 72 (7):585–596, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Campbell:2006:IPI

[CH06a] Randall C. Campbell and R. Carter Hill. Imposing parameter inequality restrictions using the principle of maximum entropy. Journal of Statistical Computation and Simulation, 76(11):985– 1000, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Coad:2006:SPC

[CH06b] D. S. Coad and A. A. Halimeh. Sequential procedures for comparing several normal means. Journal of Statistical Computation and Simulation, 76(6):519–537, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chalmers:1975:GCM

[Cha75a] C. P. Chalmers. Generation of correlation matrices with a given eigen-structure. Journal of Statistical Computation and Simulation, 4(2):133–139, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chand:1975:ETD

[Cha75b] Nanak Chand. Effect of truncating the distribution on a sequential test. Journal of Statistical Computation and Simulation, 3 (4):379–385, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chaubey:1979:GAB

[Cha79] Yogendra P. Chaubey. Gaussian approximation to a bivariate quadratic form distribution. Journal of Statistical Computation and Simulation, 9(4):253–260, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chattamvelli:1994:ADT

[Cha94] R. Chattamvelli. Another derivation of two algorithms for the noncentral χ^{2} and f distributions. Journal of Statistical Computation and Simulation, 49(3–4):207–214, 1994. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cao:2003:NML

[CHáS03] Ricardo Cao, Jeffrey Hart, and ángeles Saavedra. Nonparametric maximum likelihood estimators for AR and MA time series. Journal of Statistical Computation and Simulation, 73(5):347– 360, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:1978:AIS

[Che78] Hubert J. Chen. On approximations to the inverse Student’st distribution function. Journal of Statistical Computation and Simulation, 7(3–4):167–180, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:1980:TSF

[Che80] William Chen. On the tests of separate families of hypotheses with small sample size. Journal of Statistical Computation and Simulation, 11(3–4):183–187, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:1982:SPP

[Che82] William Chen. Simulation on probability points for testing of lognormal or Weibull distribution with a small sample. Journal of Statistical Computation and Simulation, 15(2-3):201–210, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cheng:1985:GMN

[Che85] R. C. H. Cheng. Generation of multivariate normal samples with given sample mean and covariance matrix. Journal of Statistical Computation and Simulation, 21(1):39–49, 1985. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chengalur:1987:BR

[Che87] Indushobha N. Chengalur. Book review. Journal of Statistical Computation and Simulation, 28(3):273–274, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:1994:CNS

[Che94] Ling Chen. C422. A note on Sutton’s composite test for the mean of asymmetrical distributions. Journal of Statistical Computation and Simulation, 49(3–4):244–249, 1994. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:1997:ECI

[Che97] Zhenmin Chen. Exact confidence interval for the shape parameter of a log-logistic distribution. Journal of Statistical Computation and Simulation, 56(3):193–211, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2001:SCS

[Che01] Z. Y. Chen. On the statistical computation of the sample multiple correlation coefficient. Journal of Statistical Computation and Simulation, 70(4):299–324, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2003:CNT

[Che03] Ling Chen. C470. Normality tests based on weighted order statistics. Journal of Statistical Computation and Simulation, 73(8):603–606, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2005:SCN

[Che05] Zen-Yi Chen. The S-system computation of non-central gamma distribution. Journal of Statistical Computation and Simulation, 75(10):813–829, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2006:CEM

[Che06] Yi-Hau Chen. Computationally efficient Monte Carlo EM algorithms for generalized linear mixed models. Journal of Statistical Computation and Simulation, 76(9):817–828, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chang:1991:IER

[CHH91] Kuang-Chao Chang, Chien-Pan Han, and D. L. Hawkins. Improved estimators of regression coefficients in measurement error models. Journal of Statistical Computation and Simulation, 39 (1–2):21–35, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chi:1979:QFU

[Chi79] P. Y. Chi. Quadratic forms used as non-parametric two-sample test criteria. Journal of Statistical Computation and Simulation, 8(3–4):267–280, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chiu:2000:TSO

[Chi00] Shean-Tsong Chiu. Test of significance in order selection. Journal of Statistical Computation and Simulation, 65(1-4):23–42, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chiu:2007:CKC

[Chi07] Sung Nok Chiu. Correction to Koen’s critical values in testing spatial randomness. Journal of Statistical Computation and Simulation, 77(11):1001–1004, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chiu:2008:RPD

[Chi08] Sung Nok Chiu. Reduced Palm distribution generating function for a spatial point process. Journal of Statistical Computation and Simulation, 78(12):1241–1250, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Castillo:2009:DNL

[CHM09] E. Castillo, A. S. Hadi, and R. Mínguez. Diagnostics for nonlinear regression. Journal of Statistical Computation and Simulation, 79(9):1109–1128, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Choudhury:1998:CIG

[Cho98] Askar H. Choudhury. C450. Improving a GLS approach for estimation of autoregressive moving-average models. Journal of Statistical Computation and Simulation, 60(2):173–176, 1998. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Choi:2008:IGF

[Cho08] Byungjin Choi. Improvement of goodness-of-fit test for normal distribution based on entropy and power comparison. Journal of Statistical Computation and Simulation, 78(9):781–788, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Carapeto:2003:MCS

[CHR03] M. Carapeto, W. Holt, and P. N. Refenes. On model complexity and selection. Journal of Statistical Computation and Simulation, 73(1):45–57, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Church:1980:GSE

[Chu80] Curtis K. Church. On grouping sets of exponential populations. Journal of Statistical Computation and Simulation, 10(3–4):195– 204, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chunfeng:2001:BCC

[Chu01] Huang Chunfeng. Boundary corrected cubic smoothing splines. Journal of Statistical Computation and Simulation, 70(2):107– 121, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Christie:1981:URL

[CI81] Theodore Christie, Jr. and Boris Iglewicz. Utilization of run lengths in bonus play-the-winner rules for clinical trials. Journal of Statistical Computation and Simulation, 12(2):107–119, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cicchitelli:1989:ROS

[Cic89] G. Cicchitelli. On the robustness of the one-sample t test. Journal of Statistical Computation and Simulation, 32(4):249–258, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Custer:1973:RPL

[CJ73] S. W. Custer and R. R. Jakhotia. Robustness and power of linear regression when variables are Poisson-distributed. Journal of Statistical Computation and Simulation, 2(1):39–43, 1973. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chattamvelli:1995:RRN

[CJ95] R. Chattamvelli and M. C. Jones. Recurrence relations for noncentral density, distribution functions and inverse moments. Journal of Statistical Computation and Simulation, 52(3):289– 299, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Christopher:2000:ABS

[CJ00] Assaid A. Christopher and Birch B. Jeffrey. Automatic bandwidth selection in robust nonparametric regression. Journal of Statistical Computation and Simulation, 66(4):259–272, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:1996:ADW

[CK96] Hanfeng Chen and Samuel Kotz. An asymptotic distribution of Wright’s process capability index sensitive to skewness. Journal of Statistical Computation and Simulation, 55(1–2):147–158, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2001:FDJ

[CK01] Hanfeng Chen and Grazyna Kamburowska. Fitting data to the Johnson system. Journal of Statistical Computation and Simulation, 70(1):21–32, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Caroni:2004:DFW

[CK04] C. Caroni and A. Kimber. Detection of frailty in Weibull lifetime data using outlier tests. Journal of Statistical Computation and Simulation, 74(1):15–23, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Choudhari:2001:LRT

[CKM01] Pankaj Choudhari, Debasis Kundu, and Neeraj Misra. Likelihood ratio test for simultaneous testing of the mean and the variance of a normal distribution. Journal of Statistical Computation and Simulation, 71(4):313–333, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Constantine:1989:BEG

[CKT89] Kenneth Constantine, Marvin Karson, and Siu-Keung Tse. Bootstrapping estimators of P (Y < X) in the gamma case. Journal of Statistical Computation and Simulation, 33(4):217– 231, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Clemm:1973:TER

[CKW73] D. S. Clemm, P. R. Krishnaiah, and V. B. Waikar. Tables for the extreme roots of the Wishart matrix. Journal of Statistical Computation and Simulation, 2(1):65–92, 1973. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chatillon:1991:PSC

[CL91] Guy Châtillon and Louis Laurencelle. Percentiles from a set of classes or from the raw data? An argument based on the trinomial distribution. Journal of Statistical Computation and Simulation, 38(1-4):127–138, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chaiyapong:1997:AIR

[CL97] Y. Chaiyapong and C. J. Lloyd. Accurate inference for recapture experiments with behavioural response. Journal of Statistical Computation and Simulation, 56(2):97–115, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cheung:1998:PAD

[CL98] Yin-Wong Cheung and Kon S. Lai. Power of the augmented Dickey–Fuller test with information-based lag selection. Journal of Statistical Computation and Simulation, 60(1):57–65, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Clark:1986:ECH

[Cla86] D. I. Clark. Efficient calculation of Hodges–Lehmann estimators of location. Journal of Statistical Computation and Simulation, 25(1–2):127–136, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Clayton:1996:EBM

[Cla96] Howard R. Clayton. The effectiveness of bootstrap methods in evaluating skewed auditing populations: a simulation study. Journal of Statistical Computation and Simulation, 54(1-3):255– 264, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2004:BET

[CLLN04] Cathy W. S. Chen, Jack C. Lee, Hsiang-Yu Lee, and W. F. Niu. Bayesian estimation for time-series regressions improved with exact likelihoods. Journal of Statistical Computation and Simulation, 74(10):727–740, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chang:1993:IJS

[CLP93] Ching-Hui (Charlie) Chang, Jyh-Jiuan (Grace) Lin, and Nabendu Pal. Improvements over the James–Stein estimator: A risk analysis. Journal of Statistical Computation and Simulation, 48(1–2):117–126, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Crain:1976:MCE

[CM76] I. K. Crain and R. E. Miles. Monte Carlo estimates of the distributions of the random polygons determined by random lines in a plane. Journal of Statistical Computation and Simulation, 4(4):293–325, 1976. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Czado:1998:ASD

[CM98] Claudia Czado and Axel Munk. Assessing the similarity of distributions — finite sample performance of the empirical Mallows distance. Journal of Statistical Computation and Simulation, 60 (4):319–346, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Czado:2001:BMN

[CM01] Claudia Czado and Axel Munk. Bootstrap methods for the nonparametric assessment of population bioequivalence and similarity of distributions. Journal of Statistical Computation and Simulation, 68(3):243–280, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cook:2004:SDA

[CM04] Steven Cook and Neil Manning. Size distortion of asymmetric unit root tests in the presence of level shifts. Journal of Statistical Computation and Simulation, 74(11):811–819, 2004. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Clarke:2005:RIM

[CM05a] Brenton R. Clarke and Peter L. McKinnon. Robust inference and modelling for the single ion channel. Journal of Statistical Computation and Simulation, 75(7):513–529, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cook:2005:UHM

[CM05b] Steven Cook and Neil Manning. Unobserved heterogeneity in Markovian analysis of the size distortion of unit root tests. Journal of Statistical Computation and Simulation, 75(9):709–729, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Clarke:2006:CSC

[CM06] Judith A. Clarke and Sadaf Mirza. A comparison of some common methods for detecting Granger noncausality. Journal of Statistical Computation and Simulation, 76(3):207–231, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chiou:2007:SED

[CM07a] Paul Chiou and Weiwen Miao. Shrinkage estimation for the difference between a control and treatment mean. Journal of Statistical Computation and Simulation, 77(8):651–662, 2007. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Comas:2007:SHP

[CM07b] C. Comas and J. Mateu. On soft and hard particle motions for stochastic marked point processes. Journal of Statistical Computation and Simulation, 77(12):1091–1121, 2007. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chow:1974:EMC

[CMD74] Bryant Chow, James E. Miller, and Peter C. Dickinson. Extensions of a Monte-Carlo comparison of some properties of two rank correlation coefficients in small samples. Journal of Statistical Computation and Simulation, 3(2):189–195, 1974. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cruz-Medina:2004:NES

[CMH04] I. R. Cruz-Medina and T. P. Hettmansperger. Nonparametric estimation in semi-parametric univariate mixture models. Journal of Statistical Computation and Simulation, 74(7):513–524, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chiu:2003:MLE

[CMQ03] S. N. Chiu, I. S. Molchanov, and M. P. Quine. Maximum likelihood estimation for germination-growth processes with application to neurotransmitters data. Journal of Statistical Computation and Simulation, 73(10):725–732, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cribari-Neto:2009:HCI

[CNdGAL09] Francisco Cribari-Neto and Maria da Glória A. Lima. Heteroskedasticityconsistent interval estimators. Journal of Statistical Computation and Simulation, 79(6):787–803, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cribari-Neto:2005:NET

[CNFO05] Francisco Cribari-Neto, Silvia L. P. Ferrari, and Waldemar A. S. C. Oliveira. Numerical evaluation of tests based on different heteroskedasticity-consistent covariance matrix estimators. Journal of Statistical Computation and Simulation, 75(8):611– 628, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cribari-Neto:2002:NUM

[CNV02] Francisco Cribari-Neto and Klaus L. P. Vasconcellos. Nearly unbiased maximum likelihood estimation for the beta distribution. Journal of Statistical Computation and Simulation, 72(2):107– 118, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cribari-Neto:2001:HCC

[CNZ01] Francisco Cribari-Neto and Spyros G. Zarkos. Heteroskedasticityconsistent covariance matrix estimation: White’s estimator and the bootstrap. Journal of Statistical Computation and Simulation, 68(4):391–411, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cribari-Neto:2004:LAH

[CNZ04] Francisco Cribari-Neto and Spyros G. Zarkos. Leverage-adjusted heteroskedastic bootstrap methods. Journal of Statistical Computation and Simulation, 74(3):215–232, 2004. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Coad:1992:DDA

[Coa92] D. S. Coad. Data-dependent allocation rules a comparative study of some for Bernoulli data. Journal of Statistical Computation and Simulation, 40(3–4):219–231, 1992. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Coad:1995:SAR

[Coa95] D. S. Coad. Sequential allocation rules for multi-armed clinical trials. Journal of Statistical Computation and Simulation, 52 (3):239–251, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cobb:1989:EWS

[Cob89] E. Benton Cobb. Estimation of the Weibull shape parameter in small-sample bioassay. Journal of Statistical Computation and Simulation, 31(2):93–101, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Conway:1995:ISE

[Con95] Mark R. Conway. Importance sampling to evaluate bivariate probabilities with an application to sequential phase II trials.

Journal of Statistical Computation and Simulation, 51(2-4):249– 262, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Congdon:2006:BMC

[Con06] Peter Congdon. Bayesian model comparison via parallel model output. Journal of Statistical Computation and Simulation, 76 (2):149–165, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cook:2007:CIT

[Coo07] Steven Cook. Co-integration testing using local-to-unity detrending: the impact of structural change under the null. Journal of Statistical Computation and Simulation, 77(3):265–270, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:1995:PBT

[Cor95] Gauss M. Cordeiro. Performance of a Bartlett-type modification for the deviance. Journal of Statistical Computation and Simulation, 51(2-4):385–403, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Corduas:1997:IIF

[Cor97] Marcella Corduas. Indirect inference for fractional time series models. Journal of Statistical Computation and Simulation, 59 (3):221–232, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cornell:2000:DMM

[Cor00] John A. Cornell. Developing mixture models, are we done. Journal of Statistical Computation and Simulation, 66(2):127–144, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:2004:CLR

[Cor04] Gauss M. Cordeiro. Corrected likelihood ratio tests in symmetric nonlinear regression models. Journal of Statistical Computation and Simulation, 74(8):609–620, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Carter:1976:EAO

[CP76] Michael C. Carter and Rudolph S. Parrish. An empirical analysis of overshoot duration times in a stationary Gaussian process. Journal of Statistical Computation and Simulation, 4(4):279– 292, 1976. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Caroni:1991:MOT

[CP91] C. Caroni and P. Prescott. Multivariate outlier tests with structured covariance matrices. Journal of Statistical Computation and Simulation, 38(1-4):165–179, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Caroni:1995:UIT

[CP95] C. Caroni and P. Prescott. Union-intersection testing for outliers in multivariate normal data. Journal of Statistical Computation and Simulation, 51(2-4):185–196, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cleroux:1973:AML

[CR73] R. Cléroux and P. Robert. An algorithm for maximum likelihood estimation in incomplete block variance component models. Journal of Statistical Computation and Simulation, 2(3): 237–251, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chaffin:1993:ESK

[CR93] Wilkie W. Chaffin and Steven G. Rhiel. The effect of skewness and kurtosis on the one-sample T test and the impact of knowledge of the population standard deviation. Journal of Statistical Computation and Simulation, 46(1–2):79–90, 1993. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cheang:2003:FSP

[CR03] Wai-Kwong Cheang and Gregory Reinsel. Finite sample properties of ML and REML estimators in time series regression models with long memory noise. Journal of Statistical Computation and Simulation, 73(4):233–259, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Craigmile:2005:AWB

[Cra05] Peter F. Craigmile. Approximate wavelet-based simulation of long memory processes. Journal of Statistical Computation and Simulation, 75(5):363–379, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cressie:1989:CET

[Cre89] Noel Cressie. C326. Ergodicity for time series and spatial processes. Journal of Statistical Computation and Simulation, 32 (1–2):61–63, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cade:2006:RSP

[CRM06] Brian S. Cade, Jon D. Richards, and Paul W. Mielke, Jr. Rank score and permutation testing alternatives for regression quantile estimates. Journal of Statistical Computation and Simulation, 76(4):331–355, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cross:1974:TFM

[Cro74] M. J. Cross. Tables of finite-mean nonsymmetric stable distributions as computed from their convergent and asymptotic series. Journal of Statistical Computation and Simulation, 3(1):1–27, 1974. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cowles:1999:PBI

[CRR99] Mary Kathryn Cowles, Gareth O. Roberts, and Jeffrey S. Rosenthal. Possible biases induced by MCMC convergence diagnostics. Journal of Statistical Computation and Simulation, 64(1): 87–104, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Comte:2007:FSP

[CRT07] F. Comte, Y. Rozenholc, and M.-L. Taupin. Finite sample penalization in adaptive density deconvolution. Journal of Statistical Computation and Simulation, 77(11):977–1000, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Canavos:1986:SBI

[CS86] George C. Canavos and Charles H. Smith. Sensitivity of a Bayesian inference to prior assumptions. Journal of Statistical Computation and Simulation, 25(1–2):25–36, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chattamvelli:1994:EAN

[CS94] R. Chattamvelli and R. Shanmugam. An enhanced algorithm for noncentral t-distribution. Journal of Statistical Computation and Simulation, 49(1–2):77–83, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:1995:MCE

[CS95a] Huifen Chen and Bruce W. Schmeiser. Monte Carlo estimation for guaranteed-coverage non-normal tolerance intervals. Journal of Statistical Computation and Simulation, 51(2-4):223–238, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:1995:ATN

[CS95b] Ling Chen and Samuel S. Shapiro. An alternative test for normality based on normalized spacings. Journal of Statistical Computation and Simulation, 53(3–4):269–287, 1995. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cornelius:1997:EGL

[CS97] P. L. Cornelius and M. S. Seyedsadr. Estimation of general linear-bilinear models for two-way tables. Journal of Statistical Computation and Simulation, 58(4):287–322, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Corcoran:2005:PPA

[CS05] J. N. Corcoran and U. Schneider. Pseudo-perfect and adaptive variants of the Metropolis–Hastings algorithm with an independent candidate density. Journal of Statistical Computation and Simulation, 75(6):459–475, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:2008:MSI

[CS08] Gauss M. Cordeiro and Borko D. Stosić. Maple script for improving test statistics. Journal of Statistical Computation and Simulation, 78(11):1045–1053, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chengalur-Smith:1993:MMS

[CSAR93] I. N. Chengalur-Smith, C. Jesse Arnold, and R. Marion Reynolds, Jr. Multiple means Shewhart charts using variable sampling intervals. Journal of Statistical Computation and Simulation, 45(1–2):1–21, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Colosimo:2000:BEP

[CSC00] E. A. Colosimo, A. F. Silva, and F. R. B. Cruz. Bias evaluation in the proportional hazards model. Journal of Statistical Computation and Simulation, 65(1-4):191–201, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chandra:1983:SPS

[CSS83] Mahesh Chandra, Nozer D. Singpurwalla, and Michael A. Stephens. Some problems in simulating the quantiles of the maxima and other functionals of Gaussian processes. Journal of Statistical Computation and Simulation, 18(1):45–57, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:1982:CCI

[CT82] Hubert J. Chen and Paul J. Tsai. A class of confidence intervals for the largest component mean of a multivariate normal population. Journal of Statistical Computation and Simulation, 14 (2):133–148, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cheng:1988:SPP

[CT88] R. C. H. Cheng and K. M. Thornton. Selected percentage points of the Moran statistic. Journal of Statistical Computation and Simulation, 30(3):189–194, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Crainiceanu:2007:NPT

[CV07] Ciprian M. Crainiceanu and Timothy J. Vogelsang. Nonmonotonic power for tests of a mean shift in a time series. Journal of Statistical Computation and Simulation, 77(6):457–476, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cook:2008:WSC

[CV08] Steven Cook and Dimitrios Vougas. A weighted symmetric cointegration test. Journal of Statistical Computation and Simulation, 78(6):559–565, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:1998:SOB

[CVS98] Gauss M. Cordeiro, Klaus L. P. Vasconcellos, and Maria Luiza F. Santos. on the second-order bias of parameter estimates in nonlinear regression models with Student t errors. Journal of Statistical Computation and Simulation, 60(4):363–378, 1998. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Carter:1972:SCA

[CW72] Michael C. Carter and William O. Williford. Some computationai aspects of Bayesian analysis. Journal of Statistical Computation and Simulation, 1(2):157–172, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Crosby:1974:EIA

[CW74] D. S. Crosby and M. P. Weinreb. Effect of incorrect atmospheric statistics on the accuracy of temperature profiles derived from satellite measurements. Journal of Statistical Computation and Simulation, 3(1):41–51, 1974. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Correa:1999:LBS

[CW99] Jose Andres Correa and David B. Wolfson. Length-bias: some characterizations and applications. Journal of Statistical Computation and Simulation, 64(3):209–219, 1999. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2001:GFT

[CW01] Cathy W. S. Chen and Yu-Wen Wen. On goodness of fit for time series regression models. Journal of Statistical Computation and Simulation, 69(3):239–256, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Crabbe:1989:MRI

[CY89] Lealand E. Crabbe and Brian B. Young. Model reduction via the internally balanced state space representation. Journal of Statistical Computation and Simulation, 33(4):233–248, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comment [Aok91].

Crabbe:1991:CNT

[CY91] Leland E. Crabbe and Brian B. Young. C380. A note on time series model reduction algorithms. Journal of Statistical Computation and Simulation, 39(3):178–185, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chatterjee:1992:UEF

[CY92] Sangit Chatterjee and Mustafa Yilmaz. Use of estimated fractal dimension in model identification for time series. Journal of Statistical Computation and Simulation, 41(3–4):129–141, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:1999:CSS

[CY99] Huifen Chen and Tsu-Kuang Yang. Computation of the sample size and coverage for guaranteed-coverage nonnormal tolerance intervals. Journal of Statistical Computation and Simulation, 63 (4):299–320, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cornell:1990:CAG

[CYB90] J. E. Cornell, D. M. Young, and T. L. Bratcher. C376. An algorithm for generating covariance matrices with specified departures from sphericity. Journal of Statistical Computation and Simulation, 37(3–4):240–243, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Campbell:1999:ADC

[CYC99] Nancy L. Campbell, Linda J. Young, and George A. Capuano. Analyzing over-dispersed count data in two-way crossclassification problems using generalized linear models. Journal of Statistical Computation and Simulation, 63(3):263–281, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Camarinopoulos:2002:FRR

[CZ02] L. Camarinopoulos and G. Zioutas. Formulating robust regression estimation as an optimum allocation problem. Journal of Statistical Computation and Simulation, 72(9):687–705, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2007:CEL

[CZ07] Kun Chen and Mai Zhou. Computation of the empirical likelihood ratio from censored data. Journal of Statistical Computation and Simulation, 77(12):1033–1042, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

David:1987:CAS

[DA87] H. A. David and D. M. Andrews. Closed adaptive sequential paired-comparison selection procedures. Journal of Statistical Computation and Simulation, 27(2):127–141, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dagpunar:1978:SVT

[Dag78] J. S. Dagpunar. Sampling of variates from a truncated gamma distribution. Journal of Statistical Computation and Simulation, 8(1):59–64, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dagenais:1989:SSP

[Dag89] Marcel G. Dagenais. Small sample performance of parameter estimators for Tobit models with serial correlation. Journal of Statistical Computation and Simulation, 33(1):11–26, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dagenais:1995:PAE

[Dag95] Marcel G. Dagenais. The performance of alternative estimators for the probit model with first-order serial correlation. Journal of Statistical Computation and Simulation, 51(2-4):343–351, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Damsleth:1982:EMA

[Dam82] Eivind Damsleth. Estimation in moving average models, why does it fail? Journal of Statistical Computation and Simulation, 16(2):109–128, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Declerck:1998:BLR

[DAM98] Lieven Declerck, Marc Aerts, and Geert Molenberghs. Behaviour of the likelihood ratio test statistic under a Bahadur model for exchangeable binary data. Journal of Statistical Computation and Simulation, 61(1–2):15–38, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Daniel:1979:CNA

[Dan79] Cuthbert Daniel. C27. Numerical adjustment of sums of squares. Journal of Statistical Computation and Simulation, 8(3–4):314– 315, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comment [Dan80b].

Daniel:1980:CCF

[Dan80a] Cuthbert Daniel. C60. A convenient formula for inverting a frequently occurring matrix. Journal of Statistical Computation and Simulation, 10(2):160, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Daniel:1980:CES

[Dan80b] Cuthbert Daniel. C61. Extension of s. s. (y_{1}, y_{2}) = (1/2)(y_{1} − y_{2})^{2} to n = 3 and 4 (see C27). Journal of Statistical Computation and Simulation, 10(2):160–161, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Dan79].

Daniel:1987:EPP

[Dan87] Zelterman Daniel. Estimating percentage points by simulation. Journal of Statistical Computation and Simulation, 27(2):107– 125, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

David:1979:CDS

[Dav79] H. A. David. C35. Does stronger opposition make it harder to win? Journal of Statistical Computation and Simulation, 9(1): 71–72, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Davis:1993:SLB

[Dav93] Linda J. Davis. Selection of load balancing parameters. Journal of Statistical Computation and Simulation, 44(3–4):209–230, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Day:1987:NSS

[Day87] J. W. R. Day. A new simple straightforward and practical asymptotic expansion for the incomplete beta function. Journal of Statistical Computation and Simulation, 28(3):227–243, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Deligonul:1984:SVE

[DB84a] Z. S. Deligonul and S. Bilgen. Solution of the Volterra equation of renewal theory with the Galerkin technique using cubic splines. Journal of Statistical Computation and Simulation, 20 (1):37–45, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Doris:1984:ERC

[DB84b] Binkley Doris and Jeffrey B. Birch. Estimation of regression coefficients under model misspecification and non-normality of errors. Journal of Statistical Computation and Simulation, 19 (2):155–170, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DeVleeschauwer:2002:EEQ

[DBVK02] D. De Vleeschauwer, M. J. C. Büchli, A. Van Moffaert, and R. E. Kooij. End-to-end queuing delay assessment in multi-service IP networks. Journal of Statistical Computation and Simulation, 72 (10):803–824, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dykstra:1999:DAT

[DC99] Richard Dykstra and Chris Carolan. The distribution of the argmax of two-sided Brownian motion with quadratic drift. Journal of Statistical Computation and Simulation, 63(1):47– 58, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

David:2000:ABS

[DC00] Burt A. David and Clint W. Coakley. Automatic bandwidth selection for modified m-smoother. Journal of Statistical Computation and Simulation, 66(4):295–319, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dasgupta:2002:SRI

[DC02] Nairanjana Dasgupta and Gongwei Chen. Some robustness issues for comparing multiple logistic regression slopes to a control for small samples. Journal of Statistical Computation and Simulation, 72(12):925–935, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dyer:2003:SSE

[DCA03] John Dyer, Michael Conerly, and B. Michael Adams. A simulation study and evaluation of multivariate forecast based control charts applied to ARMA processes. Journal of Statistical Computation and Simulation, 73(10):709–724, 2003. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dayal:1977:NEI

[DD77] Hari H. Dayal and James M. Dickey. Numerical evaluation of integrals involving products of t-densities. Journal of Statistical Computation and Simulation, 6(1):19–27, 1977. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dachs:1978:CTT

[DD78] J. Norberto Walter Dachs and Hélio Dos Santos Migon. Comparison of tests for the two-way layout. Journal of Statistical Computation and Simulation, 7(1):1–11, 1978. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Demirhan:2009:SSP

[DDB09] Y. Parlak Demirhan, H. Demirhan, and S. Bacanli. A simulation study on power comparisons for group sequential tests of nonparametric statistics. Journal of Statistical Computation and Simulation, 79(6):769–785, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DeGooijer:1981:IMS

[De 81] J. G. De Gooijer. An investigation of the moments of the sample autocovariances and autocorrelations for general ARMA processes. Journal of Statistical Computation and Simulation, 12 (3–4):175–192, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DeWit:1997:IIP

[De 97] M. S. De Wit. Identification of the important parameters in thermal building simulation models. Journal of Statistical Computation and Simulation, 57(1-4):305–320, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Davis:2006:ODE

[DE06] G. M. Davis and K. B. Ensor. Outlier detection in environmental monitoring network data: an application to ambient ozone measurements for Houston, Texas. Journal of Statistical Computation and Simulation, 76(5):407–422, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Deak:1980:FPG

[Deá80] I. Deák. Fast procedures for generating stationary normal vectors. Journal of Statistical Computation and Simulation, 10(3– 4):225–242, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Deak:1986:CPR

[Dea86] I. Deak. Computing probabilities of rectangles in case of multinormal distribution. Journal of Statistical Computation and Simulation, 26(1–2):101–114, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Deegan:1976:TNA

[Dee76] John Deegan, Jr. A test of the numerical accuracy of some matrix inversion algorithms commonly used in least squares programs. Journal of Statistical Computation and Simulation, 4 (4):269–278, 1976. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Deligonul:1983:CIC

[Del83] Z. Seyda Deligönul. C156. On the independence condition for the classical variance estimator in an autoregressive moving-average process. Journal of Statistical Computation and Simulation, 17 (3):233–236, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dembski:1990:CRD

[Dem90] William A. Dembski. C373. Reverse diffusion-limited aggregation. Journal of Statistical Computation and Simulation, 37(3– 4):231–234, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Demirtas:2006:MMO

[Dem06] Hakan Demirtas. A method for multivariate ordinal data generation given marginal distributions and correlations. Journal of Statistical Computation and Simulation, 76(11):1017–1025, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dent:1977:CEL

[Den77] Warren Dent. Computation of the exact likelihood function of an Arima process. Journal of Statistical Computation and Simulation, 5(3):193–206, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Devroye:1982:NAR

[Dev82] Luc Devroye. A note on approximations in random variate generation. Journal of Statistical Computation and Simulation, 14 (2):149–158, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Devroye:1984:UPI

[Dev84] Luc Devroye. On the use of probability inequalities in random variate generation. Journal of Statistical Computation and Simulation, 20(2):91–100, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Devroye:1992:RVG

[Dev92] Luc Devroye. Random variate generation for the digamma and trigamma distributions. Journal of Statistical Computation and Simulation, 43(3–4):197–216, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dey:1984:SEP

[Dey84] Dipak K. Dey. On simultaneous estimation of parametric functions in a contingency table. Journal of Statistical Computation and Simulation, 20(3):261–269, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Deaton:1980:CPA

[DF80] M. L. Deaton and R. V. Foutz. C75. Path analysis for complexvalued random variables with an application to spectral analysis of time series. Journal of Statistical Computation and Simulation, 12(1):61–65, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dupuis:1998:CCI

[DF98] D. J. Dupuis and C. A. Field. A comparison of confidence intervals for generalized extreme-value distributions. Journal of Statistical Computation and Simulation, 61(4):341–360, 1998. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Demirtas:2008:PMN

[DFY08] Hakan Demirtas, Sally A. Freels, and Recai M. Yucel. Plausibility of multivariate normality assumption when multiply imputing non-Gaussian continuous outcomes: a simulation assessment. Journal of Statistical Computation and Simulation, 78 (1):69–84, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dods:1995:ASA

[DG95] Johannah L. Dods and David E. A. Giles. Alternative strategies for ‘augmenting’ the Dickey–Fuller test: size-robustness in the face of pre-testing. Journal of Statistical Computation and Simulation, 53(3–4):243–258, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dias:2002:BAH

[DG02] Ronaldo Dias and Dani Gamerman. A Bayesian approach to hybrid splines non-parametric regression. Journal of Statistical Computation and Simulation, 72(4):285–297, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dunning:1977:GPD

[DH77] Kenneth A. Dunning and James N. Hanson. Generalized Pearson distributions and nonlinear programming. Journal of Statistical Computation and Simulation, 6(2):115–128, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dutter:1981:NMN

[DH81a] Rudolf Dutter and Peter J. Huber. Numerical methods for the nonlinear robust regression problem. Journal of Statistical Computation and Simulation, 13(2):79–113, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dyer:1981:EPS

[DH81b] Danny Dyer and Mickie Sue Harbin. An empirical power study of multi-sample tests for exponentiality. Journal of Statistical Computation and Simulation, 12(3–4):277–291, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dharan:1985:PSS

[Dha85] Bala G. Dharan. A priori sample size evaluation and information matrix computation for time series models. Journal of Statistical Computation and Simulation, 21(2):171–177, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DiNardo:2005:CBO

[Di 05] E. Di Nardo. On the connection between orthant probabilities and the first passage time problem. Journal of Statistical Computation and Simulation, 75(6):437–445, 2005. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dias:1998:DEH

[Dia98] Ronaldo Dias. Density estimation via hybrid splines. Journal of Statistical Computation and Simulation, 60(4):277–293, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dickinson:1978:TRE

[Dic78] Bradley W. Dickinson. Two recursive estimates of autoregressive models based on maximum likelihood. Journal of Statistical Computation and Simulation, 7(2):85–92, 1978. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dietz:1983:SSP

[Die83] E. Jacqueun Dietz. Small sample power of multivariate nonparametric tests for monotone trend. Journal of Statistical Computation and Simulation, 18(2-3):159–177, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dietrich:1993:CEC

[Die93] C. R. Dietrich. Computationally efficient Cholesky factorization of a covariance matrix with block Toeplitz structure. Journal of Statistical Computation and Simulation, 45(3–4):203–218, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dietrich:1994:NCI

[Die94] C. R. Dietrich. A note on computational issues associated with restricted maximum likelihood estimation of covariance parameters. Journal of Statistical Computation and Simulation, 49 (1–2):11–20, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dielman:2005:LAV

[Die05] Terry E. Dielman. Least absolute value regression: recent contributions. Journal of Statistical Computation and Simulation, 75(4):263–286, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dielman:2006:VEH

[Die06] Terry E. Dielman. Variance estimates and hypothesis tests in least absolute value regression. Journal of Statistical Computation and Simulation, 76(2):103–114, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dinwoodie:2008:SEI

[Din08] Ian H. Dinwoodie. Statistical estimation of internal available bandwidth. Journal of Statistical Computation and Simulation, 78(7):639–652, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dionne:1981:SSP

[Dio81] Louise Dionne. Small-sample properties of universally efficient nonparametric estimators of shift. Journal of Statistical Computation and Simulation, 12(3–4):259–276, 1981. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dette:2005:FSP

[DK05] Holger Dette and Robert Kwiecien. Finite sample performance of sequential designs for model identification. Journal of Statistical Computation and Simulation, 75(6):477–495, 2005. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dutt:1975:STC

[DL75] J. E. Dutt and T. K. Lin. A short table for computing normal orthant probabilities of dimensions 4 and 5. Journal of Statistical Computation and Simulation, 4(2):95–120, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dinse:1980:APC

[DL80] Gregg E. Dinse and S. W. Lagakos. The analysis of partiallycensored date from a first-order semi-Markov model. Journal of Statistical Computation and Simulation, 11(3–4):209–222, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dagenais:1981:NAM

[DL81] Marcel G. Dagenais and Tran Cong Liem. Numerical approximations of marginals from “well-behaved” joint posteriors. Journal of Statistical Computation and Simulation, 12(3–4):157–173, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Durilleul:1992:LRT

[DL92] Pierre Durilleul and Pierre Legendre. Lack of robustness in two tests of normality against autocorrelation in sample data. Journal of Statistical Computation and Simulation, 42(1–2):79–91, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Derouen:1978:AMT

[DM78] T. A. Derouen and T. J. Mitchell. An application of G-minimax techniques to the problem of fixed precision estimation of the binomial. Journal of Statistical Computation and Simulation, 7 (3–4):189–197, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dodge:1979:AFL

[DM79] Yadolah Dodge and Dibyen Majumdar. An algorithm for finding least square generalized inverses for classification models with arbitrary patterns. Journal of Statistical Computation and Simulation, 9(1):1–17, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Decoster:1991:ECD

[DM91] Gregory P. Decoster and Douglas W. Mitchell. The efficacy of the correlation dimension technique in detecting determinism in small samples. Journal of Statistical Computation and Simulation, 39(4):221–229, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Donatos:1993:SSL

[DM93] George S. Donatos and George C. Michailidis. A simulation study of least squares and ridge estimators for normal and nonnormal autocorrelated disturbances. Journal of Statistical Computation and Simulation, 47(1–2):49–66, 1993. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dodge:2005:RNG

[DM05] Yadolah Dodge and Giuseppe Melfi. Random number generators and rare events in the continued fraction of π. Journal of Statistical Computation and Simulation, 75(3):189–197, 2005. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic),

1563-5163. URL http://www.tandfonline.com/doi/abs/10.
1080/00949650410001687181.

Dykstara:1983:NLR

[DMF83] Richard L. Dykstara, Richard W. Madsen, and Kenneth Fairbanks. A nonparametric likelihood ratio test. Journal of Statistical Computation and Simulation, 18(4):247–264, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dietrich:1994:GCG

[DN94] C. R. Dietrich and G. N. Newsam. Generating correlated Gaussian random fields by orthogonal polynomial approximations to the square root of the covariance matrix. Journal of Statistical Computation and Simulation, 50(1–2):91–109, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DeGiovanni:2006:CES

[DN06] Livia De Giovanni and Maurizio Naldi. Comparative evaluation of semiparametric long-memory estimators. Journal of Statistical Computation and Simulation, 76(7):645–654, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Do:1992:SSB

[Do92] Kim-Anh Do. A simulation study of balanced and antithetic bootstrap resampling methods. Journal of Statistical Computation and Simulation, 40(3–4):153–166, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dodgson:1983:MPS

[Dod83] J. H. Dodgson. The moments and power of the skewness statistic \sqrt{b_{1}} in sampling from the Gram–Charlier distribution. Journal of Statistical Computation and Simulation, 18(1):7–14, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dodson:2000:IGG

[Dod00] C. T. J. Dodson. Information geodesics for gamma models of communication clustering. Journal of Statistical Computation and Simulation, 65(1-4):133–146, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Doguwa:1989:CSE

[Dog89] Sani I. Doguwa. A comparative study of the edge-corrected kernel-based nearest neighbour density estimators for point processes. Journal of Statistical Computation and Simulation, 33 (2):83–100, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Doguwa:1992:EPO

[Dog92] Sani I. Doguwa. On the estimation of the point-object nearest neighbour distribution F (y) for point processes. Journal of Statistical Computation and Simulation, 41(1–2):95–107, 1992. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Doguwa:1993:CNS

[Dog93] Sani I. Doguwa. On correlations of neuronal spike discharges. Journal of Statistical Computation and Simulation, 47(1–2):11– 24, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Doge:2001:PSR

[Dög01] Gunter Döge. Perfect simulation for random sequential adsorption of d-dimensional spheres with random radii. Journal of Statistical Computation and Simulation, 69(2):141–156, 2001. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dongsheng:1997:TOS

[Don97] Tu Dongsheng. Two one-sided tests procedures in establishing therapeutic equivalence with binary clinical endpoints: fixed sample performances and sample size derformination. Journal of Statistical Computation and Simulation, 59(3):271–290, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dorfman:1988:FSC

[Dor88] Alan Dorfman. A finite sample comparison of White’s direct and simple tests for heteroscedasticity. Journal of Statistical Computation and Simulation, 29(4):299–308, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dorvlo:2001:CRL

[Dor01] Atsu S. S. Dorvlo. Component reduction in linear discriminant analysis. Journal of Statistical Computation and Simulation, 71 (4):335–344, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Doros:2008:IDT

[Dor08] Gheorghe Doros. Improving dispersion tests by symmetrization. Journal of Statistical Computation and Simulation, 78(5):417– 428, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Downer:2002:PAO

[Dow02] Robert Downer. A permutation alternative and other test procedures for spatially correlated data in one-way Anova. Journal of Statistical Computation and Simulation, 72(9):747–757, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dey:1993:CPB

[DP93] Dipak K. Dey and Fengchun Peng. On the choice of prior for the Bayes estimation in accelerated life testing. Journal of Statistical Computation and Simulation, 48(3–4):207–217, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Dette:2006:CSM

[DP06] Holger Dette and Kay F. Pilz. A comparative study of monotone nonparametric kernel estimates. Journal of Statistical Computation and Simulation, 76(1):41–56, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dasgupta:2001:SST

[DPS01] Nairanjana Dasgupta, Francis G. Pascual, and John D. Spurrier. Small-sample techniques for comparing several logistic regression slopes to a standard. Journal of Statistical Computation and Simulation, 71(2):141–161, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dielman:1994:ELA

[DR94] Terry E. Dielman and Elizabeth L. Rose. Estimation in least absolute value regression with autocorrelated errors. Journal of Statistical Computation and Simulation, 50(1–2):29–43, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dielman:2002:BVT

[DR02] Terry Dielman and Elizabeth Rose. Bootstrap versus traditional hypothesis testing procedures for coefficients in least absolute value regression. Journal of Statistical Computation and Simulation, 72(8):665–675, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Drescher:2008:TPL

[Dre08] Daniel Drescher. Testing for presence of a latent process in count series. Journal of Statistical Computation and Simulation, 78 (7):595–607, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Desmond:2008:EPB

[DRYL08] Anthony F. Desmond, Gabriel A. Rodríguez-Yam, and Xuewen Lu. Estimation of parameters for a Birnbaum–Saunders regression model with censored data. Journal of Statistical Computation and Simulation, 78(11):983–997, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DeBeer:1989:MDW

[DS89] C. F. De Beer and J. W. H. Swanepoel. A modified Durbin– Watson test for serial correlation in multiple regression under nonnormality using the bootstrap. Journal of Statistical Computation and Simulation, 33(2):75–81, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dryver:2009:VRM

[DS09] Arthur L. Dryver and Jantra Sukkasem. Validating risk models with a focus on credit scoring models. Journal of Statistical Computation and Simulation, 79(2):181–193, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dasgupta:2006:UNM

[DSL06] N. Dasgupta, E. Solorzano, and N. A. Lazar. Using numerical methods to find the least favorable configuration when comparing k test treatments with both positive and negative controls. Journal of Statistical Computation and Simulation, 76(3):251– 265, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dougherty:2006:NAU

[DSS06] Daniel P. Dougherty, Eric A. Stahlberg, and Wolfgang Sadee. Network analysis using transitive closure: New methods for exploring networks. Journal of Statistical Computation and Simulation, 76(6):539–551, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dupuis:1996:EPOb

[Dup96a] D. J. Dupuis. Estimating the probability of obtaining nonfeasible parameter estimates of the generalized extreme-value distribution. Journal of Statistical Computation and Simulation, 56(1):23–38, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dupuis:1996:EPOa

[Dup96b] D. J. Dupuis. Estimating the probability of obtaining nonfeasible parameter estimates of the generalized Pareto distribution. Journal of Statistical Computation and Simulation, 54(1-3):197– 209, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dutter:1977:NSR

[Dut77] Rudolf Dutter. Numerical solution of robust regression problems: computational aspects, a comparison. Journal of Statistical Computation and Simulation, 5(3):207–238, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dutilleul:1999:MAM

[Dut99] Pierre Dutilleul. The MLE algorithm for the matrix normal distribution. Journal of Statistical Computation and Simulation, 64(2):105–123, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DeWet:1986:BCI

[DV86] T. De Wet and J. W. J. Van Wyk. Bootstrap confidence intervals for regression coefficients when the residuals are dependent. Journal of Statistical Computation and Simulation, 23(4):317– 327, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DuPlessis:1995:BCM

[DV95] J. L. Du Plessis and A. J. Van Der Merwe. Bayesian calibration in meat science research-estimation of the pH in postmortem muscle. Journal of Statistical Computation and Simulation, 53 (1–2):1–11, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dishon:1980:SSC

[DW80] Menachem Dishon and George H. Weiss. Small sample comparison of estimation methods for the beta distribution. Journal of Statistical Computation and Simulation, 11(1):1–11, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Drezner:1990:CBN

[DW90] Zvi Drezner and G. O. Wesolowsky. On the computation of the bivariate normal integral. Journal of Statistical Computation and Simulation, 35(1–2):101–107, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dupuis:2001:MFP

[DW01] D. J. Dupuis and C. Winchester. More on the four-parameter kappa distribution. Journal of Statistical Computation and Simulation, 71(2):99–113, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DeAngelis:1992:BCC

[DY92] Daniela De Angelis and Alastair Young. Bootstrapping the correlation coefficient: a comparison of smoothing strategies. Journal of Statistical Computation and Simulation, 40(3–4):167–176, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dykstra:1985:CAP

[Dyk85] Richard L. Dykstra. Computational aspects of l-projections. Journal of Statistical Computation and Simulation, 21(3–4):265– 274, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Easterling:1978:EPN

[EA78] R. G. Easterling and H. E. Anderson. The effect of preliminary normality goodness of fit tests on subsequent inference. Journal of Statistical Computation and Simulation, 8(1):1–11, 1978. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Engeman:1990:AED

[EB90] Richard M. Engeman and Jeffrey F. Bromaghin. An approach to estimating density from line transect data where the animals move in response to the observer. Journal of Statistical Computation and Simulation, 34(2-3):145–154, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

El-Bassiouni:2000:IEM

[EBA00] M. Y. El-Bassiouni and M. E. M. Abdelhafez. Interval estimation of the mean in a two-stage nested model. Journal of Statistical Computation and Simulation, 67(4):333–350, 2000. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ebbeler:1973:NEL

[Ebb73] Donald H. Ebbeler. A note on estimation in lognormal linear models. Journal of Statistical Computation and Simulation, 2 (3):225–231, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ebrahimi:1993:EBS

[Ebr93] Nader Ebrahimi. Estimating a bivariate survival function and its marginals under positive quadrant dependence. Journal of Statistical Computation and Simulation, 47(1–2):25–35, 1993. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Engelhart:1986:SAC

[EBS86] Max Engelhart, Lee J. Bain, and Wei-Kei Shuie. Statistical analysis of a compound exponential failure model. Journal of Statistical Computation and Simulation, 23(4):299–315, 1986. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Edelman:1994:CCI

[Ede94] David Edelman. C414. A confidence interval for an unknown unimodal density function based on a sample of size 1. Journal of Statistical Computation and Simulation, 49(3–4):231–235, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Editor:1983:CMB

[Edi83] Good I. J. Section Editor. C183. A modification, based on probability density estimation, of the Sherman–Foutz one-sample test. Journal of Statistical Computation and Simulation, 18(4):322, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Editor:2004:CCC

[Edi04] I. J. Good Section Editor. Comments, conjectures, and conclusions. Journal of Statistical Computation and Simulation, 74 (12):913–922, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Edwards:1985:ESB

[Edw85] Don Edwards. Exact simulation-based inference: A survey, with additions. Journal of Statistical Computation and Simulation, 22 (3–4):307–326, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Edward:1989:SES

[Edw89] Brent Edward. Statistical expert systems: an example. Journal of Statistical Computation and Simulation, 31(2):103–110, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Edwards:1995:EME

[Edw95] Lloyd J. Edwards. Effect of measurement error in testing separate linear models. Journal of Statistical Computation and Simulation, 52(3):205–216, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Emad-Eldin:1988:MOS

[EEÖ88] Aly A. A. Emad-Eldin and Aydin Öztürk. A modified onesample Q–Q plot and a test for normality. Journal of Statistical Computation and Simulation, 29(1):1–15, 1988. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Elston:1989:SME

[EG89] D. A. Elston and C. A. Glasbey. Simulating from a mixture of exponential distributions with some negatively weighted components. Journal of Statistical Computation and Simulation, 33(1): 1–9, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Eskridge:2004:LSD

[EGM^{+}04] K. M. Eskridge, S. G. Gilmour, R. Mead, N. A. Butler, and D. A. Travnicek. Large supersaturated designs. Journal of Statistical Computation and Simulation, 74(7):525–542, 2004. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ebrahimi:1992:TDU

[EH92] Nader Ebrahimi and Mohamed Habibullah. Testing to determine the underlying distribution using incomplete observations when the life time determine is proportionally related to the censoring time distribution. Journal of Statistical Computation and Simulation, 40(1–2):109–118, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Elfessi:2001:SSL

[EH01] Abdulaziz Elfessi and Robert H. Hoar. Simulation study of a linear relationship between two variables affected by errors. Journal of Statistical Computation and Simulation, 71(1):29– 40, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Erasto:2007:BAF

[EH07] Panu Erästö and Lasse Holmström. Bayesian analysis of features in a scatter plot with dependent observations and errors in predictors. Journal of Statistical Computation and Simulation, 77(5):421–431, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Eubank:1998:SSO

[EK98] R. L. Eubank and J. T. Kim. A simple second order smoother. Journal of Statistical Computation and Simulation, 61(3):271– 285, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Einbeck:2003:OML

[EK03] Jochen Einbeck and Göran Kauermann. Online monitoring with local smoothing methods and adaptive ridging. Journal of Statistical Computation and Simulation, 73(12):913–929, 2003. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

ElLozy:1982:ECD

[El 82] Mohamed El Lozy. Efficient computation of the distribution functions of Student’s t chi-squared and f to moderate accuracy. Journal of Statistical Computation and Simulation, 14(3–4):179– 189, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ellingsen:1975:LRR

[EL75] Walte R. Ellingsen and James F. Leathrum. On-line ridge regression; sequential biased estimation for nonorthogonal problems. Journal of Statistical Computation and Simulation, 3(3): 249–264, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ekberg:1997:UAS

[ELB97] C. Ekberg and I. Lundén-Burö. Uncertainty analysis for some actinides under groundwater conditions. Journal of Statistical Computation and Simulation, 57(1-4):271–284, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Ellenberg:2000:CSC

[Ell00] Jonas H. Ellenberg. Communication of statistical concepts: examples in medical collaboration. Journal of Statistical Computation and Simulation, 66(3):189–207, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Escobar:1986:PAL

[EM86] Luis A. Escobar and William Q. Meeker, Jr. Planning accelerated life tests with Type II censored data. Journal of Statistical Computation and Simulation, 23(4):273–297, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ensor:1990:ROA

[EN90] Katheerine B. Ensor and H. Joseph Newton. A recursive in order algorithm for least squares estimates of an autoregressive process. Journal of Statistical Computation and Simulation, 37 (1–2):115–126, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Engeman:1986:SSC

[EOD86] Richard M. Engeman, David L. Otis, and William E. Dusenberry. Small sample comparison of Thompson’s estimator to some common bioassay estimators. Journal of Statistical Computation and Simulation, 25(3–4):237–250, 1986. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Emrich:1992:SSS

[EP92] Lawrence J. Emrich and Marion R. Piedmonte. On some small sample properties of generalized estimating equation estimates for multivariate dichotomous outcomes. Journal of Statistical Computation and Simulation, 41(1–2):19–29, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Eriksson:1983:CCL

[Eri83] Bo Eriksson. On the construction of confidence limits for the regression coefficient when the residuals are dependent. Journal of Statistical Computation and Simulation, 17(4):297–309, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Eric:1997:MBM

[Éri97] Marchand Éric. On moments of beta mixtures, the noncentral beta distribution, and the coefficient of determination. Journal of Statistical Computation and Simulation, 59(2):161–178, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Epps:1986:OTT

[ES86] T. W. Epps and Kenneth J. Singleton. An omnibus test for the two-sample problem using the empirical characteristic function. Journal of Statistical Computation and Simulation, 26(3–4):177– 203, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Evans:1988:MCC

[ES88] M. Evans and T. Swartz. Monte Carlo computation of some multivariate normal probabilities. Journal of Statistical Computation and Simulation, 30(2):117–128, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Escalante-Sanboval:1998:MEF

[ESRV98] Carlos A. Escalante-Sanboval and Jose A. Raynal-Villasenor. Multivariate estimation of floods: the trivariate Gumbel distribution. Journal of Statistical Computation and Simulation, 61 (4):313–340, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Etiksson:1981:CLU

[Eti81] Bo Etiksson. Confidence levels using normal approximation to modified t-statistics for dependent variables. Journal of Statistical Computation and Simulation, 13(2):131–148, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Everitt:1988:TMN

[Eve88] B. S. Everitt. A test of multivariate normality against the alternative that the distribution is a mixture. Journal of Statistical Computation and Simulation, 30(2):103–115, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Everson:2001:EBI

[Eve01] Philip J. Everson. Exact Bayesian inference for normal hierarchical models. Journal of Statistical Computation and Simulation, 68(3):223–241, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Eslinger:1991:MHD

[EW91] Paul W. Eslinger and Wayne A. Woodward. Minimum Hellinger distance estimation for normal models. Journal of Statistical Computation and Simulation, 39(1–2):95–114, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Famoye:1998:CGG

[Fam98] Felix Famoye. Computer generation of generalized negative binomial deviates. Journal of Statistical Computation and Simulation, 60(2):107–122, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Famoye:1999:ETG

[Fam99] Felix Famoye. EDF tests for the generalized Poisson distribution. Journal of Statistical Computation and Simulation, 63(2): 159–168, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fan:1995:BBC

[Fan95] Tsai-Hung Fan. Bayesian bootstrap clones for finite state Markov chains. Journal of Statistical Computation and Simulation, 53(3–4):289–298, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fan:2003:IRA

[Fan03] Shu-Kai S. Fan. Implementation of ridge analysis on quadratic response surfaces. Journal of Statistical Computation and Simulation, 73(9):675–684, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Farebrother:1978:CTS

[Far78] R. W. Farebrother. The computation of three stage least squares estimates. Journal of Statistical Computation and Simulation, 6 (3–4):183–187, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Faraway:1990:BSB

[Far90a] Julian J. Faraway. Bootstrap selection of bandwidth and confidence bands for nonparametric regression. Journal of Statistical Computation and Simulation, 37(1–2):37–44, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Farebrother:1990:CLG

[Far90b] R. W. Farebrother. The cumulants of the logarithm of a gamma variable. Journal of Statistical Computation and Simulation, 36 (4):243–245, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Forsythe:1982:SUI

[FAS82] Alan B. Forsythe, Clauder Archer, and Jan Schwartz. Selective use of an instrumental variable. Journal of Statistical Computation and Simulation, 15(2-3):193–200, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fan:1990:ECD

[FB90] Tsai-Hung Fan and James O. Berger. Exact convolution of t distributions, with applications to Bayesian inference for a normal mean with t prior distributions. Journal of Statistical Computation and Simulation, 36(4):209–228, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fan:2009:BAH

[FBC09] Tsai-Hung Fan, N. Balakrishnan, and Chia-Chen Chang. The Bayesian approach for highly reliable electro-explosive devices using one-shot device testing. Journal of Statistical Computation and Simulation, 79(9):1143–1154, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fok:1989:VEM

[FC89] D. S. K. Fok and D. Crevier. Volume estimation by Monte Carlo methods. Journal of Statistical Computation and Simulation, 31 (4):223–235, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ferrari:1994:MFC

[FC94] Silvia L. P. Ferrari and Gauss M. Cordeiro. Matrix formulae for computing improved score tests. Journal of Statistical Computation and Simulation, 49(3–4):195–206, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fai:1996:AFT

[FC96] Alex Hrong-Tai Fai and Paul L. Cornelius. Approximate f-tests of multiple degree of freedom hypotheses in generalized least squares analyses of unbalanced split-plot experiments. Journal of Statistical Computation and Simulation, 54(4):363–378, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fan:2005:BCP

[FcC05] Tsai-Hung Fan and Wei chen Chen. Bayesian change points analysis on the seismic activity in northeastern Taiwan. Journal of Statistical Computation and Simulation, 75(11):857–868, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ferrari:1999:RAB

[FCN99] Silvia L. P. Ferrari and Francisco Cribari-Neto. On the robustness of analytical and bootstrap corrections to score tests in regressions models. Journal of Statistical Computation and Simulation, 64(2):177–191, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fearn:1979:ACI

[Fea79] T. Fearn. Aiding the choice of inverse chi prior distributions. Journal of Statistical Computation and Simulation, 10(1):25– 29, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Feltz:2002:CGK

[Fel02] Carol J. Feltz. Customizing generalizations of the Kolmogorov– Smirnov goodness-of-fit test. Journal of Statistical Computation and Simulation, 72(2):179–186, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Friedman:1984:TSC

[FF84] Linda Weiser Friedman and Hershey H. Friedman. Tutorial statistical considerations in computer simulation: the state of the art. Journal of Statistical Computation and Simulation, 19(3): 237–263, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ferrari:2007:APL

[FFCN07] Silvia L. P. Ferrari, Michel Ferreira Da Silva, and Francisco Cribari-Neto. Adjusted profile likelihoods for the Weibull shape parameter. Journal of Statistical Computation and Simulation, 77(7):531–548, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

FerreiraDaSilva:2008:ILI

[FFCN08] Michel Ferreira Da Silva, Silvia L. P. Ferrari, and Francisco Cribari-Neto. Improved likelihood inference for the shape parameter in Weibull regression. Journal of Statistical Computation and Simulation, 78(9):789–811, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fleischer:2009:STP

[FGSV09] Frank Fleischer, Catherine Gloaguen, Volker Schmidt, and Florian Voss. Simulation of the typical Poisson–Voronoi–Cox– Voronoi cell. Journal of Statistical Computation and Simulation, 79(7):939–957, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fairclough:1986:MLM

[FH86] Diane L. Fairclough and Ronald W. Helms. A mixed linear model with linear covariance structure: a sensitivity analysis of maximum likelihood estimators. Journal of Statistical Computation and Simulation, 25(3–4):205–236, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Feng:2009:SBB

[FH09] Yuanhua Feng and Siegfried Heiler. A simple bootstrap bandwidth selector for local polynomial fitting. Journal of Statistical Computation and Simulation, 79(12):1425–1439, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Field:1993:TAL

[Fie93] Christopher Field. Tail areas of linear combinations of chisquares and non-central chi-squares. Journal of Statistical Computation and Simulation, 45(3–4):243–248, 1993. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Firinguetti:1997:RRC

[Fir97] Luis Firinguetti. Ridge regression in the context of a system of seemingly unrelated regression equations. Journal of Statistical Computation and Simulation, 56(2):145–162, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fishman:1972:VRS

[Fis72] George S. Fishman. Variance reduction in simulation studies. Journal of Statistical Computation and Simulation, 1(2):173– 182, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fisher:1973:NEI

[Fis73a] N. I. Fisher. A note on the evaluation of the incomplete gamma function. Journal of Statistical Computation and Simulation, 2 (4):325–332, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fishman:1973:VRN

[Fis73b] George S. Fishman. Variance reduction for normal variates in Monte Carlo studies. Journal of Statistical Computation and Simulation, 2(4):365–374, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Franke:1983:EIP

[FJ83] Richard Franke and Toke Jayakchandran. An empirical investigation of the properties of a new goodness-of-fit test. Journal of Statistical Computation and Simulation, 16(3–4):175–187, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Franke:1984:TNM

[FJ84] R. Franke and T. Jayachandran. Tables for a new multivariate goodness-of-fit test. Journal of Statistical Computation and Simulation, 20(2):101–114, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Foutz:1982:CSI

[FJT82] R. V. Foutz, D. R. Jensen, and Y. L. Tong. C120. Sharper inequalities for the joint distributions of quadratic forms. Journal of Statistical Computation and Simulation, 14(2):159–163, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fang:2004:FSC

[FK04] Yue Fang and Sergio G. Koreisha. Forecasting with serially correlated regression models. Journal of Statistical Computation and Simulation, 74(9):625–649, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fenstad:1980:RES

[FKW80] Grete U. Fenstad, Morten Kjaernes, and Lars WallØe. Robust estimation of standard deviation. Journal of Statistical Computation and Simulation, 10(2):113–132, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fiskeaux:1982:TSP

[FL82] Charlie D. Fiskeaux and Robert F. Ling. Tukey smoothers as preprocessors for positive ar(1) parameter estimation in the presence of additive contamination. Journal of Statistical Computation and Simulation, 15(4):315–331, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fleischer:1995:TTP

[Fle95] Karlheinz Fleischer. Two tests of pseudo random number generators for independence and uniform distribution. Journal of Statistical Computation and Simulation, 52(4):311–322, 1995. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fu:2009:WDW

[FMC09] Wenjiang J. Fu, Bani Mallick, and Raymond J. Carroll. Why do we observe misclassification errors smaller than the Bayes error? Journal of Statistical Computation and Simulation, 79 (5):717–722, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fang:2006:PPS

[FMOR06] Yue Fang, Michael A. Martin, Terence J. O’neill, and Steven Roberts. On the power of portmanteau serial correlation tests. Journal of Statistical Computation and Simulation, 76(7):593– 604, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fokianos:2007:DRM

[Fok07] Konstantinos Fokianos. Density ratio model selection. Journal of Statistical Computation and Simulation, 77(9):805–819, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fong:1990:RER

[Fon90] Duncan K. H. Fong. Ranking and estimation of related means in two-way models — a Bayesian approach. Journal of Statistical Computation and Simulation, 34(2-3):107–117, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fong:1992:BAE

[Fon92] Duncan K. H. Fong. A Bayesian approach to the estimation of the largest normal mean. Journal of Statistical Computation and Simulation, 40(1–2):119–133, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Formann:1997:SSC

[For97] Anton K. Formann. Small-sample comparison of the exact and asymptotic upper tail probabilities of chi-squared goodness-of-fit statistics: the binomial and the mixture binomial. Journal of Statistical Computation and Simulation, 56(3):229–249, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Foster:1995:CSS

[Fos95] Peter Foster. A comparative study of some bias correction techniques for kernel- based density estimators. Journal of Statistical Computation and Simulation, 51(2-4):137–152, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Foutz:1980:MCE

[Fou80] Robert V. Foutz. A method for constructing exact tests from test statistics that have unknown null distributions. Journal of Statistical Computation and Simulation, 10(3–4):187–193, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Foutz:1981:CSM

[Fou81] Robert V. Foutz. C83. On the superiority of Monte Carlo tests. Journal of Statistical Computation and Simulation, 12(2):135– 137, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Felipe:2007:NFE

[FP07] A. Felipe and L. Pardo. New family of estimators for the loglinear model of quasi-independence based on power-divergence measures. Journal of Statistical Computation and Simulation, 77(5):407–420, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fairclough:1992:MCC

[FPRS92] Diane L. Fairclough, William C. Pierni, Gary J. Ridgway, and Neil C. Schwertman. A Monte Carlo comparison of the smoothing, scoring and em algorithms for dispersion matrix estimation with incomplete growth curve data. Journal of Statistical Computation and Simulation, 43(1–2):77–92, 1992. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Frazier:1974:ABI

[Fra74] Larry T. Frazier. An analysis of a Bayes inverse regression method of confidence intervals in linear calibration. Journal of Statistical Computation and Simulation, 3(2):99–103, 1974. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Frank:1979:EGT

[Fra79] Ove Frank. Estimating a graph from triad counts. Journal of Statistical Computation and Simulation, 9(1):31–46, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Franklin:1988:CEN

[Fra88] L. A. Franklin. The complete exact null distribution of Spearman’s rho for n = 12(1)18. Journal of Statistical Computation and Simulation, 29(3):578–580, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Franco:2006:URT

[FRB06] Glaura C. Franco, Valdério A. Reisen, and Paula A. Barros. Unit root tests using semi-parametric estimators of the long-memory parameter. Journal of Statistical Computation and Simulation, 76(8):727–735, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Frey:2007:NPI

[Fre07] J. Frey. A note on a probability involving independent order statistics. Journal of Statistical Computation and Simulation, 77 (11):969–975, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Frey:2009:ACR

[Fre09] Jesse Frey. An algorithm for computing rectangular multinomial probabilities. Journal of Statistical Computation and Simulation, 79(12):1483–1489, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Frisen:1979:NAE

[Fri79] Marianne Frisen. A note on alternating estimation in non-linear regression. Journal of Statistical Computation and Simulation, 9(1):19–23, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Frisen:1985:CFA

[Fri85] M. Frisén. C231. The forgotten assumption in the superpopulation model. Journal of Statistical Computation and Simulation, 21(3–4):331–334, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fried:2007:RLE

[Fri07] Roland Fried. Robust location estimation under dependence. Journal of Statistical Computation and Simulation, 77(2):131– 147, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

From:1989:PSS

[Fro89] Steven G. From. On the performances of some selection procedures for the best population assuming parent distributions with varying tail weight. Journal of Statistical Computation and Simulation, 33(1):27–35, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

From:1995:TSP

[Fro95] Steven G. From. A two-stage procedure for estimating the parameters of some discrete distributions. Journal of Statistical Computation and Simulation, 53(1–2):23–39, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

From:2004:ADR

[Fro04] Steven G. From. Approximating the distribution of a renewal process using generalized Poisson distributions. Journal of Statistical Computation and Simulation, 74(9):667–681, 2004. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Farrell:2006:CST

[FRS06] Patrick J. Farrell and Katrina Rogers-Stewart. Comprehensive study of tests for normality and symmetry: extending the Spiegelhalter test. Journal of Statistical Computation and Simulation, 76(9):803–816, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Franklin:1975:CEA

[FS75] Mark A. Franklin and Amitava Sen. Comparison of exact and approximate variate generation methods for the Erlang distribution. Journal of Statistical Computation and Simulation, 4(1): 1–18, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Frank:1981:PDS

[FS81] Ove Frank and Klas Svensson. On probability distributions of single-linkage dendrograms. Journal of Statistical Computation and Simulation, 12(2):121–131, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fenstad:1990:TSS

[FS90] Grete U. Fenstad and Eva Skovlund. A two-sample sequential rank test by sen investigated by stochastic simulation. Journal of Statistical Computation and Simulation, 36(2-3):129–137, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fairclough:1994:EPD

[FS94] Diane L. Fairclough and Neil C. Schwertman. Estimation of a positive define general dispersion matrix from repeated measures with. Journal of Statistical Computation and Simulation, 50 (1–2):21–27, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Farrell:2003:RBR

[FS03] Patrick Farrell and T. W. F. Stroud. Random balanced resampling: A new method for estimating variance components in unbalanced designs. Journal of Statistical Computation and Simulation, 73(3):181–194, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Farrell:2007:TMN

[FSBN07] Patrick J. Farrell, Matias Salibian-Barrera, and Katarzyna Naczk. On tests for multivariate normality and associated simulation studies. Journal of Statistical Computation and Simulation, 77(12):1065–1080, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Firat:1996:IBL

[FT96] Mehmet Z. Firat and R. Thompson. Investigation of bimodality in likelihoods and posterior densities. Journal of Statistical Computation and Simulation, 54(4):379–386, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fu:2009:BEM

[FTS09] Zhi-Hui Fu, Jian Tao, and Ning-Zhong Shi. Bayesian estimation in the multidimensional three-parameter logistic model. Journal of Statistical Computation and Simulation, 79(6):819–835, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fung:1979:MCS

[Fun79] Karen Yuen Fung. A Monte Carlo study of the Studentized Wilcoxon statistic for the Behrens–Fisher problem. Journal of Statistical Computation and Simulation, 10(1):15–24, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fung:1988:CVT

[Fun88] Wing-Kam Fung. Critical values for testing in multivariate statistical outliers. Journal of Statistical Computation and Simulation, 30(3):195–212, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ferrari:1997:SOA

[FUOCN97] Silvia L. P. Ferrari, Miguel A. Uribe-Opazo, and Francisco Cribari-Neto. Second order asymptotics for score tests in exponential family nonlinear models. Journal of Statistical Computation and Simulation, 59(2):179–194, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Furno:1991:CEH

[Fur91] Marilena Furno. Comparison of estimators for heteroskedastic models. Journal of Statistical Computation and Simulation, 38 (1-4):99–107, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Furno:1996:SSB

[Fur96] Marilena Furno. Small sample behavior of a robust heteroskedasticity consistent covariance matrix estimator. Journal of Statistical Computation and Simulation, 54(1-3):115–128, 1996. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Furno:2004:ATQ

[Fur04] Marilena Furno. ARCH tests and quantile regressions. Journal of Statistical Computation and Simulation, 74(4):277–292, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Furno:2007:NLT

[Fur07] Marilena Furno. Non-linearity tests based on order statistics and quantile regressions. Journal of Statistical Computation and Simulation, 77(3):189–211, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Furno:2008:MEE

[Fur08] Marilena Furno. Misspecification and estimation effect in the Lagrange multiplier tests for heteroskedasticity. Journal of Statistical Computation and Simulation, 78(4):299–313, 2008. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Feig:1988:REC

[FW88] Douglas G. Feig and Timothy A. Wittig. Ridge estimated confidence intervals: a Monte Carlo evaluation. Journal of Statistical Computation and Simulation, 29(2):127–142, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Griffith:1995:TOA

[GA95] Daniel A. Griffith and Sone Akio. Trade-offs associated with normalizing constant computional simplifications for estimating spatial statistical models. Journal of Statistical Computation and Simulation, 51(2-4):165–183, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gil-Alana:2000:ERT

[GA00] L. A. Gil-Alana. Evaluation of Robinson’s (1994) tests in finite samples. Journal of Statistical Computation and Simulation, 68 (1):39–63, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gil-Alana:2001:JTF

[GA01a] L. A. Gil-Alana. A joint test of fractional cyclic integration and a linear time trend. Journal of Statistical Computation and Simulation, 71(4):269–285, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gil-Alana:2001:PTR

[GA01b] Luis A. Gil-Alana. The power of the tests of Robinson (1994) in the context of fractionally integrated moving average models. Journal of Statistical Computation and Simulation, 70(3):267– 272, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gil-Alana:2002:LPT

[GA02] Luis Gil-Alana. Logarithmic and power transformations in the context of fractionally integrated processes. Journal of Statistical Computation and Simulation, 72(12):949–957, 2002. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gil-Alana:2009:BFC

[GA09] Luis A. Gil-Alana. A bivariate fractionally cointegrated relationship in the context of cyclical structures. Journal of Statistical Computation and Simulation, 79(11):1355–1370, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gallagher:2002:OIG

[Gal02] Colin Gallagher. Order identification for Gaussian moving averages using the covariation. Journal of Statistical Computation and Simulation, 72(4):279–283, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ghitany:2009:EMD

[GAM09] M. E. Ghitany and D. K. Al-Mutairi. Estimation methods for the discrete Poisson–Lindley distribution. Journal of Statistical Computation and Simulation, 79(1):1–9, 2009. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gans:1984:SSD

[Gan84] Daniel J. Gans. The search for significance: different tests on the same data. Journal of Statistical Computation and Simulation, 19(1):1–21, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gan:1989:CCS

[Gan89] F. F. Gan. Combined cumulative sum and Shewhart variance charts. Journal of Statistical Computation and Simulation, 32 (3):149–163, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gan:1990:MOG

[Gan90] F. F. Gan. Monitoring observations generated from a binomial distribution using modified exponentially weighted moving average control chart. Journal of Statistical Computation and Simulation, 37(1–2):45–60, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gan:1991:ECC

[Gan91] F. F. Gan. Ewma control chart under linear drift. Journal of Statistical Computation and Simulation, 38(1-4):181–200, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gan:1993:EWM

[Gan93a] F. F. Gan. Exponentially weighted moving average control charts with reflecting boundaries. Journal of Statistical Computation and Simulation, 46(1–2):45–67, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gan:1993:ODE

[Gan93b] F. F. Gan. An optimal design of EWMA control charts based on median run length. Journal of Statistical Computation and Simulation, 45(3–4):169–184, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gao:2004:CBD

[Gao04] Sujuan Gao. Combining binomial data using the logistic normal model. Journal of Statistical Computation and Simulation, 74 (4):293–306, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gates:1991:EMC

[Gat91] John Gates. Exact Monte Carlo tests using several statistics. Journal of Statistical Computation and Simulation, 38(1-4):211– 218, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gatto:2000:MST

[Gat00] Riccardo Gatto. Multivariate saddlepoint test for the wrapped normal model. Journal of Statistical Computation and Simulation, 65(1-4):271–285, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gatto:2009:AAD

[Gat09] Riccardo Gatto. Asymptotic approximations to the distribution of Kendall’s sample τ. Journal of Statistical Computation and Simulation, 79(5):671–679, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Guerard:1986:CFA

[GB86] John B. Guerard, Jr. and Carl R. Beidleman. Composite forecasting of annual earnings: an application of biased regression techniques. Journal of Statistical Computation and Simulation, 24(1):1–16, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Genz:1999:NCM

[GB99] Alan Genz and Frank Bretz. Numerical computation of multivariate t-probabilities with application to power calculation of multiple contrasts. Journal of Statistical Computation and Simulation, 63(4):103–117, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CPS

[GBSS84] I. J. Good, K. O. Bowman, L. R. Shenton, and G. Szekeres. C212. A Poisson sum up to the mean and a Ramanujan problem. Journal of Statistical Computation and Simulation, 20(2):167– 173, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gbur:1981:ECR

[Gbu81] Edward E. Gbur. The effect of constraints on the response variable in regression. Journal of Statistical Computation and Simulation, 13(2):115–130, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CGR

[GC90] I. J. Good and I. N. Chengalur. C340. A generalization of Rissanen’s distribution. Journal of Statistical Computation and Simulation, 34(2-3):155–158, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gupta:2003:SCC

[GC03] Arjun Gupta and Tuhao Chen. On the sample characterization criterion for normal distributions. Journal of Statistical Computation and Simulation, 73(3):155–163, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gefeller:1992:NNK

[GD92] Olaf Gefeller and Holger Dette. Nearest neighbour kernel estimation of the hazard function from censored data. Journal of Statistical Computation and Simulation, 43(1–2):93–101, 1992. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Genizi:1976:EIC

[Gen76] A. Genizi. On estimating the interaction component of variance from unbalanced data. Journal of Statistical Computation and Simulation, 5(1):61–64, 1976. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gerig:1978:NA

[Ger78] Thomas M. Gerig. News and abstracts. Journal of Statistical Computation and Simulation, 6(3–4):317–318, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CFP

[GF89] I. J. Good and R. V. Foutz. C329. The “Fisher–Poisson trick” extended to contingency tables. Journal of Statistical Computation and Simulation, 32(1–2):76–81, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gerig:1975:CML

[GG75] Thomas M. Gerig and A. Ronald Gallant. Computing methods for linear models subject to linear parametric constraints. Journal of Statistical Computation and Simulation, 3(3):283– 296, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gaskins:1977:CCD

[GG77] R. A. Gaskins and I. J. Good. C4. Compenis on the distribution of the likelihood ratio statistic for the equiprobable multinomial. Journal of Statistical Computation and Simulation, 6(2):150– 152, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CBO

[GG78] I. J. Good and R. A. Gaskins. C19. Bayesian oedemology. Journal of Statistical Computation and Simulation, 7(3–4):299–301, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gaskins:1980:CCB

[GG80] R. A. Gaskins and I. J. Good. C62. The covariance between the chi-squareds for the margins of a contingency table. Journal of Statistical Computation and Simulation, 10(2):161–162, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CSC

[GG90] I. J. Good and N. L. Gilinsky. C346. A simple comment concerning the Galton–Watson–Bienayme process. Journal of Statistical Computation and Simulation, 34(2-3):165–166, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gutierrez:2009:TPS

[GGSNR09] R. Gutiérrez, R. Gutiérrez-Sánchez, A. Nafidi, and E. Ramos. Three-parameter stochastic lognormal diffusion model: statistical computation and simulating annealing — application to real case. Journal of Statistical Computation and Simulation, 79(1):25–38, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Godambe:1976:EAD

[GH76] A. V. Godambe and W. L. Harkness. Exact and asymptotic distributions of vectors of column totals relate to random matrices and their approximations — I. Journal of Statistical Computation and Simulation, 5(1):23–35, 1976. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CFR

[GH82a] I. J. Good and Golde I. Holtzman. C141. On s-fold repeats and their additive and multiflicative properties. Journal of Statistical Computation and Simulation, 16(1):66–69, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CMN

[GH82b] I. J. Good and Golde I. Hotzman. C142. The moments of the number of s-fold repeats. Journal of Statistical Computation and Simulation, 16(1):69–75, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CED

[GH83a] I. J. Good and C. I. Holtzman. C179. The exact distribution of χ^{2} for a non-equiprobable multinomial. Journal of Statistical Computation and Simulation, 18(2-3):238–239, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Good:1983:CCCb

[GH83b] I. J. Good and C. I. Holtzman. C180. Conjectures concerning the cumulants of the multivariate lognormal distribution. Journal of Statistical Computation and Simulation, 18(2-3):239–241, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gunning:2007:ILH

[GH07] P. Gunning and J. Horgan. Improving the Lavallée and Hidiroglou algorithm for stratification of skewed populations. Journal of Statistical Computation and Simulation, 77(4):277– 291, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ghazali:2001:MJM

[Gha01] Syed Shakir Ali Ghazali. Modification of Jessen’s method 2. Journal of Statistical Computation and Simulation, 70(1):45– 54, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Garcia-Hiernaux:2009:FEM

[GHCJ09] Alfredo García-Hiernaux, José Casals, and Miguel Jerez. Fast estimation methods for time-series models in state-space form. Journal of Statistical Computation and Simulation, 79(2):121– 134, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CDH

[GHDB89] I. J. Good, G. I. Holtzman, M. L. Deaton, and L. H. Bernstein. C328. Diagnosis of heart attack from two enzyme measurements by means of bivariate probability density estimation: statistical details. Journal of Statistical Computation and Simulation, 32 (1–2):68–76, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ghorai:1995:LSC

[Gho95] J. K. Ghorai. Large sample confidence interval for the quantiles in the proportional hazards model of random censorship. Journal of Statistical Computation and Simulation, 53(1–2):91– 102, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gao:1999:MCI

[GHW99] Andre H. Gao, Michael Hazilla, and George H. K. Wang. A Monte Carlo investigation of the BDS statistic. Journal of Statistical Computation and Simulation, 62(4):319–356, 1999. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gilberto:1995:IRR

[Gil95] Paula A. Gilberto. Influence and residuals in restricted generalized linear models. Journal of Statistical Computation and Simulation, 51(2-4):315–331, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gill:2007:ECV

[Gil07] Peter M. W. Gill. Efficient calculation of p-values in linearstatistic permutation significance tests. Journal of Statistical Computation and Simulation, 77(1):55–61, 2007. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gupta:1977:BBT

[GJ77] R. P. Gupta and G. C. Jain. A bivariate Borel–Tanner distribution and its approximation. Journal of Statistical Computation and Simulation, 6(2):129–135, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gupta:1979:MBD

[GJ79] R. P. Gupta and G. C. Jain. Modified Borel distributions. Journal of Statistical Computation and Simulation, 9(2):117–125, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CDN

[GJ82] I. J. Good and D. R. Jensen. C131. Determining the noncentral Wishart characteristic function via quadratic expressions. Journal of Statistical Computation and Simulation, 15(1):85– 87, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CNM

[GJH90] I. J. Good, D. R. Jensen, and Y. V. Hui. C345. Numerology for the masses of elementary verticles [sic]: an updating. Journal of Statistical Computation and Simulation, 34(2-3):164–165, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gutierrez-Jaimez:2002:EPL

[GJJL02] R. Gutiérrez-Jáimez and J. D. Jiménez-López. On the efficiency property of the linear estimators in the bivariate Pearson Type VII distribution. Journal of Statistical Computation and Simulation, 72(12):937–948, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gutierrez-Jaimez:2006:ELE

[GJJLGS06] R. Gutiérrez-Jáimez, J. D. Jiménez-López, and R. Gutiérrez-Sánchez. Efficient linear estimation problem in the bivariate Kotz distribution under dependence assumptions. Journal of Statistical Computation and Simulation, 76(2):115–130, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ganeshanandam:1990:ERE

[GK90] S. Ganeshanandam and W. J. Krzanowski. Error-rate estimation in two-group discriminant analysis using the linear discriminant function. Journal of Statistical Computation and Simulation, 36 (2-3):157–175, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Genz:2000:NES

[GK00] Alan Genz and Koon-Shlng Kwong. Numerical evaluation of singular multivariate normal distributions. Journal of Statistical Computation and Simulation, 68(1):1–21, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gupta:2001:GED

[GK01] Rameshwar D. Gupta and Debasis Kundu. Generalized exponential distribution: different method of estimations. Journal of Statistical Computation and Simulation, 69(4):315–337, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gupta:2004:DBG

[GK04] Rameshwar D. Gupta and Debasis Kundu. Discriminating between gamma and generalized exponential distributions. Journal of Statistical Computation and Simulation, 74(2):107–121, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Goh:2005:PCN

[GK05] Kim-Leng Goh and Maxwell King. Pre-checking for nonmonotonicity of the Wald statistic. Journal of Statistical Computation and Simulation, 75(9):751–759, 2005. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CRB

[GL83] I. J. Good and Byron C. Lewis. C147. The relationship between an informational measure of association and a generalization of one of Yule’s measures for 2 × s tables, and the relevance of generalized determinants and permanents. Journal of Statistical Computation and Simulation, 17(1):63–66, 1983. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CGP

[GL84] I. J. Good and B. C. Lewis. C185. Generation of pairs of related probability vectors. Journal of Statistical Computation and Simulation, 19(1):82–85, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:MLM

[GL85] I. J. Good and B. C. Lewis. A maximum-likelihood/minimum-discrepancy method for estimation of multinomial probabilities when there are related multinomials. Journal of Statistical Computation and Simulation, 21(3–4):227–240, 1985. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CBT

[GL86] I. J. Good and B. C. Lewis. The choice between techniques for choosing between methods of estimation. Journal of Statistical Computation and Simulation, 25(1–2):1–8, 1986. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1988:CFN

[GL88] I. J. Good and Byron C. Lewis. C304. A further note on probability estimation for pairs of related multinomials. Journal of Statistical Computation and Simulation, 29(4):353–357, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gupta:1990:MOM

[GL90] A. K. Gupta and T. P. Logan. On a multiple observations model in discriminant analysis. Journal of Statistical Computation and Simulation, 34(2-3):119–132, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Giles:1992:SPD

[GL92] David Giles and Offer Lieberman. Some properties of the Durbin–Watson test after a preliminary t-test. Journal of Statistical Computation and Simulation, 41(3–4):219–227, 1992. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gu:1996:SAS

[GL96] Minggao Gu and Shaolin Li. A sampling approach to selfconsistent estimators with arbitrary censored data. Journal of Statistical Computation and Simulation, 55(1–2):101–110, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gupta:1999:VST

[GLC99] A. K. Gupta, T. P. Logan, and Jie Chen. A variable selection technique in discriminant analysis with application in marketing data. Journal of Statistical Computation and Simulation, 63(2): 187–199, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Glendinning:1989:EIA

[Gle89a] R. H. Glendinning. An evaluation of the ICM algorithm for image reconstruction. Journal of Statistical Computation and Simulation, 31(3):169–185, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Glendinning:1989:CHZ

[Gle89b] Richard Glendinning. The convex hull of the zeros of a random polynomial. Journal of Statistical Computation and Simulation, 31(2):111–118, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Glendinning:1991:CHD

[Gle91] R. H. Glendinning. The convex hull of a dependent vector-valued process. Journal of Statistical Computation and Simulation, 38 (1-4):219–237, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Glendinning:1997:TJP

[Gle97] R. H. Glendinning. Testing for a jump in the periodogram. Journal of Statistical Computation and Simulation, 56(2):117– 144, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gray:1972:EDA

[GLP72] H. L. Gray, T. O. Lewis, and J. L. Poirot. Effects of digital approximation in linear models. Journal of Statistical Computation and Simulation, 1(3):287–294, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Glynn:1984:SAF

[Gly84] Peter W. Glynn. Some asymptotic formulas for Markov chains with applications to simulation. Journal of Statistical Computation and Simulation, 19(2):97–112, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Greenstreet:1977:DCS

[GM77] R. L. Greenstreet and R. B. Madden. Design considerations for stratified finite populations using the linear least-squares prediction approach. Journal of Statistical Computation and Simulation, 6(2):83–100, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ganesalingam:1979:SSR

[GM79] S. Ganesalingam and G. J. McLachlan. Small sample results for a linear discriminant function estimated from a mixture of normal populations. Journal of Statistical Computation and Simulation, 9(2):151–158, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gulati:2006:TSF

[GM06] Sneh Gulati and Jie Mi. Testing for scale families using total variation distance. Journal of Statistical Computation and Simulation, 76(9):773–792, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Green:2008:ASP

[GM08] D. H. Green and J. S. Morgan. Asymptotic spatial properties of stochastic cell growth on a regular tessellation. Journal of Statistical Computation and Simulation, 78(2):115–124, 2008. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gonzalez-Manteiga:2008:BMS

[GMLM^{+}08] W. González-Manteiga, M. J. Lombardía, I. Molina, D. Morales, and L. Santamaría. Bootstrap mean squared error of a smallarea EBLUP. Journal of Statistical Computation and Simulation, 78(5):443–462, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Giguere:1982:MAE

[GMR82] J. C. M. Giguere, E. B. Manoukian, and J. M. A. Roy. Maximum absolute error for Bartlett’s chi-square approximation. Journal of Statistical Computation and Simulation, 15(2-3):109–117, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

George:1995:CCI

[GMSS95] Karen J. George, Joseph W. McKean, William R. Schucany, and Simon J. Sheather. A comparison of confidence intervals from R-estimators in regression. Journal of Statistical Computation and Simulation, 53(1–2):13–22, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gupta:1989:AND

[GN89] A. K. Gupta and D. K. Nagar. Asymptotic nonnull distribution of likelihood ratio statistic for testing homogeneity of complex multivariate Gaussian populations. Journal of Statistical Computation and Simulation, 31(2):83–91, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gneiting:1997:NSM

[Gne97] Tilmann Gneiting. Normal scale mixtures and dual probability densities. Journal of Statistical Computation and Simulation, 59 (4):375–384, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

George:1990:ECL

[GO90] Karen George and M. R. Osborne. The efficient computation of linear rank statistics. Journal of Statistical Computation and Simulation, 35(3–4):227–237, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gomes:2003:MLR

[GO03] M. Ivette Gomes and Orlando Oliveira. Maximum likelihood revisited under a semi-parametric context — estimation of the tail index. Journal of Statistical Computation and Simulation, 73(4):285–301, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Golhar:1972:EFS

[Gol72] M. B. Golhar. The errors of first and second kinds in Welch– Aspin’s solution of the Behrens–Fisher problem. Journal of Statistical Computation and Simulation, 1(3):209–224, 1972. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Goldsmith:1977:EDS

[Gol77] Heather Goldsmith. The exact distributions of the serial correlation coefficients and an evaluation on some approximate distributions. Journal of Statistical Computation and Simulation, 5(2):115–134, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1977:CDL

[Goo77] I. J. Good. C2. Does −2 log(likelihoodratio) exceed x^{2}? Journal of Statistical Computation and Simulation, 6(1):70, 1977. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CML

[Goo78a] I. J. Good. C10. Are maximum-likelihood estimates invariant? Journal of Statistical Computation and Simulation, 7(1):80–81, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CCN

[Goo78b] I. J. Good. C11. The complexity of nutrition advice. Journal of Statistical Computation and Simulation, 7(1):81, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CMK

[Goo78c] I. J. Good. C14. Moments of the Kolmogorov–Smirnov onesample statistic. Journal of Statistical Computation and Simulation, 7(3–4):289–290, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CST

[Goo78d] I. J. Good. C15. A signs test when estimating probability densities. Journal of Statistical Computation and Simulation, 7(3–4): 290–292, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CET

[Goo78e] I. J. Good. C16. Ethical treatments. Journal of Statistical Computation and Simulation, 7(3–4):292–295, 1978. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CHC

[Goo78f] I. J. Good. C17. Histograms with 2^{n} classes. Journal of Statistical Computation and Simulation, 7(3–4):295, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CFB

[Goo78g] I. J. Good. C18. A fuzzy Bayesian method for estimating probabilities given two related multinomial distributions. Journal of Statistical Computation and Simulation, 7(3–4):296–299, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CPA

[Goo78h] I. J. Good. C22. Path analysis and correlation for power functions. Journal of Statistical Computation and Simulation, 8(1): 80, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CMC

[Goo78i] I. J. Good. C23. Monozygotic criminals. Journal of Statistical Computation and Simulation, 8(2):161–162, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CMA

[Goo78j] I. J. Good. C24. A method for avoiding overflows and underflows. Journal of Statistical Computation and Simulation, 8(2): 162–163, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CIT

[Goo78k] I. J. Good. C5. Interpolation in tables of the normal distribution. Journal of Statistical Computation and Simulation, 6(3–4):313, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CMV

[Goo78l] I. J. Good. C6. A minimal-vocabulary algorithm? Journal of Statistical Computation and Simulation, 6(3–4):314, 1978. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CAA

[Goo78m] I. J. Good. C7. Adenine arabinoside therapy. Journal of Statistical Computation and Simulation, 6(3–4):314–315, 1978. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CIC

[Goo78n] I. J. Good. C9. An improvement in the chi-squared goodnessof-fit test. Journal of Statistical Computation and Simulation, 7(1):79–81, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CCRa

[Goo79a] I. J. Good. C25. A comment on runs of signs. Journal of Statistical Computation and Simulation, 8(3–4):311–312, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CCS

[Goo79b] I. J. Good. C26. A comparison of some statistical estimates for the numbers of contingency tables. Journal of Statistical Computation and Simulation, 8(3–4):312–314, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CIM

[Goo79c] I. J. Good. C28. The importance of mathematical consistency, at least in a discriminiation problem. Journal of Statistical Computation and Simulation, 8(3–4):315–316, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CCCa

[Goo79d] I. J. Good. C29. A comment on Cuthbert Daniel’s adjustment of sums of squares. Journal of Statistical Computation and Simulation, 8(3–4):316, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CPF

[Goo79e] I. J. Good. C38. “Proper fees” in multiple-choice examinations. Journal of Statistical Computation and Simulation, 9(1):73–74, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CPX

[Goo79f] I. J. Good. C40. Partitioning of X^{2} and of the likelihood ratio for contingency tables. Journal of Statistical Computation and Simulation, 9(1):74–77, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CCJ

[Goo79g] I. J. Good. C41. On the combination of judgements concerning quantiles of a distribution with potential application to the estimation of mineral resources. Journal of Statistical Computation and Simulation, 9(1):77–79, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comment [Goo79h].

Good:1979:CCCb

[Goo79h] I. J. Good. C42. Comment on C41. Journal of Statistical Computation and Simulation, 9(2):159, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo79g].

Good:1979:CBM

[Goo79i] I. J. Good. C43. The Bayesian meaning of invariant estimation. Journal of Statistical Computation and Simulation, 9(2):160– 161, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CBB

[Goo79j] I. J. Good. C44. Bayes’s billiard-table argument extended to multinomials. Journal of Statistical Computation and Simulation, 9(2):161–163, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CII

[Goo79k] I. J. Good. C45. An index of influentiality of a scientist. Journal of Statistical Computation and Simulation, 9(2):163, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CVP

[Goo79l] I. J. Good. C46. A voting proposal: power-summing. Journal of Statistical Computation and Simulation, 9(2):164, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CCP

[Goo79m] I. J. Good. C47. A comment on “proper fees” (c38). Journal of Statistical Computation and Simulation, 9(2):164–165, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CBE

[Goo79n] I. J. Good. C49. A Boolean error? Journal of Statistical Computation and Simulation, 9(3):236–238, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comments [Goo80e, Goo81p].

Good:1979:CPS

[Goo79o] I. J. Good. C50. Predictive sample re-use and the estimation of probabilities. Journal of Statistical Computation and Simulation, 9(3):238–239, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CCRb

[Goo79p] I. J. Good. C52. The clustering of random variables. Journal of Statistical Computation and Simulation, 9(3):241–243, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CPC

[Goo79q] I. J. Good. C53. Partial correlation and spherical trigonometry. Journal of Statistical Computation and Simulation, 9(3):243– 245, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CMG

[Goo79r] I. J. Good. C54. Moments of a generalized gamma density. Journal of Statistical Computation and Simulation, 9(4):311, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CFM

[Goo79s] I. J. Good. C56. Fractional moments and cumulants: some unsolved problems. Journal of Statistical Computation and Simulation, 9(4):314–315, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CFG

[Goo79t] I. J. Good. C57. Further generallzations of X^{2}. Journal of Statistical Computation and Simulation, 9(4):315, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CGX

[Goo79u] I. J. Good. C58. Generalizations of X^{2} for contingency tables. Journal of Statistical Computation and Simulation, 9(4):315– 316, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CNO

[Goo80a] I. J. Good. C59. The number of orderings of n candidates when ties and omissions are both allowed. Journal of Statistical Computation and Simulation, 10(2):159, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comment [Goo80g].

Good:1980:CPD

[Goo80b] I. J. Good. C64. The probability distribution of the gaps between the Mersenne primes. Journal of Statistical Computation and Simulation, 10(2):165, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CAR

[Goo80c] I. J. Good. C67. Another relationship between weight of evidence and errors of the first and second kinds. Journal of Statistical Computation and Simulation, 10(3–4):315–316, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CGC

[Goo80d] I. J. Good. C68. The grading of competitors. Journal of Statistical Computation and Simulation, 10(3–4):316–318, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CCCa

[Goo80e] I. J. Good. C70. Comment concerning C49. Journal of Statistical Computation and Simulation, 11(1):74, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo79n].

Good:1980:CDC

[Goo80f] I. J. Good. C71. Degrees of causation in regression analysis. Journal of Statistical Computation and Simulation, 11(2):153– 155, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CCCb

[Goo80g] I. J. Good. C72. Comment concerning C59. Journal of Statistical Computation and Simulation, 11(2):155–156, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo80a].

Good:1980:CDS

[Goo80h] I. J. Good. C73. The diminishing significance of a p-value as the sample size increases. Journal of Statistical Computation and Simulation, 11(3–4):307–313, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CTV

[Goo80i] I. J. Good. C74. Terminology for various kinds of Bayesian methods. Journal of Statistical Computation and Simulation, 11 (3–4):309–313, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comment [Goo83-29].

Good:1980:CFD

[Goo80j] I. J. Good. C77. Functions of distinct random vectors having identical distributions. Journal of Statistical Computation and Simulation, 12(1):68–70, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CGP

[Goo80k] I. J. Good. C78. On Godambe’s paradox. Journal of Statistical Computation and Simulation, 12(1):70–72, 1980. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CVL

[Goo80l] I. J. Good. C79. Vinograde’s lemma, singular decompositions and k-frames. Journal of Statistical Computation and Simulation, 12(1):72–74, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CFP

[Goo80m] I. J. Good. C80. Feynman’s path integrals and Sewall Wright’s path analysis. Journal of Statistical Computation and Simulation, 12(1):74–77, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CSC

[Goo80n] I. J. Good. C81. A simplification in “the causal calculus”. Journal of Statistical Computation and Simulation, 12(1):77– 78, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:BR

[Goo80o] Phillip Good. Book review. Journal of Statistical Computation and Simulation, 11(2):157–160, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CCS

[Goo81a] I. J. Good. C100. Can scientific induction be meaningfully questioned? Journal of Statistical Computation and Simulation, 13 (2):154, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CDP

[Goo81b] I. J. Good. C101. Dynamic probability does not contradict reductio ad absurdum. Journal of Statistical Computation and Simulation, 13(2):154–155, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CEPa

[Goo81c] I. J. Good. C102. An error by Peirce concerning “weight of evidence”. Journal of Statistical Computation and Simulation, 13(2):155–157, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CEPb

[Goo81d] I. J. Good. C103. The effect of permutations of rows and columns on measures of association. Journal of Statistical Computation and Simulation, 13(3–4):309–312, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CICa

[Goo81e] I. J. Good. C104. Informational correlations between two distributions. Journal of Statistical Computation and Simulation, 13 (3–4):312–314, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CWS

[Goo81f] I. J. Good. C105. When should marginal cases be dropped in the interests of simplicity? Journal of Statistical Computation and Simulation, 13(3–4):314–315, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CICb

[Goo81g] I. J. Good. C106. International clinical trials. Journal of Statistical Computation and Simulation, 13(3–4):315, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Good:1981:CSCa

[Goo81h] I. J. Good. C107. Some comments about the Sherman–Foutz test. Journal of Statistical Computation and Simulation, 13(3– 4):316–319, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CAC

[Goo81i] I. J. Good. C108. Analogue of chi-squared that is powerful against bumpy alternatives. Journal of Statistical Computation and Simulation, 13(3–4):319–323, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CNM

[Goo81j] I. J. Good. C109. A new Monte Carlo computation for the characteristic function of Pearson’s chi-squared. Journal of Statistical Computation and Simulation, 13(3–4):323–324, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CRC

[Goo81k] I. J. Good. C110. Randomly connected genetic nets. Journal of Statistical Computation and Simulation, 13(3–4):324–327, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CSD

[Goo81l] I. J. Good. C114. On the statistical determination of authorship. Journal of Statistical Computation and Simulation, 14(1):55–60, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CAE

[Goo81m] I. J. Good. C115. Assessment of eye-ball descriptions. Journal of Statistical Computation and Simulation, 14(1):60–63, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CSCb

[Goo81n] I. J. Good. C116. A supplement to chi-squared when testing homogeneity. Journal of Statistical Computation and Simulation, 14(1):63–67, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CMCb

[Goo81o] I. J. Good. C117. A Monte Carlo method that will be impracticable before the year 2001. Journal of Statistical Computation and Simulation, 14(1):67–69, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CCC

[Goo81p] I. J. Good. C84. A comment on C49, “A Boolean error”. Journal of Statistical Computation and Simulation, 12(2):137–138, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo79n].

Good:1981:CRP

[Goo81q] I. J. Good. C87. Roughness penalties, invariant under rotation for multidimensional probability density estimation. Journal of Statistical Computation and Simulation, 12(2):142–144, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CAV

[Goo81r] I. J. Good. C88. An approximation of value in the Bayesian analyses of contingency tables. Journal of Statistical Computation and Simulation, 12(2):145–147, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CTC

[Goo81s] I. J. Good. C89. The total cost of collecting income tax. Journal of Statistical Computation and Simulation, 12(2):147–148, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CAF

[Goo81t] I. J. Good. C90. The autocorrelation functions. Journal of Statistical Computation and Simulation, 12(2):148–152, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CN

[Goo81u] I. J. Good. C92. Neologisms. Journal of Statistical Computation and Simulation, 12(3–4):313, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CWP

[Goo81v] I. J. Good. C94. When is g positive in the mixed Dirichlet approach to contingency tables? Journal of Statistical Computation and Simulation, 13(1):49–52, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CMCa

[Goo81w] I. J. Good. C95. The Monte Carlo computation of Bayes factors for contingency tables. Journal of Statistical Computation and Simulation, 13(1):52–56, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CWE

[Goo81x] I. J. Good. C96. The weight of evidence provided by uncertain testimony or from an uncertain event. Journal of Statistical Computation and Simulation, 13(1):56–60, 1981. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CGD

[Goo81y] I. J. Good. C97. Generalized determinants and generalized Jacobians. Journal of Statistical Computation and Simulation, 13 (1):60–62, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CFP

[Goo81z] I. J. Good. C98. Flexible priors for estimating a normal distribution. Journal of Statistical Computation and Simulation, 13 (2):151–153, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CPS

[Goo81-27] I. J. Good. C99. The probability of a scientific law. Journal of Statistical Computation and Simulation, 13(2):153, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:EA

[Goo81-28] I. J. Good. Errata and annenda. Journal of Statistical Computation and Simulation, 14(1):79, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:FCE

[Goo81-29] I. J. Good. The fast calculation of the exact distribution of Pearson’s chi-squared and of the number of repeats within the cells of a multinomial by using a Fast Fourier Transform. Journal of Statistical Computation and Simulation, 14(1):71–78, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CPN

[Goo82a] I. J. Good. C122. A paradox in the Neyman–Pearson theory. Journal of Statistical Computation and Simulation, 14(3–4):322, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CMS

[Goo82b] I. J. Good. C123. Are the methods of statistical mechanics applicable to society? Journal of Statistical Computation and Simulation, 14(3–4):322–323, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CPD

[Goo82c] I. J. Good. C124. Probability densities versus discriminant functions. Journal of Statistical Computation and Simulation, 14(3– 4):321–323, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CPP

[Goo82d] I. J. Good. C126. Protom: a projection-tomographic method for density estimation from multivariate data. Journal of Statistical Computation and Simulation, 15(1):73–75, 1982. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CWM

[Goo82e] I. J. Good. C127. When is maximum likelihood estimation O.K. for a Bayesian? Journal of Statistical Computation and Simulation, 15(1):75–77, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CISa

[Goo82f] I. J. Good. C129. An index of separateness of clusters and a permutation test for its statistical significance. Journal of Statistical Computation and Simulation, 15(1):81–84, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CBL

[Goo82g] I. J. Good. C130. Bayesian likelihood: a point of terminology. Journal of Statistical Computation and Simulation, 15(1):84– 85, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CMR

[Goo82h] I. J. Good. C133. On the measurement of regret. Journal of Statistical Computation and Simulation, 15(2-3):237–238, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CPI

[Goo82i] I. J. Good. C134. A proposal for the improvement of the refereeing process. Journal of Statistical Computation and Simulation, 15(2-3):238, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CIB

[Goo82j] I. J. Good. C135. Interactions between indexes of separateliess. Journal of Statistical Computation and Simulation, 15(2-3):239– 242, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CISb

[Goo82k] I. J. Good. C136. The irregular shapes of polypeptide chains. Journal of Statistical Computation and Simulation, 15(2-3):243– 247, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CFC

[Goo82l] I. J. Good. C138. The fast calculation of the exact distribution of Pearson’s chi-squared and of the number of repeats within the cells of a multinomial by using a Fast Fourier Transform: an addendum. Journal of Statistical Computation and Simulation, 15(4):336–337, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CSE

[Goo82m] I. J. Good. C139. The standard error of the estimated “coverage” of a sample of species or vocabulary. Journal of Statistical Computation and Simulation, 15(4):337, 1982. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CST

[Goo82n] I. J. Good. C140. Standardized tail-area probabilities. Journal of Statistical Computation and Simulation, 16(1):65–66, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CAP

[Goo82o] I. J. Good. The cumulants of an analogue of Pearson’s chisquared. Journal of Statistical Computation and Simulation, 15 (2-3):171–181, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CSIa

[Goo83a] I. J. Good. C143. Scientific induction: “universal and predictive”. Journal of Statistical Computation and Simulation, 16(3– 4):311–312, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CDS

[Goo83b] I. J. Good. C144. The diminishing significance of a fixed pvalue as the sample size increases: a discrete model. Journal of Statistical Computation and Simulation, 16(3–4):312–3 14, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CMA

[Goo83c] I. J. Good. C145. A measure of adhockery. Journal of Statistical Computation and Simulation, 16(3–4):314, 1983. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CCB

[Goo83d] I. J. Good. C146. Correlation between power functions. Journal of Statistical Computation and Simulation, 16(3–4):314–316, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CPE

[Goo83e] I. J. Good. C148. Probability estimation by maximum penalized likelihood for large contingency tables and for other categorical data. Journal of Statistical Computation and Simulation, 17(1): 66–67, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CSC

[Goo83f] I. J. Good. C149. A simple consequence of Daniel Bernoulli’s logarithmic formula for utility. Journal of Statistical Computation and Simulation, 17(1):67–68, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CDP

[Goo83g] I. J. Good. C150. “Density” a point of terminology. Journal of Statistical Computation and Simulation, 17(1):68, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CA

[Goo83h] I. J. Good. C151. Antisurprise. Journal of Statistical Computation and Simulation, 17(1):69–71, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CND

[Goo83i] I. J. Good. C152. On the number of duplicated fingerprints. Journal of Statistical Computation and Simulation, 17(1):71– 74, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CCA

[Goo83j] I. J. Good. C154. A conjecture about the history of errors of the first and second kinds. Journal of Statistical Computation and Simulation, 17(2):155, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CIA

[Goo83k] I. J. Good. C157. An improved algorithm for the fast calculation of the exact distribution of Pearson’s chi-squared. Journal of Statistical Computation and Simulation, 17(3):236–242, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CNE

[Goo83l] I. J. Good. C158. A note on the extraction of coefficients from power series. Journal of Statistical Computation and Simulation, 17(3):243–244, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CFCa

[Goo83m] I. J. Good. C160. A further comment on multiple-choice examinations. Journal of Statistical Computation and Simulation, 17 (4):311–313, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CWF

[Goo83n] I. J. Good. C161. When are free observations of positive expected value. Journal of Statistical Computation and Simulation, 17(4):313–315, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CMC

[Goo83o] I. J. Good. C162. Moments and cumulants of weights of evidence. Journal of Statistical Computation and Simulation, 17 (4):315–319, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CME

[Goo83p] I. J. Good. C163. The moments of an exponential. Journal of Statistical Computation and Simulation, 17(4):320–321, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CCL

[Goo83q] I. J. Good. C164. The cumulants of the lognormal distribution, including some conjectures. Journal of Statistical Computation and Simulation, 17(4):321–328, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CCCa

[Goo83r] I. J. Good. C165. A correction concerning my interpretation of Peirce, and the Bayesian interpretation of Neyman–Pearson “hypothesis determination. Journal of Statistical Computation and Simulation, 18(1):71–84, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CMW

[Goo83s] I. J. Good. C166. The moments of weight of evidence gives intermediate hypotheses. Journal of Statistical Computation and Simulation, 18(1):74–79, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CPO

[Goo83t] I. J. Good. C167. The probability of odd partial quotients in continued fractions. Journal of Statistical Computation and Simulation, 18(1):79–80, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CFD

[Goo83u] I. J. Good. C168. Fractional dimensions, continued fractions, and Markovity. Journal of Statistical Computation and Simulation, 18(1):80–84, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See erratum [Goo83-27].

Good:1983:CBNa

[Goo83v] I. J. Good. C169. Barndorff–Nielsen’s plausibility function and a principle of least surprise. Journal of Statistical Computation and Simulation, 18(2-3):215–218, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CPM

[Goo83w] I. J. Good. C170. A potential misstatement of Bayes’s theorem. Journal of Statistical Computation and Simulation, 18(2-3):218– 219, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CFCb

[Goo83x] I. J. Good. C171. Further combinatorial identities related to the cumulants of a lognormal distribution. Journal of Statistical Computation and Simulation, 18(2-3):219–226, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CPR

[Goo83y] I. J. Good. C172. Probabilities of residues for partial quotients of continued fractions, and a question about moments. Journal of Statistical Computation and Simulation, 18(2-3):226–227, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See erratum [Goo89o].

Good:1983:CGD

[Goo83z] I. J. Good. C173. A geometry of distributions. Journal of Statistical Computation and Simulation, 18(2-3):227–229, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CSIb

[Goo83-27] I. J. Good. C174. Scientific induction, Jeffreys’s and my approaches compared. Journal of Statistical Computation and Simulation, 18(2-3):229–231, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comments [Goo83-28, Goo83-30].

Good:1983:CHB

[Goo83-28] I. J. Good. C175. Hierarchical Bayes, mixed Dirichlet priors, penalized likelihood, etc. (see the first paragraph of C174.). Journal of Statistical Computation and Simulation, 18(2-3):231–234, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo83-27]. This note was my response to comments by Tom Leonard.

Good:1983:CBNb

[Goo83-29] I. J. Good. C176. The Bayes/non-Bayes compromise or synthesis and Box’s current philosophy (see the first paragraph of C74.). Journal of Statistical Computation and Simulation, 18(2-3):234– 236, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo80i].

Good:1983:CDI

[Goo83-30] I. J. Good. C177. The device of imaginary results, “WHAT IF”, and preposterior analysis (see the first paragraph of C174). Journal of Statistical Computation and Simulation, 18(2-3):236, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo83-27].

Good:1983:CTT

[Goo83-31] I. J. Good. C178. Tail trouble when sampling. Journal of Statistical Computation and Simulation, 18(2-3):237, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Good:1984:CRR

[Goo84a] I. J. Good. C186. The repeat rate with respect to a given multinomial population, and the Dirichlet–multinomial, distribution. Journal of Statistical Computation and Simulation, 19(1):85– 89, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CLE

[Goo84b] I. J. Good. C187. Laplace’s estimate of the mass of Jupiter. Journal of Statistical Computation and Simulation, 19(1):89– 90, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CCD

[Goo84c] I. J. Good. C189. A commonsense design for multiple comparisons. Journal of Statistical Computation and Simulation, 19(1): 94, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CHS

[Goo84d] I. J. Good. C191. How should tail-area probabilities be standardized for sample size in unpaired comparisons? Journal of Statistical Computation and Simulation, 19(2):174, 1984. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:COT

[Goo84e] I. J. Good. C192. One tail versus two-tails, and the harmonicmean rule of thumb. Journal of Statistical Computation and Simulation, 19(2):174–176, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CPV

[Goo84f] I. J. Good. C193. Paired versus unpaired comparisons and the Hakvunic-mean rule of thumb. Journal of Statistical Computation and Simulation, 19(2):176–177, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CEC

[Goo84g] I. J. Good. C195. Editorial comment on Kempthorne’s C194. Journal of Statistical Computation and Simulation, 19(2):182– 184, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CBE

[Goo84h] I. J. Good. C197. The best explicatum for weight of evidence. Journal of Statistical Computation and Simulation, 19(4):294– 299, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CPT

[Goo84i] I. J. Good. C198. A probability too small to achieve using a computer. Journal of Statistical Computation and Simulation, 19(4):299–300, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CST

[Goo84j] I. J. Good. C199. Standardized tail-area probabilities: a possible interpretation. Journal of Statistical Computation and Simulation, 19(4):299–300, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CTB

[Goo84k] I. J. Good. C200. The tolerant Bayesian’s interpretation of a tail-area probability. Journal of Statistical Computation and Simulation, 19(4):300–302, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CBI

[Goo84l] I. J. Good. C201. A Bayesian interpretation of ancillarity. Journal of Statistical Computation and Simulation, 19(4):302–308, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CPA

[Goo84m] I. J. Good. C202. Probability applied to Diophantine equations. Journal of Statistical Computation and Simulation, 19(4):308– 313, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See errata [Ano90a].

Good:1984:CQD

[Goo84n] I. J. Good. C203. A quick and dirty method for nonparametric estimation of a density function. Journal of Statistical Computation and Simulation, 19(4):313–314, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CEH

[Goo84o] I. J. Good. C204. The early history of the Fisher–Yates–Irwin formula and Fisher’s ’exact test’. Journal of Statistical Computation and Simulation, 19(4):315–319, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CMP

[Goo84p] I. J. Good. C205. Monotonic properties of the moments of a Bayes factor and the relationship to measures of divergence. Journal of Statistical Computation and Simulation, 19(4):320– 325, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CFN

[Goo84q] I. J. Good. C208. A further note on the early history of the Fisher–Yates–Irwin formula. Journal of Statistical Computation and Simulation, 20(2):155–159, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CEN

[Goo84r] I. J. Good. C209. An error by Neyman noticed by Dickey. Journal of Statistical Computation and Simulation, 20(2):159–160, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:COJ

[Goo84s] I. J. Good. C210. The only-just effect in nonparametric procedures, and the achievement of objectivity by using a statistician’s stooge. Journal of Statistical Computation and Simulation, 20(2):160–164, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CSH

[Goo84t] I. J. Good. C213. A sharpening of the harmonic-mean rule of thumb for combining tests “in parallel”. Journal of Statistical Computation and Simulation, 20(2):173–176, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CHM

[Goo84u] I. J. Good. C214. The harmonic-mean rule of thumb: some classes of applications. Journal of Statistical Computation and Simulation, 20(2):176–179, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CPL

[Goo84v] I. J. Good. C215. Penalized least surprise. Journal of Statistical Computation and Simulation, 20(2):179–180, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CCC

[Goo84w] I. J. Good. Comment concerning the conjectured value of a determinant occurring in statistics. Journal of Statistical Computation and Simulation, 19(1):95, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CWE

[Goo84x] J. J. Good. C190. Weights of evidence in medical diagnosis. Journal of Statistical Computation and Simulation, 19(2):171– 173, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CPI

[Goo85a] I. J. Good. C216. Probabilistic induction is inevitable. Journal of Statistical Computation and Simulation, 20(4):323–324, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CTA

[Goo85b] I. J. Good. C217. Tail-area probabilities and Bayes factors as “distances” from the null hypothesis. Journal of Statistical Computation and Simulation, 20(4):325, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CCC

[Goo85c] I. J. Good. C218. A correction concerning degrees of causation. Journal of Statistical Computation and Simulation, 20(4):326, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CNN

[Goo85d] I. J. Good. C220. The number of numerical multiplications when multiplying polynomials by recursive bisection. Journal of Statistical Computation and Simulation, 20(4):329–331, 1985. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CAS

[Goo85e] I. J. Good. C223. The accuracy of Stirling’s formula for x! when x is small. Journal of Statistical Computation and Simulation, 21(1):84–86, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CHN

[Goo85f] I. J. Good. C224. A historical note regarding the mass of Jupiter. Journal of Statistical Computation and Simulation, 21(1):86–87, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CWD

[Goo85g] I. J. Good. C225. Who discovered the God–Koheleth law? Journal of Statistical Computation and Simulation, 21(1):87– 88, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CNM

[Goo85h] I. J. Good. C226. A new measure of surprise. Journal of Statistical Computation and Simulation, 21(1):88–89, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CHC

[Goo85i] I. J. Good. C227. A historical comment concerning the index of separateness. Journal of Statistical Computation and Simulation, 21(1):89–90, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CNF

[Goo85j] I. J. Good. C228. The number of functionals: a combinatorial curiosity. Journal of Statistical Computation and Simulation, 21 (1):90, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CPDa

[Goo85k] I. J. Good. C229. Probability density estimation over subintervals of the range. Journal of Statistical Computation and Simulation, 21(1):90–91, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CSW

[Goo85l] I. J. Good. C232. Sufftctmcy and weight of evidence. Journal of Statistical Computation and Simulation, 21(3–4):334–336, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CFS

[Goo85m] I. J. Good. C233. Fisher and subjectivity. Journal of Statistical Computation and Simulation, 21(3–4):336–337, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CDL

[Goo85n] I. J. Good. C234. In defense of the law of the excluded middle, or all logic is probabilistic. Journal of Statistical Computation and Simulation, 21(3–4):337–339, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CAA

[Goo85o] I. J. Good. C235. Another application of the Wilcoxon signedrank test. Journal of Statistical Computation and Simulation, 21(3–4):339, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CMN

[Goo85p] I. J. Good. C236. Mathematically natural generalizations of some measures of association and dependence for contingency tables. Journal of Statistical Computation and Simulation, 22 (1):83–91, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CEP

[Goo85q] I. J. Good. C238. Explanatory power depends on more than probabilities. Journal of Statistical Computation and Simulation, 22(2):184–186, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CED

[Goo85r] I. J. Good. C241. The exact distribution of a statistic for matched n-tuples allowing for ties. Journal of Statistical Computation and Simulation, 23(1–2):142–146, 1985. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CPDb

[Goo85s] I. J. Good. C244. The probability distribution of the gaps between the Mersenne primes, II. Journal of Statistical Computation and Simulation, 23(1–2):154–156, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comment [Goo86b].

Good:1986:CVS

[Goo86a] I. J. Good. C246. Very small tails of the t distribution, and significance tests for clustering. Journal of Statistical Computation and Simulation, 23(3):243–248, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CCCa

[Goo86b] I. J. Good. C248. Comment on C244. Journal of Statistical Computation and Simulation, 23(3):250, 1986. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo85s].

Good:1986:CSM

[Goo86c] I. J. Good. C250. “Saturated model” or “quasimodel”, a point of terminology. Journal of Statistical Computation and Simulation, 24(2):168–169, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CSB

[Goo86d] I. J. Good. C260. Some Bayes/non-Bayes compromises based on covert Bayesianism. Journal of Statistical Computation and Simulation, 25(3–4):298–300, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CSS

[Goo86e] I. J. Good. C261. The species-sam pling problem again. Journal of Statistical Computation and Simulation, 25(3–4):301–304, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CGA

[Goo86f] I. J. Good. C264. A geometrical approach to the amalgamation paradox. Journal of Statistical Computation and Simulation, 26 (1–2):129–132, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CCI

[Goo86g] I. J. Good. C265. The computer-intensive form of a Bayes/ non-Bayes compromise. Journal of Statistical Computation and Simulation, 26(1–2):132–133, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CRP

[Goo86h] I. J. Good. C266. Recognizability, permutability, and homogeneity. Journal of Statistical Computation and Simulation, 26 (1–2):133–134, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CME

[Goo86i] I. J. Good. C267. The 3d monocular effect. Journal of Statistical Computation and Simulation, 26(1–2):134–139, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CLR

[Goo86j] I. J. Good. C268. The likelihood-ratio approach to the generalized Behrens–Fisher problem: one-way Anova under heteroscedasticity. Journal of Statistical Computation and Simulation, 26(1–2):139–145, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CND

[Goo86k] I. J. Good. C269. The newly discovered poem, possibly by Shakespeare. Journal of Statistical Computation and Simulation, 26(1–2):145–146, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CNP

[Goo86l] I. J. Good. C270. Newcomb’s problem and the foundations of probability. Journal of Statistical Computation and Simulation, 26(1–2):127–148, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CAE

[Goo86m] I. J. Good. C271. Age of earliest memory: an informal sampling experiment. Journal of Statistical Computation and Simulation, 26(3–4):299–301, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CFB

[Goo86n] I. J. Good. C272. A flexible Bayesian model for comparing two treatments. Journal of Statistical Computation and Simulation, 26(3–4):301–305, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CHC

[Goo86o] I. J. Good. C273. A historical comment concerning the Monte Carlo estimation of p-values. Journal of Statistical Computation and Simulation, 26(3–4):306, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CMP

[Goo86p] I. J. Good. C274. Multinomial prediction and improved approximations to the shrinking parameter that maximizes a Dirichlet– multinomial likelihood. Journal of Statistical Computation and Simulation, 26(3–4):306–311, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CMB

[Goo86q] I. J. Good. C275. The maximum of a Bayes factor against “independence” in a contingency table, and generalizations to higher dimensions. Journal of Statistical Computation and Simulation, 26(3–4):312–316, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CTM

[Goo86r] I. J. Good. C276. On terminology for matrices of special forms. Journal of Statistical Computation and Simulation, 26(3–4):316– 317, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CCCb

[Goo86s] I. J. Good. Comments, conjectures, and conclusions C258. Editorial note on C257 regarding the t-test. Journal of Statistical Computation and Simulation, 25(3–4):296–297, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [LS86].

Good:1987:CSC

[Goo87a] I. J. Good. C277. Some checks on the consistency of probability judgments and a conjectured method for improving those judgments. Journal of Statistical Computation and Simulation, 28(1):63–71, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:CIE

[Goo87b] I. J. Good. C278. Indices of economic concentration. Journal of Statistical Computation and Simulation, 28(1):71–72, 1987. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:CCN

[Goo87c] I. J. Good. C283. The calculation of the normal or error function with normally acceptable proportional error. Journal of Statistical Computation and Simulation, 28(1):83–84, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:CC

[Goo87d] I. J. Good. C284. Confidence. Journal of Statistical Computation and Simulation, 28(1):85–87, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:CSU

[Goo87e] I. J. Good. C285. A survey of the use of the Fast Fourier Transform for computing distributions. Journal of Statistical Computation and Simulation, 28(1):87–93, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:CCL

[Goo87f] I. J. Good. C286. Correlation and linearity: some lapses in exposition. Journal of Statistical Computation and Simulation, 28(1):93–94, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:CCC

[Goo87g] I. J. Good. C288. A correction concerning ’ancillarity’ for significance tests. Journal of Statistical Computation and Simulation, 28(2):182–183, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:CMP

[Goo87h] I. J. Good. C289. Maximum penalized likelihood for somewhat parametric estimation. Journal of Statistical Computation and Simulation, 28(3):261–264, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:CED

[Goo87i] I. J. Good. C294. An example of dynamic probability in number theory. Journal of Statistical Computation and Simulation, 28 (4):349–350, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1988:CBR

[Goo88a] I. J. Good. C297. Bayes’s red billiard ball is also a herring, and why Bayes withheld publication. Journal of Statistical Computation and Simulation, 29(4):335–340, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1988:CIS

[Goo88b] I. J. Good. C299. An inequality for the smallest singular value of a matrix products. Journal of Statistical Computation and Simulation, 29(4):343–346, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1988:CWS

[Goo88c] I. J. Good. C300. What should be the usual shape of a loss function? Journal of Statistical Computation and Simulation, 29(4):346–347, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1988:CER

[Goo88d] I. J. Good. C302. An early risk assessment for an atomic reactor. Journal of Statistical Computation and Simulation, 29(4):351, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1988:CPI

[Goo88e] I. J. Good. C303. The pragmatic irrelevance of De Finetti’s theorem. Journal of Statistical Computation and Simulation, 29 (4):351–353, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See apology [Goo89r].

Good:1988:CNC

[Goo88f] I. J. Good. C305. The number of correct significant digits in a piece of physical numerology. Journal of Statistical Computation and Simulation, 29(4):358–359, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See errata [Ano90a].

Good:1988:EN

[Goo88g] Phillip I. Good. Editor’s note. Journal of Statistical Computation and Simulation, 29(4):333, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1988:SSP

[Goo88h] Phillip I. Good. Statistics software for the professional statistician. Journal of Statistical Computation and Simulation, 30 (2):149–156, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CCP

[Goo89a] I. J. Good. C311. On the combination of pieces of evidence. Journal of Statistical Computation and Simulation, 31(1):54– 58, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See errata [Ano90a].

Good:1989:CYA

[Goo89b] I. J. Good. C312. Yet another argument for the explicatum of weight of evidence. Journal of Statistical Computation and Simulation, 31(1):58–59, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CRD

[Goo89c] I. J. Good. C313. The rate of diagnosing a fault, for example, in a communication channel. Journal of Statistical Computation and Simulation, 31(1):60–61, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CCN

[Goo89d] I. J. Good. C314. Chaotics and number theory. Journal of Statistical Computation and Simulation, 31(1):61–62, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CSF

[Goo89e] I. J. Good. C315. A suspicious feature of the Popper/Miller argument. Journal of Statistical Computation and Simulation, 31(1):62–63, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CIL

[Goo89f] I. J. Good. C316. Interpretation of a large coefficient of determination. Journal of Statistical Computation and Simulation, 31(1):63–64, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CTC

[Goo89g] I. J. Good. C317. The theorem of corroboration and undermining, and Popper’s demarcation rule. Journal of Statistical Computation and Simulation, 31(2):119–120, 1989. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CPC

[Goo89h] I. J. Good. C318. The paradox of confirmation and stoogian observations. Journal of Statistical Computation and Simulation, 31(2):121, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CWE

[Goo89i] I. J. Good. C319. Weight of evidence and a compelling metaprinciple. Journal of Statistical Computation and Simulation, 31(2): 121–123, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CWA

[Goo89j] I. J. Good. C320. Weak attacks on transitivity of preference. Journal of Statistical Computation and Simulation, 31(2):123– 124, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CNN

[Goo89k] I. J. Good. C321. The need for a normative theory of rationality. Journal of Statistical Computation and Simulation, 31(2):124– 126, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CMV

[Goo89l] I. J. Good. C322. Medialized p-values. Journal of Statistical Computation and Simulation, 31(2):126–127, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CTS

[Goo89m] I. J. Good. C323. A test for a specified Markov model or branching process. Journal of Statistical Computation and Simulation, 31(2):127–128, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CEC

[Goo89n] I. J. Good. C324. Erratum to C168 on fractional dimensions, continued fractions and Markovity. Journal of Statistical Computation and Simulation, 31(2):128, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo83u].

Good:1989:CEA

[Goo89o] I. J. Good. C325. Erratum and addendum to C172. Journal of Statistical Computation and Simulation, 31(2):128–129, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo83y].

Good:1989:CAFa

[Goo89p] I. J. Good. C327. Addenda for “The fractional dimensional theory of continued fractions”, Proc. Cambridge Philos. Soc. 37 (1941), 199–228. Journal of Statistical Computation and Simulation, 32(1–2):64–68, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CVE

[Goo89q] I. J. Good. C330. The variability of evaluations of speculative ideas. Journal of Statistical Computation and Simulation, 32 (1–2):81–88, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CAFb

[Goo89r] I. J. Good. C331. An apology to De Finetti. Journal of Statistical Computation and Simulation, 32(1–2):88–90, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo88e].

Good:1989:CSI

[Goo89s] I. J. Good. C332. Surprise indexes and p-values. Journal of Statistical Computation and Simulation, 32(1–2):90–92, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CJN

[Goo89t] I. J. Good. C333. On the judgement of “numerological” assertions. Journal of Statistical Computation and Simulation, 32 (1–2):92–103, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CNP

[Goo89u] I. J. Good. C334. On the neutron and proton masses: a numerological case study. Journal of Statistical Computation and Simulation, 32(1–2):103–111, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See errata [Ano90a].

Good:1989:CSN

[Goo89v] I. J. Good. C335. Some numerology for the mass ratio of proton to electron, and a partially Bayesian evaluation. Journal of Statistical Computation and Simulation, 32(1–2):111–115, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:EDS

[Goo89w] Phillip Good. Editorial developing statistical software II: Choosing the statistic. AI in statistical software. Journal of Statistical Computation and Simulation, 31(2):71–72, 1989. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CVT

[Goo90a] I. J. Good. C342.5. The p-value trap, an editorial comment on C342. Journal of Statistical Computation and Simulation, 34(2-3):162, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Joh90b].

Good:1990:CUD

[Goo90b] I. J. Good. C344. The use of divergent series in statistical problems. Journal of Statistical Computation and Simulation, 34 (2-3):164, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CRR

[Goo90c] I. J. Good. C347. Random-response questionnaires and liedetectors. Journal of Statistical Computation and Simulation, 34 (2-3):166–168, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CMT

[Goo90d] I. J. Good. C356. Multivariate tests after dimensionality reduction. Journal of Statistical Computation and Simulation, 36(1): 45, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CI

[Goo90e] I. J. Good. C357. Isolikelihoods. Journal of Statistical Computation and Simulation, 36(1):45–47, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CFD

[Goo90f] I. J. Good. C359. Factorial designs and the programming of games. Journal of Statistical Computation and Simulation, 36 (1):49–50, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CEC

[Goo90g] I. J. Good. C361. Editorial concerning Terrell’s conjecture. Journal of Statistical Computation and Simulation, 36(1):53– 54, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CBF

[Goo90h] I. J. Good. C363. A Bessel-function inequality derived by a statistical argument. Journal of Statistical Computation and Simulation, 36(1):56–57, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CMV

[Goo90i] I. J. Good. C367. Medialized p-values, an earlier reference. Journal of Statistical Computation and Simulation, 36(2-3):177, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CEDa

[Goo90j] I. J. Good. C368. On the exact distribution of Pearson’s χ^{2} for the lady tasting beer. Journal of Statistical Computation and Simulation, 36(2-3):177–179, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CID

[Goo90k] I. J. Good. C369. Invariant distance in statistics: some miscellaneous comments. Journal of Statistical Computation and Simulation, 36(2-3):179–186, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CCB

[Goo90l] I. J. Good. C370. A compromise between credibility and subjective probability. Journal of Statistical Computation and Simulation, 36(2-3):186–193, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CFY

[Goo90m] I. J. Good. C372. Fifty years of errors. Journal of Statistical Computation and Simulation, 36(2-3):193–194, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CEDb

[Goo90n] I. J. Good. C377. Exploratory data analysis. Journal of Statistical Computation and Simulation, 37(3–4):243–245, 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CCC

[Goo90o] I. J. Good. C378. Comments concerning the hadron quantal hypothesis. Journal of Statistical Computation and Simulation, 37 (3–4):245–247, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1991:CCC

[Goo91] I. J. Good. C383. A comment concerning optional stopping. Journal of Statistical Computation and Simulation, 39(3):191– 192, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1992:CRR

[Goo92a] I. J. Good. C386. Ridge regression and the harmonic-mean rule of thumb. Journal of Statistical Computation and Simulation, 40(3–4):262, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1992:CFC

[Goo92b] I. J. Good. C387. Further comments concerning the lady tasting tea or beer: p-values and restricted randomization. Journal of Statistical Computation and Simulation, 40(3–4):263–267, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1992:CFN

[Goo92c] I. J. Good. C388. A further note on physical numerology. Journal of Statistical Computation and Simulation, 40(3–4):267–268, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1992:CPCa

[Goo92d] I. J. Good. C389. Partial correlation and spherical trigonometry, II. Journal of Statistical Computation and Simulation, 43(3–4): 253, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1992:CPCb

[Goo92e] I. J. Good. C390. Partial correlation coefficients are analogous to partial differential coefficients. Journal of Statistical Computation and Simulation, 43(3–4):254–256, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1993:CTS

[Goo93a] I. J. Good. C394. Tendencies to be sufficient or necessary causes. Journal of Statistical Computation and Simulation, 44(3–4):261– 262, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1993:CRR

[Goo93b] I. J. Good. C397. Refutation and rejection versus inexactification, and other comments concerning terminology. Journal of Statistical Computation and Simulation, 47(1–2):91–92, 1993.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1993:CVL

[Goo93c] I. J. Good. C398. P -values, luck, and surprise. Journal of Statistical Computation and Simulation, 47(1–2):92–94, 1993. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1993:CTF

[Goo93d] I. J. Good. C399. Trigonometrical factorization of the Fibonacci and Lucas numbers, and sums of powers of cotangents. Journal of Statistical Computation and Simulation, 47(1–2):94–99, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1993:CTM

[Goo93e] I. J. Good. C400. A tentative measure of probabilistic causation relevant to the philosophy of the law. Journal of Statistical Computation and Simulation, 47(1–2):99–105, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CCE

[Goo94a] I. J. Good. C407. A comment on the evaluation of numerology. Journal of Statistical Computation and Simulation, 49(3–4):216– 217, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CFA

[Goo94b] I. J. Good. C415. A fourth argument for the explication of causal tendency. Journal of Statistical Computation and Simulation, 49 (3–4):235–237, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CFL

[Goo94c] I. J. Good. C416. Fibonacci and Lucas numbers and trigonometry. Journal of Statistical Computation and Simulation, 49(3–4): 237, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CFE

[Goo94d] I. J. Good. C417. The flybutter effect: the obverse of the butterfly effect. Journal of Statistical Computation and Simulation, 49 (3–4):237–238, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CFP

[Goo94e] I. J. Good. C418. Fuzzy probability and hierarchical Bayes. Journal of Statistical Computation and Simulation, 49(3–4):238– 239, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CKT

[Goo94f] I. J. Good. C419. A key to the topic-independent philosophy of science. Journal of Statistical Computation and Simulation, 49 (3–4):239–240, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CES

[Goo94g] I. J. Good. C420. The existence of sharp null hypotheses. Journal of Statistical Computation and Simulation, 49(3–4):241–242, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CTL

[Goo94h] I. J. Good. C421. Turing’s little theorem is not really paradoxical. Journal of Statistical Computation and Simulation, 49 (3–4):242–244, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CCW

[Goo94i] I. J. Good. C423. A comment on the Wilcoxon sign-rank test and a conjecture concerning hypothyroidism. Journal of Statistical Computation and Simulation, 49(3–4):249–250, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CST

[Goo94j] I. J. Good. C424. Significance tests: are they more important in science than in technology? Journal of Statistical Computation and Simulation, 49(3–4):250–251, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CRV

[Goo94k] I. J. Good. C426. Rationality versus minimax. Journal of Statistical Computation and Simulation, 49(3–4):253, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1995:CCC

[Goo95a] I. J. Good. C435. A comment concerning physical numerology. Journal of Statistical Computation and Simulation, 52(2):190– 191, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1995:CAR

[Goo95b] I. J. Good. C436. An apology related to blurred hypotheses. Journal of Statistical Computation and Simulation, 52(2):191– 192, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1995:CHR

[Goo95c] I. J. Good. C437. Haphazardness is not randomness, and the defeat of Hitler. Journal of Statistical Computation and Simulation, 52(2):192, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1995:CRA

[Goo95d] I. J. Good. C438. Reliability always depends on probability of course. Journal of Statistical Computation and Simulation, 52 (2):192–193, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1995:CDD

[Goo95e] I. J. Good. C439. Dual density functions. Journal of Statistical Computation and Simulation, 52(2):193–194, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1996:CCI

[Goo96a] I. J. Good. C443. Can the idea of the QH test for normality be used for testing the Weibull distribution? Journal of Statistical Computation and Simulation, 55(3):257–263, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1996:CRC

[Goo96b] I. J. Good. C445. A reply to C444. Journal of Statistical Computation and Simulation, 55(3):265, 1996. CODEN JSCSAJ.

ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [BH96].

Good:1998:CEE

[Goo98a] I. J. Good. C446. The excess of even numbers of carbon atoms in organic compounds, and Eddington’s fisherman’s fallacy: dogs, hot dogs and astronomers. Journal of Statistical Computation and Simulation, 60(2):157–160, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1998:CCC

[Goo98b] I. J. Good. Comments, conjectures and conclusions. Journal of Statistical Computation and Simulation, 61(4):287–312, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:2000:TAE

[Goo00] I. J. Good. Turing’s anticipation of empirical Bayes in connection with the cryptanalysis of the Naval Enigma. Journal of Statistical Computation and Simulation, 66(2):101–111, 2000. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:2003:CSI

[Goo03] I. J. Good. C473. Straddling the interquantile range. Journal of Statistical Computation and Simulation, 73(8):612–616, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:2006:CCCa

[Goo06a] I. J. Good. Comments, conjectures and conclusions. Journal of Statistical Computation and Simulation, 76(8):737–740, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:2006:CCCb

[Goo06b] I. J. Good. Comments, conjectures and conclusions. Journal of Statistical Computation and Simulation, 76(11):1027–1030, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Giulietti:2009:TSH

[GOS09] Monica Giulietti, Jesús Otero, and Jeremy Smith. Testing for stationarity in heterogeneous panel data in the presence of crosssection dependence. Journal of Statistical Computation and Simulation, 79(2):195–203, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Goutis:1993:REB

[Gou93] Constantinos Goutis. Recovering extra-binomial variation. Journal of Statistical Computation and Simulation, 45(3–4):233–242, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Graves:1995:PTA

[GP95] Thomas S. Graves and James L. Pazdan. A permutation test analogue to Tarone’s test for trend in survival analysis. Journal of Statistical Computation and Simulation, 53(1–2):79–89, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gomes:2007:SSO

[GP07] M. Ivette Gomes and Dinis Pestana. A simple second-order reduced bias’ tail index estimator. Journal of Statistical Computation and Simulation, 77(6):487–502, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gardner:1980:MCE

[GPW80] R. H. Gardner, J. E. Pinder III, and R. S. Wood. Monte Carlo estimation of percentiles for the multi-Smirnov test. Journal of Statistical Computation and Simulation, 10(3–4):243–249, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ghoral:1979:CPM

[GR79a] Jugal Ghoral and Herman Rubin. Computational procedure for maximum penalized likelihood estimate. Journal of Statistical Computation and Simulation, 10(1):65–78, 1979. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gupta:1979:ECP

[GR79b] A. K. Gupta and V. K. Rohatgi. Estimation of the common positive mean of a multivariate population. Journal of Statistical Computation and Simulation, 8(3–4):253–265, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gordon:1983:CDF

[GR83] N. H. Gordon and P. F. Ramig. Cumulative distribution function of the sum of correlated chi- squared random variables the sum of correlated chi- squared random variables. Journal of Statistical Computation and Simulation, 17(1):1–9, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gadre:2008:MAU

[GR08] M. P. Gadre and R. N. Rattihalli. Multi-attribute unit and group runs control chart to identify the process deterioration. Journal of Statistical Computation and Simulation, 78(11):999– 1008, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gray:1986:SGA

[Gra86] J. Brian Gray. A simple graphic for assessing influence in regression. Journal of Statistical Computation and Simulation, 24 (2):121–134, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gray:1988:CIM

[Gra88] J. Brian Gray. A classification of influence measures. Journal of Statistical Computation and Simulation, 30(3):159–171, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Grier:1992:ESO

[Gri92] David Alan Grier. An extended sweep operator for the cross validation of variable selection in linear regression. Journal of Statistical Computation and Simulation, 43(1–2):117–126, 1992. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Grigoletto:2002:WTH

[Gri02] Matteo Grigoletto. Weighted transformed hazard Anova for censored and truncated data. Journal of Statistical Computation and Simulation, 72(9):759–773, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Griffith:2004:FML

[Gri04] Daniel A. Griffith. Faster maximum likelihood estimation of very large spatial autoregressive models: an extension of the Smirnov–Anselin result. Journal of Statistical Computation and Simulation, 74(12):855–866, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Grillenzoni:2009:RNP

[Gri09] Carlo Grillenzoni. Robust non-parametric smoothing of nonstationary time series. Journal of Statistical Computation and Simulation, 79(4):379–393, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gomes:2008:HAC

[GRVV08] M. Ivette Gomes, Lígia Rodrigues, Björn Vandewalle, and Clara Viseu. A heuristic adaptive choice of the threshold for biascorrected Hill estimators. Journal of Statistical Computation and Simulation, 78(2):133–150, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Golub:1973:NCU

[GS73] Gene H. Golub and George P. H. Styan. Numerical computations for univariate linear models. Journal of Statistical Computation and Simulation, 2(3):253–274, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gordon:1978:GSM

[GS78] Gilbert R. Gordon and Julius Surkis. A general simulation methodology for optical bar code evaluation. Journal of Statistical Computation and Simulation, 8(2):85–101, 1978. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CEH

[GS83] I. J. Good and Eric P. Smith. C155. The early history of the quadratic index of diversity or repeat rate. Journal of Statistical Computation and Simulation, 17(3):231–233, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CPB

[GS84] I. J. Good and Eric P. Smith. C184. The possible bias of the Pearson chi-squared test in non-equiprobable cases. Journal of Statistical Computation and Simulation, 19(1):79–82, 1984. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CCP

[GS85] I. J. Good and Eric P. Smith. C240. A continuum of principal component methods. Journal of Statistical Computation and Simulation, 23(1–2):136–142, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See correction [Mar90b].

Good:1986:CPS

[GS86] I. J. Good and Eric P. Smith. C247. A power series for the tail-area probability of Student’s t distribution. Journal of Statistical Computation and Simulation, 23(3):248–250, 1986. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:CSG

[GS87] I. J. Good and Eric P. Smith. C280. Some general measures of similarity and their values under multinormality. Journal of Statistical Computation and Simulation, 28(1):75–79, 1987. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Grabarnik:2001:INP

[GS01] Pavel Grabarnik and Aila Särkkä. Interacting neighbour point processes: Some models for clustering. Journal of Statistical Computation and Simulation, 68(2):103–125, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gargallo:2007:MCE

[GS07] P. Gargallo and M. Salvador. Monitoring changes in exponential family models: a two-sided Bayesian decision approach. Journal of Statistical Computation and Simulation, 77(9):781–796, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Goodlow:1987:SSC

[GSC87] Janis L. Goodlow, Donald M. Stablein, and Walter H. Carter, Jr. Sample sizes for cancer therapy survival studies based on response surface methods. Journal of Statistical Computation and Simulation, 28(1):31–41, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gonzales:1978:EAA

[GSF78] Teofilo Gonzales, Sartaj Sahni, and W. R. Franta. An efficient approximate algorithm for the Kolmogorov–Smirnov and Lilliefors tests. Journal of Statistical Computation and Simulation, 6(3–4):257–263, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gu:2004:UPP

[GSH04] Ming Gao Gu, Liuquan Sun, and Changquan Huang. A universal procedure for parametric frailty models. Journal of Statistical Computation and Simulation, 74(1):1–13, 2004. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Groggel:1987:EWC

[GSS87] David J. Groggel, Robert L. Schaefer, and John H. Skillings. The effect of withdrawal censoring on two-sample tests. Journal of Statistical Computation and Simulation, 28(4):275–302, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gang:2001:TNM

[GsW01] Li Gang, Peter Y. S. shao, and Yining Wang. Tests of a normal mean vector with hypotheses determined by linear inequalities. Journal of Statistical Computation and Simulation, 70(4):289– 298, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gupta:1978:HPT

[GT78] A. K. Gupta and D. S. Tracy. Hermite polynomials and truncated trivariate normal distributions. Journal of Statistical Computation and Simulation, 7(3–4):269–286, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CRI

[GT81] I. J. Good and T. N. Tideman. C93. The relevance of imaginary alternatives. Journal of Statistical Computation and Simulation, 12(3–4):313–315, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CJE

[GT85] I. J. Good and Gordon Tullock. C221. On judicial errors. Journal of Statistical Computation and Simulation, 21(1):75–76, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Geurts:1990:CPS

[GT90] Michael D. Geurts and H. Dennis Tolley. Causal partitioning and sales forecasting accuracy. Journal of Statistical Computation and Simulation, 37(1–2):1–12, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Guenther:1978:EPN

[Gue78] William C. Guenther. Evaluation of probabilities for the noncentral distributions and the difference of two T -variables with a desk calculator. Journal of Statistical Computation and Simulation, 6(3–4):199–206, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Guenther:1982:CTF

[Gue82] William C. Guenther. C132. Trigonometric functions and combinations from probability distributions. Journal of Statistical Computation and Simulation, 15(2-3):233–237, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Guerard:1989:ATS

[Gue89] John B. Guerard, Jr. Automatic time series modeling, intervention analysis, and effective forecasting. Journal of Statistical Computation and Simulation, 34(1):43–49, 1989. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gupta:1973:TRC

[Gup73] A. K. Gupta. On a test for reality of the covariance matrix in a complex Gaussian distribution. Journal of Statistical Computation and Simulation, 2(4):333–342, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Guptat:1984:ACN

[Gup84] S. C. Guptat. An algorithm for constructing nests of 2 factor designs with orthogonal factorial structure. Journal of Statistical Computation and Simulation, 20(1):59–79, 1984. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gravel:1981:BLU

[GV81] Raymond Gravel and Constance Van Eeden. Best linear unbiased estimators of the location of a double quadratic distribution based on order statistics. Journal of Statistical Computation and Simulation, 13(3–4):225–243, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Giummole:2002:PPE

[GV02] Federica Giummole and Laura Ventura. Practical point estimation from higher-order pivots. Journal of Statistical Computation and Simulation, 72(5):419–430, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gunst:1973:DFB

[GW73] Richard F. Gunst and John T. Webster. Density functions of the bivariate chi-square distribution. Journal of Statistical Computation and Simulation, 2(3):275–288, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Geurts:1994:PTS

[GW94] Michael Geurts and David Whitlark. Partioning a time series to increase accuracy in marketing response models. Journal of Statistical Computation and Simulation, 51(1):49–55, 1994. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Geoffroy:2001:PGM

[GW01] Pedro Geoffroy and Govinda Weerakkody. A Poisson-gamma model for two-stage cluster sampling data. Journal of Statistical Computation and Simulation, 68(2):161–172, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gluhovsky:2005:SCI

[GZP05] Alexander Gluhovsky, Martin Zihlbauer, and Dimitris N. Politis. Subsampling confidence intervals for parameters of atmospheric time series: block size choice and calibration. Journal of Statistical Computation and Simulation, 75(5):381–389, 2005. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Helms:1978:NAb

[HA78] Ronald W. Helms and O. D. Anderson. News and abstracts. Journal of Statistical Computation and Simulation, 8(2):165– 168, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haber:1980:CSS

[Hab80] Michael Haber. A comparative simulation study of the small sample powers of several goodness fit test. Journal of Statistical Computation and Simulation, 11(3–4):241–250, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haber:1984:CTH

[Hab84] Michael Haber. A comparison of tests for the hypothesis of no three-factor interaction in 2 × 2 × 2 contingency tables. Journal of Statistical Computation and Simulation, 20(3):205–215, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haber:1992:ESP

[Hab92] Michael Haber. On the expected significance probabilities and Bahadur efficiencies of tests for comparing two binomial proportions. Journal of Statistical Computation and Simulation, 43 (3–4):243–251, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hall:1972:SCT

[Hal72] I. J. Hall. Some comparisons of tests for equality of variances. Journal of Statistical Computation and Simulation, 1(2):183– 194, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hall:1982:RET

[Hal82] A. D. Hall. The relative efficiency of time and frequency domain estimators in SUR systems. Journal of Statistical Computation and Simulation, 16(2):81–96, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hamilton:1977:ELC

[Ham77] Martin A. Hamilton. Estimating the logistic curve from grouped data. Journal of Statistical Computation and Simulation, 5(4):

279–301, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hanumara:1978:CEP

[Han78] R. Choudary Hanumara. A comparison of estimators of the precision of two instruments. Journal of Statistical Computation and Simulation, 7(3–4):213–223, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Han:1979:CNC

[Han79] Chien-Pai Han. On the computation of noncentral chi-squared distributions with even degrees of freedom. Journal of Statistical Computation and Simulation, 9(1):25–29, 1979. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hanson:1986:FCD

[Han86] James N. Hanson. Fitting a curve to data when both variables are subject to error. Journal of Statistical Computation and Simulation, 24(2):147–158, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hanumara:1994:NEI

[Han94] Choudary R. Hanumara. A note on the estimation of imprecisions of two instruments. Journal of Statistical Computation and Simulation, 51(1):1–5, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hendi:1998:ETH

[HANMA98] M. Hendi, H. Al-Nachawati, M. Montasser, and I. Alwasel. An exact test for HNBUE class of life distributions. Journal of Statistical Computation and Simulation, 60(3):261–275, 1998. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hartigan:1975:PGC

[Har75] J. A. Hartigan. Printer graphics for clustering. Journal of Statistical Computation and Simulation, 4(3):187–213, 1975. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hart:1985:CTP

[Har85] Jeffrey D. Hart. On the choice of a truncation point in Fourier series density estimation. Journal of Statistical Computation and Simulation, 21(2):95–116, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Harris:1987:CEQ

[Har87a] Carl M. Harris. C291. Estimating quantiles by simulation: a Brownian motion example. Journal of Statistical Computation and Simulation, 28(3):266–272, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hartfiel:1987:CLC

[Har87b] D. J. Hartfiel. Computing limits of convex sets of distribution. Journal of Statistical Computation and Simulation, 27(4):277– 291, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hartfiel:1989:SBL

[Har89] D. J. Hartfiel. Stochastic box limits for nonhomogeneous Markov chains. Journal of Statistical Computation and Simulation, 32 (1–2):17–29, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hartfiel:1991:CBM

[Har91] D. J. Hartfiel. Component bounds on Markov set-chain limiting sets. Journal of Statistical Computation and Simulation, 38(1-4): 15–24, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hartcfiel:1993:CMS

[Har93] D. J. Hartcfiel. Cyclic Markov set-chains. Journal of Statistical Computation and Simulation, 46(3–4):145–167, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Harville:2003:CEV

[Har03] David Harville. C472. The expected value of a conditional variance: An upper bound. Journal of Statistical Computation and Simulation, 73(8):609–612, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Harville:2004:MRC

[Har04] David A. Harville. Making REML computationally feasible for large data sets: use of the Gibbs sampler. Journal of Statistical Computation and Simulation, 74(2):135–153, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hovav:1989:ESE

[HAS89] Talpaz Hovav, William P. Alexander, and Richard C. Shumway. Estimation of systems of equations subject to curvature constraints. Journal of Statistical Computation and Simulation, 32 (4):201–214, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hathaway:1986:CEA

[Hat86] Richard J. Hathaway. A constrained em algorithm for univariate normal mixtures. Journal of Statistical Computation and Simulation, 23(3):211–230, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hawkins:1979:FEM

[Haw79] Douglas M. Hawkins. Fractiles of an extended multiple outlier test. Journal of Statistical Computation and Simulation, 8(3– 4):227–236, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hayes:1997:CTV

[Hay97] Andrew F. Hayes. Cautions in testing variance equality with randomization tests. Journal of Statistical Computation and Simulation, 59(1):25–31, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hanumara:1978:TSP

[HB78] R. Choudary Hanumara and Edward P. Barry. A table of some percentage points of the distribution of a difference between independent chi-square variables. Journal of Statistical Computation and Simulation, 6(3–4):169–181, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hernandez:1993:CIT

[HB93] Ramon P. Hernandez and Richard K. Burdick. Confidence intervals and tests of hypotheses on variance components in an unbalanced two-factor crossed design with interaction. Journal of Statistical Computation and Simulation, 47(1–2):67–77, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hofmann:2006:NTT

[HB06] Glenn Hofmann and N. Balakrishnan. A nonparametric test for trend based on initial ranks. Journal of Statistical Computation and Simulation, 76(9):829–837, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hitchcock:2007:EPS

[HBC07] David B. Hitchcock, James G. Booth, and George Casella. The effect of pre-smoothing functional data on cluster analysis. Journal of Statistical Computation and Simulation, 77(12): 1043–1055, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Helmut:1992:CRF

[HBK92] Schneider Helmut, Shan Lin Bin, and Tang Kwei. Comparison of renewal function estimators for censored data. Journal of Statistical Computation and Simulation, 42(3–4):125–135, 1992. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hanscom:1975:BRP

[HC75] M. A. Hanscom and R. Cléroux. The block replacement problem. Journal of Statistical Computation and Simulation, 3(3):233– 248, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ho:2006:TCS

[HC06] L. P. Ho and S. N. Chiu. Testing the complete spatial randomness by Diggle’s test without an arbitrary upper limit. Journal of Statistical Computation and Simulation, 76(7):585–591, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hill:1996:BES

[HCA96] R. Carter Hill, Phillip A. Cartwright, and Julia F. Arbaugh. Bootstrapping estimators for the seemingly unrelated regressions model. Journal of Statistical Computation and Simulation, 54(1-3):177–196, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hsieh:2007:CSP

[HCW07] Wen-Ping Hsieh, Tzu-Ming Chu, and Russ Wolfinger. Comparison of statistical performance of univariate and bivariate mixed models for Affymetrix^{(}R) probe level data. Journal of Statistical Computation and Simulation, 77(3):251–264, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hosmer:1977:IMT

[HD77] David W. Hosmer, Jr. and Nathan P. Dick. Information and mixtures of two normal distributions. Journal of Statistical Computation and Simulation, 6(2):137–148, 1977. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Halawa:2000:TRC

[HE00] A. M. Halawa and M. Y. El Bassiouni. Tests of regression coefficients under ridge regression models. Journal of Statistical Computation and Simulation, 65(1-4):341–356, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Helms:1978:NAa

[Hel78] Ronald W. Helms. News and abstracts. Journal of Statistical Computation and Simulation, 8(1):81–82, 1978. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Helms:1986:NA

[Hel86] Ronald W. Helms. News and abstracts. Journal of Statistical Computation and Simulation, 25(3–4):307–311, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Helton:1997:USA

[Hel97] J. C. Helton. Uncertainty and sensitivity analysis in the presence of stochastic and subjective uncertainty. Journal of Statistical Computation and Simulation, 57(1-4):3–76, 1997. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Henderson:1981:CIC

[Hen81] Harold V. Henderson. C85. On the inverse of a completely symmetric matrix. Journal of Statistical Computation and Simulation, 12(2):138–139, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hendrickson:1995:ERM

[Hen95] John A. Hendrickson, Jr. On the enumeration of rectangular (0, 1)-matrices. Journal of Statistical Computation and Simulation, 51(2-4):291–313, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Herson:1975:FTV

[Her75] Jay Herson. Fieller’s theorem vs. the delta method for significance intervals for ratios. Journal of Statistical Computation and Simulation, 3(3):265–274, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Holtzman:1985:CEU

[HG85] Golde I. Holtzman and I. J. Good. C242. Eeaviside’s unit function applied to midrank procedures. Journal of Statistical Computation and Simulation, 23(1–2):146–148, 1985. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haseman:1974:TGG

[HH74] J. K. Haseman and D. G. Hoel. Tables of Gehan’s generalized Wilcoxon test with fixed point censoring. Journal of Statistical Computation and Simulation, 3(2):117–135, 1974. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hayter:1992:PCB

[HH92] Anthony J. Hayter and Merrilee Hurn. Power comparisons between the F -test, the Studentised range test, and an optimal test of the equality of several normal means. Journal of Statistical Computation and Simulation, 42(3–4):173–185, 1992. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hsuan:1993:MSA

[HH93] Francis Hsuan and Burt Holland. A modified sweep algorithm for interchanging between overparameterized and cell means linear models. Journal of Statistical Computation and Simulation, 48(3–4):135–151, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hossain:1996:ULS

[HH96] Anwar Hossain and Hatem A. Howlader. Unweighted least squares estimation of Weibull parameters. Journal of Statistical Computation and Simulation, 54(1-3):265–271, 1996. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haughton:1990:ICH

[HHI90] Dominique Haughton, Jonathan Haughton, and Alan J. Izenman. Information criteria and harmonic models in time series analysis. Journal of Statistical Computation and Simulation, 35 (3–4):187–207, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:2002:IAI

[HHKD02] Y. Huang, P. Harris, S. P. J. Kirby, and J. C. Dearden. Improved approaches to interval estimation of effective doses for a logistic dose-response curve. Journal of Statistical Computation and Simulation, 72(7):565–584, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hwu:2002:CTM

[HHR02] Tian-Jong Hwu, Chien-Pai Han, and Katherine Rogers. The combination test for multivariate normality. Journal of Statistical Computation and Simulation, 72(5):379–390, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hietikko:1981:CBL

[Hie81] Harri Hietikko. Comparison between the least squares, maximum likelihood and Bayes estimation of simple ARMA models. Journal of Statistical Computation and Simulation, 12(3–4):225– 235, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Higuchi:1997:MCF

[Hig97] Tomoyuki Higuchi. Monte Carlo filter using the genetic algorithm operators. Journal of Statistical Computation and Simulation, 59(1):1–23, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hill:1987:CCL

[Hil87] I. D. Hill. C281. 95% Confidence limits for the median. Journal of Statistical Computation and Simulation, 28(1):80–81, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hines:1997:ARS

[Hin97a] R. J. O’Hara Hines. An application of retrospective sampling in the analysis of a very large clustered data set. Journal of Statistical Computation and Simulation, 59(1):63–81, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hines:1997:CTO

[Hin97b] R. J. Ohara Hines. A comparison of tests for overdispersion in generalized linear models. Journal of Statistical Computation and Simulation, 58(4):323–342, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hoel:1985:LTA

[HJ85] David G. Hoel and Robert I. Jennrich. Life table analysis with small numbers of cases: an example-multiple myeloma in Hiroshima and Nagasaki. Journal of Statistical Computation and Simulation, 20(4):311–322, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jun:1993:HSE

[hJ93] Chi hyuck Jun. Heuristic shrinkage estimators of Poisson means and a Monte Carlo comparison. Journal of Statistical Computation and Simulation, 47(3–4):181–193, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Humble:1984:SST

[HK84] S. Humble and M. J. Knapp. Small sample testing of one-shot devices with exponentially distributed lifetimes. Journal of Statistical Computation and Simulation, 20(1):21–36, 1984. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ho:1992:SMM

[HK92] Shu-Yen Ho and Jerome Klotz. Sparse matrix methods for unbalanced multifactor analysis of variance and covariance. Journal of Statistical Computation and Simulation, 41(1–2):55–72, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hartung:2000:CIB

[HK00] Joachim Hartung and Guido Knapp. Confidence intervals for the between group variance in the unbalanced one-way random effects model of analysis of variance. Journal of Statistical Computation and Simulation, 65(1-4):311–323, 2000. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Headrick:2007:PMT

[HK07] Todd C. Headrick and Rhonda K. Kowalchuk. The power method transformation: its probability density function, distribution function, and its further use for fitting data. Journal of Statistical Computation and Simulation, 77(3):229–249, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hsu:2008:RRS

[HKL08] Ying-Lin Hsu, Jau-Chuan Ke, and Ssu-Lang Lee. On a redundant repairable system with switching failure: Bayesian approach. Journal of Statistical Computation and Simulation, 78(12):1163–1180, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Horvath:2004:BMT

[HKT04] Lajos Horváth, Piotr Kokoszka, and Gilles Teyssière. Bootstrap misspecification tests for ARCH based on the empirical process of squared residuals. Journal of Statistical Computation and Simulation, 74(7):469–485, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hutcheson:1984:CDG

[HL84] K. Hutcheson and N. I. Lyons. C196. Diversity: a glance at the Federalist Papers. Journal of Statistical Computation and Simulation, 19(4):291–294, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Horswell:1992:CTM

[HL92a] Ronald L. Horswell and Stephen W. Looney. A comparison of tests for multivariate normality that are based on measures of multivariate skewness and kurtosis. Journal of Statistical Computation and Simulation, 42(1–2):21–38, 1992. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:1992:GRM

[hL92b] Kim hung Li. Generation of random matrices with orthonormal columns and multivariate normal variates with given sample mean and covariance. Journal of Statistical Computation and Simulation, 43(1–2):11–18, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hurd:1992:SEN

[HL92c] Michael D. Hurd and Hoe Kyung Lee. Simplified estimation of a nonlinear simultaneous equations model with censoring: a Monte Carlo study. Journal of Statistical Computation and Simulation, 43(1–2):55–75, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hamilton:1997:CEP

[HL97] David C. Hamilton and Mary L. Lesperance. A comparison of estimators of the proportion nonconforming in univariate and multivariate normal samples. Journal of Statistical Computation and Simulation, 59(4):333–348, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hamilton:1998:CMF

[HL98] David C. Hamilton and Mary L. Lesperance. Calculating the MVUE of the fraction nonconforming for the bivariate normal. Journal of Statistical Computation and Simulation, 61(3):259– 270, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hamza:2005:ECE

[HL05] Mounir Hamza and Denis Larocque. An empirical comparison of ensemble methods based on classification trees. Journal of Statistical Computation and Simulation, 75(8):629–643, 2005. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hussein:2009:PEM

[HL09] A. Hussein and J. Liu. Parametric estimation of mixtures of two uniform distributions. Journal of Statistical Computation and Simulation, 79(4):395–410, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hinde:1980:MCE

[HM80] A. L. Hinde and R. E. Miles. Monte Carlo estimates of the distributions of the random polygons of the Voronoi tessellation with respect to a Poisson process. Journal of Statistical Computation and Simulation, 10(3–4):205–223, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hewett:1985:RTC

[HM85] John E. Hewett and Richard Madsen. A robust test for comparing three populations relative to an ordered alternative with bivariate data. Journal of Statistical Computation and Simulation, 21(3–4):195–212, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haccou:1988:TNC

[HM88] Patsy Haccou and Evert Meelis. Testing for the number of change points in a sequence of exponential random variables. Journal of Statistical Computation and Simulation, 30(4):285– 298, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hamel:1995:MCS

[HM95] Denis Hamel and Jean-Claude Massé. A Monte Carlo study of the robustness of prediction in growth curve models. Journal of Statistical Computation and Simulation, 52(1):33–53, 1995. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hsu:1998:CAT

[HM98] Yu-Sheng Hsu and Hong Mei. Comparisons among three estimation methods in linear models when observations are pairwise correlated. Journal of Statistical Computation and Simulation, 60(3):223–236, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hefang:2000:BTS

[HMK00] Lin Hefang, Raymond H. Myers, and Ye. Keying. Bayesian twostage optimal design for mixture models. Journal of Statistical Computation and Simulation, 66(3):209–231, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hall:1989:BNB

[HMS89] Peter Hall, Michael A. Martin, and William R. Schucany. Better nonparametric bootstrap confidence intervals for the correlation coefficient. Journal of Statistical Computation and Simulation, 33(3):161–172, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hettmansperger:2002:FSP

[HMS02] Thomas Hettmansperger, Joseph Mckean, and Simon Sheather. Finite sample performance of tests for symmetry of the errors in a linear model. Journal of Statistical Computation and Simulation, 72(11):863–879, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hui:2005:BCI

[HMZ05] Terrence P. Hui, Reza Modarres, and Gang Zheng. Bootstrap confidence interval estimation of mean via ranked set sampling linear regression. Journal of Statistical Computation and Simulation, 75(7):543–553, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hall:1980:EEM

[HN80] A. D. Hall and D. F. Nicholls. The evaluation of exact maximum likelihood estimates for Varma models. Journal of Statistical Computation and Simulation, 10(3–4):251–262, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Helms:1983:NA

[HN83] Ranald W. Helms and M. Novák. News and abstracts. Journal of Statistical Computation and Simulation, 17(3):249–250, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ho:1993:FBT

[Ho93] Chih-Hsiang Ho. Forward and backward tests for an abrupt change in the intensity of a Poisson process. Journal of Statistical Computation and Simulation, 48(3–4):245–252, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hoeschele:1989:NLM

[Hoe89] I. Hoeschele. A note on local maxima in maximum likelihood, restricted maximum likelihood, and Bayesian estimation of variance components. Journal of Statistical Computation and Simulation, 33(3):149–160, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hogben:1973:LE

[Hog73] David Hogben. Letter to the editor. Journal of Statistical Computation and Simulation, 2(3):289–290, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Holtzman:1985:CMA

[Hol85a] Golde I. Holtzman. C222. The measurement of the accuracy of approximate p-values. Journal of Statistical Computation and Simulation, 21(1):76–84, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Holtzman:1985:CCD

[Hol85b] Golde I. Holtzman. C243. Computation of the distribution of Page’s criterion, given any configuration of ties, by a Fast Fourier Transform. Journal of Statistical Computation and Simulation, 23(1–2):149–154, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Honds:1990:CPE

[Hon90] Masayuki Honds. Comparison of procedures for estimation of error rates in discriminant analysis under nonnormal population. Journal of Statistical Computation and Simulation, 36(2-3):105– 115, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hormann:1993:GBR

[Hör93] Wolfgang Hörmann. The generation of binomial random variates. Journal of Statistical Computation and Simulation, 46(1– 2):101–110, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hora:1997:SUD

[Hor97] Stephen C. Hora. Sensitivity, uncertainty, and decision analyses in the prioritization of research. Journal of Statistical Computation and Simulation, 57(1-4):175–196, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hosmer:1978:UMT

[Hos78] David W. Hosmer, Jr. A use of mixtures of two normal distributions in a classification problem. Journal of Statistical Computation and Simulation, 6(3–4):281–294, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hosmane:1987:EIC

[Hos87] Balakrishna Hosmane. An empirical investigation of chi-square tests for the hypothesis of no three-factor interaction in I × J × K contingency tables. Journal of Statistical Computation and Simulation, 28(2):167–178, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hossain:1991:IOA

[Hos91] Anwar Hossain. Influential observations in Anova and Manova models. Journal of Statistical Computation and Simulation, 39 (1–2):9–19, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Houston:1985:ETA

[Hou85] Ben F. Houston. Extended tables on an approximate test for outliers in linear models. Journal of Statistical Computation and Simulation, 21(2):179–185, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Howell:1985:APD

[How85] Leonard W. Howell. An APL program for the distribution of energy deposition by charged particles passing through thin absorbers. Journal of Statistical Computation and Simulation, 22 (3–4):265–284, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Howlader:1988:LSE

[How88] H. Howlader. Linear scale estimation. Journal of Statistical Computation and Simulation, 29(2):117–126, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hazen:1980:FVS

[HP80] Gordon B. Hazen and A. Alan B. Pritsker. Formulas for the variance of the sample mean in finite state Markov processes. Journal of Statistical Computation and Simulation, 12(1):25– 40, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hall:1990:SBC

[HP90] Peter Hall and Y. E. Pittelkow. Simultaneous bootstrap confidence bands in regression. Journal of Statistical Computation and Simulation, 37(1–2):99–113, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Houshmand:1995:AAN

[HP95] Ali A. Houshmand and Srinivasarao Panganamamula. An analytical approximation and a neural network model for optimal sample size in vendor selection. Journal of Statistical Computation and Simulation, 53(1–2):65–78, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hammou:2000:MDE

[HP00] Brami El. Hammou and Nelson I. Paul. The monotone density estimators from selection biased samples. Journal of Statistical Computation and Simulation, 67(3):203–217, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hashemi-Parast:1979:DFS

[HPY79] S. M. Hashemi-Parast and D. H. Young. Distribution free slippage tests for populations following a Lehmann model. Journal of Statistical Computation and Simulation, 8(3–4):237–251, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hayen:2000:CPT

[HQ00] Andrew Hayen and Malcolm Quine. Calculating the proportion of triangles in a Poisson–Voronoi tessellation of the plane. Journal of Statistical Computation and Simulation, 67(4):351–358, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hafner:2007:ETA

[HR07] Christian M. Hafner and Jeroen V. K. Rombouts. Estimation of temporally aggregated multivariate GARCH models. Journal of Statistical Computation and Simulation, 77(8):629–650, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hall:1973:OST

[HS73] I. J. Hall and C. B. Sampson. One-sided tolerance limits for a normal population based on censored samples. Journal of Statistical Computation and Simulation, 2(4):317–324, 1973. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hazilla:1977:SSN

[HS77] Michael Hazilla and V. Kerry Smith. Sampling studies with nonlinear simultaneous equations: a preliminary appraisal. Journal of Statistical Computation and Simulation, 6(2):101–114, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Horn:1986:GSO

[HS86a] Paul S. Horn and John S. Schlipf. Generating subsets of order statistics with applications to trimmed means and means of trimmings. Journal of Statistical Computation and Simulation, 24(2):83–97, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Horn:1986:WBS

[HS86b] Paul S. Horn and John S. Schlipf. Who’s being swindled?: a case for enlightened naïveté. Journal of Statistical Computation and Simulation, 26(1–2):21–36, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huesemann:1991:COP

[HS91] J. A. Huesemann and R. R. Stevens. Computation of optimal parameters for bivariate histograms. Journal of Statistical Computation and Simulation, 38(1-4):109–125, 1991. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Holgersson:2004:TMH

[HS04] H. E. T. Holgersson and Ghazi Shukur. Testing for multivariate heteroscedasticity. Journal of Statistical Computation and Simulation, 74(12):879–896, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hassanein:1985:BLU

[HSB85] K. M. Hassanein, A. K. Md. Ehsanes Saleh, and E. F. Brown. Best linear unbiased quantile estimators of the logistic distribution using order statistics. Journal of Statistical Computation and Simulation, 23(1–2):123–131, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Holt:2004:NTI

[HSSY04] Melinda A. Holt, James D. Stamey, John W. Seaman, Jr., and Dean M. Young. A note on tests for interaction in quantal response data. Journal of Statistical Computation and Simulation, 74(9):683–690, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hsu:2002:GLL

[Hsu02] John Hsu. A generalization of the logistic linear model. Journal of Statistical Computation and Simulation, 72(7):549–564, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hoel:1975:CSM

[HSW75] David G. Hoel, Milton Sobel, and George H. Weiss. Comparison of sampling methods for choosing the best binomial population with delayed observations. Journal of Statistical Computation and Simulation, 3(4):299–313, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Husby:2007:CNU

[HSWF07] Chad E. Husby, Elizabeth A. Stasny, Douglas A. Wolfe, and Jesse Frey. Cautionary note on unbalanced ranked-set sampling. Journal of Statistical Computation and Simulation, 77(10):869– 878, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hannagan:1983:NSE

[HT83] G. M. Hannagan and P. J. Thomson. A note on spectral estimation for ARMA processes. Journal of Statistical Computation and Simulation, 18(2-3):87–92, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Herzberg:1985:MCC

[HT85] Agnes M. Herzberg and A. V. Tsukanov. The Monte-Carlo comparison of two criteria for the selection of models. Journal of Statistical Computation and Simulation, 22(2):113–126, 1985. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hamilton:1993:CBC

[HT93] A. Scott Hamilton and M. G. Jeremy Taylor. A comparison of the Box–Cox transformation method and nonparametric methods for estimating quantiles in clinical data with repeated measures. Journal of Statistical Computation and Simulation, 45 (3–4):185–201, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Holst:1999:STM

[HT99] Erik Holst and Poul Thyregod. A statistical test for the mean squared error. Journal of Statistical Computation and Simulation, 63(4):321–347, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hsueh:2007:INT

[HTC07] Huey-Miin Hsueh, Chen-An Tsai, and James J. Chen. Incorporating the number of true null hypotheses to improve power in multiple testing: application to gene microarray data. Journal of Statistical Computation and Simulation, 77(9):757–767, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:1985:SOP

[Hua85] Wen-Tao Huang. Some optimization problems in selection procedures. Journal of Statistical Computation and Simulation, 22 (1):83–91, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:1991:ECP

[Hua91] J. S. Huang. Efficient computation of the performance of bootstrap and jackknife estimators of the variance of L-statistics. Journal of Statistical Computation and Simulation, 38(1-4):45– 56, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:2001:TAL

[Hua01] Li-Shan Huang. Testing the adequacy of a linear model VIA critical smoothing. Journal of Statistical Computation and Simulation, 68(3):281–294, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hudson:1983:REL

[Hud83] H. M. Hudson. Random effects in log-linear models. Journal of Statistical Computation and Simulation, 17(1):43–50, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hunt:1987:SRM

[Hun87] D. N. Hunt. Successive over-relaxation methods in the analysis of variance. Journal of Statistical Computation and Simulation, 27(4):321–337, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hunt:2002:SIE

[Hun02] Daniel Hunt. Simulations to improve experimental designs for u-shaped dose-response modeling. Journal of Statistical Computation and Simulation, 72(9):737–746, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hutcheson:1977:CSE

[Hut77] K. Hutcheson. C3. Some exact moments of minus twice the logarithm of the likelihood ratio for multinomial distributions. Journal of Statistical Computation and Simulation, 6(2):149, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hutchinson:1993:CSA

[Hut93a] T. P. Hutchinson. C102. Second attempts at multiple-choice test items. Journal of Statistical Computation and Simulation, 47(1– 2):108–112, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hutchinson:1993:CCM

[Hut93b] T. P. Hutchinson. C403. Comments on the multivariate normal integral. Journal of Statistical Computation and Simulation, 47 (1–2):112–114, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hariharan:1993:EDP

[HV93] Hari S. Hariharan and Raja P. Velu. On estimating Dirichlet parameters — a comparison of initial values. Journal of Statistical Computation and Simulation, 48(1–2):47–58, 1993. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

He:2005:ECS

[HV05] Qin He and Eberhard O. Voit. Estimation and completion of survival data with piecewise linear models and S-distributions.

Journal of Statistical Computation and Simulation, 75(4):287– 305, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Helms:1988:NA

[HZ88] Ronald W. Helms and Wei-Min Zheng. News and abstracts. Journal of Statistical Computation and Simulation, 29(2):179, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Harville:1996:PDF

[HZ96] David A. Harville and Alan G. Zimmermann. The posterior distribution of the fixed and random effects in a mixed-effects linear model. Journal of Statistical Computation and Simulation, 54 (1-3):211–229, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hossain:2003:CEM

[HZ03] Anwar Hossain and William Zimmer. Comparison of estimation methods for Weibull parameters: Complete and censored samples. Journal of Statistical Computation and Simulation, 73 (2):145–153, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Iglarsh:1980:WER

[IC80] Harvey J. Iglarsh and David C. Cheng. Weighted estimators in regression with multicollinearity. Journal of Statistical Computation and Simulation, 10(2):103–112, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ines:2000:OWW

[IC00] Fortin Ines and Kuzmics Christoph. Optimal window width choice in spectral density estimation review and simulation. Journal of Statistical Computation and Simulation, 67(2):109– 131, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Iglewicz:1978:ALR

[IHM78] Boris Iglewicz, David Huang, and Gaston Mendoza. Another look at the robustness of Colton’s model for comparing two medical treatments. Journal of Statistical Computation and Simulation, 7(3–4):259–267, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ismail:1986:EGD

[IJP86] R. Ismail, S. Jeyaratnam, and S. Panchapakesan. Estimation of pr[X > Y] for gamma distributions. Journal of Statistical Computation and Simulation, 26(3–4):253–267, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Inglot:2003:DDC

[IJW03] Tadeusz Inglot and Alicja Janic-Wróblewska. Data driven chisquare test for uniformity with unequal cells. Journal of Statistical Computation and Simulation, 73(8):545–561, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Iliopoulos:2003:SBD

[IK03] George Iliopoulos and Dimitris Karlis. Simulation from the Bessel distribution with applications. Journal of Statistical Computation and Simulation, 73(7):491–506, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Iglewicz:1982:ODU

[IM82] Boris Iglewicz and Jorge Martinez. Outlier detection using robust measures of scale. Journal of Statistical Computation and Simulation, 15(4):285–293, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Izraelevitz:2009:RSO

[IMLG09] David Izraelevitz, Harry F. Martz, Deborah A. Leishman, and William L. Gibson. On representing second-order uncertainty in multi-state systems via moments of mixtures of Dirichlet distributions. Journal of Statistical Computation and Simulation, 79 (7):869–880, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Insua:1997:SAS

[IMP^{+}97] David Rios Insua, Jacinto Martin, Les Proll, Simon French, and Abdellah Salhi. Sensitivity analysis in statistical decision theory: A decision analytic view. Journal of Statistical Computation and Simulation, 57(1-4):197–218, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ironyt:1986:ETE

[IP86] Telba Z. Ironyt and Carlos A. B. Pereirat. Exact tests for equality of two proportions: Fisher v. Bayes. Journal of Statistical Computation and Simulation, 25(1–2):93–114, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

J:1983:CML

[J.83] Good I. J. C182. The multivariate lognormal distribution and colored trees. Journal of Statistical Computation and Simulation, 18(4):316–322, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jaheen:2005:RSM

[Jah05] Zeinhum F. Jaheen. On record statistics from a mixture of two exponential distributions. Journal of Statistical Computation and Simulation, 75(1):1–11, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jain:1985:ODL

[Jai85] Sudha Jain. D-optimum design for logistic model. Journal of Statistical Computation and Simulation, 22(3–4):303–306, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Josephy:1993:BES

[JAK93] Norman H. Josephy, Amir D. Aczel, and Hans R. Künsch. Bootstrap estimates of the sample bivariate autocorrelation and partial autocorrelation distributions. Journal of Statistical Computation and Simulation, 46(3–4):235–249, 1993. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jamshidian:2001:NPS

[Jam01] Mortaza Jamshidian. A note on parameter and standard error estimation in adaptive robust regression. Journal of Statistical Computation and Simulation, 71(1):11–27, 2001. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Janssen:1991:MCR

[JB91] Arnold Janssen and Stephan Brenner. Monte Carlo results for conditional survival tests under randomly censored data. Journal of Statistical Computation and Simulation, 39(1–2):47–62, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jang:2009:BMA

[JCKS09] Junghoon Jang, Younshik Chung, Chansoo Kim, and Seongho Song. Bayesian meta-analysis using skewed elliptical distributions. Journal of Statistical Computation and Simulation, 79 (5):691–704, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jowett:1972:OSL

[JCM72] D. Jowett, R. L. Chamberlain, and A. G. Mexas. Omnitab — a simple language for statistical computations. Journal of Statistical Computation and Simulation, 1(2):129–147, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jan:2007:SIM

[JCS07] Show-Li Jan, Yuh-Ing Chen, and Gwowen Shieh. Simultaneous identifications of the minimum effective dose in each of several groups. Journal of Statistical Computation and Simulation, 77 (2):149–161, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1976:GAB

[Jen76] D. R. Jensen. Gaussian approximation to bivariate Rayleigh distributions. Journal of Statistical Computation and Simulation, 4(4):259–267, 1976. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1978:CCC

[Jen78] Donald R. Jensen. C13. On consistent Cauchy averages. Journal of Statistical Computation and Simulation, 7(3–4):287–289, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1980:CSC

[Jen80] D. R. Jensen. C66. Sharpened Chebychev inequalities. Journal of Statistical Computation and Simulation, 10(3–4):312–315, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1982:CRW

[Jen82a] D. R. Jensen. C125. Recovering the Wishart distribution from its sections. Journal of Statistical Computation and Simulation, 15(1):69–72, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1982:COP

[Jen82b] D. R. Jensen. C128. An optimal property of s^{2} charts. Journal of Statistical Computation and Simulation, 15(1):77–81, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1983:COH

[Jen83] D. R. Jensen. C153. The optimality of Hotelling’s T^{2} in spherical processes. Journal of Statistical Computation and Simulation, 17(2):151–154, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1989:CMLa

[Jen89a] D. R. Jensen. C336. Misspecified location and scale. I. Student’s test. Journal of Statistical Computation and Simulation, 32(1– 2):115–118, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1989:CMLb

[Jen89b] D. R. Jensen. C337. Misspecified location and scale. II. Independence of Ȳ and s^{2}. Journal of Statistical Computation and Simulation, 32(1–2):119–121, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1989:CMLc

[Jen89c] D. R. Jensen. C338. Misspecified location and scale. III. The use of S^{2}. Journal of Statistical Computation and Simulation, 32(1– 2):121–124, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1991:CMA

[Jen91] D. R. Jensen. C382. Matrix approximations of given rank. Journal of Statistical Computation and Simulation, 39(3):187–191, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1993:CPS

[Jen93a] D. R. Jensen. C393. Power of Student’s tests under spherical errors. Journal of Statistical Computation and Simulation, 44 (3–4):258–261, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1993:CBC

[Jen93b] D. R. Jensen. C405. Bivariate correlation in translation-scale models. Journal of Statistical Computation and Simulation, 47 (1–2):118–121, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1993:CLU

[Jen93c] D. R. Jensen. C406. Lower and upper bounds for Gaussian measures on R^{k}. Journal of Statistical Computation and Simulation, 47(1–2):121–124, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1994:CUC

[Jen94] D. R. Jensen. C408. Uniform Chebychev bounds on R^{n}. Journal of Statistical Computation and Simulation, 49(3–4):218–221, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1995:CEC

[Jen95a] D. R. Jensen. C432. Efficiency comparisons of central composite designs. Journal of Statistical Computation and Simulation, 52 (2):177–183, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1995:CPH

[Jen95b] D. R. Jensen. C433. Power of Hotelling’s T^{2} in Gaussian mixtures. Journal of Statistical Computation and Simulation, 52 (2):184–187, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1995:CLB

[Jen95c] D. R. Jensen. C434. Local bounds on power of T^{2} in scale families. Journal of Statistical Computation and Simulation, 52 (2):187–190, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jeng:2003:ESS

[Jen03] Shuen-Lin Jeng. Exact sample size determination in a Weibull test plan when there is time censoring. Journal of Statistical Computation and Simulation, 73(6):389–408, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1980:CRS

[JF80] D. R. Jensen and R. V. Foutz. C65. On resolving statistical hypotheses. Journal of Statistical Computation and Simulation, 10 (3–4):309–312, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jain:1980:CMR

[JG80] G. C. Jain and R. P. Gupta. On certain moment ratios of a class of discrete distributions. Journal of Statistical Computation and Simulation, 10(3–4):263–270, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Johannesson:1983:CPF

[JG83a] B. Jóhannesson and N. Giri. Comparison of power functors of some step-down tests for means with additional observations. Journal of Statistical Computation and Simulation, 17(3):173– 194, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Johannesson:1983:CVL

[JG83b] Benedikt Jóhannesson and Narayan Giri. Critical values of the locally optimum combination of two independent test statistics. Journal of Statistical Computation and Simulation, 16(3–4):251– 286, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jen:1987:THV

[JG87] Tang Jen and A. K. Gupta. On testing homogeneity of variances for Gaussian models. Journal of Statistical Computation and Simulation, 27(2):155–173, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jones:1992:ICR

[JG92] Peter N. Jones and Mclachlan J. Geoffrey. Improving the convergence rate of the em algorithm for a mixture model fitted to grouped truncated data. Journal of Statistical Computation and Simulation, 43(1–2):31–44, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jiang:1996:AEW

[JG96] Qihao Jiang and Alan J. Gross. Asymptotic evaluation of weighted nonparametric combination procedures. Journal of Statistical Computation and Simulation, 55(3):213–237, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jones:1972:SAV

[JH72] R. D. H. Jones and K. Hinkelmann. Sequential analysis of variance for irregular designs: some Monte Carlo results. Journal of Statistical Computation and Simulation, 1(3):273–280, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1990:CPCb

[JH90a] D. R. Jensen and Y. V. Hui. C343. Properties of R and S^{2} charts. II. Nonstationarity. Journal of Statistical Computation and Simulation, 34(2-3):172–175, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1990:CPCa

[JH90b] D. R. Jensen and Y. V. Hui. C348. Properties of R and S^{2} charts. I. Dependent statistics. Journal of Statistical Computation and Simulation, 34(2-3):168–172, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Javaheri:2001:ARL

[JH01] Amir Javaheri and Ali A. Houshmand. Average run length comparison of multivariate control charts. Journal of Statistical Computation and Simulation, 69(2):125–140, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jung:2009:STO

[JJH09] Byoung Cheol Jung, Myoungshic Jhun, and Sang Moon Han. Score tests for overdispersion in the bivariate negative binomial models. Journal of Statistical Computation and Simulation, 79 (1):11–24, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Johnson:1989:CAA

[JK89] Norman L. Johnson and Samuel Kotz. C310. Axiomatic approaches to formulas for combining likelihoods or evidence. Journal of Statistical Computation and Simulation, 31(1):49– 54, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See errata [Ano90a].

Jach:2008:WBC

[JK08] Agnieszka Jach and Piotr Kokoszka. Wavelet-based confidence intervals for the self-similarity parameter. Journal of Statistical Computation and Simulation, 78(12):1181–1200, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jones:1983:EIC

[JL83] M. C. Jones and H. W. Lotwick. On the errors involved in computing the empirical characteristic function. Journal of Statistical Computation and Simulation, 17(2):133–149, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jhun:1988:MCM

[JL88] Myoungshic Jhun and Pui Lam Leung. A Monte Carlo method for Foutz’s goodness-of-fit test. Journal of Statistical Computation and Simulation, 29(4):309–316, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jones:1996:TRV

[JL96] M. C. Jones and A. D. Lunn. Transformations and random variate generation: generalised ratio-of-uniforms methods. Journal of Statistical Computation and Simulation, 55(1–2):49–55, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:2005:CNV

[jLP05] Jyh jiuan Lin and Nabendu Pal. Comparison of normal variance estimators under multiple criteria and towards a compromise estimator. Journal of Statistical Computation and Simulation, 75(8):645–665, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jolliffe:1996:SSU

[JMY96] Ian T. Jolliffe, Byron J. T. Morgan, and Philip J. Young. A simulation study of the use of principal components in linear discriminant analysis. Journal of Statistical Computation and Simulation, 55(4):353–366, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jeyaratnam:1985:THO

[JO85] S. Jeyaratnam and Abdul Rahman Othman. Test of hypothesis in one-way random effects model with unequal error variances. Journal of Statistical Computation and Simulation, 21(1):51– 57, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jockel:1981:CCC

[Jöc81] K.-H. Jöckel. C82. A comment on the construction of exact tests from test statistics that have unknown null distributions. Journal of Statistical Computation and Simulation, 12(2):133– 134, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Johnson:1978:SSR

[Joh78] Lester W. Johnson. Some simulation results for the linear expenditure system with random marginal budget shares. Journal of Statistical Computation and Simulation, 7(3–4):225–236, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Johnson:1979:CCG

[Joh79] Mark E. Johnson. C51. Computer generation of the exponential power distribution. Journal of Statistical Computation and Simulation, 9(3):239–240, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Johnson:1982:CCG

[Joh82] Mark E. Johnson. C137. Computer generation of the generalized Eyraud distribution. Journal of Statistical Computation and Simulation, 15(4):333–335, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Johnstone:1990:CLL

[Joh90a] David Johnstone. C341. Likelihood logic in orthodox significance tests. Journal of Statistical Computation and Simulation, 34(2-3):158–160, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Johnstone:1990:CCR

[Joh90b] David Johnstone. C342. Consistency and relative plausibility. Journal of Statistical Computation and Simulation, 34(2-3):160– 162, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comment [Goo90a].

Johnstone:1994:CER

[Joh94] D. J. Johnstone. C412. The evidence required by auditors. Journal of Statistical Computation and Simulation, 49(3–4):228–229, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Johansson:1995:TSC

[Joh95] Per Johansson. Tests for serial correlation and overdispersion in a count data regression model. Journal of Statistical Computation and Simulation, 53(3–4):153–164, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jones:1975:PEU

[Jon75] Richard H. Jones. Probability estimation using a multinominal logistic function. Journal of Statistical Computation and Simulation, 3(4):315–329, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jones:1986:CIM

[Jon86] M. C. Jones. C245. Inverse moments of negative- binomial distributions. Journal of Statistical Computation and Simulation, 23(3):241–243, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jonsson:2007:FSD

[Jön07] Kristian Jönsson. Finite-sample distribution of a recursively mean-adjusted panel data unit root test. Journal of Statistical Computation and Simulation, 77(4):293–303, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Joseph:2001:SSK

[Jos01] Tahsoh T. Joseph. Smoothing spline and kernel estimation of a probit function. Journal of Statistical Computation and Simulation, 68(4):373–389, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

John:1983:SLC

[JR83] R. D. John and J. Robinson. Significance levels and confidence intervals for permutation tests. Journal of Statistical Computation and Simulation, 16(3–4):161–173, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1996:CCEa

[JR96a] D. R. Jensen and D. E. Ramirez. C441. Complexity in experimental design I. Relation to efficiency. Journal of Statistical Computation and Simulation, 55(3):245–250, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:1996:CCEb

[JR96b] D. R. Jensen and D. E. Ramirez. C442. Complexity in experimental design II. Comparing second-order designs. Journal of Statistical Computation and Simulation, 55(3):251–257, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:2009:AAC

[JR09] D. R. Jensen and D. E. Ramirez. Anomalies in the analysis of calibrated data. Journal of Statistical Computation and Simulation, 79(3):299–314, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jeong:2000:NED

[JS00] Dong-Bin Jeong and Sahadeb Sarkar. Negative exponential disparity based family of goodness-of-fit tests for multinomial models. Journal of Statistical Computation and Simulation, 65(1-4): 43–61, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jiang:2002:CTE

[JS02] Guoyong Jiang and Sanat Sarkar. Combination tests for the equality of the covariance matrices of two dependent bivariate normals. Journal of Statistical Computation and Simulation, 72 (6):495–505, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jowaheer:2009:FPM

[JSS09] Vandna Jowaheer, Brajendra C. Sutradhar, and Gary Sneddon. On familial Poisson mixed models with multi-dimensional random effects. Journal of Statistical Computation and Simulation, 79(8):1043–1062, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jodrey:1980:RSP

[JT80] W. S. Jodrey and E. M. Tory. Random sequential packing in R^{n}. Journal of Statistical Computation and Simulation, 10(2): 87–93, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jung:2008:SSC

[Jun08] Sin-Ho Jung. Sample size calculation for paired survival data: a simulation method. Journal of Statistical Computation and Simulation, 78(1):85–92, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Johnson:1986:MCC

[JUP86] Wesley Johnson, Jessica Utts, and Larry M. Pearson. A Monte Carlo comparison of Bayesian estimators and trimmed means. Journal of Statistical Computation and Simulation, 25(3–4):167– 192, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Johnson:1980:GPR

[JW80] D. G. Johnson and W. J. Welcht. The generation of pseudorandom correlation matrices. Journal of Statistical Computation and Simulation, 11(1):55–69, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jones:1997:RRA

[JW97] L. Allison Jones and William H. Woodall. A runs rule alternative to level crossings in statistical process control. Journal of Statistical Computation and Simulation, 59(4):315–331, 1997. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kohn:1982:NOT

[KA82] Robert Kohn and Craig Ansley. A note on obtaining the theoretical autocovariances of an ARMA process. Journal of Statistical Computation and Simulation, 15(4):273–283, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kohn:1985:CLD

[KA85] Robert Kohn and Craig F. Ansley. Computing the likelihood and its derivatives for a Gaussian ARMA model. Journal of Statistical Computation and Simulation, 22(3–4):229–263, 1985. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kohn:1993:AEA

[KA93] Robert Kohn and Craig F. Ansley. Accuracy and efficiency of alternative spline smoothing algorithms. Journal of Statistical Computation and Simulation, 46(1–2):1–18, 1993. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Knoth:2003:ATL

[KA03] Sven Knoth and Raid Amin. Autocorrelation and tolerance limits. Journal of Statistical Computation and Simulation, 73(7): 467–489, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kahn:1993:SSM

[Kah93] Margaret Kahn. A state space method for estimating frequencies and dampings. Journal of Statistical Computation and Simulation, 48(1–2):101–116, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kanji:1975:RPA

[Kan75] G. K. Kanji. Robustness of power in the analysis of variance. Journal of Statistical Computation and Simulation, 4(1):19–30, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kappenman:1981:EWL

[Kap81] Russell F. Kappenman. Estimation of Weibull location. Journal of Statistical Computation and Simulation, 13(3–4):245–254, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kappenman:1983:PES

[Kap83] Russell F. Kappenman. Parameter estimation via sample reuse. Journal of Statistical Computation and Simulation, 16(3–4):213– 222, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kappenman:1987:NED

[Kap87] Russell F. Kappenman. Nonparametric estimation of doseresponse curves with application to ED 50 estimation. Journal of Statistical Computation and Simulation, 28(1):1–13, 1987. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kappenman:1993:RPT

[Kap93] Russell F. Kappenman. A robust power transformation to symmetry. Journal of Statistical Computation and Simulation, 48 (1–2):11–18, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Karrison:1990:BCD

[Kar90] Theodore Karrison. Bootstrapping censored data with covariates. Journal of Statistical Computation and Simulation, 36(4): 195–207, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Karlsson:2009:BMB

[Kar09] Andreas Karlsson. Bootstrap methods for bias correction and confidence interval estimation for nonlinear quantile regression of longitudinal data. Journal of Statistical Computation and Simulation, 79(10):1205–1218, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kulatunga:1996:SSS

[KAS96] D. D. Sarath Kulatunga, Mikiko Asai, and Syoichi Sasbuchi. A simulation study of several tests of the equality of binomial probabilities against ordered alternatives. Journal of Statistical Computation and Simulation, 56(1):53–75, 1996. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kirisci:2005:EGB

[KAST05] Levent Kirisci, Ali A. Al-Subaihi, and Ralph Tarter. Effects of the generalized Box–Cox transformation on Type I error rate and power of Hotelling’s T^{2}. Journal of Statistical Computation and Simulation, 75(3):199–206, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Katholi:1978:CVP

[Kat78] Charles R. Katholi. On the computation of values of the psi function from rapidly converging power series expansions. Journal of Statistical Computation and Simulation, 8(1):25–42, 1978. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Khazenie:1987:BVB

[KB87] Nahid Khazenie and Ronald W. Butler. Bounds on the P -value of the best fitting regressor variable calculated to a prespecified accuracy. Journal of Statistical Computation and Simulation, 27 (4):293–306, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Koval:1996:EKE

[KB96] John J. Koval and Nicole J. M. Blackman. Estimators of kappaexact small sample properties. Journal of Statistical Computation and Simulation, 55(4):315–336, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kemp:1973:RPT

[KC73] K. E. Kemp and W. J. Conover. Robustness and power of the ttest compared with some nonparametric alternatives when sampling from a Poisson distribution. Journal of Statistical Computation and Simulation, 2(4):293–307, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Keating:1989:CJS

[KC89] Jerome P. Keating and Veronica Czitrom. A comparison of James–Stein regression with least squares in the Pitman nearness sense. Journal of Statistical Computation and Simulation, 34(1):1–9, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:1996:TPP

[KC96] Seock-Ho Kim and Allan S. Cohen. Table of percentage points of the Behrens–Fisher distribution. Journal of Statistical Computation and Simulation, 55(3):181–187, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Koutrouvelis:1997:ETP

[KC97a] Ioannis A. Koutrouvelis and George C. Canavos. Estimation in the three-parameter gamma distribution based on the empirical moment generation function. Journal of Statistical Computation and Simulation, 59(1):47–62, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kraan:1997:PPT

[KC97b] Bernd Kraan and Roger Cooke. Post-processing techniques for the joint CEC/USNRC uncertainty analysis of accident consequence codes. Journal of Statistical Computation and Simulation, 57(1-4):243–259, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kuk:1997:MCN

[KC97c] Anthony Y. C. Kuk and Yuk W. Cheng. The Monte Carlo Newton–Raphson algorithm. Journal of Statistical Computation and Simulation, 59(3):233–250, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2002:CEM

[KC02] Keeyoung Kim and Byungjin Choi. A comparison of estimators for the mean in the inverse Gaussian distribution with a known coefficient of variation. Journal of Statistical Computation and Simulation, 72(12):899–908, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kachitvichyanukul:1988:FPB

[KCS88] Voratas Kachitvichyanukul, Shiow-Wen J. Cheng, and Bruce Schemeiser. Fast Poisson and binomial algorithms for correlation induction. Journal of Statistical Computation and Simulation, 29(1):17–33, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kadilar:2003:MAI

[KE03] Cem Kadilar and Cenap Erdemìr. Modification of the Akaike information criterion to account for seasonal effects. Journal of Statistical Computation and Simulation, 73(2):135–143, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kelton:1994:EEM

[Kel94] Christina M. L. Kelton. Entry and exit in Markov-process models estimated from macro data, with applications to shifts between advertising media. Journal of Statistical Computation and Simulation, 50(3–4):213–233, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kempthorne:1984:CNF

[Kem84] Oscar Kempthorne. C194. A nearly flat prior, some unsolved problems, and some comments about improper priors. Journal of Statistical Computation and Simulation, 19(2):177–182, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kendall:1979:CLR

[Ken79a] M. G. Kendall. C30. Linear relations with all variables subject to error: rough and ready methods. Journal of Statistical Computation and Simulation, 9(1):69, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kendall:1979:CMC

[Ken79b] M. G. Kendall. C31. Monte Carlo for tails. Journal of Statistical Computation and Simulation, 9(1):69, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kendall:1979:CTV

[Ken79c] M. G. Kendall. C32. Tests for validity of regression models. Journal of Statistical Computation and Simulation, 9(1):70, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kendall:1979:CCC

[Ken79d] M. G. Kendall. C33. Contributing cells for high X^{2} in contingency tables. Journal of Statistical Computation and Simulation, 9(1):70, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kendall:1979:CPB

[Ken79e] M. G. Kendall. C34. L. S. Penrose’s billiard-table problem. Journal of Statistical Computation and Simulation, 9(1):70, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kendall:1979:CRV

[Ken79f] M. G. Kendall. C39. Regression with variable variances. Journal of Statistical Computation and Simulation, 9(1):74, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kershaw:1987:CEE

[Ker87] C. D. Kershaw. A comparison of estimators of the ED50 in up-and-down experiments. Journal of Statistical Computation and Simulation, 27(2):175–184, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kotze:1980:CPL

[KG80] T. J. V. W. Kotze and D. V. Gokhale. A comparison of the Pearson-χ^{2} and log-likelihood-ratio statistics for small samples by means of probability ordering. Journal of Statistical Computation and Simulation, 12(1):1–13, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Komori:2004:EEN

[KH04] Yoshio Komori and Hideo Hirose. Easy estimation by a new parameterization for the three-parameter lognormal distribution. Journal of Statistical Computation and Simulation, 74(1):63– 74, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2009:ESP

[KH09] Chansoo Kim and Keunhee Han. Estimation of the scale parameter of the Rayleigh distribution with multiply Type-II censored sample. Journal of Statistical Computation and Simulation, 79 (8):965–976, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kane:2004:RCG

[KHA04] Mark Kane, John Holt, and Brian Allen. Results concerning the generalized partially linear single-index model. Journal of Statistical Computation and Simulation, 74(12):897–912, 2004. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Khedhiri:2008:SEP

[Khe08] Sami Khedhiri. Simulation experiments on the performance of structural change tests in cointegration. Journal of Statistical Computation and Simulation, 78(7):585–593, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kodell:1983:CSP

[KHSG83] Ralph L. Kodell, M. Gary Haskin, Gray W. Shaw, and David W. Gaylor. Chronic: A SAS procedure for statistical analysis of carcinogenesis studies. Journal of Statistical Computation and Simulation, 16(3–4):287–310, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kibria:2004:PSP

[Kib04] B. M. Golam Kibria. Performance of the shrinkage preliminary test ridge regression estimators based on the conflicting of w, LR and LM tests. Journal of Statistical Computation and Simulation, 74(11):793–810, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kimber:1983:CSR

[Kim83] A. C. Kimber. Comparison of some robust estimators of scale in gamma samples with known shape. Journal of Statistical Computation and Simulation, 18(4):273–286, 1983. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:1992:MIC

[Kim92] Honggie Kim. Measures of influence in correspondence analysis. Journal of Statistical Computation and Simulation, 40(3–4):201– 217, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:1993:DPA

[Kim93] In-Moo Kim. A dynamic programming approach to the estimation of Markov switching regression models. Journal of Statistical Computation and Simulation, 45(1–2):61–76, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:1998:BOS

[Kim98] Jae H. Kim. Bootstrap order selection for autoregressive models. Journal of Statistical Computation and Simulation, 61(3):219– 235, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2000:APS

[Kim00] Sung-Soo Kim. All possible subset regressions using the triangular decomposition. Journal of Statistical Computation and Simulation, 65(1-4):81–94, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2007:APC

[Kim07] Yang-Jin Kim. Analysis of panel count data with measurement errors in the covariates. Journal of Statistical Computation and Simulation, 77(2):109–117, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krewski:1981:SSD

[KJ81] Daniel Krewski and Bruce Junkins. Sample size determination for the interval estimation of the mean or median of a distribution. Journal of Statistical Computation and Simulation, 13 (3–4):169–179, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ko:2009:URD

[KJ09] Young-Hyun Ko and Chi-Hyuck Jun. Use of reference distributions when dealing with unknown regression errors. Journal of Statistical Computation and Simulation, 79(10):1195–1204, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kelton:1985:DSH

[KK85] Christina M. L. Kelton and W. David Kelton. Development of specific hypothesis tests for estimated Markov chains. Journal of Statistical Computation and Simulation, 23(1–2):15–39, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kemp:1990:CSA

[Kk90] Adrienne W. Kemp and C. David kemp. A composition-search algorithm for low-parameter Poisson generation. Journal of Statistical Computation and Simulation, 35(3–4):239–244, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kelton:1991:CMV

[KK91] Christina M. L. Kelton and W. David Kelton. A comparison of micro versus macro point estimators for Markov-process models. Journal of Statistical Computation and Simulation, 38(1-4):201– 210, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kelton:1993:CSS

[KK93] Christina M. L. Kelton and W. David Kelton. Closure-state specification for Markov-process models with incomplete micro data. Journal of Statistical Computation and Simulation, 47(3– 4):195–217, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kang:1999:UEV

[KK99] Seung-Ho Kang and Jerome H. Klotz. Updating exact P -values for the conditional likelihood ratio test of independence. Journal of Statistical Computation and Simulation, 63(3):201–215, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2005:FEE

[KK05] Seong Ho Kim and Sung-Ho Kim. Fluctuation of estimates in an EM procedure for categorical data. Journal of Statistical Computation and Simulation, 75(12):941–957, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kocherlakota:1985:ENS

[KKB85] S. Kocherlakota, K. Kocherlakota, and N. Balakrishnan. Effects of nonnormality on the SPRT for the correlation coefficient: bivariate Edgeworth series distribution. Journal of Statistical Computation and Simulation, 23(1–2):41–51, 1985. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kapanoglu:2007:GAP

[KKE07] Muzaffer Kapanoglu, Ilker Ozan Koc, and Senol Erdogmus. Genetic algorithms in parameter estimation for nonlinear regression models: an experimental approach. Journal of Statistical Computation and Simulation, 77(10):851–867, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Koolaard:1993:EER

[KL93] J. P. Koolaard and C. R. O. Lawoko. Estimating error rates in discriminant analysis with correlated training observations: a simulation study. Journal of Statistical Computation and Simulation, 48(1–2):81–99, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Koo:1994:BBS

[KL94] Ja-Yong Koo and Youngjo Lee. Bivariate B-splines in generalised linear models. Journal of Statistical Computation and Simulation, 50(1–2):119–129, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krishnamoorthy:2005:CCE

[KL05] K. Krishnamoorthy and Yong Lu. On combining correlated estimators of the common mean of a multivariate normal distribution. Journal of Statistical Computation and Simulation, 75 (5):333–345, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kwok:2009:NDC

[KL09] Simon Sai Man Kwok and Wai Keung Li. A note on diagnostic checking of the double autoregressive model. Journal of Statistical Computation and Simulation, 79(5):705–715, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kleijnen:1997:SAR

[Kle97] Jack P. C. Kleijnen. Sensitivity analysis and related analyses: A review of some statistical techniques. Journal of Statistical Computation and Simulation, 57(1-4):111–142, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kemp:1983:CEA

[KM83] K. E. Kemp and Mervyn G. Marasinghe. Computational efficiency of Anova programs based on a reparameterization. Journal of Statistical Computation and Simulation, 17(2):79–89, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Koutrouvelis:1999:TSB

[KM99] Ioannis A. Koutrouvelis and Simos G. Meintanis. Testing for stability based on the empirical characteristic function with applications to financial data. Journal of Statistical Computation and Simulation, 64(4):275–300, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Khaniyev:2006:SCE

[KM06] Tahir Khaniyev and Zulfiyya Mammadova. On the stationary characteristics of the extended model of type (s,S) with Gaussian distribution of summands. Journal of Statistical Computation and Simulation, 76(10):861–874, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Keller-McNulty:1987:EFC

[KMK87] Sallie Keller-McNulty and W. J. Kennedy. Error-free computation of the Moore–Penrose inverse with application to linear least squares analysis. Journal of Statistical Computation and Simulation, 27(1):45–64, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Korhonen:1989:PDR

[KN89] Pekka J. Korhonen and Subhash C. Narula. The probability distribution of the ratio of the absolute values of two normal variables. Journal of Statistical Computation and Simulation, 33(3):173–182, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kohn:1981:NDL

[Koh81] R. Kohn. On nesting do loops to any depth. Journal of Statistical Computation and Simulation, 14(1):41–45, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kossler:2006:SSR

[Kös06] Wolfgang Kössler. Some c-sample rank tests of homogeneity against umbrella alternatives with unknown peak. Journal of Statistical Computation and Simulation, 76(1):57–74, 2006. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Koffler:1979:NDP

[KP79] Stephen L. Koffler and Douglas A. Penfield. Nonparametric discrimination procedures for non-normal distributions. Journal of Statistical Computation and Simulation, 8(3–4):281–299, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Koffler:1982:NCB

[KP82] Stephen L. Koffler and Douglas A. Penfield. Nonparametric classification based upon inverse normal scores and rank transformations. Journal of Statistical Computation and Simulation, 15(1):51–68, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kopecky:1987:IMC

[KP87a] Kenneth J. Kopecky and Dale L. Preston. Improved Monte Carlo estimation of statistical significance for tests of trend in rates proportions. Journal of Statistical Computation and Simulation, 27(4):339–361, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Koreisha:1987:EPM

[KP87b] Sergio Koreisha and Tarmo Pukkila. Estimation of the polynomial matrices of vector moving average processes. Journal of Statistical Computation and Simulation, 28(4):313–343, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Koo:1996:BSD

[KP96a] Ja-Yong Koo and Byeong U. Park. B-spline deconvolution based on the EM algorithm. Journal of Statistical Computation and Simulation, 54(4):275–288, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kuhn:1996:CDB

[KP96b] J. W. Kuhn and W. J. Padgett. C440. On data- based bandwidth selection for density estimation from censored samples: a simulation. Journal of Statistical Computation and Simulation, 55(3):239–245, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Koreisha:1999:SOI

[KP99] Sergio G. Koreisha and Tarmo Pukkila. The selection of the order and identification of nonzero elements in the polynomial matrices of vector autoregressive processes. Journal of Statistical Computation and Simulation, 62(3):207–235, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Koreisha:2004:SVA

[KP04] Sergio G. Koreisha and Tarmo Pukkila. The specification of vector autoregressive moving average models. Journal of Statistical Computation and Simulation, 74(8):547–565, 2004. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kakai:2009:ECE

[KP09] R. L. Glele Kakaï and R. Palm. Empirical comparison of error rate-estimators in logistic discriminant analysis. Journal of Statistical Computation and Simulation, 79(2):111–120, 2009. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2005:NVS

[KPH05] Tae Yoon Kim, Cheolyong Park, and S. Y. Hwang. New view on smoothing parameter selector in function estimation. Journal of Statistical Computation and Simulation, 75(1):55–64, 2005. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Koo:2009:CSM

[KPJ09] Ja-Yong Koo, Changyi Park, and Myoungshic Jhun. A classification spline machine for building a credit scorecard. Journal of Statistical Computation and Simulation, 79(5):681–689, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kamgar-Parsi:1995:DMW

[KPKPB95] Behzad Kamgar-Parsi, Behrooz Kamgar-Parsi, and Menashe Brosh. Distribution and moments of the weighted sum of uniforms random variables, with applications in reducing Monte Carlo simulations. Journal of Statistical Computation and Simulation, 52(4):399–414, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kuhn:2001:AEC

[KPS01] J. W. Kuhn, W. J. Padgett, and J. G. Surles. Absolute error criteria for bandwidth selection in density estimation from censored data. Journal of Statistical Computation and Simulation, 70(3):215–230, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Keilson:1983:SPS

[KPSW83] J. Keilson, D. Petrondas, U. Sumita, and J. Wellner. Significance points for some tests of uniformity on the sphere. Journal of Statistical Computation and Simulation, 17(3):195–218, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krishnamoorthy:1989:ECM

[KR89] K. Krishnamoorthy and Vijay K. Rohatgi. Estimation of the common mean of a bivariate normal distribution. Journal of Statistical Computation and Simulation, 31(3):187–194, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kunst:2002:DSU

[KR02] Robert Kunst and Michael Reutter. Decisions on seasonal unit roots. Journal of Statistical Computation and Simulation, 72 (5):403–418, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krikelis:1977:BDR

[Kri77] N. J. Krikelis. Bayes decision rule for a finite mixed population. Journal of Statistical Computation and Simulation, 6(1):11–18, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krutchkoff:1972:CPC

[Kru72] Richard G. Krutchkoff. The calibration problem and closeness. Journal of Statistical Computation and Simulation, 1(1):87–95, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krutchkoff:1973:BR

[Kru73] Richard G. Krutchkoff. Book reviews. Journal of Statistical Computation and Simulation, 2(2):191, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krutchkoff:1979:CPC

[Kru79] Richard G. Krutchkoff. C36. Partial censoring. Journal of Statistical Computation and Simulation, 9(1):72, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krutchkoff:1986:CCC

[Kru86a] R. G. Krutchkoff. Comments, conjectures, and conclusions C259. The mean likelihood estimator. Journal of Statistical Computation and Simulation, 25(3–4):297–298, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krutchkoff:1986:CPS

[Kru86b] Richard G. Krutchkoff. C262. Percentiles by simulation: reducing time and storage. Journal of Statistical Computation and Simulation, 25(3–4):304–305, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krutchkoff:1988:OWF

[Kru88] R. G. Krutchkoff. One-way fixed effects analysis of variance when the error variances may be unequal. Journal of Statistical Computation and Simulation, 30(4):259–271, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krutchkoff:1989:TWF

[Kru89a] R. S. Krutchkoff. Two-way fixed effects analysis of variance when the error variances may be unequal. Journal of Statistical Computation and Simulation, 32(3):177–183, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krutchkoff:1989:CKP

[Kru89b] Richard G. Krutchkoff. C308. K-philosophy. Journal of Statistical Computation and Simulation, 31(1):44–46, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krutchkoff:1990:CCU

[Kru90a] Richard G. Krutchkoff. C358. Comparisons using simulation. Journal of Statistical Computation and Simulation, 36(1):47– 48, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krutchkoff:1990:CPF

[Kru90b] Richard G. Krutchkoff. C362. Percentiles: fixed width estimation. Journal of Statistical Computation and Simulation, 36(1): 54–56, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krutchkoff:2000:SSS

[Kru00] Richard G. Krutchkoff. Are sufficient statistics sufficient. Journal of Statistical Computation and Simulation, 66(4):321–329, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krutchkoff:2003:CDL

[Kru03] Richard Krutchkoff. C471. Does the law of large numbers really say that? Journal of Statistical Computation and Simulation, 73(8):606–608, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krzanowski:1982:BGC

[Krz82] W. J. Krzanowski. Between-group comparison of principal components — some sampling results. Journal of Statistical Computation and Simulation, 15(2-3):141–154, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krzanowski:1983:CVC

[Krz83] W. J. Krzanowski. Cross-validatory choice in principal component analysis; some sampling results. Journal of Statistical Computation and Simulation, 18(4):299–314, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kachitvichyanukul:1985:CGH

[KS85] Voratas Kachitvichyanukul and Burce Schmeiser. Computer generation of hypergeometric random variates. Journal of Statistical Computation and Simulation, 22(2):127–145, 1985. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kelton:1987:EEM

[KS87a] Christina M. L. Kelton and Marlene A. Smith. Estimation efficiency for Markov chain models. Journal of Statistical Computation and Simulation, 28(2):145–165, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Klonias:1987:NTN

[KS87b] V. K. Klonias and Nash G. Stephen. Numerical techniques in nonparametric estimation. Journal of Statistical Computation and Simulation, 28(2):97–126, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kelton:1991:SIN

[KS91] Christina M. L. Kelton and Marlene A. Smith. Statistical inference in nonstationary Markov models with embedded explanatory variables. Journal of Statistical Computation and Simulation, 38(1-4):25–44, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kelley:1997:NAS

[KS97] D. Lynn Kelley and Shlomo S. Sawilowsky. Nonparametric alternatives to the f statistic in analysis of variance. Journal of Statistical Computation and Simulation, 58(4):343–359, 1997. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kumbhar:2004:TLL

[KS04] R. R. Kumbhar and D. T. Shirke. Tolerance limits for lifetime distribution of k-unit parallel system. Journal of Statistical Computation and Simulation, 74(3):201–213, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shiue:1990:SAI

[kSB90] Wei kei Shiue and Lee J. Bain. Simple approximate inference procedures for the mean of the gamma distribution. Journal of Statistical Computation and Simulation, 34(2-3):67–73, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kihlbergt:1973:EIS

[KSH73] Jaakko K. Kihlbergt, William E. Schotz, and Jay Herson. The effect of ignoring a simple dichotomy in sequential analysis. Journal of Statistical Computation and Simulation, 2(2):149–157, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kulkarni:1993:CAF

[KSK93] Pandurang M. Kulkarni, Arvind K. Shah, and Debashis Kushary. C392. On averaging Y ’s at fixed X’s in regression. Journal of Statistical Computation and Simulation, 44(3–4):255– 258, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shiue:1993:BCI

[kSwXR93] Wei kei Shiue, Chong wei Xu, and B. Charles Rea. Bootstrap confidence intervals for simulation outputs. Journal of Statistical Computation and Simulation, 45(3–4):249–255, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kutas:1985:SMP

[KT85] Tibor Kutas and János Tóth. A stochastic model of phytoplankton dynamics in Lake Balaton. Journal of Statistical Computation and Simulation, 21(3–4):241–264, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kwon:1994:CCM

[KT94] Chimyung Kwon and Jeffrey D. Tew. Combined correlation methods for metamodel estimation in multipopulation simulation experiments. Journal of Statistical Computation and Simulation, 49(1–2):49–75, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kallenberg:1997:DDS

[KT97] Wilbert C. M. Kallenberg and Ledwina Teresa. Data driven smooth tests for composite hypotheses comparison of powers. Journal of Statistical Computation and Simulation, 59(2):101– 121, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kahaner:1982:GQF

[KTB82] David Kahaner, Gary Tietjen, and Richard Beckman. Gaussianquadrature formulas for ∫_{0}^{∞} exp(−x^{2})g(x) dx. Journal of Statistical Computation and Simulation, 15(2-3):155–160, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kumar:2005:JST

[KTM05] Somesh Kumar, Yogesh Mani Tripathi, and Neeraj Misra. James–Stein type estimators for ordered normal means. Journal of Statistical Computation and Simulation, 75(7):501–511, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kuk:1987:BEV

[Kuk87] Anthony Y. C. Kuk. Bootstrap estimators of variance under sampling with probability proportional to aggregate size. Journal of Statistical Computation and Simulation, 28(4):303–311, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kuk:1989:DBE

[Kuk89] Anthony Y. C. Kuk. Double bootstrap estimation of variance under systematic sampling with probability proportional to size. Journal of Statistical Computation and Simulation, 31(2):73–82, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kuk:1999:LIS

[Kuk99] Anthony Y. C. Kuk. Laplace importance sampling for generalized linear mixed models. Journal of Statistical Computation and Simulation, 63(2):143–158, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kulperger:1990:DMB

[Kul90] R. J. Kulperger. On the distribution of the maximum of Brownian bridges with application to regression with correlated errors. Journal of Statistical Computation and Simulation, 34(2-3):97– 106, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kundu:1992:DNS

[Kun92] Debasis Kundu. Detecting the number of signals for undamped exponential models using information theoretic criteria. Journal of Statistical Computation and Simulation, 44(1–2):117–131, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kunst:1993:ASI

[Kun93] Robert M. Kunst. Apparently stable increments in finance data: could arch effects be the cause? Journal of Statistical Computation and Simulation, 45(1–2):121–127, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kundu:1998:ENS

[Kun98] Debasis Kundu. Estimating the number of sinusoids and its performance analysis. Journal of Statistical Computation and Simulation, 60(4):347–362, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kahaner:1978:CAG

[KW78] D. K. Kahaner and M. S. Waterman. C8. Asymptotics for gamma and beta percentiles. Journal of Statistical Computation and Simulation, 6(3–4):315–316, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krus:1994:MDS

[KW94] David J. Krus and Jmaes M. Webb. Maxwell’s demons: simulating probability distributions with functions of propositional calculus. Journal of Statistical Computation and Simulation, 51(1):71–77, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kao:1996:EAP

[KW96] Chiang Kao and J. Y. Wong. An exhaustive analysis of prime modulus multiplicative congruential random number generators with modulus smaller than 2^{15}. Journal of Statistical Computation and Simulation, 54(1-3):29–35, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kuhl:2000:LSE

[KW00] Michael E. Kuhl and James R. Wilson. Least squares estimation of nonhomogeneous Poisson processes. Journal of Statistical Computation and Simulation, 67(1):699–712, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

King:2004:ODP

[KW04] Joy King and Weng Kee Wong. Optimal designs for the power logistic model. Journal of Statistical Computation and Simulation, 74(11):779–791, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kwong:1995:EOS

[Kwo95] Koon-Shing Kwong. Evaluation of the one-sided percentage points of the singular multivariate normal distribution. Journal of Statistical Computation and Simulation, 51(2-4):121–135, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kwong:1996:NSC

[Kwo96] Koon-Shing Kwong. A note on simultaneous confidence intervals for multinomial proportions. Journal of Statistical Computation and Simulation, 55(1–2):111–120, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Karlis:2003:ZFA

[KX03] Dimitris Karlis and Evdokia Xekalaki. A zero frequency alternative method to the moment method of estimation in finite Poisson mixtures. Journal of Statistical Computation and Simulation, 73(6):409–427, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kuo:1993:EBR

[KY93] Lynn Kuo and Constantin Yiannoutsos. Empirical Bayes risk evaluation with Type II censored data. Journal of Statistical Computation and Simulation, 48(3–4):195–206, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2008:ISL

[KYF08] Hyune-Ju Kim, Binbing Yu, and Eric J. Feuer. Inference in segmented line regression: a simulation study. Journal of Statistical Computation and Simulation, 78(11):1087–1103, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lachenbruch:1990:CNG

[Lac90] Peter A. Lachenbruch. C355. A note on goodness-of-link tests. Journal of Statistical Computation and Simulation, 36(1):43– 45, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lachenbruch:1992:PTW

[Lac92] Peter A. Lachenbruch. The performance of tests when observations have different variances. Journal of Statistical Computation and Simulation, 40(1–2):83–92, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Laird:1982:CVC

[Lai82a] Nan M. Laird. Computation of variance components using the em algorithm. Journal of Statistical Computation and Simulation, 14(3–4):295–303, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Laird:1982:EBE

[Lai82b] Nan M. Laird. Empirical Bayes estimates using the nonparametric maximum likelihood estimate for the prior. Journal of Statistical Computation and Simulation, 15(2-3):211–220, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lakhan:1981:GAP

[Lak81] V. Chris Lakhan. Generating autocorrelated pseudo-random numbers with specific distributions. Journal of Statistical Computation and Simulation, 12(3–4):303–309, 1981. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lamotte:2005:NSD

[Lam05] Lynn R. Lamotte. A note on separated data and exact likelihood-ratio p-values in logistic regression. Journal of Statistical Computation and Simulation, 75(8):667–672, 2005. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2009:AET

[LAR09] Ming-Hsien Caleb Li and Abbas Al-Refaie. The alpha error of the Taguchi method with L_{16} array for the LTB response variable using simulation. Journal of Statistical Computation and Simulation, 79(5):645–656, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Latorre:1982:EPD

[Lat82] Giovanni Latorre. The exact posterior distribution of the crossratio of a 2 × 2 contingency table. Journal of Statistical Computation and Simulation, 16(1):19–24, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lau:1979:SVR

[Lau79] William W. Lau. Sequential vs. random retrieval of a large indexed file. Journal of Statistical Computation and Simulation, 10(1):53–64, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lawal:1992:MTW

[Law92] Bayo H. Lawal. A modified χ^{2} test when some cells have small expectations in the multinomial distribution. Journal of Statistical Computation and Simulation, 40(1–2):15–27, 1992. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lawal:2001:MSM

[Law01] H. Bayo Lawal. Modeling symmetry models in square contingency tables with ordered categories. Journal of Statistical Computation and Simulation, 71(1):59–83, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:1996:ACB

[LB96] Xihong Lin and Norman E. Breslow. Analysis of correlated binomial data in logistic-normal models. Journal of Statistical Computation and Simulation, 55(1–2):133–146, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:2001:ELI

[LB01] C.-T. Lin and N. Balakrishnan. Exact linear inference for scaled exponential distribution based on doubly Type-II censored samples. Journal of Statistical Computation and Simulation, 71(3): 183–199, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Low:2007:OSB

[LB07] C. K. Low and U. Balasooriya. Order statistics based sampling design for reliability sampling. Journal of Statistical Computation and Simulation, 77(8):709–715, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lebre:2008:EAE

[LB08a] Sophie Lèbre and Pierre-Yves Bourguignon. An EM algorithm for estimation in the mixture transition distribution model. Journal of Statistical Computation and Simulation, 78(8):713– 729, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2008:BLU

[LB08b] Tao Li and Narayanaswamy Balakrishnan. Best linear unbiased estimators of parameters of a simple linear regression model based on ordered ranked set samples. Journal of Statistical Computation and Simulation, 78(12):1267–1278, 2008. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Leeper:1992:CGM

[LC92] James D. Leeper and Suming W. Chang. Comparison of general multivariate and random-effects models for growth curve analysis: incomplete-data smallsample situations. Journal of Statistical Computation and Simulation, 44(1–2):93–104, 1992. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lai:2002:TCB

[LC02] Dejian Lai and Guanrong Chen. Testing chaos based on empirical distribution function: A simulation study. Journal of Statistical Computation and Simulation, 72(1):77–85, 2002. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Loader:1987:CBC

[LD87] C. R. Loader and J. J. Deely. Computations of boundary crossing probabilities for the Wiener process. Journal of Statistical Computation and Simulation, 27(2):95–105, 1987. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lohr:1997:CCI

[LD97] Sharon L. Lohr and Michelle Divan. Comparison of confidence intervals for variance components with unbalanced data. Journal of Statistical Computation and Simulation, 58(1):83–97, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lodell:1973:TLD

[LDC73] P. L. Lodell, B. S. Duran, and W. A. Coberly. On the table lookup in discriminate analysis. Journal of Statistical Computation and Simulation, 2(2):171–188, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lad:1992:NAF

[LDR92] Frank Lad, James M. Dickey, and Mohammad A. Rahman. Numerical application of the fundamental theorem of prevision. Journal of Statistical Computation and Simulation, 40(3–4):135– 151, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

LEcuyer:1997:TBS

[L’E97] Piette L’Ecuyer. Tests based on sum-functions of spacings for uniform random numbers. Journal of Statistical Computation and Simulation, 59(3):251–269, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:1981:MMC

[Lee81] Sik-Yum Lee. The multiplier method in constrained estimation of covariance structure models. Journal of Statistical Computation and Simulation, 12(3–4):247–257, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:1985:MLE

[Lee85] Sik-Yum Lee. Maximum likelihood estimation of polychoric correlations in r × s × t contingency tables. Journal of Statistical Computation and Simulation, 23(1–2):53–67, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:1990:CVS

[Lee90] David Kuo Chuen Lee. Corss-validation in semiparametric models: some Monte Carlo results. Journal of Statistical Computation and Simulation, 37(3–4):171–187, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:1992:CCN

[Lee92] Carl M. S. Lee. On the computation of central and noncentral beta probabilities. Journal of Statistical Computation and Simulation, 43(1–2):1–10, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:1998:CCM

[Lee98a] Sauchi Stephen Lee. C447. Combining models to improve classification rates. Journal of Statistical Computation and Simulation, 60(2):160–164, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:1998:CRQ

[Lee98b] Sauchi Stephen Lee. C451. Regularized quadratic discriminant methods for skewed binary classification. Journal of Statistical Computation and Simulation, 60(2):176–180, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:1999:NEP

[Lee99] S. J. Lee. A note on the exact prediction limits for future observations from the inverse Gaussian distribution. Journal of Statistical Computation and Simulation, 63(1):1–12, 1999. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2001:ATL

[Lee01] S. J. Lee. On the assessment of tolerance limits under inverse Gaussian distribution. Journal of Statistical Computation and Simulation, 71(3):245–265, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2002:ISS

[Lee02a] Jae Lee. Importance sampling for signal extraction model in time series analysis. Journal of Statistical Computation and Simulation, 72(11):881–894, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2002:AOL

[Lee02b] Thomas Lee. On algorithms for ordinary least squares regression spline fitting: A comparative study. Journal of Statistical Computation and Simulation, 72(8):647–663, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lewis:1985:CNO

[LEG85] Jerry W. Lewis, Luis A. Escobar, and James P. Geeghan. A consequence of non-orthogonal data. Journal of Statistical Computation and Simulation, 22(1):51–66, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Legendre:2000:CPM

[Leg00] Pierre Legendre. Comparison of permutation methods for the partial correlation and partial Mantel tests. Journal of Statistical Computation and Simulation, 67(1):37–73, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Leish:1983:NNP

[Lei83] D. L. Mc Leish. A note on non-parametric censored regression. Journal of Statistical Computation and Simulation, 18(1):1–6, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lemmer:1977:DFO

[Lem77] H. H. Lemmer. A distribution-free one-sample test for dispersion. Journal of Statistical Computation and Simulation, 5(4): 303–320, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lemmer:1980:ES

[Lem80] H. H. Lemmer. An estimator for spread. Journal of Statistical Computation and Simulation, 11(2):107–117, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lemmer:1987:MMW

[Lem87] H. H. Lemmer. A modified Mann–Whitney–Wilcoxon test for the two-sample location problem. Journal of Statistical Computation and Simulation, 27(4):307–319, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Levy:1978:ECA

[Lev78a] Kenneth J. Levy. An empirical comparison of the ANOVA F -test with alternatives which are more robust against heterogeneity of variance. Journal of Statistical Computation and Simulation, 8 (1):49–57, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Levy:1978:ESC

[Lev78b] Kenneth J. Levy. An empirical study of the cube-root test for homogeneity of variance with respect to the effects of nonnormality and power. Journal of Statistical Computation and Simulation, 7(1):71–78, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Levy:1978:PTE

[Lev78c] Kenneth J. Levy. A procedure for testing the equality of k dependent correlation coefficients. Journal of Statistical Computation and Simulation, 7(3–4):181–187, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Levy:1978:SEP

[Lev78d] Kenneth J. Levy. Some empirical power results associated with Welch’s robust analysis of variance technique. Journal of Statistical Computation and Simulation, 8(1):43–48, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Levin:1982:ANA

[Lev82] Bruce Levin. On the accuracy of a normal approximation to the power of the Mantel–Haenszel procedure. Journal of Statistical Computation and Simulation, 14(3–4):201–218, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Levine:2005:NMC

[Lev05] Richard A. Levine. A note on Markov chain Monte Carlo sweep strategies. Journal of Statistical Computation and Simulation, 75(4):253–262, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:1980:MCS

[LF80] Hyunshik Lee and Karen Yuen Fung. A Monte Carlo study on the robustness of the two-sample sequential t test. Journal of Statistical Computation and Simulation, 10(3–4):297–307, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lim:1982:STB

[LF82] T. K. Lim and Karen Yung Fung. Sequential t tests based on m estimators. Journal of Statistical Computation and Simulation, 14(3–4):271–281, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:1983:RPM

[LF83] Hyunshik Lee and Karen Yuen Fung. Robust procedures for multi-sample location problems with unequal group variances. Journal of Statistical Computation and Simulation, 18(2-3):125– 143, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:1997:NSM

[LF97] Hoonja Lee and Robert V. Foutz. A new spectral model for binary or categorical-valued time series data. Journal of Statistical Computation and Simulation, 58(3):217–235, 1997. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Levine:2004:AMC

[LF04] Richard A. Levine and Juanjuan Fan. An automated Markov chain Monte Carlo EM algorithm. Journal of Statistical Computation and Simulation, 74(5):349–360, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Laurencelle:1992:DBM

[LFC92] Louis Laurencelle, Jean-Yves Frigon, and Guy Châtillon. Difference between the mean calculated from raw data and from a frequency distribution: a new approach. Journal of Statistical Computation and Simulation, 41(3–4):201–218, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Looney:1984:RTF

[LG84] Stephen W. Looney and Thomas R. Gulledge, Jr. Regression tests of fit and probability plotting positions. Journal of Statistical Computation and Simulation, 20(2):115–127, 1984. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2009:PIC

[LG09] Hyeyoung Lee and Sujit K. Ghosh. Performance of information criteria for spatial models. Journal of Statistical Computation and Simulation, 79(1):93–106, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lawrence:1990:FMS

[LGP90] Kenneth Lawrence, Michael Geurts, and I. Robert Parket. Forecasting market share using a combination of time series data and explanatory variables. Journal of Statistical Computation and Simulation, 36(4):247–253, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Luceno:1996:COA

[LGPP96] Albeto Luceño, Francisco J. Gonzalez, and Jaime Puig-Pey. Computing optimal adjustment schemes for the general toolwear problem. Journal of Statistical Computation and Simulation, 54(1-3):87–113, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lauer:1972:JDF

[LH72] G. N. Lauer and C. P. Han. Joint distribution function of certain ratios of chi-square variates. Journal of Statistical Computation and Simulation, 1(3):255–260, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lyons:1978:CCD

[LH78] N. I. Lyons and K. Hutcheson. C20. Comparing diversities: Gini’s index. Journal of Statistical Computation and Simulation, 8(1):75–78, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Layne:1979:STA

[LH79] Benjamin H. Layne and Schuyler W. Huck. Subdivision techniques in applying the Bartlett–Kendall test for equal variances. Journal of Statistical Computation and Simulation, 9(2):83–99, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:1993:ODT

[LH93] H. Andy Lee and Yer Van Hui. Outliers detection in time series. Journal of Statistical Computation and Simulation, 45(1–2):77– 95, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:1994:RRC

[LH94] W. K. Li and Y. V. Hui. Robust residual cross correlation tests for lagged relations in time series. Journal of Statistical Computation and Simulation, 49(1–2):103–109, 1994. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lakshminarayan:1997:EMB

[LH97] Choudur K. Lakshminarayan and Chien-Pai Han. On estimating the mean in a bivariate normal distribution with equal or unequal variances. Journal of Statistical Computation and Simulation, 58(2):155–170, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Leybourne:2005:MPM

[LhKN05] Stephen J. Leybourne, Tae hwan Kim, and Paul Newbold. More powerful modifications of unit root tests allowing structural change. Journal of Statistical Computation and Simulation, 75 (11):869–888, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:1988:IUM

[Li88] Kim-Hung Li. Imputation using Markov chains. Journal of Statistical Computation and Simulation, 30(1):57–79, 1988. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:1994:CIY

[Li94] Kim-Hung Li. A computer implementation of the Yates–Grundy draw by draw procedure. Journal of Statistical Computation and Simulation, 50(3–4):147–151, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2001:TLG

[Li01] Chin-Shang Li. Testing for linearity in generalized linear models using kernel smoothing. Journal of Statistical Computation and Simulation, 69(4):339–351, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liang:2011:CBP

[Lia11] TaChen Liang. Comments on ‘Life prediction under random censorship’. Journal of Statistical Computation and Simulation, 81(6):795–797, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Wan08c].

Lillegaard:2001:TBM

[Lil01] Magnar Lillegård. Tests based on Monte Carlo simulations conditioned on maximum likelihood estimates of nuisance parameters. Journal of Statistical Computation and Simulation, 71(1): 1–10, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:1995:HAC

[LiT95] Shang P. Lin and Dei in Tang. A hierarchical active constraints search algorithm for optimal scaling of ordered categorical responses. Journal of Statistical Computation and Simulation, 53 (3–4):197–209, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:1981:ERC

[Liu81] Lon-Mu Liu. Estimation of random coefficient regression models. Journal of Statistical Computation and Simulation, 13(1):27–39, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:1986:CEA

[Liu86] Shu-Ing Liu. Computational efficiency in all possible regressions. Journal of Statistical Computation and Simulation, 26(3–4):283– 298, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2008:PES

[Liu08] Xuefeng Liu. Parameter expansion for sampling a correlation matrix: an efficient GPX–RPMH algorithm. Journal of Statistical Computation and Simulation, 78(11):1065–1076, 2008. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lund:2000:EBT

[LJ00] Ulric Lund and S. Rao Jammalamadaka. An entropy-based test for goodness of fit of the von Mises distribution. Journal of Statistical Computation and Simulation, 67(4):319–332, 2000. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2002:CIF

[LJDK02] J. W. Lee, S. J. Jo, D. L. DeMets, and K. Kim. Confidence intervals following group sequential tests in clinical trails with multivariate observations. Journal of Statistical Computation and Simulation, 72(3):247–259, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2005:CWS

[LJZB05] W. Liu, M. Jamshidian, Y. Zhang, and F. Bretz. Constant width simultaneous confidence bands in multiple linear regression with predictor variables constrained in intervals. Journal of Statistical Computation and Simulation, 75(6):425–436, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:1983:CGT

[LK83a] C. Charles Lin and W. J. Kennedy. Computer generated test problems for L_{p} estimates in linear models. Journal of Statistical Computation and Simulation, 17(3):219–222, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Loukas:1983:CST

[LK83b] S. Loukas and C. D. Kemp. On computer sampling from trivariate and multivariate discrete distributions multivariate discrete distributions. Journal of Statistical Computation and Simulation, 17(2):113–123, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:1988:NAM

[LK88] Jang Taek Lee and Byung Chun Kim. A new approach for the W -matrix. Journal of Statistical Computation and Simulation, 29(3):241–254, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2007:STO

[LKKL07] Juneyoung Lee, André I. Khuri, Kee Whan Kim, and Sangkon Lee. On the size of the F -test for the one-way random model with heterogeneous error variances. Journal of Statistical Computation and Simulation, 77(6):443–455, 2007. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Learmonth:1973:NPS

[LL73] G. P. Learmonth and P. A. W. Lewis. Naval Postgraduate School random number generator package LLRANDOM. Report NP555LW73061A, Naval Postgraduate School, Monterey, CA, USA, 1973. The shuffling algorithm proposed in this report does not lengthen the period, and only marginally reduces the lattice structure of linear congruential generators, despite the apparently tiny difference with the [BD76] algorithm: see [Bay90] for a comparison, both mathematical, and graphical.

Lemeshow:1979:EVR

[LL79] Stanley Lemeshow and Paul S. Levy. Estimating the variances of ratio estimates in complex sample surveys with two primary sampling units per stratum — a comparison of balanced replication and jackknife techniques. Journal of Statistical Computation and Simulation, 8(3–4):191–205, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:1988:EPC

[LL88] Sik-Yum Lee and Ming-Long Lam. Estimation of polychoric correlation with elliptical latent variables. Journal of Statistical Computation and Simulation, 30(3):173–188, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lo:1990:TNA

[LL90] Chan-Lam Lo and Wai-Keung Li. Two new approaches to robust estimation in time series. Journal of Statistical Computation and Simulation, 34(4):209–255, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lau:1991:EPE

[LL91] Hon-Shiang Lau and Amy Hing-Ling Lau. Effective procedures for estimating beta distribution’s parameters and their confidence intervals. Journal of Statistical Computation and Simulation, 38(1-4):139–150, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lau:1993:RSU

[LL93] Hon-Shiang Lau and Amy Hing-Ling Lau. The reliability of the stability-under-addition test for the stable-Paretian hypothesis. Journal of Statistical Computation and Simulation, 48(1–2):67– 80, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:1999:NBE

[LL99] Jack Lee and Y. L. Lio. A note on Bayesian estimation and prediction for the beta-binomial model. Journal of Statistical Computation and Simulation, 63(1):73–91, 1999. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Love:2003:WPF

[LL03] Thomas Love and Kevin Linderman. A Weibull process failure mechanism for the economic design of MEWMA control charts. Journal of Statistical Computation and Simulation, 73(3):195– 202, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2006:ELF

[LL06] Seung-Hwan Lee and Eun-Joo Lee. On the extended lack-of-fit tests for the censored regression model. Journal of Statistical Computation and Simulation, 76(9):793–801, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2007:CCI

[LL07] Xinmin Li and Guoying Li. Comparison of confidence intervals on the among group variance in the unbalanced variance component model. Journal of Statistical Computation and Simulation, 77(6):477–486, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2008:ESS

[LL08a] Jaechoul Lee and Robert Lund. Equivalent sample sizes in time series regressions. Journal of Statistical Computation and Simulation, 78(4):285–297, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:2008:TDN

[LL08b] Jin-Guan Lin and Yong Li. Testing for departures from nominal dispersion in generalized nonlinear models with varying dispersion and/or additive random effects. Journal of Statistical Computation and Simulation, 78(10):925–939, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Lee:2009:TET

[LL09a] Seung-Hwan Lee and Eun-Joo Lee. On testing equality of two censored samples. Journal of Statistical Computation and Simulation, 79(2):135–143, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2009:HDT

[LL09b] Yong Li and Jin-Guan Lin. Heteroscedasticity diagnostics in two-phase linear regression models. Journal of Statistical Computation and Simulation, 79(7):923–938, 2009. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2009:EMC

[LLKJ09] Sanghan Lee, Johan Lim, Seung-Jean Kim, and Yongsung Joo. Estimating monotone convex functions via sequential shape modification. Journal of Statistical Computation and Simulation, 79(8):989–1000, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lloyd:1993:TRM

[Llo93] Chris J. Lloyd. Testing recapture model M_{T} for departure from homogeneity. Journal of Statistical Computation and Simulation, 46(1–2):35–44, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lloyd:2005:ETP

[Llo05] Chris J. Lloyd. Estimating test power adjusted for size. Journal of Statistical Computation and Simulation, 75(11):921–933, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Levy:1974:CUE

[LM74] Paul S. Levy and Nathan Mantel. Combining unbiased estimates — a further examination of some old estimators. Journal of Statistical Computation and Simulation, 3(2):147–160, 1974. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lal:1984:PIB

[LM84] C. D. Lal and T. Moore. Probability integrals of a bivariate gamma distribution. Journal of Statistical Computation and Simulation, 19(3):205–213, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lachos:2008:ILI

[LMB08] V. H. Lachos, L. C. Montenegro, and H. Bolfarine. Inference and local influence assessment in skew-normal null intercept measurement error model. Journal of Statistical Computation and Simulation, 78(3):395–419, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Levy:1977:MCC

[LN77] Kenneth J. Levy and Subhash C. Narula. A Monte Carlo comparison of several methods for analyzing small sample binomial data in a disproportionate two- factor experiment without replication. Journal of Statistical Computation and Simulation, 5 (3):239–244, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lohrding:1973:TKS

[Loh73] Ronald K. Lohrding. Three Kolmogorov–Smirnov-type onesample tests with improved power properties. Journal of Statistical Computation and Simulation, 2(2):139–148, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lohrding:1975:TST

[Loh75] Ronald K. Lohrding. A two sample test of equality of coefficients of variation or relative errors. Journal of Statistical Computation and Simulation, 4(1):31–36, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Loh:1986:TMN

[Loh86] Wei-Yin Loh. Testing multivariate normality by simulation. Journal of Statistical Computation and Simulation, 26(3–4):243– 252, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lopes:2004:CEM

[LOR04] S. R. C. Lopes, B. P. Olbermann, and V. A. Reisen. A comparison of estimation methods in non-stationary ARFIMA processes. Journal of Statistical Computation and Simulation, 74 (5):339–347, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lotwick:1982:SSS

[Lot82] H. W. Lotwick. Simulation of some spatial hard core models, and the complete packing problem. Journal of Statistical Computation and Simulation, 15(4):295–314, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Loukas:1984:SMC

[Lou84] S. Loukas. Simple methods for computer generation of bivariate beta random variables. Journal of Statistical Computation and Simulation, 20(2):145–152, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Loughin:1995:RBR

[Lou95] Thomas M. Loughin. A residual bootstrap for regression parameters in proportional hazards models. Journal of Statistical Computation and Simulation, 52(4):367–384, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Loukas:1985:EBN

[LP85] S. Loukas and H. Papageorgiou. Estimation for the bivariate negative binomial distribution. Journal of Statistical Computation and Simulation, 22(1):67–82, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lenth:1989:RCI

[LP89] R. V. Lenth and A. R. Padmanabhan. Robust confidence intervals for the center of a symmetric distribution. Journal of Statistical Computation and Simulation, 33(3):183–196, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Littell:1992:RPC

[LP92] Ramon C. Littell and Kenneth M. Portier. Robust parametric confidence bands for distribution functions. Journal of Statistical Computation and Simulation, 41(3–4):187–200, 1992. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lyons:2000:ASM

[LP00] Ben Lyons and Dawn Peters. Applying Skovgaard’s modified directed likelihood statistic to mixed linear models. Journal of Statistical Computation and Simulation, 65(1-4):225–242, 2000.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2009:BST

[LP09] Jialiang Li and Mari Palta. Bandwidth selection through crossvalidation for semi-parametric varying-coefficient partially linear models. Journal of Statistical Computation and Simulation, 79(11):1277–1286, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lydersen:2005:CET

[LPSL05] Stian Lydersen, Vivek Pradhan, Pralay Senchaudhuri, and Petter Laake. Comparison of exact tests for association in unordered contingency tables using standard, mid-p, and randomized test versions. Journal of Statistical Computation and Simulation, 75 (6):447–458, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lipsitz:2002:DFA

[LPZ02] Stuart Lipsitz, Michael Parzen, and Lue Ping Zhao. A degreesof-freedom approximation in multiple imputation. Journal of Statistical Computation and Simulation, 72(4):309–318, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lefante:1986:CNO

[LS86] John J. Lefante, Jr. and Arvind K. Shah. C257. A note on the one-sample t-test. Journal of Statistical Computation and Simulation, 25(3–4):295–296, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comment [Goo86s].

Lesage:1988:CSR

[LS88] James P. Lesage and Stephen D. Simon. Centering and scaling of regression algorithms in the face of ill-conditioning. Journal of Statistical Computation and Simulation, 30(4):273–283, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

lee:1996:TVS

[lS96] S. J. lee and Ping Sa. Testing the variance of symmetric heavytailed distributions. Journal of Statistical Computation and Simulation, 56(1):39–52, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liao:1999:NGF

[LS99] Min Liao and Toshiyuki Shimokawa. A new goodness-of-fit test for Type-I extreme-value and 2-parameter Weibull distributions with estimated parameters. Journal of Statistical Computation and Simulation, 64(1):23–48, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lianfen:2001:BCH

[LS01] Qian Lianfen and Wang Suojin. Bias-corrected heteroscedasticity robust covariance matrix (sandwich) estimators. Journal of Statistical Computation and Simulation, 70(2):161–174, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lemonte:2008:BBI

[LSCN08] Artur J. Lemonte, Alexandre B. Simas, and Francisco Cribari-Neto. Bootstrap-based improved estimators for the twoparameter Birnbaum–Saunders distribution. Journal of Statistical Computation and Simulation, 78(1):37–49, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

LeRoux:1997:VSE

[LSL97] N. J. Le Roux, S. J. Steel, and N. Louw. Variable selection and error rate estimation in discriminant analysis. Journal of Statistical Computation and Simulation, 59(3):195–219, 1997. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:1993:CLZ

[LSS93] A. J. Lee, A. J. Scott, and S. C. Soo. Comparing Liang–Zeger estimates with maximum likelihood in bivariate logistic regression. Journal of Statistical Computation and Simulation, 44(3– 4):133–148, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Leiva:2008:RNG

[LSSP08] Víctor Leiva, Antonio Sanhueza, Pranab K. Sen, and Gilberto A. Paula. Random number generators for the generalized Birnbaum–Saunders distribution. Journal of Statistical Computation and Simulation, 78(11):1105–1118, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

5163. URL http://www.tandfonline.com/doi/abs/10.1080/
00949650701550242.

Liu:1990:SCN

[LT90] Ming Chung Liu and R. Y. Taylor. Simulations and computations of nonparametric density estimates for the deconvolution problem. Journal of Statistical Computation and Simulation, 35 (3–4):145–167, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lyon:1995:ITF

[LT95] John Lyon and Chin-Ling Tsai. Improved tests for the firstorder autoregressive model with heteroscedasticity. Journal of Statistical Computation and Simulation, 52(1):71–83, 1995. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lyambabaje:1990:CSE

[LTV90] A. Lyambabaje, S. Tardif, and C. Van Eeden. A comparison of several estimators of the center of a symmetric distribution. Journal of Statistical Computation and Simulation, 35(1–2):19– 30, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lu:1997:STO

[Lu97] Wang-Shu Lu. Score tests for overdispersion in Poisson regression models. Journal of Statistical Computation and Simulation, 56(3):213–228, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lu:1999:TEB

[Lu99] Wang-Shu Lu. Testing extra-binomial variations. Journal of Statistical Computation and Simulation, 63(1):93–103, 1999. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Luceno:2008:MLV

[Luc08] Alberto Luceño. Maximum likelihood vs. maximum goodness of fit estimation of the three-parameter Weibull distribution.

Journal of Statistical Computation and Simulation, 78(10):941– 949, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lund:2006:SAR

[Lun06] Robert Lund. A seasonal analysis of riverflow trends. Journal of Statistical Computation and Simulation, 76(5):397–405, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Leeper:1982:THG

[LW82] James D. Leeper and Robert F. Woolson. Testing hypotheses for the growth curve model when the data are incomplete. Journal of Statistical Computation and Simulation, 15(2-3):97–107, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:1995:CNN

[LW95] W. Z. Liu and A. P. White. A comparison of nearest neighbour and tree-based methods of non-parametric discriminant analysis. Journal of Statistical Computation and Simulation, 53(1–2): 41–50, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2004:HTS

[LW04] Anna Liu and Yuedong Wang. Hypothesis testing in smoothing spline models. Journal of Statistical Computation and Simulation, 74(8):581–597, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2008:SIL

[LWH08] Wei Liu, H. P. Wynn, and A. J. Hayter. Statistical inferences for linear regression models when the covariates have functional relationships: polynomial regression. Journal of Statistical Computation and Simulation, 78(4):315–324, 2008. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Latham:1996:SSM

[LY96] Geoff A. Latham and S. Yu. N-stage splitting for maximum penalized likelihood estimation. Journal of Statistical Computation and Simulation, 55(1–2):159–168, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Latham:1999:SSM

[LY99] G. A. Latham and S. Yu. N-stage splitting for maximum penalized likelihood estimation with Gaussian data and stationary linear iterative methods. Journal of Statistical Computation and Simulation, 62(4):375–393, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lye:1991:FSP

[Lye91] J. N. Lye. Finite-sample properties of limited-information estimators in misspecified simultaneous equation models. Journal of Statistical Computation and Simulation, 39(4):241–257, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Le:1992:GTE

[LZ92] Nhu D. Le and Ruben H. Zamar. A global test for effects in 2^{k} factorial design without replicates. Journal of Statistical Computation and Simulation, 41(1–2):41–54, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2008:MPD

[LZNL08] Jialiang Li, Chunming Zhang, Erik V. Nordheim, and Charles E. Lehner. On the multivariate predictive distribution of multidimensional effective dose: a Bayesian approach. Journal of Statistical Computation and Simulation, 78(5):429–442, 2008. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mccabe:1974:SSP

[MA74] George P. Mccabe, Jr. and James N. Arvesen. A subset selection procedure for regression variables. Journal of Statistical Computation and Simulation, 3(2):137–146, 1974. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ma:1997:SAM

[Ma97] Chunsheng Ma. Saddlepoint approximations for multinomial distributions, with applications to the joint. Journal of Statistical Computation and Simulation, 59(1):33–45, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Ma:1999:RBA

[Ma99] Chunsheng Ma. Representations, bounds and approximations for tail probabilities of multivariate non-central hypergeometric and negative hypergeometric distributions. Journal of Statistical Computation and Simulation, 62(3):237–258, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Macleod:1983:RMS

[Mac83] D. A. Macleod. Reconstructed mean squares: A technique for factorial experiments with many missing cells. Journal of Statistical Computation and Simulation, 17(1):51–61, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maclaren:1992:LUL

[Mac92] N. M. Maclaren. A limit on the usable length of a pseudorandom sequence. Journal of Statistical Computation and Simulation, 42 (1–2):47–54, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Madsen:1977:NSS

[Mad77] R. W. Madsen. A note on a sequential signed-rank test for symmetry. Journal of Statistical Computation and Simulation, 5(4):259–265, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Madsen:1978:STP

[Mad78] Richard W. Madsen. A k-stage test procedure using p-values. Journal of Statistical Computation and Simulation, 8(2):117– 131, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maeshiro:1987:FSS

[Mae87] Asatoshi Maeshiro. Fatal shortcomings of stationary AR (1) exogeneous variables in econometric Monte Carlo tudies of estimators. Journal of Statistical Computation and Simulation, 28 (3):191–212, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maghsoodloo:1975:EWE

[Mag75a] S. Maghsoodloo. Eccentricities for which ellipsoidal probabilities are good approximations to spherical probabilities. Journal of Statistical Computation and Simulation, 3(4):369–378, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maghsoodloo:1975:EQK

[Mag75b] S. Maghsoodloo. Estimates of the quantiles of Kendall’s partial rank correlation coefficient. Journal of Statistical Computation and Simulation, 4(2):155–164, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maghsoodloo:1981:CCC

[Mag81] Saeed Maghsoodloo. C111. The continuity correction for chisquared for 2 × 2 tables. Journal of Statistical Computation and Simulation, 14(1):47–51, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maharaj:1996:STC

[Mah96] Elizabeth Ann Maharaj. A significance test for classifying ARMA models. Journal of Statistical Computation and Simulation, 54(4):305–331, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maindonald:1977:LSC

[Mai77] J. H. Maindonald. Least squares computations based on the Cholesky decomposition of the correlation matrix. Journal of Statistical Computation and Simulation, 5(4):247–258, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maire:2003:RVU

[Mai03] Sylvain Maire. Reducing variance using iterated control variates. Journal of Statistical Computation and Simulation, 73(1):1–30, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mak:2000:HRM

[Mak00] Tak K. Mak. Heteroscedastic regression models with nonnormally distributed errors. Journal of Statistical Computation and Simulation, 67(1):21–36, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mallows:1979:CHM

[Mal79] C. L. Mallows. C48. How many independence hypotheses are there? Journal of Statistical Computation and Simulation, 9 (3):235–236, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Malik:2006:NAC

[Mal06] Malik Beshir Malik. A numerical algorithm for computing exact distribution of a truncated random variable when truncation points are assigned point masses. Journal of Statistical Computation and Simulation, 76(11):957–965, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marcus:1976:PST

[Mar76] Ruth Marcus. The powers of some tests of the equality of exponential means against an ordered alternative. Journal of Statistical Computation and Simulation, 5(1):37–42, 1976. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Markham:1981:SSE

[Mar81] R. Markham. Small sample efficiencies of tests based on the method of n rankings. Journal of Statistical Computation and Simulation, 12(3–4):193–207, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marrero:1983:PSS

[Mar83] Osvaldo Marrero. The performance of several statistical tests for seasonality in monthly data. Journal of Statistical Computation and Simulation, 17(4):275–296, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marcucci:1986:CPS

[Mar86a] Mark Marcucci. A comparison of the power of some tests for repeated measurements. Journal of Statistical Computation and Simulation, 26(1–2):37–53, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marsaglia:1986:CIG

[Mar86b] John C. W. Marsaglia. C249. The incomplete gamma function and Ramanujan’s rational approximation to e^{x}. Journal of Statistical Computation and Simulation, 24(2):163–168, 1986. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Markowski:1987:AES

[Mar87] Edward Markowski. Accurate estimation of standard errors when comparing simulated power functions. Journal of Statistical Computation and Simulation, 27(4):269–275, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marx:1990:CPC

[Mar90a] Brian D. Marx. 2374. A continuum of principal component regressions. Journal of Statistical Computation and Simulation, 37 (3–4):234–236, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marx:1990:CCC

[Mar90b] Brian D. Marx. C371. A correction for C240, Good and Smith, JSCS 22, pp. 136–142. Journal of Statistical Computation and Simulation, 36(2-3):193, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [GS85].

Marx:1992:SST

[Mar92] Brian D. Marx. A simulation study of two-sample inference based on one-sample nonparametric confidence intervals. Journal of Statistical Computation and Simulation, 42(3–4):161–171, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marschner:1995:CAS

[Mar95] Ian C. Marschner. Computation of age-specific HIV incidence estimates using the EM algorithm. Journal of Statistical Computation and Simulation, 53(3–4):299–312, 1995. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marchand:1996:CMTa

[Mar96a] Eric Marchand. Computing the moments of a truncated noncentral chi-square distribution. Journal of Statistical Computation and Simulation, 54(4):387–391, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marchand:1996:CMTb

[Mar96b] Eric Marchand. Computing the moments of a truncated noncentral chi-square distribution. Journal of Statistical Computation and Simulation, 55(1–2):23–29, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marchand:1998:CNE

[Mar98] Eric Marchand. C448. A note on the equality E(X) = E[E(X/Y)] where E(XIY = y) is constant. Journal of Statistical Computation and Simulation, 60(2):165–168, 1998. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maswadah:2003:CCI

[Mas03] M. Maswadah. Conditional confidence interval estimation for the inverse Weibull distribution based on censored generalized order statistics. Journal of Statistical Computation and Simulation, 73(12):887–898, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mathai:1979:EDE

[Mat79] A. M. Mathai. The exact distributions and the exact percentage points for testing equality of variances in independent normal populations. Journal of Statistical Computation and Simulation, 9(3):169–182, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mays:1999:NST

[May99] Darcy Mays. Near-saturated two-stage designs with dispersion effects. Journal of Statistical Computation and Simulation, 63 (1):13–35, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mays:2006:BAT

[May06] Darcy P. Mays. Bayesian application to the two-stage nearsaturated experimental design method with dispersion effects. Journal of Statistical Computation and Simulation, 76(5):459– 473, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Moazzami:1986:SSP

[MB86] B. Moazzami and A. Busse. The small sample performance of some limited information estimators of a dynamic structural equation with autocorrelated errors. Journal of Statistical Computation and Simulation, 26(1–2):55–78, 1986. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Moazzami:1990:LIE

[MB90] B. Moazzami and A. Buse. Limited information estimator of a dynamic structural equation with autocorrelated errors: a correction and new evidence. Journal of Statistical Computation and Simulation, 34(4):195–208, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Meeuwissen:1997:MID

[MB97] A. M. H. Meeuwissen and Tim Bedford. Minimally informative distributions with given rank correlation for use in uncertainty analysis. Journal of Statistical Computation and Simulation, 57 (1-4):143–174, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mays:2000:OMR

[MBE00] James E. Mays, Jeffrey B. Birch, and Richard L. Einsporn. An overview of model-robust regression. Journal of Statistical Computation and Simulation, 66(1):79–100, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Meelis:1991:NPT

[MBH91] Evert Meelis, Mientje Bressers, and Patsy Haccou. Nonparametric testing for the number of change points in a sequence of independent random variables. Journal of Statistical Computation and Simulation, 39(3):129–137, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

McGilchrist:1991:RER

[MC91] C. A. McGilchrist and B. R. Cullis. Reml estimation for repeated measures analysis. Journal of Statistical Computation and Simulation, 38(1-4):151–163, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Montenegro:2004:BCC

[MCCC04] L. C. C. Montenegro, E. A. Colosimo, G. M. Cordeiro, and F. R. B. Cruz. Bias correction in the Cox regression model. Journal of Statistical Computation and Simulation, 74(5):379– 386, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mcdonald:1974:ADF

[Mcd74] James B. Mcdonald. An approximation of the distribution function of the LIML identifiability test statistic using the method of moments. Journal of Statistical Computation and Simulation, 3 (1):53–66, 1974. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

McGilchrist:1989:BMR

[McG89] C. A. McGilchrist. Bias of ML and REML estimators in regression models with ARMA errors. Journal of Statistical Computation and Simulation, 32(3):127–136, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

McKenzie:1986:NDT

[McK86] Ed McKenzie. A note on the derivation of theoretical autocovariances for ARMA models. Journal of Statistical Computation and Simulation, 24(2):159–162, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

McLeish:1978:VRM

[McL78] D. L. McLeish. Variations of the Robbins–Monro procedure for estimating ED(p) in the logit model. Journal of Statistical Computation and Simulation, 8(1):13–24, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

McLachlan:1980:EEB

[McL80] G. J. McLachlan. The efficiency of Efron’s “Bootstrap” approach applied to error rate estimation in discriminant analysis. Journal of Statistical Computation and Simulation, 11(3–4):273– 279, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

McNichols:1987:CML

[McN87] D. T. McNichols. C279. Maximum likelihood estimation when sampling without replacement. Journal of Statistical Computation and Simulation, 28(1):73–74, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

McNeil:2008:SNA

[McN08] Alexander J. McNeil. Sampling nested Archimedean copulas. Journal of Statistical Computation and Simulation, 78(6):567– 581, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

McSweeney:2006:CPT

[McS06] Laura A. McSweeney. Comparison of periodogram tests. Journal of Statistical Computation and Simulation, 76(4):357–369, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Metzler:1972:CSN

[ME72] Carl M. Metzler and Gary L. Elfring. A computer study of nonlinear standard curves in chemical and biological assays. Journal of Statistical Computation and Simulation, 1(3):261–271, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mee:1993:CFU

[Mee93] Robert W. Mee. Computing factors for uniform simultaneous confidence bands and prediction bands and linear regression. Journal of Statistical Computation and Simulation, 47(3–4):161– 165, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Meintanis:2008:TGE

[Mei08] Simos G. Meintanis. Tests for generalized exponential laws based on the empirical Mellin transform. Journal of Statistical Computation and Simulation, 78(11):1077–1085, 2008. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Meintanis:2009:GFT

[Mei09] Simos G. Meintanis. Goodness-of-fit testing by transforming to normality: comparison between classical and characteristic function-based methods. Journal of Statistical Computation and Simulation, 79(2):205–212, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Melnick:1980:MRD

[Mel80] Edward L. Melnick. Moments of ranked discrete variables with an application to independent Poisson variates. Journal of Statistical Computation and Simulation, 12(1):51–60, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mendes:2000:ABM

[Men00] Beatriz Vaz De. Melo Mendes. Assessing the bias of maximum likelihood estimates of contaminated GARCH models. Journal of Statistical Computation and Simulation, 67(4):359–376, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Manly:2002:TMV

[MF02] Bryan Manly and R. I. C. Chris Francis. Testing for mean and variance differences with samples from distributions that may be non-normal with unequal variances. Journal of Statistical Computation and Simulation, 72(8):633–646, 2002. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marsaglia:1980:CGN

[MG80] George Marsaglia and I. J. Good. C69. Generating a normal sample with given sample mean and variance. Journal of Statistical Computation and Simulation, 11(1):71–74, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Magis:2006:NIP

[MG06] David Magis and Paul Gérard. A note on the influence of pairwise misclassification in two-way tables. Journal of Statistical Computation and Simulation, 76(10):875–887, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mohberg:1978:LMA

[MGG78] N. R. Mohberg, M. Ghosh, and J. E. Grizzle. Linear models analysis of small samples of categorized ordinal response data. Journal of Statistical Computation and Simulation, 7(3–4):237– 251, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mejias:2009:MCC

[MGG09] R. Pino Mejías, M. D. Jiménez Gamero, and A. Enguix González. A Monte Carlo comparison of three consistent bootstrap procedures. Journal of Statistical Computation and Simulation, 79(4):323–334, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Martins:1999:SRB

[MGN99] M. João Martins, M. Ivette Gomes, and Manuela Nevas. Some results on the behaviour of Hill’s estimator. Journal of Statistical Computation and Simulation, 63(3):283–297, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Martinson:1972:MLS

[MH72] E. O. Martinson and M. A. Hamdan. Maximum likelihood and some other asymptotically efficient estimators of correlation in two way contingency tables. Journal of Statistical Computation and Simulation, 1(1):45–54, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Madsen:1978:MST

[MH78] Richard W. Madsen and John E. Hewett. Multi-stage tests based on sequential ranks. Journal of Statistical Computation and Simulation, 7(2):93–105, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Millar:1999:MOD

[MH99] Anne Michèle Millar and David C. Hamilton. Modern outlier detection methods and their effect on subsequent inference. Journal of Statistical Computation and Simulation, 64(2):125–150, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Molefe:2007:TLM

[MH07] Ayrin C. Molefe and Balakrishna Hosmane. Test for link misspecification in dependent binary regression using generalized estimating equations. Journal of Statistical Computation and Simulation, 77(2):95–107, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Minhajuddin:2004:SMD

[MHS04] Abu T. M. Minhajuddin, Ian R. Harris, and William R. Schucany. Simulating multivariate distributions with specific correlations. Journal of Statistical Computation and Simulation, 74 (8):599–607, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mihram:1972:GST

[Mih72] G. Arthur Mihram. A glossary of simulation terminology. Journal of Statistical Computation and Simulation, 1(1):35–44, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mihram:1974:BSE

[Mih74] G. Arthur Mihram. Blockinq in simular experimental designs. Journal of Statistical Computation and Simulation, 3(1):29–32, 1974. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Miller:1979:MLE

[Mil79] John J. Miller. Maximum likelihood estimation of variance components — a Monte Carlo study. Journal of Statistical Computation and Simulation, 8(3–4):175–190, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Modarres:1993:RTC

[MJ93] Reza Modarres and Robert W. Jernigan. A robust test for comparing correlation matrices. Journal of Statistical Computation and Simulation, 46(3–4):169–181, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mara:2008:CSE

[MJ08] Thierry Alex Mara and Onimihamina Rakoto Joseph. Comparison of some efficient methods to evaluate the main effect of computer model factors. Journal of Statistical Computation and Simulation, 78(2):167–178, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Martz:1974:SEB

[MK74] H. F. Martz, Jr. and S. J. Kamat. A smooth empirical Bayes estimator for the mean of a multivariate normal distribution. Journal of Statistical Computation and Simulation, 3(2):179– 188, 1974. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mitra:1997:CME

[MK97a] Amit Mitra and Debasis Kundu. Consistent method for estimating sinusoidal frequencies: a non-iterative approach. Journal of Statistical Computation and Simulation, 58(2):171–194, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Moore:1997:CIN

[MK97b] Brett Moore and K. Krishnamoorthy. Combining independent normal sample means by weighting with their standard errors. Journal of Statistical Computation and Simulation, 58(2):145– 153, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mitra:2008:ALC

[MK08] Sharmishtha Mitra and Debasis Kundu. Analysis of left censored data from the generalized exponential distribution. Journal of Statistical Computation and Simulation, 78(7):669–679, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mitra:1994:ERS

[MKDM94] Amit Mitra, Debasis Kundu, I. D. Dhariyal, and Neeraj Misra. Estimating the ratio of the smaller and the larger of two uniform scale parameters. Journal of Statistical Computation and Simulation, 50(3–4):197–211, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Martz:1974:SCS

[ML74] H. F. Martz and M. G. Lian. A survey and comparison of several empirical Bayes estimators for the binomial parameter. Journal of Statistical Computation and Simulation, 3(2):165–178, 1974. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mitra:1979:MCC

[ML79] Amitava Mitra and Robert F. Ling. A Monte Carlo comparison of some ridge and other biased estimators. Journal of Statistical Computation and Simulation, 9(3):195–215, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

McDonald:1980:NCC

[MM80] James B. McDonald and Dennis D. Maynes. A note on the computation of k-class estimators using a least squares algorithm.

Journal of Statistical Computation and Simulation, 11(3–4):189– 196, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Markowski:1993:CES

[MM93] Edward P. Markowski and Carol A. Markowski. C39.5. Evaluation of some alternatives to the Welch test. Journal of Statistical Computation and Simulation, 44(3–4):262–264, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Moussa:1999:MCE

[MM99] S. Moussa and M. Moore. Monte Carlo estimates of the distributions of the Poisson Voronoi tessellation. Journal of Statistical Computation and Simulation, 64(3):251–273, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mecklin:2005:MCC

[MM05] Christopher J. Mecklin and Daniel J. Mundfrom. A Monte Carlo comparison of the Type I and Type II error rates of tests of multivariate normality. Journal of Statistical Computation and Simulation, 75(2):93–107, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mitchell:2008:CAT

[MM08] Heather Mitchell and Michael D. Mckenzie. A comparison of alternative techniques for selecting an optimum ARCH model. Journal of Statistical Computation and Simulation, 78(1):51– 67, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Martin:2008:DBC

[MMP08] N. Martín, D. Morales, and L. Pardo. Divergence-based confidence intervals in false-positive misclassification model. Journal of Statistical Computation and Simulation, 78(6):527–540, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marhuenda:2005:RUT

[MMPP05] Y. Marhuenda, D. Morales, J. A. Pardo, and M. C. Pardo. Rukhin’s uniformity test based on sample quantiles. Journal of Statistical Computation and Simulation, 75(3):217–235, 2005.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Melloy:1992:DPL

[MMR92] Brian J. Melloy, John I. Mccool, and Matthew Rosenshine. Determining the production lot size for precision components. Journal of Statistical Computation and Simulation, 41(1–2):1– 18, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mingxian:1991:PML

[MMW91] Xu Mingxian, John J. Miller, and Edward J. Wegman. Parallelizing multiple linear regression for speed and redundancy: an empirical study. Journal of Statistical Computation and Simulation, 39(4):205–214, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mehrotra:1983:SCL

[MMWM83] Kishan G. Mehrotra, Joel E. Michalek, Thomas White, and Daniel Mihalko. Score computation for linear rank procedures. Journal of Statistical Computation and Simulation, 16(3–4):201– 211, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Miller:1991:SAG

[MO91] Klaus Miller and Walter Olbricht. On the statistical analysis of grouped Markov chains. Journal of Statistical Computation and Simulation, 39(1–2):69–89, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mohle:2005:SAI

[Möh05] M. Möhle. Simulation algorithms for integrals of a class of sampling distributions arising in population genetics. Journal of Statistical Computation and Simulation, 75(9):731–749, 2005. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Molenaar:1979:SAB

[Mol79] Ivo W. Molenaar. Simple approximations to the Behrens–Fisher distribution. Journal of Statistical Computation and Simulation, 9(4):283–288, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Molchanov:1999:DLA

[Mol99] I. S. Molchanov. Diffusion-limited aggregation with jumps and flights. Journal of Statistical Computation and Simulation, 64 (4):357–381, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Moon:1990:SSP

[Moo90] Choon-Geol Moon. Small sample performance of large sample tests for Tobit versus p-Tobit. Journal of Statistical Computation and Simulation, 37(1–2):89–97, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mackisack:1994:MPA

[MOS94] M. S. Mackisack, M. R. Osborne, and G. K. Smyth. A modified Prony algorithm for estimating sinusoidal frequencies. Journal of Statistical Computation and Simulation, 49(1–2):111–124, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mousa:1995:EBE

[Mou95] M. A. M. Ali Mousa. Empirical Bayes estimators for the Burr Type XII accelerated life testing model based on type-2 censored data. Journal of Statistical Computation and Simulation, 52(2): 95–103, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mousa:2001:IPP

[Mou01] M. A. M. Ali Mousa. Inference and prediction for Pareto progressively censored data. Journal of Statistical Computation and Simulation, 71(2):163–181, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Moustafa:2005:END

[Mou05] Hanem M. Moustafa. The efficiency of a nonlinear discriminant function based on unclassified initial samples from a mixture of two Weibull populations. Journal of Statistical Computation and Simulation, 75(1):65–73, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maghsoodloo:1981:ABK

[MP81] S. Maghsoodloo and L. László Pallos. Asymptotic behavior of Kendall’s partial rank correlation coefficient and additional quantile estimates. Journal of Statistical Computation and Simulation, 13(1):41–48, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Muller:1996:DBT

[MP96] Werner G. Müller and Antonio C. M. Ponce De Leon. Discrimination between two binary data models: sequentially designed experiments. Journal of Statistical Computation and Simulation, 55(1–2):87–100, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Menendez:2005:TSM

[MPPZ05] M. L. Menéndez, J. A. Pardo, L. Pardo, and K. Zografos. On tests of symmetry, marginal homogeneity and quasi-symmetry in two-way contingency tables based on minimum ϕ-divergence estimator with constraints. Journal of Statistical Computation and Simulation, 75(7):555–580, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Moustafa:2003:MND

[MR03] H. M. Moustafa and S. G. Ramadan. On MLE of a nonlinear discriminant function from a mixture of two Gompertz distributions based on small sample size. Journal of Statistical Computation and Simulation, 73(12):867–885, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Muller:2009:STI

[MR09] Samuel Müller and Kaspar Rufibach. Smooth tail-index estimation. Journal of Statistical Computation and Simulation, 79(9): 1155–1167, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Matloff:1984:CTM

[MRR84] Norman Matloff, Robin Rose, and Tai Raymond. A comparison of two methods for estimating optimal weights in regression analysis. Journal of Statistical Computation and Simulation, 19 (4):265–274, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Meyer:1980:UFM

[MS80] Carl D. Meyer, Jr. and James M. Shoaf. Updating finite Markov chains by using techniques of group matrix inversion. Journal of Statistical Computation and Simulation, 11(3–4):163–181, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

McLachian:1995:ARE

[MS95] G. J. McLachian and D. Scot. Asymptotic relative efficiency of the linear discriminant function under partial nonrandom classification of the training data. Journal of Statistical Computation and Simulation, 52(4):415–426, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mays:1998:TSC

[MS98] Darcy P. Mays and Karen R. Schwartz. Two-stage central composite designs with dispersion effects. Journal of Statistical Computation and Simulation, 61(3):191–218, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

McDonald:1999:EPS

[MS99] David A. McDonald and Leif K. Sandal. Estimating the parameters of stochastic differential equations using a criterion function. Journal of Statistical Computation and Simulation, 64(3):235– 250, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mantalos:2001:BJT

[MS01] Panagiotis Mantalos and Ghazi Shukur. Booststrapped Johansen tests for cointegration relationships: a graphical analysis. Journal of Statistical Computation and Simulation, 68(4): 351–371, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mihalko:1980:SRM

[MT80] Daniel Mihalko and Y. L. Tong. A sequential regression method in Monte Carlo studies. Journal of Statistical Computation and Simulation, 12(1):41–50, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Morgenthaler:2001:TWP

[MT01] Stephan Morgenthaler and John W. Tukey. Two-way plots: I. The case of straight line segments. Journal of Statistical Computation and Simulation, 71(4):287–312, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mammadov:2008:ABU

[MTO08] M. Mammadov, E. Tas, and R. E. Omay. Accelerating backpropagation using effective parameters at each step and an experimental evaluation. Journal of Statistical Computation and Simulation, 78(11):1055–1064, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Munzel:2006:NPC

[Mun06] Ullrich Munzel. Non-parametric confidence intervals for shift effects based on paired ranks. Journal of Statistical Computation and Simulation, 76(9):765–772, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Murphy:1978:NCM

[Mur78] B. P. Murphy. A note on current methods of partitioning the contigency x^{2} statistic. Journal of Statistical Computation and Simulation, 6(3–4):157–167, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Murtagh:1983:PTH

[Mur83] F. Murtagh. A probability theory of hierarchic clustering using random dendrograms. Journal of Statistical Computation and Simulation, 18(2-3):145–157, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Magel:1984:REO

[MW84] Rhonda Magel and F. T. Wright. Robust estimates of ordered parameters. Journal of Statistical Computation and Simulation, 20(1):47–58, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maragah:1992:EAR

[MW92] Hazem D. Maragah and William H. Woodall. The effect of autocorrelation on the retrospective X-chart. Journal of Statistical Computation and Simulation, 40(1–2):29–42, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marschner:1994:IEA

[MW94] Ian C. Marschner and Lyndsey F. Watson. An improved EMS algorithm for back-projection of AIDS incidence data. Journal of Statistical Computation and Simulation, 50(1–2):1–20, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mena:2004:DFC

[MW04] Ramsés H. Mena and Stephen G. Walker. A density function connected with a non-negative self-decomposable random variable. Journal of Statistical Computation and Simulation, 74 (10):765–775, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mielke:1990:SEC

[MY90] P. W. Mielke, Jr. and Y. C. Yao. On g-sample empirical coverage tests: exact and simulated null distributions of test statistics with small and moderate sizes. Journal of Statistical Computation and Simulation, 35(1–2):31–39, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Myers:1979:BR

[Mye79] Sharon L. Myers. Book reviews. Journal of Statistical Computation and Simulation, 9(4):317–319, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Myers:1983:BR

[Mye83] Sharon L. Myers. Book review. Journal of Statistical Computation and Simulation, 17(1):75–77, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Myers:1998:CJS

[Mye98] Leann Myers. Comparability of the James’ second-order approximation test and the Alexander and Govern A statistic for non-normal heteroscedastic data. Journal of Statistical Computation and Simulation, 60(3):207–222, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Meaux:2001:CPC

[MYS01] Laurie M. Meaux, Dean M. Young, and John W. Seaman, Jr. A comparison of parametric conditional error-rate estimators for the two-group linear discriminant function. Journal of Statistical Computation and Simulation, 69(3):277–291, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Mardia:1977:TML

[MZ77] K. V. Mardia and P. J. Zemroch. Table of maximum likelihood estimates for the Bingham distribution. Journal of Statistical Computation and Simulation, 6(1):29–34, 1977. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mantalos:2008:IEB

[MZ08] Panagiotis Mantalos and Konstantinos Zografos. Interval estimation for a binomial proportion: a bootstrap approach. Journal of Statistical Computation and Simulation, 78(12):1251– 1265, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marsaglia:1989:CAA

[MZZ89] George Marsaglia, Arif Zaman, and Youlu Zheng. C309. An algorithm for the area of the union of a collection of convex sets. Journal of Statistical Computation and Simulation, 31(1):46– 49, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nudurupati:2009:NAS

[NA09] Sai V. Nudurupati and Asheber Abebe. A nonparametric allocation scheme for classification based on transvariation probabilities. Journal of Statistical Computation and Simulation, 79 (8):977–987, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nabeya:1983:AES

[Nab83] Seiji Nabeya. Asymptotic expansions for the sum of the series used in sequential analysis. Journal of Statistical Computation and Simulation, 16(3–4):223–240, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nadarajah:1999:SME

[Nad99] S. Nadarajah. Simulation of multivariate extreme values. Journal of Statistical Computation and Simulation, 62(4):395–410, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nagarsenker:1975:EDL

[Nag75] B. N. Nagarsenker. The exact distribution of the likelihood ratio criterion for testing equality of means, variances and covariances.

Journal of Statistical Computation and Simulation, 4(3):225– 233, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nandram:1998:BAT

[Nan98] Balgobin Nandram. A Bayesian analysis of the three-stage hierarchical multinomial model. Journal of Statistical Computation and Simulation, 61(1–2):97–126, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Narayanan:1990:CGD

[Nar90] A. Narayanan. Computer generation of Dirichlet random vectors. Journal of Statistical Computation and Simulation, 36(1): 19–30, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Naranjo:1996:MEE

[Nar96] Joshua D. Naranjo. More efficient L-estimates. Journal of Statistical Computation and Simulation, 54(4):355–361, 1996. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nath:1982:IRT

[Nat82] Ravinder Nath. An investigation of the rank transform in the multivariate one- sample location problem. Journal of Statistical Computation and Simulation, 16(2):139–155, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nottingham:2000:LLR

[NBB00] Quintion J. Nottingham, Jeffrey B. Birch, and Barry A. Bodt. Local logistic regression an application to army penetration data. Journal of Statistical Computation and Simulation, 66 (1):35–50, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Norman:1972:CPG

[NC72] J. E. Norman and L. E. Cannon. A computer program for the generation of random variables from any discrete distribution. Journal of Statistical Computation and Simulation, 1(4):331– 348, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nandram:1996:RGL

[NC96] Balgobin Nandram and Ming-Hui Chen. Reparameterizing the generalized linear model to accelerate Gibbs sampler convergence. Journal of Statistical Computation and Simulation, 54 (1-3):129–144, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nandram:2006:BAT

[NC06] Balgobin Nandram and Jai Won Choi. Bayesian analysis of a two-way categorical table incorporating intraclass correlation. Journal of Statistical Computation and Simulation, 76(3):233– 249, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Negassa:2000:TSP

[NCA^{+}00] Abdissa Negassa, Antonio Ciampi, Michal Abrahamowicz, Stanley Shapiro, and Jean-François Boivin. Tree-structured prognostic classification for censored survival data: validation of computationally inexpensive model selection criteria. Journal of Statistical Computation and Simulation, 67(4):289–317, 2000. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nakamori:2009:SEN

[NCÁHC^{+}09] S. Nakamori, R. Caballero-Águila, A. Hermoso-Carazo, J. D. Jiménez-López, and J. Linares-Pérez. Signal estimation with nonlinear uncertain observations using covariance information. Journal of Statistical Computation and Simulation, 79(1):55– 66, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nath:1984:MCC

[ND84] Ravinder Nath and Benjamin S. Duran. A Monte Carlo comparison of some t-tests. Journal of Statistical Computation and Simulation, 20(3):235–260, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Neuhauser:2007:CSN

[Neu07] Markus Neuhäuser. A comparative study of nonparametric twosample tests after Levene’s transformation. Journal of Statistical Computation and Simulation, 77(6):517–526, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ntzoufras:2000:SSV

[NFD00] Ioannis Ntzoufras, Jonathan J. Forster, and Petros Dellaportas. Stochastic search variable selection for log-linear models. Journal of Statistical Computation and Simulation, 68(1):23– 37, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ng:1996:APC

[Ng96] C. K. Ng. All pairwise comparisons of exponential location parameters under Type I censoring — equal scale parameter case. Journal of Statistical Computation and Simulation, 54(1-3):19– 27, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nickerson:1996:SAS

[Nic96] David M. Nickerson. Sequential, accelerated-sequential and three-stage estimation of the mean of a first-order stationary autoregressive process: A Monte Carlo study. Journal of Statistical Computation and Simulation, 54(4):333–353, 1996. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nicolau:2005:MSN

[Nic05] João Nicolau. Method for simulating non-linear stochastic differential equations in R^{1}. Journal of Statistical Computation and Simulation, 75(8):595–609, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Niermann:2006:EEP

[Nie06] Stefan Niermann. Evolutionary estimation of parameters of Johnson distributions. Journal of Statistical Computation and Simulation, 76(3):185–193, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nagar:1988:TCS

[NJG88] D. K. Nagar, S. K. Jain, and A. K. Gupta. On testing circular stationary and related models. Journal of Statistical Computation and Simulation, 29(3):225–239, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nath:1992:CCL

[NJJ92] Ravinder Nath, Wade M. Jackson, and Thomas W. Jones. A comparison of the classical and the linear programming approaches to the classification problem in discriminant analysis. Journal of Statistical Computation and Simulation, 41(1–2):73– 93, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nandram:2002:MLC

[NK02] Balgobin Nandram and Hyunjoong Kim. Marginal likelihood for a class of Bayesian generalized linear models. Journal of Statistical Computation and Simulation, 72(4):319–340, 2002. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nadarajah:2007:JBG

[NK07] Saralees Nadarajah and Samuel Kotz. Jensen’s bivariate gamma distribution: ratios of components. Journal of Statistical Computation and Simulation, 77(4):349–358, 2007. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nikoloulopoulos:2008:MCD

[NK08] Aristidis K. Nikoloulopoulos and Dimitris Karlis. On modeling count data: a comparison of some well-known discrete distributions. Journal of Statistical Computation and Simulation, 78 (3):437–457, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Narula:1977:MCC

[NL77a] Subhash C. Narula and Kenneth J. Levy. A Monte Carlo comparison of several methods for analyzing small sample binomial data in a two-factor experiment without replication. Journal of Statistical Computation and Simulation, 5(2):135–143, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Narula:1977:ACS

[NL77b] Subhash C. Narula and Franz S. Li. Approximations to the chi-square distribution. Journal of Statistical Computation and Simulation, 5(4):267–277, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nelson:2008:PCE

[NL08] Kerrie P. Nelson and Brian G. Leroux. Properties and comparison of estimation methods in a log-linear generalized linear mixed model. Journal of Statistical Computation and Simulation, 78(3):367–384, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ng:1998:MLT

[NM98] S. K. Ng and G. J. McLachlan. On modifications to the longterm survival mixture model in the presence of competing risks. Journal of Statistical Computation and Simulation, 61(1–2):77– 96, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nance:1975:IFR

[NO75] Richard E. Nance and Claude Overstreet, Jr. Implementation of Fortran random number generators on computers with one’s complement arithmetic. Journal of Statistical Computation and Simulation, 4(3):235–243, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Norden:1984:NEM

[Nor84] R. H. Norden. On the numerical evaluation of the moments of the distribution of states at time t in the stochastic logistic process. Journal of Statistical Computation and Simulation, 20 (1):1–20, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nord:1991:IRS

[Nor91] R. S. Nord. Irreversible random sequential filling of lattices by Monte Carlo simulation. Journal of Statistical Computation and Simulation, 39(4):231–240, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Newton:1981:SMA

[NP81] H. Joseph Newton and Marcello Pagano. On the stationarity of multiple autoregressive approximants: theory and algorithms. Journal of Statistical Computation and Simulation, 13(3–4):195– 208, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nelson:1998:IPE

[NP98] Ray D. Nelson and Rulon D. Pope. Improvements in the power of empirical stochastic dominance comparisons through kernel density estimation: a Monte Carlo study. Journal of Statistical Computation and Simulation, 60(4):295–317, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nardon:2009:STG

[NP09] Martina Nardon and Paolo Pianca. Simulation techniques for generalized Gaussian densities. Journal of Statistical Computation and Simulation, 79(11):1317–1329, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nelson:1986:DSW

[NS86] Barry L. Nelson and Bruce W. Schmeiser. Decomposition of some well-known variance reduction techniques. Journal of Statistical Computation and Simulation, 23(3):183–209, 1986. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Narayann:1989:TAD

[NS89] A. Narayann and Thomas W. Sager. Table for the asymptotic distribution of univariate mode estimators. Journal of Statistical Computation and Simulation, 33(1):37–51, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nguyen:2008:MAR

[NS08a] D. V. Nguyen and D. Sentürk. Multicovariate-adjusted regression models. Journal of Statistical Computation and Simulation, 78(9):813–827, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nosek:2008:PSL

[NS08b] K. Nosek and Z. Szkutnik. A power study of k-linear-r-ahead recursive residuals test for change-point in finite sequences. Journal of Statistical Computation and Simulation, 78(12):1201– 1213, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Neuhauser:2009:AMC

[NS09] Markus Neuhäuser and Roswitha Senske. The analysis of multicentre clinical trials when there is heterogeneity between centres. Journal of Statistical Computation and Simulation, 79(11): 1381–1387, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Narula:1991:CCP

[NSG91] Subhash C. Narula, V. A. Sposito, and James E. Gentle. Comparison of computer programs for simple linear L_{1} regression. Journal of Statistical Computation and Simulation, 39(1–2):63– 68, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nandram:2009:BIS

[NTK09] Balgobin Nandram, Maria C. S. Toto, and Myron Katzoff. Bayesian inference for a stratified categorical variable allowing all possible category choices. Journal of Statistical Computation and Simulation, 79(2):161–179, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nuel:2008:CDF

[Nue08] Grégory Nuel. Cumulative distribution function of a geometric Poisson distribution. Journal of Statistical Computation and Simulation, 78(3):385–394, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Narula:1983:SVL

[NW83] Subhash C. Narula and John F. Wellington. Selection of variables in linear regression: a pragmatic approach. Journal of Statistical Computation and Simulation, 17(3):159–172, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Narula:1986:AFA

[NW86] Subhash C. Narula and John F. Wellington. An algorithm to find all regression quantiles. Journal of Statistical Computation and Simulation, 26(3–4):269–281, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ng:2009:SEP

[NW09] Hon Keung Tony Ng and Zhu Wang. Statistical estimation for the parameters of Weibull distribution based on progressively Type-I interval censored sample. Journal of Statistical Computation and Simulation, 79(2):145–159, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Oh:1992:AIS

[OB92] Man-Suk Oh and James O. Berger. Adaptive importance sampling in Monte Carlo integration. Journal of Statistical Computation and Simulation, 41(3–4):143–168, 1992. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ozturk:2005:EPM

[OBW05] Omer Ozturk, Omer C. Bilgin, and Douglas A. Wolfe. Estimation of population mean and variance in flock management: a ranked set sampling approach in a finite population setting. Journal of Statistical Computation and Simulation, 75(11):905– 919, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Edlund:1989:ITF

[oE89] Per olov Edlund. On identification of transfer function models by biased regression methods. Journal of Statistical Computation and Simulation, 31(3):131–148, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

OGorman:2006:ATS

[O’G06] Thomas W. O’Gorman. An adaptive test of a subset of regression coefficients using permutations of residuals. Journal of Statistical Computation and Simulation, 76(12):1095–1103, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

OGorman:2009:CSA

[O’G09] Thomas W. O’Gorman. A comparison of several adaptive tests for paired data. Journal of Statistical Computation and Simulation, 79(9):1083–1093, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ogasawara:2007:AEA

[Oga07] H. Ogasawara. Asymptotic expansion and asymptotic robustness of the normal-theory estimators in the random regression model. Journal of Statistical Computation and Simulation, 77 (10):821–838, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

OHara:1982:CDF

[OHLH82] Thoamas F. O’Hara, David W. Hosmer, Jr., Stanley Lemshow, and Stuart C. Hartz. A comparison of discriminant function and maximum likelihood estimates of logistic coefficients for categorical-scaled data. Journal of Statistical Computation and Simulation, 14(3–4):169–178, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ogunyemi:1993:ABE

[OHN93] O. T. Ogunyemi, J. E. Hutton, and P. I. Nelson. Approximate Bayes estimators for stochastic processes. Journal of Statistical Computation and Simulation, 45(3–4):219–231, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ohtani:1998:MPM

[Oht98] Kazuhiro Ohtani. Mes performance of the minimum mean squared error estimators in a linear regression model when relevant regressors are omitted. Journal of Statistical Computation and Simulation, 61(1–2):61–75, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ou:1997:SCI

[OHW97] Shyh-Tyan Ou, Ronald R. Hocking, and Suojin Wang. Saddlepoint confidence intervals for variance components. Journal of Statistical Computation and Simulation, 58(3):255–268, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ostrouchov:1988:AAC

[OM88] George Ostrouchov and William Q. Meeker, Jr. Accuracy of approximate confidence bounds computed from interval censored Weibull and lognormal data. Journal of Statistical Computation and Simulation, 29(1):43–76, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

ONeill:1982:CAM

[O’N82] M. E. O’Neill. A comparison of the additive and multiplicative definitions of second-order interaction in 2 × 2 × 2 contingency tables. Journal of Statistical Computation and Simulation, 15 (1):33–50, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ong:1995:CBG

[Ong95] S. H. Ong. Computation of bivariate gamma and inverted beta distribution functions. Journal of Statistical Computation and Simulation, 51(2-4):153–163, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Onar:2000:IGA

[OP00a] A. Onar and W. J. Padgett. Inverse Gaussian accelerated test models based on cumulative damage. Journal of Statistical Computation and Simulation, 66(3):233–247, 2000. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Orme:2000:AER

[OP00b] Chris D. Orme and Simon A. Peters. Asymptotic expansions and the reliability of tests in accelerated failure time models. Journal of Statistical Computation and Simulation, 65(1-4):109– 132, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Orasch:2004:TWS

[OP04] Markus Orasch and William Pouliot. Tabulating weighted supnorm functionals used in change-point analysis. Journal of Statistical Computation and Simulation, 74(4):249–276, 2004. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ortega:2008:DRG

[OPB08] Edwin M. M. Ortega, Gilberto A. Paula, and Heleno Bolfarine. Deviance residuals in generalised log-gamma regression models with censored observations. Journal of Statistical Computation and Simulation, 78(8):747–764, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

OBrien:1982:MCS

[OPS82] Patricia N. O’Brien, Fraderick J. Parenté, and Charles J. Schmitt. A Monte Carlo study on the robustness of four MANOVA criterion tests. Journal of Statistical Computation and Simulation, 15(2-3):183–192, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ogwang:1992:SSS

[OR92] Tomso Ogwang and U. L. Gouranga Rao. A simulation study of structural estimators of dynamic simultaneous equations models with autoregressive disturbances. Journal of Statistical Computation and Simulation, 44(1–2):27–46, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Osborne:2002:LCM

[OS02] Jason A. Osborne and Thomas A. Severini. The Lorenz curve for model assessment in exponential order statistic models. Journal of Statistical Computation and Simulation, 72(1):87–97, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ouysse:2006:AFM

[Ouy06] Rachida Ouysse. Approximate factor models: Finite sample distributions. Journal of Statistical Computation and Simulation, 76(4):279–303, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ojeda:2002:PRR

[OVL02] Silvia M. Ojeda, Ronny O. Vallejos, and María M. Lucini. Performance of robust RA estimator for bidimensional autoregressive models. Journal of Statistical Computation and Simulation, 72(1):47–62, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Owen:1977:CMS

[Owe77] A. D. Owen. On the calculation of the moments of several econometric estimators. Journal of Statistical Computation and Simulation, 6(1):1–10, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Owent:1981:PAT

[Owe81] A. D. Owent. The power of alternative tests for serial correlation in dynamic models estimated by instrumental variables. Journal of Statistical Computation and Simulation, 12(2):81– 91, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Oman:1981:MAD

[OZ81] Samuel D. Oman and S. Zacks. A mixture approximation to the distribution of a weighted sum of chi-squared variables. Journal of Statistical Computation and Simulation, 13(3–4):215–224, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Padgett:1978:CCI

[Pad78] W. J. Padgett. C21. Comment on inverse Gaussian random number generation. Journal of Statistical Computation and Simulation, 8(1):78–79, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Padgett:1982:API

[Pad82] W. J. Padgett. An approximate prediction interval for the mean of future observations from the inverse Gaussian distribution. Journal of Statistical Computation and Simulation, 14(3–4):191– 199, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pak:1999:RTS

[Pak99] Ro Jin Pak. Robust test statistics induced from the minimum L_{2} distance estimation. Journal of Statistical Computation and Simulation, 64(3):193–207, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pan:1998:RIT

[Pan98] Wei Pan. Rank invariant tests with left truncated and interval censored data. Journal of Statistical Computation and Simulation, 61(1–2):163–174, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pan:1999:LTT

[Pan99a] Guohua Pan. On a Levene type test for equality of two variances. Journal of Statistical Computation and Simulation, 63(1):59–71, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pang:1999:MBD

[Pan99b] Wan-Kai Pang. Modelling binary data: a Gibbs sampling approach. Journal of Statistical Computation and Simulation, 64 (1):73–85, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Papageorgiou:1980:CPG

[Pap80] Haralambos Papageorgiou. C76. On probability generating functions for the absolute difference of two random variables. Journal of Statistical Computation and Simulation, 12(1):65–68, 1980.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Papageorgiou:1983:CCN

[Pap83] H. Papageorgiou. C159. A characterization of the negative multinomial as a Poisson mixture. Journal of Statistical Computation and Simulation, 17(3):245–247, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Paparoditis:1993:CSA

[Pap93] Efstathios Paparoditis. A comparison of some autocovariancebased methods of ARMA model selection: a simulation study. Journal of Statistical Computation and Simulation, 45(1–2):97– 120, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Parr:1980:CPA

[Par80] William C. Parr. A conditional property of adaptive estimators. Journal of Statistical Computation and Simulation, 10(3–4):289– 295, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Parrish:1981:EBC

[Par81] Rudolph S. Parrish. Evaluation of bivariate cumulative probabilities using moments to fourth order. Journal of Statistical Computation and Simulation, 13(3–4):181–194, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Parrish:1987:EAM

[Par87] Rudolph S. Parrish. Evaluation and approximation of multivariate cumulative joint probabilities. Journal of Statistical Computation and Simulation, 27(1):1–33, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Parrish:1992:NTC

[Par92] Rudolph S. Parrish. New tables of coefficients and percentage points for the w test for normality. Journal of Statistical Computation and Simulation, 41(3–4):169–185, 1992. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Park:1999:CTR

[Par99] Dongryeon Park. Comparison of two response curve estimators. Journal of Statistical Computation and Simulation, 62(3):259– 269, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pascual:2003:RFL

[Pas03] Francis Pascual. The random fatigue-limit model in multi-factor experiments. Journal of Statistical Computation and Simulation, 73(10):733–752, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Patefield:1976:VAD

[Pat76] W. M. Patefield. On the validity of approximate distributions arising in fitting a linear functional relationship. Journal of Statistical Computation and Simulation, 5(1):43–60, 1976. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Paulsen:1984:IRN

[Pau84] Jostein Paulsen. Impact of random number generators in time series Monte Carlo simulation. Journal of Statistical Computation and Simulation, 19(1):23–33, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Paula:1992:BCE

[Pau92] Gilberto A. Paula. Bias correction for exponential family nonlinear models. Journal of Statistical Computation and Simulation, 40(1–2):43–54, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pawitan:2001:TSE

[Paw01] Yudi Pawitan. Two-staged estimation of variance components in generalized linear mixed models. Journal of Statistical Computation and Simulation, 69(1):1–17, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Phillips:1972:PGG

[PB72] Don T. Phillips and Charles S. Beightler. Procedures for generating gamma variates with non-integer parameter sets. Journal of Statistical Computation and Simulation, 1(3):197–208, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Piegorsch:1993:MMS

[PB93] Walter W. Piegorsch and A. John Bailer. Minimum mean-square error quadrature. Journal of Statistical Computation and Simulation, 46(3–4):217–234, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

pace:1997:FC

[pB97] R. Kelley pace and Ronald P. Barry. Fast cars. Journal of Statistical Computation and Simulation, 59(2):123–145, 1997. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Polansky:2000:MPB

[PB00] Alan M. Polansky and Edsel R. Baker. Multistage plug-in bandwidth selection for kernel distribution function estimates. Journal of Statistical Computation and Simulation, 65(1-4):63–80, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Price:2001:EVS

[PB01] Robert M. Price and Douglas G. Bonett. Estimating the variance of the sample median. Journal of Statistical Computation and Simulation, 68(3):295–305, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Price:2002:DFC

[PB02] Robert M. Price and Douglas G. Bonett. Distribution-free confidence intervals for difference and ratio of medians. Journal of Statistical Computation and Simulation, 72(2):119–124, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Park:2003:GKL

[PB03] Chanseok Park and Ayanendranath Basu. The generalized Kullback–Leibler divergence and robust inference. Journal of Statistical Computation and Simulation, 73(5):311–332, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Prasad:2004:RMC

[PB04] Samineni S. R. Prasad and Bikash Bhadury. Robustness of measures of common cause sigma in presence of data correlation.

Journal of Statistical Computation and Simulation, 74(5):313– 338, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Petersen:1978:CMS

[PBWU78] Martin R. Petersen, Donald F. Butcher, Stanley Wearden, and Vincent A. Uthoff. A comparison of mean separation techniques. Journal of Statistical Computation and Simulation, 7(1):33–54, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pavia:2009:UIO

[PCS09] Jose Manuel Pavia, Bernardi Cabrer, and Ramon Sala. Updating input-output matrices: assessing alternatives through simulation. Journal of Statistical Computation and Simulation, 79(12):1467–1482, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pardo:2003:MPD

[PD03] Leandro Pardo and María Del Carmen Pardo. Minimum powerdivergence estimator in three-way contingency tables. Journal of Statistical Computation and Simulation, 73(11):819–831, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pepple:1993:SEG

[Pep93] Patricia A. Pepple. Simultaneous estimation of gamma means in the presence of a nuisance parameter. Journal of Statistical Computation and Simulation, 46(3–4):183–200, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Peskun:1980:TTC

[Pes80] Peter H. Peskun. Theoretical tests for choosing the parameters of the general mixed linear congruential pseudorandom number generator. Journal of Statistical Computation and Simulation, 11(3–4):281–305, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pettitt:1980:SRE

[Pet80] A. N. Pettitt. Some results on estimating a change-point using non-parametric type statistics. Journal of Statistical Computation and Simulation, 11(3–4):261–272, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Petersen:1988:PSP

[Pet88] Martin R. Petersen. Post study power estimation by computer. Journal of Statistical Computation and Simulation, 29(3):195– 210, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pettersson:2002:CSM

[Pet02] Tomas Pettersson. A comparative study of model-based tests of independence for ordinal data using the bootstrap. Journal of Statistical Computation and Simulation, 72(3):187–203, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Potzelberger:1993:FID

[PF93] Klaus Pötzelberger and Klaus Felsenstein. On the Fisher information of discretized data. Journal of Statistical Computation and Simulation, 46(3–4):125–144, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pan:2004:ADP

[PF04] Wei Pan and Kenneth A. Frank. An approximation to the distribution of the product of two dependent correlation coefficients. Journal of Statistical Computation and Simulation, 74(6):419– 443, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pham-Gia:2009:TCM

[PGT09] T. Pham-Gia and N. Turkkan. Testing a covariance matrix: exact null distribution of its likelihood criterion. Journal of Statistical Computation and Simulation, 79(11):1331–1340, 2009. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Portela:2004:IRB

[PGV04] Javier Portela and M. A. Gómez-Villegas. Implementation of a robust Bayesian method. Journal of Statistical Computation and Simulation, 74(4):235–248, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Han:1978:CNC

[pH78] Chien pai Han. On the computation of noncentral chi-squared distributions. Journal of Statistical Computation and Simulation, 6(3–4):207–210, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Papadopoulos:1995:SRI

[PH95] Alex S. Papadopoulos and Robert M. Hoekstra. Some robustness issues in the Bayesian analysis of a Weibull failure model. Journal of Statistical Computation and Simulation, 52(1):85– 94, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pawitan:2003:CCI

[PH03] Yudi Pawitan and Jian Huang. Constrained clustering of irregularly sampled spatial data. Journal of Statistical Computation and Simulation, 73(12):853–865, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

pham:1995:DEC

[pha95] Dinh Tuan pham. On the discretisation error in the computation of the empirical characteristic function. Journal of Statistical Computation and Simulation, 53(3–4):129–141, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Phillip:1991:MPT

[Phi91] Good I. Phillip. Most powerful tests for use in matched pair experiments when data may be censored. Journal of Statistical Computation and Simulation, 38(1-4):57–63, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Philippe:1997:PSO

[Phi97] Anne Philippe. Processing simulation output by Riemann sums. Journal of Statistical Computation and Simulation, 59(4):295– 314, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pasia:2005:UTS

[PHO05] Joseph M. Pasia, Augusto Y. Hermosilla, and Hernando Ombao. A useful tool for statistical estimation: genetic algorithms.

Journal of Statistical Computation and Simulation, 75(4):237– 251, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Piepho:1994:CDF

[Pie94a] Hans-Peter Piepho. C410. a distribution-free test for homoscedasticity in a two-way layout. Journal of Statistical Computation and Simulation, 49(3–4):223–225, 1994. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Piepho:1994:CRT

[Pie94b] Mans-Peter Piepho. C409. A randomization test for homoscedasticity in a balanced one-way layout. Journal of Statistical Computation and Simulation, 49(3–4):222–223, 1994. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Piepho:1997:TED

[Pie97] H. P. Piepho. Tests for equality of dispersion in divariate samples — review and empirical comparison. Journal of Statistical Computation and Simulation, 56(4):353–372, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pigeot:1991:SSE

[Pig91] Iris Pigeot. A simulation study of estimators of a common odds ratio in several 2 × 2 tables. Journal of Statistical Computation and Simulation, 38(1-4):65–82, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pinkert:1978:ECI

[Pin78] James R. Pinkert. Elimination of computer-induced errors. Journal of Statistical Computation and Simulation, 6(3–4):189– 198, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Papaioannou:1972:PTS

[PK72] Takis Papaioannou and Oscar Kempthorne. Parallel tangents and steepest descent optimization algorithm — a computer implementation with application to linear, partially linear models and qualitative data. Journal of Statistical Computation and Simulation, 1(4):349–376, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pal:1999:SEC

[PL99] Nabendu Pal and Wooi K. Lom. Shrinkage estimation of a correlation coefficient and two examples with real life data-sets. Journal of Statistical Computation and Simulation, 62(4):357– 373, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Peck:1988:CSB

[PLD88] Roger Peck, Jennings W. Linda, and Young M. Dean. A comparison of several biased estimators for improving the expected error rate of the sample quadratic discriminant function. Journal of Statistical Computation and Simulation, 29(2):143–156, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pascual:2002:MDW

[PM02] Francis Pascual and Grace Montepiedra. On minimax designs when there are two candidate models. Journal of Statistical Computation and Simulation, 72(11):841–862, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pavur:1986:PVR

[PN86] Robert Pavur and Ravinder Nath. Parametric versus rank transform procedures in the two-way factorial experiment: a comparative study. Journal of Statistical Computation and Simulation, 23(3):231–240, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Potscher:1998:DEA

[PN98] Benedikt M. Pötscher and Andreas J. Novák. The distribution of estimators after model selection: large and small sample results. Journal of Statistical Computation and Simulation, 60(1):19– 56, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pendleton:1983:MCA

[PNM83] Brian F. Pendleton, Isadore Newman, and Rodney S. Marshall. A Monte Carlo approach to correlational spuriousness and ratio variables. Journal of Statistical Computation and Simulation, 18 (2-3):93–124, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Piegorsch:2006:ERE

[PNW06] Walter W. Piegorsch, Daniela K. Nitcheva, and R. Webster West. Excess risk estimation under multistage model misspecification. Journal of Statistical Computation and Simulation, 76 (5):423–430, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Puza:2005:LBR

[PO05] Borek D. Puza and Terence J. O’neill. Length-biased, withreplacement sampling from an exponential finite population. Journal of Statistical Computation and Simulation, 75(3):159– 174, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Puza:2006:GCP

[PO06] Borek Puza and Terence O’neill. Generalised Clopper–Pearson confidence intervals for the binomial proportion. Journal of Statistical Computation and Simulation, 76(6):489–508, 2006. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Poitras:1992:TRD

[Poi92] Geoffrey Poitras. Testing regression disturbances for normality with stable alternatives: further Monte Carlo evidence. Journal of Statistical Computation and Simulation, 41(1–2):109–123, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Polansky:1994:EPF

[Pol94] Marcia Polansky. Estimation of the parameters of first order autoregressive process for balanced repeated measures designs. Journal of Statistical Computation and Simulation, 51(1):57– 69, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Polansky:1997:ICV

[Pol97] Asan M. Polansky. Inexact control variates for the iterated bootstrap. Journal of Statistical Computation and Simulation, 59(1): 83–99, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Poon:1980:MCS

[Poo80] Amy H. Poon. A Monte Carlo study of the power of some ksample tests for ordered binomial alternatives. Journal of Statistical Computation and Simulation, 11(3–4):251–259, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Popp:2008:NIO

[Pop08] Stephan Popp. New innovational outlier unit root test with a break at an unknown time. Journal of Statistical Computation and Simulation, 78(12):1145–1161, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Porteous:1988:PSR

[Por88] B. T. Porteous. Plotting squared radii: improving the χ^{2} approximation. Journal of Statistical Computation and Simulation, 29(1):35–41, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Posten:1978:RTS

[Pos78] Harry O. Posten. The robustness of the two-sample t-test over the Pearson system. Journal of Statistical Computation and Simulation, 6(3–4):295–311, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Posten:1979:ROS

[Pos79] Harry O. Posten. The robustness of the one-sample t-test over the Pearson system. Journal of Statistical Computation and Simulation, 9(2):133–149, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Posten:1982:TSW

[Pos82] Harry O. Posten. Two-sample Wilcoxon power over the Pearson system and comparison with the t-test. Journal of Statistical Computation and Simulation, 16(1):1–18, 1982. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Posten:1994:NAN

[Pos94] Harry O. Posten. A new algorithm for the noncentral t distribution function. Journal of Statistical Computation and Simulation, 51(1):79–87, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Potter:1981:OAC

[Pot81] Randall W. Potter. Optimal allocation of computer time in Monte Carlo studies of interval procedures. Journal of Statistical Computation and Simulation, 13(3–4):297–307, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pouloukas:2004:ECL

[Pou04] Stavros Pouloukas. Estimation and comparison of lognormal parameters in the presence of censored data. Journal of Statistical Computation and Simulation, 74(3):157–169, 2004. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pope:1973:GFT

[PP73] P. T. Pope and E. N. Powers. A goodness-of-fit test using randomly selected cell boundaries. Journal of Statistical Computation and Simulation, 2(4):343–347, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pfaff:1980:NAA

[PP80] Th. Pfaff and J. Pfanzagl. On the numerical accuracy of approximations based on Edgeworth expansions. Journal of Statistical Computation and Simulation, 11(3–4):223–239, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pfaff:1985:AAE

[PP85] Th. Pfaff and J. Pfanzagl. On the accuracy of asymptotic expansions for power functions. Journal of Statistical Computation and Simulation, 22(1):1–25, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Park:2003:CMG

[PP03] Sangun Park and Dongryeon Park. Correcting moments for goodness of fit tests based on two entropy estimates. Journal of Statistical Computation and Simulation, 73(9):685–694, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Park:2006:ASD

[PP06a] Chanseok Park and W. J. Padgett. Analysis of strength distributions of multi-modal failures using the EM algorithm. Journal of Statistical Computation and Simulation, 76(7):619–636, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Park:2006:PNE

[PP06b] Dongryeon Park and Sangun Park. Parametric and nonparametric estimators of ED100 α. Journal of Statistical Computation and Simulation, 76(8):661–672, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Perez:2006:CBN

[PP06c] Timothy D. Perez and Jeffrey S. Pontius. Conventional bootstrap and normal confidence interval estimation under adaptive cluster sampling. Journal of Statistical Computation and Simulation, 76(9):755–764, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Patil:1977:BTP

[PPK77] S. A. Patil, D. I. Patel, and J. L. Kovner. On bivariate truncated Poisson distribution. Journal of Statistical Computation and Simulation, 6(1):49–66, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Perera:2006:SAM

[PPRW06] D. I. Perera, M. S. Peiris, J. Robinson, and N. C. Weber. Saddlepoint approximation methods for testing of serial correlation in panel time series data. Journal of Statistical Computation and Simulation, 76(11):1001–1013, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Peterson:1984:MCI

[PQ84] Bercedis Peterson and Dana Quade. A Monte Carlo investigation into the properties of a proposed robust one-sample test of location. Journal of Statistical Computation and Simulation, 19 (2):113–127, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Plant:1989:ACN

[PQ89] Mark W. Plant and Richard E. Quandt. On the accuracy and cost of numerical integration in several variables. Journal of Statistical Computation and Simulation, 32(4):229–248, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pirie:1984:SET

[PR84] Walter R. Pirie and Howard H. Rauch. Simulated efficiencies of tests and estimators from general linear models analysis based on ranks: the two-way layout with interaction. Journal of Statistical Computation and Simulation, 20(3):197–204, 1984. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Paulson:1987:SED

[PRS87] A. S. Paulson, P. Roohan, and P. Sullo. Some empirical distribution function tests for multivariate normality. Journal of Statistical Computation and Simulation, 28(1):15–30, 1987. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pruitt:1993:SSC

[Pru93] Ronald S. Pruitt. Small sample comparison of six bivariate survival curve estimators. Journal of Statistical Computation and Simulation, 45(3–4):147–167, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Price:1985:MCS

[PS85] Thomas E. Price, Jr. and D. P. Story. Monte Carlo simulation of numerical integration. Journal of Statistical Computation and Simulation, 23(1–2):97–112, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Park:1998:AGC

[PS98] Chul Gyu Park and Dong Wan Shin. An algorithm for generating correlated random variables in a class of infinitely divisible distributions. Journal of Statistical Computation and Simulation, 61(1–2):127–139, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pace:1999:PEB

[PS99] Lungi Pace and Alessandra Salvan. Point estimation based on confidence intervals: exponential families. Journal of Statistical Computation and Simulation, 64(1):1–21, 1999. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Polzehl:2009:NSA

[PS09] Jörg Polzehl and Stefan Sperlich. A note on structural adaptive dimension reduction. Journal of Statistical Computation and Simulation, 79(6):805–818, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Paul:2003:EID

[PSB03] Sudhir Paul, Krishna Saha, and Uditha Balasooriya. An empirical investigation of different operating characteristics of several estimators of the intraclass correlation in the analysis of binary data. Journal of Statistical Computation and Simulation, 73 (7):507–523, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Paige:1975:CSD

[PSW75] C. C. Paige, George P. H. Styan, and Peter G. Wachter. Computation of the stationary distribution of a Markov chain. Journal of Statistical Computation and Simulation, 4(3):173–186, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Poston:1998:PAS

[PSW98] W. L. Poston, J. L. Solka, and E. J. Wegman. A parallel algorithm for subset selection. Journal of Statistical Computation and Simulation, 60(1):1–17, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Padadopoulos:1989:BBL

[PT89] Alex S. Padadopoulos and Ram C. Tiwari. Bayesian bootstrap lower confidence interval estimation of the reliability and failure rate. Journal of Statistical Computation and Simulation, 32(3): 185–192, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ploymenis:1999:NDL

[PT99] A. Ploymenis and D. M. Titterington. A note on the distribution of the likelihood ratio statistic for normal mixture models with known proportions. Journal of Statistical Computation and Simulation, 64(2):167–175, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pan:2002:GLM

[PT02] Guohua Pan and Winson Taam. On generalized linear model method for detecting dispersion effects in unreplicated factorial designs. Journal of Statistical Computation and Simulation, 72 (6):431–450, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Padgett:2003:LCB

[PT03] W. J. Padgett and Meredith Tomlinson. Lower confidence bounds for percentiles of Weibull and Birnbaum–Saunders distributions. Journal of Statistical Computation and Simulation, 73(6):429–443, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pullin:1979:GNV

[Pul79] D. I. Pullin. Generation of normal variates with given sample mean and variance. Journal of Statistical Computation and Simulation, 9(4):303–309, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Price:1993:RLS

[PV93] Lydia J. Price and Wilfried R. Vanhonacker. Recursive leastsquares approach to data transferability: exposition and numerical results. Journal of Statistical Computation and Simulation, 47(3–4):147–160, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Prescott:1983:MLE

[PW83] P. Prescott and A. T. Walden. Maximum likelihood estimation of the parameters of the three-parameter generalized extremevalue distribution from censored samples. Journal of Statistical Computation and Simulation, 16(3–4):241–250, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Piegorsch:1986:TSE

[PW86] Walter W. Piegorsch and Clarice R. Weinberg. Testing synergistic effects for simultaneous exposures with stratified dichotomous response. Journal of Statistical Computation and Simulation, 26(1–2):1–19, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pages:1997:SLD

[PX97] Gilles Pagès and Yi-Jun Xiao. Sequences with low discrepancy and pseudo-random numbers: theoretical results and numerical tests. Journal of Statistical Computation and Simulation, 56 (2):163–188, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Perakis:2002:PCI

[PX02] Michael Perakis and Evdokia Xekalaki. A process capability index that is based on the proportion of conformance. Journal of Statistical Computation and Simulation, 72(9):707–718, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Perakis:2005:PCI

[PX05] Michael Perakis and Evdokia Xekalaki. A process capability index for discrete processes. Journal of Statistical Computation and Simulation, 75(3):175–187, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pettit:1999:BAI

[PY99] L. I. Pettit and K. D. S. Young. Bayesian analysis for inverse Gaussian lifetime data with measures of degradation. Journal of Statistical Computation and Simulation, 63(3):217–234, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Payton:1993:CUS

[PYC93] Mark E. Payton, Linda J. Young, and J. P. Chandler. C404. Using SAS to solve a system of non-linear equations. Journal of Statistical Computation and Simulation, 47(1–2):114–117, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qian:2003:EWP

[QC03] Lianfen Qian and José Correa. Estimation of Weibull parameters for grouped data with competing risks. Journal of Statistical Computation and Simulation, 73(4):261–275, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Quesenberry:1983:APS

[QD83] C. P. Quesenberry and E. Jacquelin Dietz. Agreement probabilities for some CPIT–Neyman smooth tests. Journal of Statistical Computation and Simulation, 17(2):125–131, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Quintana:1987:LSI

[QFG87] José M. Quintana, O’Reilly J. Federico, and Susana Gómez. Least squares with inequality restrictions: a symmetric positivedefinite linear complementarity problem algorithm. Journal of Statistical Computation and Simulation, 28(2):127–143, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Quesenberry:1980:CBU

[QH80] C. P. Quesenberry and Craige Hales. Concentration bands for uniformity plots. Journal of Statistical Computation and Simulation, 11(1):41–53, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qiu:2001:RCE

[QL01] Peihua Qiu and Chap Le. ROC curve estimation based on local smoothing. Journal of Statistical Computation and Simulation, 70(1):55–69, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Quesenberry:1977:PSS

[QM77] C. P. Quesenberry and F. L. Miller, Jr. Power studies of some tests for uniformity. Journal of Statistical Computation and Simulation, 5(3):169–191, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Quesenberry:1982:DRF

[QQ82] C. P. Quesenberry and Chatles Quesenberry, Jr. On the distribution of residuals form fitted parametric models. Journal of Statistical Computation and Simulation, 15(2-3):129–140, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qian:1992:SRE

[QT92] W. Qian and D. M. Titterington. Stochastic relaxations and em algorithms for Markov random fields. Journal of Statistical Computation and Simulation, 40(1–2):55–69, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qu:2006:BWE

[Qu06] Leming Qu. Bayesian wavelet estimation of partially linear models. Journal of Statistical Computation and Simulation, 76(7):

605–617, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Quan:1995:STI

[QY95] Hui Quan and Kai F. Yu. Sequential testing with interval censored data. Journal of Statistical Computation and Simulation, 51(2-4):239–247, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Quan:1992:JPH

[QyT92] Hui Quan and Wei yann Tsai. Jackknife for the proportional hazards model. Journal of Statistical Computation and Simulation, 43(3–4):163–176, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rohatgi:1981:CTE

[RA81] V. K. Rohatgi and Rashpal Ahluwalia. C86. Testing equality of variances in sampling from two normal populations. Journal of Statistical Computation and Simulation, 12(2):140–142, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Raqab:2001:ELS

[RA01] Mohammad M. Raqab and M. Ahsanullah. Estimation of the location and scale parameters of generalized exponential distribution based on order statistics. Journal of Statistical Computation and Simulation, 69(2):109–123, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rodriguez-Avi:2005:CTS

[RACSGJ05] José Rodríguez-Avi, Antonio Conde-Sánchez, and Ramón Gutiérrez-Jáimez. Computational techniques in the study of discrete distributions generated by the function _{3}F_{2}. Journal of Statistical Computation and Simulation, 75(7):531–542, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ramsey:1989:DBT

[Ram89] Philip H. Ramsey. Determining the best test for the 2 × 2 comparative trial. Journal of Statistical Computation and Simulation, 34(1):51–65, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rayner:1985:CBP

[Ray85] J. C. W. Rayner. C230. Bias of the Pearson chi-squared test. Journal of Statistical Computation and Simulation, 21(3–4):329– 331, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rigdon:1988:EIF

[RB88] Steven E. Rigdon and Asit P. Basu. Estimating the intensity function of a Weibull process at the current time: failure truncated case. Journal of Statistical Computation and Simulation, 30(1):17–38, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ruppel:1992:CCR

[RB92] Patricia L. Ruppel and Stephen Buxser. Combining and comparing results from combining and comparing results from receptor binding assays. Journal of Statistical Computation and Simulation, 44(1–2):75–92, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Robinson:2000:MMP

[RB00a] Timothy J. Robinson and Jeffery B. Birch. Model misspecification in parametric dual modeling. Journal of Statistical Computation and Simulation, 66(2):113–126, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rojo:2000:EQF

[RB00b] Javier Rojo and Julia Bader. On estimating the quantile function for distributions constrained by star-shaped ordering. Journal of Statistical Computation and Simulation, 67(2):133–172, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Raghavan:1998:AMI

[RC98] Nandini Raghavan and Dennis D. Cox. Adaptive mixture importance sampling. Journal of Statistical Computation and Simulation, 60(3):237–259, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Russell:1981:ACO

[REK81] K. G. Russell, J. A. Eccleston, and G. J. Knudsen. Algorithms for the construction of (M, S)-optimal block designs and rowcolumn designs. Journal of Statistical Computation and Simulation, 12(2):93–105, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Russell:2009:DCS

[RELW09] K. G. Russell, J. A. Eccleston, S. M. Lewis, and D. C. Woods. Design considerations for small experiments and simple logistic regression. Journal of Statistical Computation and Simulation, 79(1):81–91, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rawlings:1986:SDA

[RFGE86] Robert R. Rawlings, Vivian B. Faden, Barry I. Graubard, and Michael J. Eckardt. A study on discriminant analysis techniques applied to multivariate lognormal data. Journal of Statistical Computation and Simulation, 26(1–2):79–100, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rahman:1993:IAD

[RG93a] M. S. Rahman and R. P. Gupta. Identification and approximation of distribution by moment ratios. Journal of Statistical Computation and Simulation, 48(1–2):59–65, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rice:1993:CVA

[RG93b] William R. Rice and Steven D. Gaines. Calculating P -values for Anova with unequal variances. Journal of Statistical Computation and Simulation, 46(1–2):19–22, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ramanayake:2002:CPL

[RG02] Asoka Ramanayake and Arjun Gupta. Change points with linear trend followed by abrupt change for the exponential distribution. Journal of Statistical Computation and Simulation, 72(4):263– 278, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ramsey:2006:ICE

[RG06] Philip H. Ramsey and W. A. Gibson. Improved communality estimation in factor analysis. Journal of Statistical Computation and Simulation, 76(2):93–101, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rudolph:1976:RNG

[RH76] E. Rudolph and D. M. Hawkins. Random number generators in cyclic queuing applications. Journal of Statistical Computation and Simulation, 5(1):65–71, 1976. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Remme:1980:SCL

[RHH80] J. Remme, J. D. F. Habbema, and J. Hermans. A simulative comparison of linear, quadratic and kernel discrimination. Journal of Statistical Computation and Simulation, 11(2):87–106, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rhiel:1986:ENN

[Rhi86] G. Steven Rhiel. The effect of non-normality on the use of the range in estimating the population standard deviation. Journal of Statistical Computation and Simulation, 24(1):71–82, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Riddell:1980:ESR

[Rid80] W. Craig Riddell. Estimating switching regressions: a computational note. Journal of Statistical Computation and Simulation, 10(2):95–101, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Riddel:2003:FSP

[Rid03] Mary Riddel. Finite sample properties of nonstationary binary response models: A Monte Carlo analysis. Journal of Statistical Computation and Simulation, 73(3):203–222, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rigdon:1995:IEC

[Rig95] Steven E. Rigdon. An integral equation for the in-control average run length of a multivariate exponentially weighted moving average control chart. Journal of Statistical Computation and Simulation, 52(4):351–365, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Remmenga:1995:CIP

[RJ95] Marta D. Remmenga and Dallas E. Johnson. A comparison of inference procedures in unbalanced split-plot designs. Journal of Statistical Computation and Simulation, 51(2-4):353–367, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rubin:2006:EGE

[RJ06] Herman Rubin and Brad C. Johnson. Efficient generation of exponential and normal deviates. Journal of Statistical Computation and Simulation, 76(6):509–518, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Reiter:2005:CDR

[RK05] Jerome P. Reiter and Christine N. Kohnen. Categorical data regression diagnostics for remote access servers. Journal of Statistical Computation and Simulation, 75(11):889–903, 2005. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

RoldanNofuentes:2008:EAC

[RL08] J. A. Roldan Nofuentes and J. D. Luna Del Castillo. EM algorithm for comparing two binary diagnostic tests when not all the patients are verified. Journal of Statistical Computation and Simulation, 78(1):19–35, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rojo:1996:ESO

[RM96] Javvier Rojo and Zheng Ma. On the estimation of stochastically ordered survival functions. Journal of Statistical Computation and Simulation, 55(1–2):1–21, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Raqab:2002:BPT

[RM02a] Mohamed Raqab and Mohamed Madi. Bayesian prediction of the total time on test using doubly censored Rayleigh data. Journal of Statistical Computation and Simulation, 72(10):781–789, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rollans:2002:EOS

[RM02b] S. Rollans and D. L. McLeish. Estimating the optimum of a stochastic system using simulation. Journal of Statistical Computation and Simulation, 72(5):357–377, 2002. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Raqab:2005:BIG

[RM05] Mohamed Z. Raqab and Mohamed T. Madi. Bayesian inference for the generalized exponential distribution. Journal of Statistical Computation and Simulation, 75(10):841–852, 2005. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Robert:1988:PTC

[RMH88] Crawford G. Robert, Geurts D. Michael, and Rinne J. Heikki. Permanent and temporary components of a time series in market analysis and forecasting. Journal of Statistical Computation and Simulation, 29(2):157–177, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rayner:1990:HLR

[RML90] J. C. W. Rayner, B. E. J. Manly, and G. F. Liddell. Hierarchic likelihood ratio tests for equality of covariance matrices. Journal of Statistical Computation and Simulation, 35(1–2):91– 99, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rodger:1976:TSN

[Rod76] R. S. Rodger. Tables of Stein’s non-central parameter Dβ; v_{1}, v_{2} required to set power for numerical alternatives to H_{0} tested by two-stage sampling Anova. Journal of Statistical Computation and Simulation, 5(1):1–22, 1976. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rodriguez:2006:FSS

[Rod06] Ricardo J. Rodriguez. Finite sequences of St. Petersburg games: inferences from a simulation study. Journal of Statistical Computation and Simulation, 76(10):925–933, 2006. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rodriguez:2007:FSB

[Rod07] Gabriel Rodríguez. Finite sample behaviour of the level shift model using quasi-differenced data. Journal of Statistical Computation and Simulation, 77(10):889–905, 2007. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rohatgi:1988:CMW

[Roh88] Vijay K. Rohatgi. C298. On the moments of f(x) when f is discrete. Journal of Statistical Computation and Simulation, 29 (4):340–343, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Roland:2001:RLM

[Rol01] Fried Roland. Recursive linear mixed models: some results on shared parameter estimation. Journal of Statistical Computation and Simulation, 70(2):135–160, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ronning:1989:MLE

[Ron89] G. Ronning. Maximum likelihood estimation of Dirichlet distributions. Journal of Statistical Computation and Simulation, 32 (4):215–221, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rost:1995:SSW

[Ros95] D. Rost. A simulation study of the weighted k-means cluster procedure. Journal of Statistical Computation and Simulation, 53(1–2):51–63, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rosadi:2006:OIG

[Ros06] Dedi Rosadi. Order identification for Gaussian moving averages using the codifference function. Journal of Statistical Computation and Simulation, 76(6):553–559, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Royston:1989:CSW

[Roy89] J. P. Royston. Correcting the Shapiro–Wilk W for ties. Journal of Statistical Computation and Simulation, 31(4):237–249, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Roy:1993:CEC

[Roy93] Tapon Roy. C396. An exact confidence interval for the ratio of means using regression methods. Journal of Statistical Computation and Simulation, 47(1–2):89–90, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rahman:2001:ETP

[RP01] Mezbahur Rahman and Larry M. Pearson. Estimation in twoparameter exponential distributions. Journal of Statistical Computation and Simulation, 70(4):371–386, 2001. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rey:1977:OSD

[RR77] G. Rey and H. Riedwyl. One-sample distribution-free test statistic V (n, k) sensitive to location differences. Journal of Statistical Computation and Simulation, 6(2):75–81, 1977. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rutemiller:1979:SSP

[RR79] Herbert Rutemiller and E. Schafer Ray. Simulation studies of the power of Birnbaum’s test. Journal of Statistical Computation and Simulation, 9(4):261–281, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Richardson:1982:EEA

[RR82] David H. Richardson and Robert J. Rohr. Experimental evidence on approximations to two-stage least-squares distributions. Journal of Statistical Computation and Simulation, 15 (2-3):221–232, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ramsey:1993:UVC

[RR93a] Philip H. Ramsey and Patricia P. Ramsey. Updated version of the critical values of the standardized fourth moment. Journal of Statistical Computation and Simulation, 44(3–4):231–241, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Razzaghi:1993:LSE

[RR93b] Mehdi Razzaghi and Mohsen Razzaghi. Legendre series estimate of a distribution function. Journal of Statistical Computation and Simulation, 48(1–2):19–27, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ramsey:2001:LST

[RR01] Patricia P. Ramsey and Philip H. Ramsey. Least-squares tests of time-series, invention effects with and without autocorrelations. Journal of Statistical Computation and Simulation, 70(1):33– 44, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ramsey:2003:CLS

[RR03] Patricia Ramsey and Philip Ramsey. Comparing least-squares lines for testing level changes in interrupted time-series analysis. Journal of Statistical Computation and Simulation, 73(1):31–44, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ramsey:2006:RTL

[RR06] Patricia P. Ramsey and Philip H. Ramsey. Robust testing of level changes in interrupted time-series analysis. Journal of Statistical Computation and Simulation, 76(10):913–923, 2006. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ramsey:2008:BIT

[RR08] Philip H. Ramsey and Patricia P. Ramsey. Brief investigation of tests of variability in the two-sample case. Journal of Statistical Computation and Simulation, 78(12):1125–1131, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rahni:1997:AGS

[RRCD97] N. Rahni, N. Ramdani, Y. Candau, and P. Dalicieux. Application of group screening to dynamic building energy simulation models. Journal of Statistical Computation and Simulation, 57 (1-4):285–304, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Reisen:2006:ESL

[RRP06] Valderio A. Reisen, Alexandre L. Rodrigues, and Wilfredo Palma. Estimating seasonal long-memory processes: a Monte Carlo study. Journal of Statistical Computation and Simulation, 76(4):305–316, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Robert:1999:CCM

[RRT99] Christian P. Robert, Tobias Rydén, and D. M. Titterington. Convergence controls for MCMC algorithms, with applications to hidden Markov chains. Journal of Statistical Computation and Simulation, 64(4):327–355, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Raghunandanan:1977:NPI

[RS77] K. Raghunandanan and R. Srinivasan. Nonparametric prediction intervals with controlled tail proportions. Journal of Statistical Computation and Simulation, 5(2):85–90, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rutemiller:1983:UE

[RS83] Herbert C. Rutemiller and Ray E. Schafer. Unbiased estimation of. Journal of Statistical Computation and Simulation, 17(2): 101–112, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rubinstein:1985:VRU

[RS85] R. Y. Rubinstein and G. Samorodnitsky. Variance reduction by the use of common and antithetic random variables. Journal of Statistical Computation and Simulation, 22(2):161–180, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rocke:1986:SPR

[RS86] David M. Rocke and David F. Shanno. The scale problem in robust regression M-estimates. Journal of Statistical Computation and Simulation, 24(1):47–69, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rayner:1988:AUC

[RS88a] J. C. W. Rayner and A. J. Seyb. Anova under categorisation. Journal of Statistical Computation and Simulation, 29(3):181– 193, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Robertazzi:1988:OMP

[RS88b] T. G. Robertazzi and S. C. Schwartz. Optimizing the measurement processing order for linear regression. Journal of Statistical Computation and Simulation, 30(4):233–258, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Robinson:1989:WSM

[RS89] Ian Robinson and Simon Sheather. Weighted selection for the multiset X + X with application to R-estimates and associated confidence limits. Journal of Statistical Computation and Simulation, 31(1):19–35, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rohatgi:1990:CUD

[RS90a] Vijay K. Rohatgi and Gabor J. Székely. C343. An unexpected decomposition of a symmetric random variable. Journal of Statistical Computation and Simulation, 34(2-3):162–164, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rom:1990:API

[RS90b] Dror Rom and Sanat K. Sarkar. Approximating probability integrals of multivariate normal using association models. Journal of Statistical Computation and Simulation, 35(1–2):109–119, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rohatgi:1992:CBS

[RS92] Vijay K. Rohatgi and Gabor J. Szekely. C385. On the background of some correlation inequalities. Journal of Statistical Computation and Simulation, 40(3–4):260–262, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rotelli:1996:BR

[RS96] Mathew D. Rotelli and Eric Smith. Book reviews. Journal of Statistical Computation and Simulation, 54(1-3):273–274, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rao:1997:SSA

[RS97] G. Prabhakar Rao and P. K. Sarkar. Sensitivity studies of air scattered neutron dose from particle accelerators. Journal of Statistical Computation and Simulation, 57(1-4):261–270, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Riaz:2009:MDB

[RS09] Muhammad Riaz and Aamir Saghir. A mean deviation-based approach to monitor process variability. Journal of Statistical Computation and Simulation, 79(10):1173–1193, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ramberg:1978:GSO

[RT78] John S. Ramberg and Pandu R. Tadikamalla. On the generation of subsets of order statistics. Journal of Statistical Computation and Simulation, 6(3–4):239–241, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Raftery:1988:CHM

[RT88] Adrian E. Raftery and Elizabeth A. Thompson. C301. How many nuclear reactor accidents? Journal of Statistical Computation and Simulation, 29(4):347–350, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Raftery:1990:CWP

[RT90] Adrian E. Raftery and Elizabeth A. Thompson. C351. What is the probability of a serious nuclear reactor accident? Journal of Statistical Computation and Simulation, 36(1):31–34, 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rudas:1986:MCC

[Rud86] T. Rudas. A Monte Carlo comparison of the small sample behaviour of the Pearson, the likelihood ratio and the Cressie–Read statistics. Journal of Statistical Computation and Simulation, 24 (2):107–120, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rufibach:2007:CML

[Ruf07] Kaspar Rufibach. Computing maximum likelihood estimators of a log-concave density function. Journal of Statistical Computation and Simulation, 77(7):561–574, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rothblum:1998:MLE

[RV98] Uriel G. Rothblum and Yehuda Vardi. Maximum likelihood estimation of cell probabilities in constrained multinomial models. Journal of Statistical Computation and Simulation, 61(1–2):141– 161, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rijmen:2008:APV

[RV08a] Frank Rijmen and Jirí Vomlel. Assessing the performance of variational methods for mixed logistic regression models. Journal of Statistical Computation and Simulation, 78(8):765–779, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rukhin:2008:TRA

[RV08b] Andrew L. Rukhin and Zeev Volkovich. Testing randomness via aperiodic words. Journal of Statistical Computation and Simulation, 78(12):1133–1144, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Romero-Villafranca:2008:OST

[RVZRZP08] R. Romero-Villafranca, L. Zúnica, R. Romero-Zúnica, and J. A. Pagura. One-sided tolerance limits for unbalanced one-way random effects models: a generalized Mee and Owen procedure. Journal of Statistical Computation and Simulation, 78(12):1215– 1227, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Roach:1977:OAS

[RW77] William Roach and Roger Wright. Optimal antithetic sampling plans. Journal of Statistical Computation and Simulation, 5(2): 99–114, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rudolfer:1993:EOP

[RW93] Stephan M. Rudolfer and Peter C. Watson. Evaluation of orthant probabilities for singular bi- and trivariate normal distributions. Journal of Statistical Computation and Simulation, 48 (3–4):219–232, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Runger:1996:UPR

[RW96] George C. Runger and Thomas R. Willemain. Using partial ranking information in the design of small-sample comparisons. Journal of Statistical Computation and Simulation, 54(1-3):75– 86, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ravishanker:1995:SET

[RWD95] Nalini Ravishanker, Lilian S. Y. Wu, and Dipak K. Dey. Shrinkage estimation in time series using a bootstrapped covariance estimate. Journal of Statistical Computation and Simulation, 53 (3–4):259–267, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rudolfer:1995:OMU

[RWL95] Stephan M. Rudolfer, Peter C. Watson, and Emmanuel Lesaffre. Are ordinal models useful for classification? A revised analysis. Journal of Statistical Computation and Simulation, 52(2):105– 132, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ryan:1980:NMG

[Rya80] Thomas P. Ryan. A new method of generating correlation matrices. Journal of Statistical Computation and Simulation, 11 (2):79–85, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schwertman:1979:SIV

[SA79] N. C. Schwertman and D. M. Allen. Smoothing an indefinite variance-covariance matrix. Journal of Statistical Computation and Simulation, 9(3):183–194, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sheffield:1983:CNR

[SA83] L. G. Sheffield and R. R. Anderson. C181. A note on the randomness of protein sequences. Journal of Statistical Computation and Simulation, 18(4):315–316, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sharma:1992:RNN

[SA92] Subhash C. Sharma and Mukhtar M. Ali. Robustness to nonnormality of the null distribution and power of the Durbin– Watson test in regressions with and without an intercept. Journal of Statistical Computation and Simulation, 42(1–2):93–105, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sklar:1995:PLA

[SA95] Michael G. Sklar and Ronald D. Armstrong. A piecewise linear approximation procedure for L_{p} norm curve fitting. Journal of Statistical Computation and Simulation, 52(4):323–335, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sievers:2004:RER

[SA04a] Gerald L. Sievers and Asheber Abebe. Rank estimation of regression coefficients using iterated reweighted least squares. Journal of Statistical Computation and Simulation, 74(11):821– 831, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Singh:2004:CEE

[SA04b] V. P. Singh and Maqsood Ahmad. A comparative evaluation of the estimators of the three-parameter generalized Pareto distribution. Journal of Statistical Computation and Simulation, 74 (2):91–106, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sherman:2006:EBA

[SAC06] Michael Sherman, Tatiyana V. Apanasovich, and Raymond J. Carroll. On estimation in binary autologistic spatial models. Journal of Statistical Computation and Simulation, 76(2):167– 179, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Samawi:2003:MPS

[SAD03] Hani Samawi and Walid Abu-Dayyeh. More powerful sign test using median ranked set sample: Finite sample power comparison. Journal of Statistical Computation and Simulation, 73(10): 697–708, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sahai:1979:SSB

[Sah79] Hardeo Sahai. A supplement to Sowey’s bibliography on random number generation and related topics. Journal of Statistical Computation and Simulation, 10(1):31–52, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Sahu:2002:BEM

[Sah02] Sujit K. Sahu. Bayesian estimation and model choice in item response models. Journal of Statistical Computation and Simulation, 72(3):217–232, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Seong:2008:NSR

[SAJ08] Byeongchan Seong, Sung K. Ahn, and Yongil Jeon. A note on spurious regression in seasonal time series. Journal of Statistical Computation and Simulation, 78(9):843–851, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Santanna:1989:GIR

[San89] A. P. Sant’anna. Generalized inverses in recursive least squares estimation with multiple additional observations. Journal of Statistical Computation and Simulation, 33(1):53–63, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sarkar:1998:CCB

[Sar98] Sanat K. Sarkar. C452. Correction to: On improving the min test for the analysis of combination drug trials, JSCS 51 (1995), pp. 197–213. Journal of Statistical Computation and Simulation, 60(2):180–181, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [SSW95].

Scholz:1973:TUC

[SB73] J. E. Scholz and D. R. Barr. Tables for unbiased confidence intervals for ratios of variances when sampling from two independent normal distributions. Journal of Statistical Computation and Simulation, 2(1):45–53, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shenton:1982:LEP

[SB82a] L. R. Shenton and K. O. Bowman. A Lagrange expansion for the parameters of Johnson’s S_{u}. Journal of Statistical Computation and Simulation, 15(2-3):89–95, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shiue:1982:ELS

[SB82b] Wei Kei Shiue and Lee J. Bain. Experiment length and sample size for negative binomial sampling. Journal of Statistical Computation and Simulation, 16(1):57–64, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Somerville:1982:CML

[SB82c] Paul N. Somerville and S. J. Bean. A comparison of maximum likelihood and least squares for the estimation of a cumulative distribution. Journal of Statistical Computation and Simulation, 14(3–4):229–239, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shenton:1985:CCF

[SB85] L. R. Shenton and K. O. Bowman. C239. The continued fraction for ln(z): a Stieltjes conjecture. Journal of Statistical Computation and Simulation, 23(1–2):133–136, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sarkar:1986:MMP

[SB86] A. Sarkar and A. K. Bhattacharji. A Markov model in process control. Journal of Statistical Computation and Simulation, 26 (1–2):115–126, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shih:1987:CMV

[SB87] W. J. Shih and Bruce Binkowitz. C282. Median versus geometric mean for lognormal samples. Journal of Statistical Computation and Simulation, 28(1):81–83, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Samaranayake:1988:CIT

[SB88] V. A. Samaranayake and Lee J. Bain. A confidence interval for treatment component-of-variance with applications to differences in means of two exponential distributions. Journal of Statistical Computation and Simulation, 29(4):317–332, 1988. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sinha:1991:CUB

[SB91] Kishore Sinha and Bhagwandas. C381. Unequal block-size designs with orthogonal factorial structure. Journal of Statistical Computation and Simulation, 39(3):185–187, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Sharples:1992:RAC

[SB92] Katrina Sharples and Norman Breslow. Regression analysis of correlated binary data: some small sample results for the estimating equation approach. Journal of Statistical Computation and Simulation, 42(1–2):1–20, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sutradhar:1993:MCC

[SB93] Brajendra C. Sutradhar and Roy F. Bartlett. Monte Carlo comparison of Wald’s, likelihood ratio and Rao’s tests. Journal of Statistical Computation and Simulation, 46(1–2):23–33, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Stewart:1996:IAD

[SB96] D. E. Stewart and M. S. Bebbington. An iterative aggregation/disaggregation procedure for modelling the long-term behaviour of continuous-time evanescent random processes. Journal of Statistical Computation and Simulation, 56(1):77–95, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Starnes:2000:ARM

[SB00] B. A. Starnes and Jeffrey B. Birch. Asymptotic results for model robust regression. Journal of Statistical Computation and Simulation, 66(1):19–33, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schoenfeld:2005:CPS

[SB05] David A. Schoenfeld and Michael Borenstein. Calculating the power or sample size for the logistic and proportional hazards models. Journal of Statistical Computation and Simulation, 75 (10):771–785, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sanjel:2008:LPA

[SB08] Deepak Sanjel and N. Balakrishnan. A Laguerre polynomial approximation for a goodness-of-fit test for exponential distribution based on progressively censored data. Journal of Statistical Computation and Simulation, 78(6):503–513, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

SantiagoMaia:2003:CLR

[SBC03] André Luis Santiago Maia, Antônio Carlos Braga Junior, and Gauss Cordeiro. Corrected likelihood ratio and score tests for the beta distribution. Journal of Statistical Computation and Simulation, 73(8):585–596, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sibson:1981:SRM

[SBO81] Robin Sibson, Adrian Bowyer, and Clive Osmond. Studies in the robustness of multidimensional scaling: Euclidean models and simulation studies. Journal of Statistical Computation and Simulation, 13(3–4):273–296, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schulman:1975:CAS

[SC75] Robert S. Schulman and Elliot M. Cramer. Comparison of the accuracies of some regression algorithms. Journal of Statistical Computation and Simulation, 4(2):85–93, 1975. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Stablein:1982:MCS

[SC82] Donald M. Stablein and Walter H. Carter, Jr. A Monte Carlo study of the effects of grouping on the analysis of censored survival data. Journal of Statistical Computation and Simulation, 14(2):113–123, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schwertman:1995:MPS

[SC95] Neil C. Schwertman and Nancy J. Carter. A more practical Scheffé-type multiple comparison procedure for commonly encountered numbers of comparisons. Journal of Statistical Computation and Simulation, 53(3–4):181–196, 1995. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Suen:2000:FRF

[SC00] Chung-Yi Suen and Hegang Chen. A family of regular fractional factorial designs with maximum resolution. Journal of Statistical Computation and Simulation, 66(1):67–78, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Simas:2009:APR

[SC09a] Alexandre B. Simas and Gauss M. Cordeiro. Adjusted Pearson residuals in exponential family nonlinear models. Journal of Statistical Computation and Simulation, 79(4):411–425, 2009. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Stosic:2009:UMM

[SC09b] Borko D. Stosić and Gauss M. Cordeiro. Using Maple and Mathematica to derive bias corrections for two parameter distributions. Journal of Statistical Computation and Simulation, 79 (6):751–767, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Seong:2007:ISC

[SCA07] Byeongchan Seong, Sinsup Cho, and Sung K. Ahn. Inference of seasonal cointegration with linear restrictions. Journal of Statistical Computation and Simulation, 77(7):593–603, 2007. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shoukri:2005:EWP

[SCAH05] M. M. Shoukri, M. A. Chaudhary, and A. Al-Halees. Estimating P (Y < X) when X and Y are paired exponential variables. Journal of Statistical Computation and Simulation, 75(1):25– 38, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sutradhar:2007:EEP

[SCB07] Brajendra C. Sutradhar, David P. T. Chu, and Wasimul Bari. Estimation effects on powers of two simple test statistics in identifying an outlier in linear models. Journal of Statistical Computation and Simulation, 77(4):305–328, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schucany:1972:OSS

[Sch72] W. R. Schucany. Order statistics in simulation. Journal of Statistical Computation and Simulation, 1(3):281–286, 1972. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schafer:1974:APS

[Sch74] R. E. Schafer. On assessing the precision of simulations. Journal of Statistical Computation and Simulation, 3(1):67–69, 1974. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schwartz:1975:DAA

[Sch75] Ronald D. Schwartz. Discriminant analysis and its application to the stock market. Journal of Statistical Computation and Simulation, 3(3):217–223, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schmeiser:1978:GMM

[Sch78] Bruce Schmeiser. Generation of the maximum (minimum) value in digital computer simulation. Journal of Statistical Computation and Simulation, 8(2):103–115, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schechtman:1983:ITP

[Sch83] Edna Schechtman. Inference in two-phase regression: A simulation study with nonnormal observations. Journal of Statistical Computation and Simulation, 17(3):223–229, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schaefer:1986:AEL

[Sch86] Robert L. Schaefer. Alternative estimators in logistic regression when the data are collinear. Journal of Statistical Computation and Simulation, 25(1–2):75–91, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schechtman:1992:ATS

[Sch92] Kenneth B. Schechtman. Adaptive two-stage sampling in clinical trials. Journal of Statistical Computation and Simulation, 42 (1–2):55–64, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schafer:2002:LAF

[Sch02] Daniel W. Schafer. Likelihood analysis and flexible structural modeling for measurement error model regression. Journal of Statistical Computation and Simulation, 72(1):33–45, 2002. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sunwoo:1997:AUL

[ScK97] Hasik Sunwoo and Byung chun Kim. Analysis of the unbalanced linear model based on the balanced model. Journal of Statistical Computation and Simulation, 56(4):373–385, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Susko:2001:DTM

[SCK01] Edward Susko, Jiahua Chen, and J. D. Kalbfleisch. A diagnostic tool for mixture models. Journal of Statistical Computation and Simulation, 69(4):293–313, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Singh:1990:CEE

[SCM90] V. P. Singh, J. F. Cruise, and Ming Ma. A comparative evaluation of the estimators of the Weibull distribution by Monte Carlo simulation. Journal of Statistical Computation and Simulation, 36(4):229–241, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Scott:2002:MLE

[Sco02] William Andrew Scott. Maximum likelihood estimation using the empirical Fisher information matrix. Journal of Statistical Computation and Simulation, 72(8):599–611, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Segreti:1979:MCE

[SCW79] Anthony C. Segreti, Walter H. Carter, Jr., and Galen W. Wampler. Monte Carlo evaluation of several sequential optimization techniques when the response is time to an event. Journal of Statistical Computation and Simulation, 9(4):289– 301, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Segreti:1981:MCE

[SCW81] Anthony C. Segreti, Walter H. Carter, Jr., and Galen L. Wampler. A Monte Carlo evaluation of the robustness of several sequential optimization techniques when the response is time to an event. Journal of Statistical Computation and Simulation, 12 (3–4):209–216, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shaddick:2007:MCC

[SCW07] G. Shaddick, L. L. Choo, and S. G. Walker. Modelling correlated count data with covariates. Journal of Statistical Computation and Simulation, 77(11):945–954, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Swanepoel:1992:SNM

[SD92] J. W. H. Swanepoel and C. F. De Beer. Some new model selection criteria in simple regression. Journal of Statistical Computation and Simulation, 42(3–4):137–149, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sahama:2001:SSC

[SD01a] Tony R. Sahama and Neil T. Diamond. Sample size considerations and augmentation of computer experiments. Journal of Statistical Computation and Simulation, 68(4):307–319, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Smith:2001:EPI

[SD01b] Peter Smith and Brian Dawkins. Estimating the permanent by importance sampling from a finite population. Journal of Statistical Computation and Simulation, 70(3):197–214, 2001. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sullivan:2002:RPA

[SD02] Lisa M. Sullivan and Ralph B. D’Agostino. Robustness and power of analysis of covariance applied to data distorted from normality by floor effects: Non-homogeneous regression slopes. Journal of Statistical Computation and Simulation, 72(2):141– 165, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schmidt:2008:ISG

[SdFdGCM08] Alexandra M. Schmidt, Ma. de Fátima da G. Conceição, and Guido A. Moreira. Investigating the sensitivity of Gaussian processes to the choice of their correlation function and prior specifications. Journal of Statistical Computation and Simulation, 78 (8):681–699, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sneh:1993:CBM

[SDP93] Gulati Sneh, S. D. Durham, and W. J. Padgett. C401. On bounds for moments of the brittle fracture distribution. Journal of Statistical Computation and Simulation, 47(1–2):105–108, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schluchter:1990:SSA

[SE90] Mark D. Schluchter and Janet T. Elashoff. Small-sample adjustments to tests with unbalanced repeated measures assuming several covariance structures. Journal of Statistical Computation and Simulation, 37(1–2):69–87, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shukur:2002:SSP

[SE02] Ghazi Shukur and David Edgerton. The small sample properties of the reset test as applied to systems of equations. Journal of Statistical Computation and Simulation, 72(12):909–924, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sefton:1992:LMT

[Sef92] Martin Sefton. A Lagrange multiplier test of uniformity against beta alternatives. Journal of Statistical Computation and Simulation, 44(1–2):17–25, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sellami:2008:CSN

[Sel08] Afef Sellami. Comparative survey on nonlinear filtering methods: the quantization and the particle filtering approaches. Journal of Statistical Computation and Simulation, 78(2):93–113, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sen:2002:BEP

[Sen02] Ananda Sen. Bayesian estimation and prediction of the intensity of the power law process. Journal of Statistical Computation and Simulation, 72(8):613–631, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Strand:1977:NA

[SF77] Rodney Strand and Mike Farrell. News and abstracts. Journal of Statistical Computation and Simulation, 6(1):71, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schneeberger:1993:DRT

[SF93] H. Schneeberger and K. Fleischer. The distribution of the ratio of two variables. Journal of Statistical Computation and Simulation, 47(3–4):227–240, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shieh:2008:BAA

[SFC08] Grace S. Shieh, Tsai-Hung Fan, and Hsueh-Ping Chu. A Bayesian approach to assessing differential expression of microarray data. Journal of Statistical Computation and Simulation, 78(2):179–191, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schwertman:1985:MCS

[SFSS85] Neil C. Schwertman, William Flynn, Sallysue Stein, and Kathryn L. Schenk. A Monte Carlo study of alternative procedures for testing the hypothesis of parallelism for complete and incomplete growth curve data. Journal of Statistical Computation and Simulation, 21(1):1–37, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schabenberger:1996:PAS

[SG96] Oliver Schabenberger and Timothy G. Gregoire. Populationaveraged and subjectspecific approaches for clustered categorical data. Journal of Statistical Computation and Simulation, 54(1-3):231–253, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Surkis:1975:NVR

[SGH75] Julius Surkis, Gilbert R. Gordon, and John Hynes. A new variance reduction technique for evaluating system reliability problems by simulation. Journal of Statistical Computation and Simulation, 4(3):215–224, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Smith:1985:VSM

[SGH85] David W. Smith, D. S. Gill, and James J. Hammond. Variable selection in multivariate multiple regression. Journal of Statistical Computation and Simulation, 22(3–4):217–227, 1985. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shalabh:2004:UKE

[SGR04] Shalabh, Leon J. Gleser, and Ori Rosen. On the usefulness of knowledge of error variances in the consistent estimation of an unreplicated ultrastructural model. Journal of Statistical Computation and Simulation, 74(6):391–417, 2004. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Small:1994:EPS

[SGW94] J. P. Small, D. E. A. Giles, and K. J. White. The exact powers of some autocorrelation tests when relevant regressors are omitted. Journal of Statistical Computation and Simulation, 50(1–2):45– 57, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sugg:1972:CAH

[SH72] Merritt N. Sugg and Kermit Hutcheson. Computational algorithm for higher order moments of multinomial type distributions. Journal of Statistical Computation and Simulation, 1(2): 149–155, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shah:1985:CCD

[Sha85] Arvind K. Shah. C237. On the coefficient of determination. Journal of Statistical Computation and Simulation, 22(2):181– 184, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shah:1987:CEE

[Sha87a] Arvind K. Shah. C296. On estimating the exact variance of products. Journal of Statistical Computation and Simulation, 28(4):353–354, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sharma:1987:EAA

[Sha87b] Subhash C. Sharma. The effects of autocorrelation among errors on the consistency property of OLS estimator. Journal of Statistical Computation and Simulation, 28(1):43–62, 1987. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2001:BR

[Sha01a] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 71(3):267–268, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shayib:2001:EPP

[Sha01b] Mohammed A. Shayib. Estimation procedures and prediction intervals. Journal of Statistical Computation and Simulation, 70 (2):123–134, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2002:BRb

[Sha02a] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 72(9):775, 2002. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2002:BRc

[Sha02b] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 72(9):777–778, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2002:BRa

[Sha02c] Ramalingam Shanmugam. Book reviews. Journal of Statistical Computation and Simulation, 72(6):517–518, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2003:BGG

[Sha03a] Ramalingam Shanmugam. Biostatistical genetics and genetic epidemiology. Journal of Statistical Computation and Simulation, 73(7):543–544, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2003:BR

[Sha03b] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 73(10):773–774, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Shanmugam:2003:BEV

[Sha03c] Ramalingam Shanmugam. Encyclopedia of Environmetrics, Volume 2. Journal of Statistical Computation and Simulation, 73 (6):465–466, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2003:CSR

[Sha03d] Ramalingam Shanmugam. Case studies in reliability and maintenance. Journal of Statistical Computation and Simulation, 73 (9):695, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2003:PFE

[Sha03e] Ramalingam Shanmugam. A package of four edited books on structural equations modeling. Journal of Statistical Computation and Simulation, 73(5):387–388, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2004:BRa

[Sha04a] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 74(1):75–76, 2004. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2004:BRb

[Sha04b] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 74(1):77, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2004:BRc

[Sha04c] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 74(2):155, 2004. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2004:BRd

[Sha04d] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 74(3):233, 2004. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2004:BRe

[Sha04e] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 74(4):307–312, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2004:BRf

[Sha04f] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 74(7):543, 2004. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2004:BRg

[Sha04g] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 74(7):545–546, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2004:BRh

[Sha04h] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 74(8):621, 2004. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2004:BRi

[Sha04i] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 74(8):623, 2004. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2004:BRj

[Sha04j] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 74(9):697–700, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2004:BRk

[Sha04k] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 74(11):851, 2004. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2004:BRl

[Sha04l] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 74(11):853–854, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Shanmugam:2004:BRm

[Sha04m] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 74(12):923, 2004. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2004:BRT

[Sha04n] Ramalingam Shanmugam. Book review two books on geometry in statistics. Journal of Statistical Computation and Simulation, 74(12):925–926, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2005:BRa

[Sha05a] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 75(1):75–78, 2005. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2005:BRb

[Sha05b] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 75(2):155–157, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2005:BRBa

[Sha05c] Ramalingam Shanmugam. Book review: Books on allied topics to statistics. Journal of Statistical Computation and Simulation, 75(8):679–685, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2005:BRR

[Sha05d] Ramalingam Shanmugam. Book review reliable books about reliability and related concepts. Journal of Statistical Computation and Simulation, 75(6):497–500, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2005:BRS

[Sha05e] Ramalingam Shanmugam. Book review: Statistics and biomathematics books with health applications. Journal of Statistical Computation and Simulation, 75(4):307–311, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Shanmugam:2005:BRT

[Sha05f] Ramalingam Shanmugam. Book review: Two operations research books with statistical thoughts. Journal of Statistical Computation and Simulation, 75(4):313–314, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2006:BRBb

[Sha06a] Ramalingam Shanmugam. Book review books about likelihood and asymptotic statistics. Journal of Statistical Computation and Simulation, 76(11):1031–1033, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2006:BRBa

[Sha06b] Ramalingam Shanmugam. Book review books about time series and/or spatial data analysis. Journal of Statistical Computation and Simulation, 76(10):935–940, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2006:BRE

[Sha06c] Ramalingam Shanmugam. Book review excellent books about statistical distributions and probability theories. Journal of Statistical Computation and Simulation, 76(8):741–753, 2006. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2006:BRS

[Sha06d] Ramalingam Shanmugam. Book review selective books on software SAS, SPSS, and Internet outsourcing. Journal of Statistical Computation and Simulation, 76(12):1107–1113, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2007:BBR

[Sha07a] Ramalingam Shanmugam. Bioinformatics books are reviewed. Journal of Statistical Computation and Simulation, 77(11):1005– 1011, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2007:BCS

[Sha07b] Ramalingam Shanmugam. A book on common statistics errors is reviewed. Journal of Statistical Computation and Simulation, 77(12):1123, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2007:BRa

[Sha07c] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 77(3):271–272, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2007:BRc

[Sha07d] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 77(3):273–275, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2007:BRd

[Sha07e] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 77(4):359–364, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2007:BRe

[Sha07f] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 77(5):435–441, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2007:BRf

[Sha07g] Ramalingam Shanmugam. Book review. Journal of Statistical Computation and Simulation, 77(6):527–530, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2007:BRR

[Sha07h] Ramalingam Shanmugam. Book review a review of design and analysis books. Journal of Statistical Computation and Simulation, 77(2):185–188, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2007:BRB

[Sha07i] Ramalingam Shanmugam. Book review best Bayesian statistics books. Journal of Statistical Computation and Simulation, 77 (1):89–93, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2007:BRO

[Sha07j] Ramalingam Shanmugam. Book review outstanding psychometric books. Journal of Statistical Computation and Simulation, 77(7):625–628, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2008:BSA

[Sha08a] Ramalingam Shanmugam. A book on the statistical analysis of cost effectiveness data is reviewed. Journal of Statistical Computation and Simulation, 78(12):1279, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2008:BRa

[Sha08b] Ramalingam Shanmugam. Book reviews. Journal of Statistical Computation and Simulation, 78(9):869–872, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2008:BRb

[Sha08c] Ramalingam Shanmugam. Book reviews. Journal of Statistical Computation and Simulation, 78(11):1119–1123, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2009:BBR

[Sha09a] Ramalingam Shanmugam. Bioinformatics books are reviewed. Journal of Statistical Computation and Simulation, 79(7):959– 964, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2009:BCS

[Sha09b] Ramalingam Shanmugam. A book on common statistics errors is reviewed. Journal of Statistical Computation and Simulation, 79(4):427, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2009:BR

[Sha09c] Ramalingam Shanmugam. Book reviews. Journal of Statistical Computation and Simulation, 79(2):213–214, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Shanmugam:2009:BDM

[Sha09d] Ramalingam Shanmugam. Books on data mining are reviewed. Journal of Statistical Computation and Simulation, 79(1):107– 110, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2009:BEM

[Sha09e] Ramalingam Shanmugam. Books on econometrics and marketing are reviewed. Journal of Statistical Computation and Simulation, 79(3):321–322, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2009:BSS

[Sha09f] Ramalingam Shanmugam. Books on SAS, SPSS, and computing are reviewed. Journal of Statistical Computation and Simulation, 79(5):747–750, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2009:EBR

[Sha09g] Ramalingam Shanmugam. Encyclopedia of biostatistics are reviewed. Journal of Statistical Computation and Simulation, 79(10):1275–1276, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2009:EGG

[Sha09h] Ramalingam Shanmugam. Encyclopedia of genetics, genomics, proteomics and bioinformatics are reviewed. Journal of Statistical Computation and Simulation, 79(9):1169–1172, 2009. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2009:ESS

[Sha09i] Ramalingam Shanmugam. Encyclopedia of statistical sciences are reviewed. Journal of Statistical Computation and Simulation, 79(11):1389–1392, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2009:IBR

[Sha09j] Ramalingam Shanmugam. Informatics books are reviewed. Journal of Statistical Computation and Simulation, 79(6):855– 858, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2009:SHR

[Sha09k] Ramalingam Shanmugam. Software and hardware reliability books are reviewed. Journal of Statistical Computation and Simulation, 79(12):1491–1495, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2009:TIB

[Sha09l] Ramalingam Shanmugam. Tomographic imaging books are reviewed. Journal of Statistical Computation and Simulation, 79 (8):1063–1066, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sheehan:1983:AEF

[She83] Dennis P. Sheehan. Approximating estimators of the first-order autoregression. Journal of Statistical Computation and Simulation, 18(1):15–43, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shephard:1991:NIR

[She91] N. G. Shephard. Numerical integration rules for multivariate inversions. Journal of Statistical Computation and Simulation, 39(1–2):37–46, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2009:SEO

[She09] Pao-Sheng Shen. Semiparametric estimation of the occurrence rate for prevalent survival data in competing risk models. Journal of Statistical Computation and Simulation, 79(7):899–908, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schmitz:1985:SSP

[SHH85] P. I. M. Schmitz, J. D. F. Habbema, and J. Hermans. A simulation study of the performance of five discriminant analysis methods for mixtures of continuous and binary variables. Journal of Statistical Computation and Simulation, 23(1–2):69–95, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shih:1985:MLE

[Shi85] W. J. Shih. Maximum likelihood estimation and likelihood ratio test with incomplete pairs. Journal of Statistical Computation and Simulation, 21(3–4):187–194, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shiraishi:2007:UBC

[Shi07] Yuichi Shiraishi. An upper bound on the convergence time of the Gibbs sampler in Ising models. Journal of Statistical Computation and Simulation, 77(10):839–850, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shore:1986:AID

[Sho86] Haim Shore. An approximation for the inverse distribution function of a combination of random variables, with an application to operating theatres. Journal of Statistical Computation and Simulation, 23(3):157–181, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shore:1995:ITP

[Sho95] Haim Shore. Identifying a two-parameter distribution by the first two sample moment (partial and complete). Journal of Statistical Computation and Simulation, 52(1):17–32, 1995. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shu:2008:AEW

[Shu08] Lianjie Shu. An adaptive exponentially weighted moving average control chart for monitoring process variances. Journal of Statistical Computation and Simulation, 78(4):367–384, 2008. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sikand:1993:PST

[SHW93] Yash Sikand, John E. Hewett, and F. T. Wright. On the performance of some tests for the equality of two discrete distributions with an ordered alternative. Journal of Statistical Computation and Simulation, 46(3–4):201–215, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Suich:1972:ESS

[SI72] Ronald Suich and Boris Iglewicz. An empirical study of the sample size distribution of a truncated sequential t-test. Journal of Statistical Computation and Simulation, 1(3):247–254, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Siemiatycki:1978:MST

[Sie78] Jack Siemiatycki. Mantel’s space-time clustering statistic: computing higher moments and a comparison of various data transforms. Journal of Statistical Computation and Simulation, 7(1): 13–31, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Simonoff:1987:ODR

[Sim87] Jeffrey S. Simonoff. Outlier detection and robust estimation of scale. Journal of Statistical Computation and Simulation, 27(1): 79–92, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sim:1993:GPG

[Sim93] C. H. Sim. Generation of Poisson and gamma random vectors with given marginals and covariance matrix. Journal of Statistical Computation and Simulation, 47(1–2):1–10, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sim:2000:CNG

[Sim00] C.-H. Sim. S-chart for non-Gaussian variables. Journal of Statistical Computation and Simulation, 65(1-4):147–156, 2000. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sinha:1990:CGC

[Sin90a] Kishore Sinha. C352. Geometrical configurations and PBIB designs. Journal of Statistical Computation and Simulation, 36(1): 34–35, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sinha:1990:CSN

[Sin90b] Kishore Sinha. C365. Some new E- and MV-optimal group divisible designs with unequal block sizes. Journal of Statistical Computation and Simulation, 36(1):59–60, 1990. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sinha:1990:CNB

[Sin90c] Kishore Sinha. C367. Nearly balanced block designs. Journal of Statistical Computation and Simulation, 36(1):63–64, 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sutradhar:2006:ALC

[SJ06] Brajendra C. Sutradhar and Vandna Jowaheer. Analyzing longitudinal count data from adaptive clinical trials: a weighted generalized quasi-likelihood approach. Journal of Statistical Computation and Simulation, 76(12):1079–1093, 2006. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Smith:2007:CAU

[SJ07] Christina D. Smith and Dallas E. Johnson. Comparing analyses of unbalanced split-plot experiments. Journal of Statistical Computation and Simulation, 77(2):119–129, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Samiuddin:1993:BDE

[SJEs93] M. Samiuddin, M. C. Jones, and G. M. El-sayyad. On binbased density estimation. Journal of Statistical Computation and Simulation, 47(3–4):241–252, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sun:2006:NLM

[SJF06] Yan Sun, Venkata K. Jandhyala, and Stergios B. Fotopoulos. Non-linear mixture models for cross-sectional financial log returns. Journal of Statistical Computation and Simulation, 76 (11):967–983, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shieh:2007:PSS

[SJR07] Gwowen Shieh, Show-Li Jan, and Ronald H. Randles. Power and sample size determinations for the Wilcoxon signed-rank test. Journal of Statistical Computation and Simulation, 77(8): 717–724, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sennetti:1980:FCU

[SK80] John T. Sennetti and Rajesh K. Kakar. Field conditions for the use of certain empirical Bayes estimators. Journal of Statistical Computation and Simulation, 11(2):119–138, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Sarkar:1987:FVT

[SK87] A. Sarkar and B. Kartikeyan. Forecasting by Volterra type models. Journal of Statistical Computation and Simulation, 28(3): 245–260, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sinha:1990:CRD

[SK90a] Kishore Sinha and Sanpei Kageyama. C354. On resolvability of duplicate designs. Journal of Statistical Computation and Simulation, 36(1):40–42, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sinha:1990:CID

[SK90b] Kishore Sinha and Sanpei Kageyama. C366. Intermediate designs from BIBD(s^{2}, s, 1) to BIBD(s^{2} +s+1, s+1, 1). Journal of Statistical Computation and Simulation, 36(1):61–63, 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Saha:2008:AWT

[SK08] Krishna K. Saha and Maxwell L. King. An alternative Wald type test for two linear restrictions with applications to nonlinear models. Journal of Statistical Computation and Simulation, 78(11):1017–1031, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schuurmann:1975:EPP

[SKC75] F. J. Schuurmann, P. R. Krishnaiah, and A. K. Chattopadhyay. Exact percentage points of the distribution of the race of a multivariate beta matrix. Journal of Statistical Computation and Simulation, 3(4):331–345, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shisong:1986:EMM

[SKJ86] Mao Shisong, Samuel Kotz, and Norman L. Johnson. Estimation by moments in a model of faulty inspection. Journal of Statistical Computation and Simulation, 23(4):265–272, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shah:1987:CNU

[SL87] Arvind K. Shah and John J. Lefante, Jr. C293. A note on using a hypothesis-dependent variance estimate. Journal of Statistical Computation and Simulation, 28(4):347–349, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Subramaniam:1988:CIS

[SL88] Murali Subramaniam and Lawrence Leemis. Confidence interval scatterplots for evaluating confidence interval performance. Journal of Statistical Computation and Simulation, 30(2):129– 143, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shorack:1989:MMW

[SL89a] Galen R. Shorack and Nhu D. Le. Modification of the Mann– Whitney test. Journal of Statistical Computation and Simulation, 31(1):9–17, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sim:1989:SNB

[SL89b] C. H. Sim and P. A. Lee. Simulation of negative binomial processes. Journal of Statistical Computation and Simulation, 34 (1):29–42, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shih:1993:VGN

[SL93] Li-Hsing Shih and Lawrence M. Leemis. Variate generation for a nonhomogeneous Poisson process with time dependent covariates. Journal of Statistical Computation and Simulation, 44(3– 4):165–186, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shi:1998:FID

[SL98a] Daoji Shi and C. D. Lai. Fisher information for Downton’s bivariate exponential distribution. Journal of Statistical Computation and Simulation, 60(2):123–127, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shih:1998:CSM

[SL98b] Neng-Hui Shih and Jsen-Shung Lin. A conditional simulation method for estimating the critical component of networks. Journal of Statistical Computation and Simulation, 60(2):97–106, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sheu:2009:MMA

[SL09] Shey-Huei Sheu and Shin-Li Lu. Monitoring the mean of autocorrelated observations with one generally weighted moving average control chart. Journal of Statistical Computation and Simulation, 79(12):1393–1406, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Steel:2000:CPS

[SLL00] S. J. Steel, N. Louw, and N. J. Le Roux. A comparison of the post selection error rate behaviour of the normal linear and quadratic discriminant rules. Journal of Statistical Computation and Simulation, 65(1-4):157–172, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Smethurst:1991:TFM

[SM91] Philip A. Smethurst and Govind S. Mudholkar. Two facets of the Moran statistic. Journal of Statistical Computation and Simulation, 39(4):215–220, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schwertman:1994:ACI

[SM94] Neil C. Schwertman and Ricardo Martinez. Approximate confidence intervals for the difference in two Poisson parameters. Journal of Statistical Computation and Simulation, 50(3–4):235– 247, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Simpsona:1996:BRR

[SM96]  James R. Simpsona and Douglas C. Montgomery.  A

biased-robust regression technique for the combined outliermulticollinearity problem. Journal of Statistical Computation and Simulation, 56(1):1–22, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Safadi:2003:BAA

[SM03] Thelma Sáfadi and Pedro Morettin. A Bayesian analysis of autoregressive models with random normal coefficients. Journal of Statistical Computation and Simulation, 73(8):563–573, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Smadi:2005:LEP

[Sma05] Abdullah A. Smadi. LS estimation of periodic autoregressive models with non-Gaussian errors: a simulation study. Journal of Statistical Computation and Simulation, 75(3):207–216, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sheather:1997:FSS

[SMH97] Simon J. Sheather, Joseph W. McKean, and Thomas P. Hettmansperger. Finite sample stability properties of the least median of squares estimator. Journal of Statistical Computation and Simulation, 58(4):371–383, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Smith:1978:NPD

[Smi78] V. Kerry Smith. A note on the power of the Durbin–Watson test with 2SLS. Journal of Statistical Computation and Simulation, 7(2):115–122, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Smith:1984:CAV

[Smi84] Eric P. Smith. C188. On the asymptotic variance of Socransky’s proportional similarity index. Journal of Statistical Computation and Simulation, 19(1):90–94, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Smith:1989:BRa

[Smi89a] Eric P. Smith. Book reviews. Journal of Statistical Computation and Simulation, 31(4):251, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Smith:1989:BRb

[Smi89b] Eric P. Smith. Book reviews. Journal of Statistical Computation and Simulation, 32(1–2):59–60, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Smith:1989:NTC

[Smi89c] P. J. Smith. A new technique for calculating Fourier integrals based on the Poisson summation formula. Journal of Statistical Computation and Simulation, 33(3):135–147, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Smith:1989:MLE

[Smi89d] Woollcott Smith. A maximum likelihood estimator for totally positive probability tables: a method for smoothing two-way tables. Journal of Statistical Computation and Simulation, 33 (2):69–73, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Smith:1994:CDF

[Smi94] Peter J. Smith. On the calculation of density functions for certain random geometric series. Journal of Statistical Computation and Simulation, 50(1–2):111–117, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sasabuchi:2003:ETS

[SMO03] Shoichi Sasabuchi, Takashi Miura, and Hitoshi Oda. Estimation and test of several multivariate normal means under an order restriction when the dimension is larger than two. Journal of Statistical Computation and Simulation, 73(9):619–641, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shi:2006:WCH

[SMP^{+}06] B. Shi, K. P. Moloney, Y. Pan, V. K. Leonard, B. Vidakovic, J. A. Jacko, and F. Sainfort. Wavelet classification of high frequency pupillary responses. Journal of Statistical Computation and Simulation, 76(5):431–445, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Stucky:1976:EPT

[SMV76] W. Stucky, E. Merck, and J. Vollmar. Exact probabilities for tied linear rank tests. Journal of Statistical Computation and Simulation, 5(1):73–81, 1976. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Singh:1983:LRT

[SN83] N. Singh and P. Narayan. The likelihood ratio test for the equality of two-parameter exponential distributions based on Type II censored samples. Journal of Statistical Computation and Simulation, 18(4):287–297, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Stepniak:1995:ICM

[SN95] C. Stepniak and M. Niezgoda. Inverting covariance matrices in unbalanced hierarchical models. Journal of Statistical Computation and Simulation, 51(2-4):215–221, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Santos:2005:ENS

[SN05] M. Isabel Reis Santos and Acácio M. O. Porta Nova. Estimation of nonlinear simulation metamodels using control variates. Journal of Statistical Computation and Simulation, 75(12):959– 973, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

See:2007:CIP

[SNB07a] K. See, R. B. Noble, and A. J. Bailer. Computing inclusion probabilities to obtain Horvitz–Thompson estimators for sampling plans excluding neighbouring units. Journal of Statistical Computation and Simulation, 77(11):955–967, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

See:2007:CRE

[SNB07b] Kyoungah See, Robert B. Noble, and A. John Bailer. Comparison of relative efficiencies of sampling plans excluding certain neighboring units: a simulation study. Journal of Statistical Computation and Simulation, 77(7):605–612, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shapla:2009:MAR

[SNC09] Tanweer J. Shapla, Truc T. Nguyen, and John T. Chen. Multilevel attributable risk in cross-sectional studies. Journal of Statistical Computation and Simulation, 79(1):39–54, 2009. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schwertman:1989:MCS

[SND89] Neil C. Schwertman, Andrew Neath, and Darien Davis. A Monte Carlo study of successive difference analysis of growth curve data at random observation times. Journal of Statistical Computation and Simulation, 34(1):11–28, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Stangenhaus:1993:BCI

[SNP93] Gabriela Stangenhaus, Subhash C. Narula, and Ferreira F. Pedro. Bootstrap confidence intervals for the minimum sum of absolute errors regression. Journal of Statistical Computation and Simulation, 48(3–4):127–133, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Snyder:1988:CAK

[Sny88] R. D. Snyder. Computational aspects of Kalman filtering with a diffuse prior distribution. Journal of Statistical Computation and Simulation, 29(1):77–86, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sohn:1992:VSL

[Soh92] So Young Sohn. Variable selection in a linear growth curve model with autoregressive within-individual errors. Journal of Statistical Computation and Simulation, 40(3–4):247–255, 1992. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sohn:1994:CSF

[Soh94] So Young Sohn. A comparative study of four estimators for analyzing the random event rate of the Poisson process. Journal of Statistical Computation and Simulation, 49(1–2):1–10, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sohn:1996:IPD

[Soh96] So Young Sohn. Influence of a prior distribution on traffic intensity estimation with covariates. Journal of Statistical Computation and Simulation, 55(3):169–180, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sohn:2000:MAP

[Soh00] So Young Sohn. Meta analysis with potentially correlated marketing study results. Journal of Statistical Computation and Simulation, 65(1-4):95–107, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Solow:1990:MAM

[Sol90] Andrew R. Solow. A method for approximating multivariate normal orthant probabilities. Journal of Statistical Computation and Simulation, 37(3–4):225–229, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Somerville:1984:MRS

[Som84] Paul N. Somerville. A multiple range subset selection procedure. Journal of Statistical Computation and Simulation, 19(3):215– 226, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Somerville:1985:NSS

[Som85] Paul N. Somerville. A new subset selection method for normal populations. Journal of Statistical Computation and Simulation, 22(1):27–50, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Song:1997:GDR

[Son97] Peter Xue-Kun Song. Generating dependent random numbers with given correlations and margins from exponential dispersion models. Journal of Statistical Computation and Simulation, 56 (4):317–335, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Steel:2004:MPV

[SOU04] S. J. Steel, S. Oosthuizen, and D. W. Uys. A measure of post variable selection error in multiple linear regression, and its estimation. Journal of Statistical Computation and Simulation, 74 (2):123–133, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Stein:1986:EEP

[SP86] William E. Stein and Roger C. Pfaffenberger. Estimation of error probabilities in stochastic dominance. Journal of Statistical Computation and Simulation, 25(1–2):137–165, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Singha:1997:SSV

[SP97] N. Singha and M. S. Peiris. A simulation study on vector ARMA processes with nonstationary innovation: a new approach to identification. Journal of Statistical Computation and Simulation, 58(1):37–58, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Silvapulle:2000:ENN

[SP00] Param Sothy Silvapulle and Jan M. Podivinsky. The effect of non-normal disturbances and conditional heteroskedasticity on multiple cointegration tests. Journal of Statistical Computation and Simulation, 65(1-4):173–189, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Soares:2001:ICD

[SP01a] Paulo Soares and Carlos Daniel Paulino. Incomplete categorical data analysis: a Bayesian perspective. Journal of Statistical Computation and Simulation, 69(2):157–170, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Svetliza:2001:DLL

[SP01b] Carolina F. Svetliza and Gilberto A. Paula. On diagnostics in log-linear negative binomial models. Journal of Statistical Computation and Simulation, 71(3):231–244, 2001. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Spezia:2006:BAN

[Spe06] Luigi Spezia. Bayesian analysis of non-homogeneous hidden Markov models. Journal of Statistical Computation and Simulation, 76(8):713–725, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Spitzep:1982:HTM

[Spi82] John J. Spitzep. Hypothesis testing in models using the Box–Cox transformation. Journal of Statistical Computation and Simulation, 15(2-3):161–170, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Spiess:1998:MAE

[Spi98] Martin Spiess. A mixed approach for the estimation of probit models with correlated responses: some finite sample results. Journal of Statistical Computation and Simulation, 61(1–2):39– 59, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Spitzner:2006:UGF

[Spi06] D. J. Spitzner. Use of goodness-of-fit procedures in high dimensional testing. Journal of Statistical Computation and Simulation, 76(5):447–457, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shalabh:2009:CER

[SPK09] Shalabh, Chandra Mani Paudel, and Narinder Kumar. Consistent estimation of regression parameters under replicated ultrastructural model with non-normal errors. Journal of Statistical Computation and Simulation, 79(3):251–274, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Salama:2002:CDS

[SQ02] Ibrahim Salama and Dana Quade. Computing the distribution of Spearman’s footrule in O(n^{4}) time. Journal of Statistical Computation and Simulation, 72(11):895–898, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sheng:2007:VCM

[SQ07] J. Sheng and P. Qiu. p-value calculation for multi-stage additive tests. Journal of Statistical Computation and Simulation, 77(12):1057–1064, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shorter:1997:RAG

[SR97] Jeffrey A. Shorter and Herschel A. Rabitz. Risk analysis by the guided Monte Carlo technique. Journal of Statistical Computation and Simulation, 57(1-4):321–336, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Stoumbos:2000:RNN

[SR00] Zachary G. B. Stoumbos and Marion R. Reynolds, Jr. Robustness to non-normality and autocorrelation of individuals control charts. Journal of Statistical Computation and Simulation, 66 (2):145–187, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Scott:1975:ETS

[SS75] C. Scott and A. K. Md. Esanes Saleh. The effect of truncation on the size of the t-test. Journal of Statistical Computation and Simulation, 3(4):345–368, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schechtman:1986:EPR

[SS86] Edna Schechtman and Gideon Schechtman. Estimating the parameters in regression with uniformly distributed errors. Journal of Statistical Computation and Simulation, 26(3–4):269–281, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Swain:1988:CVM

[SS88] James J. Swain and Bruce W. Schmeiser. Control variates for Monte Carlo analysis of nonlinear statistical models, II: raw moments and variances. Journal of Statistical Computation and Simulation, 30(1):39–56, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Singhal:1992:SDA

[SS92] R. A. Singhal and Hardeo Sahai. Sampling distribution of the Anova estimator of between variance component in samples from a non-normal universe. Journal of Statistical Computation and Simulation, 43(1–2):19–30, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schmidt:1995:RMR

[SS95] Karsten Schmidt and Peter Stahlecker. Reducing the maximum risk of regression estimators by polyhedral projection. Journal of Statistical Computation and Simulation, 52(1):1–15, 1995. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sarkar:1997:ADS

[SS97a] A. Sarkar and K. M. S. Sharma. An approach to direct selection of best subset AR model. Journal of Statistical Computation and Simulation, 56(3):273–291, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Scott:1997:E

[SS97b] Marian Scott and Andrea Saltelli. Editorial. Journal of Statistical Computation and Simulation, 57(1-4):1–2, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Solorzano:1999:OSS

[SS99] Eleanne Solorzano and John Spurrier. One-sided simultaneous comparisons with more than one control. Journal of Statistical Computation and Simulation, 63(1):37–46, 1999. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Staniswalis:1993:DBP

[SSM93] Joan G. Staniswalis, Thomas A. Severini, and Panagis G. Moschopoulos. On a data based power transformation for reducing skewness. Journal of Statistical Computation and Simulation, 46(1–2):91–100, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Selukar:1995:CSA

[SSM95] Rajesh S. Selukar, Arvind K. Shah, and Satya N. Mishra. C431. A simple approach to simultaneous comparison of normal means. Journal of Statistical Computation and Simulation, 52(2):174– 177, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

See:2000:RES

[SSSB00] K. See, J. Stufken, S. Y. Song, and A. J. Bailer. Relative efficiencies of sampling plans for selecting a small number of units from a rectangular region. Journal of Statistical Computation and Simulation, 66(4):273–294, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sarkar:1995:IMT

[SSW95] Sanat K. Sarkar, Steve Snapinn, and Wenjin Wang. On improving the min test for the analysis of combination drug trials. Journal of Statistical Computation and Simulation, 51(2-4):197– 213, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See correction [Sar98].

Sahai:1974:CAP

[ST74] Hardeo Sahai and William O. Thompson. Comparisons of approximations to the percentiles of the t, χ^{2}, and F distributions. Journal of Statistical Computation and Simulation, 3(1):81–93, 1974. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shoukri:1987:BSI

[ST87] M. M. Shoukri and D. S. Tracy. Bayesian statistical inference in the Paretian law. Journal of Statistical Computation and Simulation, 27(1):65–77, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Simonoff:1988:JBQ

[ST88] Jeffrey S. Simonoff and Chih-Ling Tsai. Jackknifing and bootstrapping quasi-likelihood estimators. Journal of Statistical Computation and Simulation, 30(3):213–232, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schmid:2001:PPM

[STG^{+}01] Christopher H. Schmid, Norma Terrin, John L. Griffith, Ralph B. D’Agostino, Sr., and Harry P. Selker. Predictive performance of missing data methods for logistic regression, classification trees and neural networks. Journal of Statistical Computation and Simulation, 71(2):115–140, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shalabh:2009:SRE

[STH09] Shalabh, H. Toutenburg, and C. Heumann. Stein-rule estimation under an extended balanced loss function. Journal of Statistical Computation and Simulation, 79(10):1259–1273, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Strelitz:1989:CNF

[Str89] Richard A. Strelitz. Computing the normalization factor in Bingham distributions. Journal of Statistical Computation and Simulation, 32(4):193–200, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schipp:1994:ACP

[STS94] Bernhard Schipp, Götz Trenkler, and Peter Stahlecker. Assessing coverage-probabilities for approximate minimax estimators with respect to interval restrictions. Journal of Statistical Computation and Simulation, 50(1–2):59–74, 1994. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sumita:1983:SIL

[Sum83] Ushio Sumita. On sums of independent logistic and folded logistic variants. Journal of Statistical Computation and Simulation, 17(4):251–274, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Swain:1988:LSE

[SVW88] James J. Swain, Sekhar Venkatraman, and James R. Wilson. Least-squares estimation of distribution functions in Johnson’s translation system. Journal of Statistical Computation and Simulation, 29(4):271–297, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Simon:1975:CAS

[SW75] Richard Simon and George H. Weiss. A class of adaptive sampling schemes for selecting the better of two binomial populations. Journal of Statistical Computation and Simulation, 4(1): 37–47, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Srinivasan:1982:CBL

[SW82] R. Srinivasan and Robert M. Wharton. Confidence bands for the Laplace distribution. Journal of Statistical Computation and Simulation, 14(2):89–99, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Srinivasan:1983:CSS

[SW83] R. Srinivasan and B. Michael Wharton. A comparative study of some methods for constructing confidence bands on a normal distribution function. Journal of Statistical Computation and Simulation, 17(2):91–99, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Srinivasan:1984:SCI

[SW84] R. Srinivasan and B. Michael Wharton. Simultaneous confidence intervals for the logistic distribution. Journal of Statistical Computation and Simulation, 20(3):183–195, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Skovlund:1988:STS

[SW88] Eva Skovlund and Lars Wallöe. A sequential two-sample test developed by stochastic simulation. Journal of Statistical Computation and Simulation, 29(2):87–104, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Soper:1990:MCE

[SW90] Keith A. Soper and Peter H. Westfall. Monte Carlo estimation of significance levels for carcinogenicity tests using univariate and multivariate models. Journal of Statistical Computation and Simulation, 37(3–4):189–209, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schuurmann:1973:PPJ

[SWK73] F. J. Schuurmann, V. B. Waikar, and P. R. Krishnaiah. Percentage points of the joint distribution of the extreme roots of the random matrix S_{1}(S_{1} + S_{2})^{−1}. Journal of Statistical Computation and Simulation, 2(1):17–38, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shayib:1989:MGF

[SY89] Mohammed A. Shayib and David H. Young. Modified goodness of fit tests in gamma regression. Journal of Statistical Computation and Simulation, 33(2):125–133, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Staniswalis:1992:LAS

[SY92] Joan G. Staniswalis and Brian S. Yandell. Locally adaptive smoothing splines. Journal of Statistical Computation and Simulation, 43(1–2):45–53, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sy:2001:ISP

[Sy01a] Bon K. Sy. Information-statistical pattern based approach for data mining. Journal of Statistical Computation and Simulation, 69(2):171–201, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sy:2001:PMS

[Sy01b] Bon K. Sy. Probability model selection using informationtheoretic optimization criterion. Journal of Statistical Computation and Simulation, 69(3):203–224, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Saw:2008:IEE

[SY08] Sutaip L. C. Saw and Jongsay Yong. Interval estimation for the exponential inverse Gaussian distribution. Journal of Statistical Computation and Simulation, 78(4):339–349, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Safi:2002:MSI

[SZ02] S. Safi and A. Zeroual. Ma system identification using higher order cumulants application to modelling solar radiation. Journal of Statistical Computation and Simulation, 72(7):533–548, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsui:1992:NCE

[TA92] Albert K. Tsui and Mukhtar M. Ali. Numerical computation of exact moments of the least squares estimator in a first-order stationary autoregressive model. Journal of Statistical Computation and Simulation, 42(1–2):65–78, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tseng:1994:MAP

[TA94] Sarah Tseng and Benjamin M. Adams. Monitoring autocorrelated processes with an exponentially weighted moving average forecast. Journal of Statistical Computation and Simulation, 50 (3–4):187–195, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tejedor:2008:NCS

[TA08] I. Herranz Tejedor and A. Martín Andrés. A numerical comparison of several unconditional exact tests in problems of equivalence based on the difference of proportions. Journal of Statistical Computation and Simulation, 78(11):969–981, 2008. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tadikamalla:1981:FDO

[Tad81] Pandu R. Tadikamalla. On a family of distributions obtained by the transformation of the gamma distribution. Journal of Statistical Computation and Simulation, 13(3–4):209–214, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tahsoh:1990:EPF

[Tah90] Joseph T. Tahsoh. Estimation of a probit function using a linear spline. Journal of Statistical Computation and Simulation, 35 (1–2):1–11, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tibaldi:2003:SHL

[TAM^{+}03] Fabián Tibaldi, José Cortiñas Abrahantes, Geert Molenberghs, Didier Renard, Tomasz Burzykowski, Marc Buyse, Max Parmar, Theo Stijnen, and Russ Wolfinger. Simplified hierarchical linear models for the evaluation of surrogate endpoints. Journal of Statistical Computation and Simulation, 73(9):643–658, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tan:1982:APD

[Tan82] W. Y. Tan. On approximating probability distributions of a statistic of Brown–Forsythe from normal and non-normal universes. Journal of Statistical Computation and Simulation, 16 (1):35–55, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tang:2001:SAS

[Tan01] Hui-Chin Tang. A statistical analysis of the screening measure of multiple recursive random number generators of orders one and two. Journal of Statistical Computation and Simulation, 71 (4):345–356, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Taylor:1972:CCC

[Tay72] Malcolm S. Taylor. Computer construction of cyclic balanced incomplete block designs. Journal of Statistical Computation and Simulation, 1(1):65–70, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Taylor:1977:PES

[Tay77] Malcolm S. Taylor. On the power of the extreme spread. Journal of Statistical Computation and Simulation, 5(2):159–166, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Taylor:1990:OSE

[Tay90] Charles C. Taylor. Orthogonal series estimators and crossvalidation. Journal of Statistical Computation and Simulation, 37(3–4):151–158, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Taylor:1994:SME

[Tay94] C. C. Taylor. Simulation methods to estimate smoothing parameters in image reconstruction. Journal of Statistical Computation and Simulation, 49(3–4):161–177, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tamhane:2002:BDL

[TAY02] Ajit Tamhane, Bruce Ankenman, and Ying Yang. The beta distribution as a latent response model for ordinal data (i): Estimation of location and dispersion parameters. Journal of Statistical Computation and Simulation, 72(6):473–494, 2002. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Taylor:1982:QMS

[TB82] Malcolm S. Taylor and Mark P. Becker. Quantiles of multisample Smirnov type statistics. Journal of Statistical Computation and Simulation, 16(1):25–34, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Titterington:1985:CSS

[TB85] D. M. Titterington and A. W. Bowman. A comparative study of smoothing procedures for ordered categorical data. Journal of Statistical Computation and Simulation, 21(3–4):291–312, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tan:1986:BIT

[TB86] W. Y. Tan and N. Balakrishnan. Bayesian insight into Tiku’s robust procedures based on asymmetric censored samples. Journal of Statistical Computation and Simulation, 24(1):17–31, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tan:1988:BAR

[TB88] W. Y. Tan and N. Balakrishnan. Bayesian approach to robust comparison of two means based on asymmetric Type-II censored samples. Journal of Statistical Computation and Simulation, 30 (2):81–102, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ting:1990:CIL

[TBG^{+}90] Naitee Ting, Richard K. Burdick, Franklin A. Graybill, S. Jeyaratnam, and Tai-Fang C. Lu. Confidence intervals on linear combinations of variance components that are unrestricted in sign. Journal of Statistical Computation and Simulation, 35(3– 4):135–143, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Turcott:1995:LDC

[TBT95] Robert G. Turcott, Peter D. R. Barker, and Malvin C. Teich. Long-duration correlation in the sequence of action potentials in an insect visual interneuron. Journal of Statistical Computation and Simulation, 52(3):253–271, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Taylor:1973:CBI

[TC73] Malcolm S. Taylor and Raymond N. Carr. Construction of balanced incomplete block designs with partial cycles of blocks. Journal of Statistical Computation and Simulation, 2(4):357– 363, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tilquin:1975:PRM

[TC75] C. Tilquin and R. Cléroux. Periodic replacement with minimal repair at failure and general cost function. Journal of Statistical Computation and Simulation, 4(1):63–77, 1975. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Thomas:1992:ARE

[TC92] Ronald G. Thomas and Michael Conlon. An algorithm for the rapid evaluation of the power function for Fisher’s exact test. Journal of Statistical Computation and Simulation, 44(1–2):63– 73, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tang:1994:TLI

[TC94] Loon Ching Tang and Doug Shang Chang. Tolerance limits for inverse Gaussian distribution. Journal of Statistical Computation and Simulation, 51(1):21–29, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Trivez:2008:SEC

[TC08] F. Javier Trívez and Beatriz Catalán. Specification error caused by level shifts and temporary changes in ARMA–GARCH models. Journal of Statistical Computation and Simulation, 78(9): 853–868, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Terrell:1987:CCS

[Ter87] George R. Terrell. C290. Chi-squared left-tail probabilities. Journal of Statistical Computation and Simulation, 28(3):264– 266, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Terrell:1990:CHS

[Ter90] George R. Terrell. C360. The half-space problem. Journal of Statistical Computation and Simulation, 36(1):50–53, 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Thomas:1973:RCC

[TG73] Marlin A. Thomas and Gary W. Gemmill. Robustness of confidence coefficients associated with tolerance limits. Journal of Statistical Computation and Simulation, 2(4):349–355, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Talwar:1978:RTC

[TG78] Prem P. Talwar and James E. Gentle. A robust test for the coincidence of regressions. Journal of Statistical Computation and Simulation, 7(2):107–113, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tang:1990:TIS

[TG90] Jen Tang and Rrjun K. Gupta. Testing independence of several groups of variables. Journal of Statistical Computation and Simulation, 37(1–2):27–35, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tutz:2005:ACO

[TH05] Gerhard Tutz and Klaus Hechenbichler. Aggregating classifiers with ordinal response structure. Journal of Statistical Computation and Simulation, 75(5):391–408, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsay:2009:GAP

[TH09] Wen-Jen Tsay and Wolfgang Karl Härdle. A generalized ARFIMA process with Markov-switching fractional differencing parameter. Journal of Statistical Computation and Simulation, 79(5):731–745, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Thomas:1974:SSP

[Tho74] Marlin A. Thomas. A simple sequential procedure for sampling termination in simulation investigations. Journal of Statistical Computation and Simulation, 3(2):161–164, 1974. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Thomas:1989:CNC

[Tho89] G. E. Thomas. C306. A note on correcting for ties with Spearman’s p. Journal of Statistical Computation and Simulation, 31(1):37–40, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Thompson:1992:IRT

[Tho92] G. L. Thompson. On intercomponent rank tests for the bivariate two-sample location model. Journal of Statistical Computation and Simulation, 43(1–2):103–116, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Thoresen:2006:CME

[Tho06] Magne Thoresen. Correction for measurement error in multiple logistic regression: A simulation study. Journal of Statistical Computation and Simulation, 76(6):475–487, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Taylor:1996:SCR

[TJ96] Paul C. Taylor and M. C. Jones. Splitting criteria for regression trees. Journal of Statistical Computation and Simulation, 55 (4):267–285, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tu:2000:BRA

[TJKB00] Xin M. Tu, Gang Jia, Jeanne Kowalski, and Silviu A. Bacanu. Bayesian regression analysis of data with censored initiating and terminating times: applications to aids. Journal of Statistical Computation and Simulation, 65(1-4):1–21, 2000. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Turnbull:1978:ASP

[TKS78] Bruce W. Turnbull, Haya Kaspi, and Richard L. Smith. Adaptive sequential procedures for selecting the best of several normal populations. Journal of Statistical Computation and Simulation, 7(2):195–200, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Thompson:1989:CNA

[TKT89] A. M. Thompson, J. W. Kay, and D. M. Titterington. A cautionary note about cross-validatory choice. Journal of Statistical Computation and Simulation, 33(4):199–216, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wan:1995:URM

[tKWY95] Alan t. K. Wan and Minxian Yang. On the use of the f ratio in a mis-specified model with an interval restriction. Journal of Statistical Computation and Simulation, 52(2):151–161, 1995. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tummala:1996:SDR

[TL96] V. M. Rao Tummala and Hong Ling. Sampling distribution of the random consistency index of the analytic hierarchy process(AHP). Journal of Statistical Computation and Simulation, 55(1–2):121–131, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tang:2007:NSI

[TL07a] Ya-Yong Tang and Yeh Lam. Numerical solution to an integral equation in geometric process. Journal of Statistical Computation and Simulation, 77(7):549–560, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Thoresen:2007:SSS

[TL07b] Magne Thoresen and Petter Laake. A simulation study of statistical tests in logistic measurement error models. Journal of Statistical Computation and Simulation, 77(8):683–694, 2007. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Takeuchi:1973:AEE

[TMW73] Kei Takeuchi, Morris Meisner, and Joseph Wanderling. Asymptotic efficiencies of estimators of location a computational study. Journal of Statistical Computation and Simulation, 2(4):375– 390, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tang:2006:ECA

[TNG^{+}06] Man-Lai Tang, Hon Keung Tony Ng, Jianhua Guo, Wai Chan, and Ben Ping-Shing Chan. Exact Cochran–Armitage trend tests: comparisons under different models. Journal of Statistical Computation and Simulation, 76(10):847–859, 2006. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Thas:2004:EAD

[TO04] Olivier Thas and Jean-Pierre Ottoy. An extension of the Anderson–Darling k-sample test to arbitrary sample space partition sizes. Journal of Statistical Computation and Simulation, 74(9):651–665, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Toothaker:1972:EJV

[Too72] Larry E. Toothaker. The effect of the joint violation of two assumptions on the size of the F -test under permutation for the randomized block design. Journal of Statistical Computation and Simulation, 1(1):55–64, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Taam:1998:VUI

[TP98] Winson Taam and Guohua Pan. The value of unranked items in the analysis of incomplete ranking survey. Journal of Statistical Computation and Simulation, 61(1–2):1–14, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Taylor:2006:SRC

[TP06] R. L. Taylor and W. J. Padgett. The summer research conferences and the SRCOS: a historical perspective. Journal of Statistical Computation and Simulation, 76(5):373–383, 2006. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Turkkan:1993:CHP

[TPG93] Noyan Turkkan and T. Pham-Gia. Computation of the highest posterior density interval in Bayesian analysis. Journal of Statistical Computation and Simulation, 44(3–4):243–250, 1993. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tadikamalla:1975:AMG

[TR75] Pandu R. Tadikamalla and John S. Ramberg. An approximate method for generating gamma and other variates. Journal of Statistical Computation and Simulation, 3(3):275–282, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Trenkler:1994:CMP

[Tre94] Götz Trenkler. C413. On the Moore–Penrose inverse of a completely symmetric matrix. Journal of Statistical Computation and Simulation, 49(3–4):230–231, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Trenkler:1995:CSS

[Tre95] Götz Trenkler. C430. On the singularity of the sample covariance matrix. Journal of Statistical Computation and Simulation, 52 (2):172–173, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tai-Shing:2000:HP

[TS00] Lau Tai-Shing. On the heterogeneity of proportions. Journal of Statistical Computation and Simulation, 67(2):173–188, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tutz:2004:SMM

[TS04] G. Tutz and T. Scholz. Semiparametric modelling of multicategorical data. Journal of Statistical Computation and Simulation, 74(3):183–200, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tagore:2009:CIL

[TS09] Vickneswary Tagore and Brajendra C. Sutradhar. Conditional inference in linear versus nonlinear models for binary time series.

Journal of Statistical Computation and Simulation, 79(7):881– 897, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsao:2006:D

[Tsa06] Min Tsao. E distribution. Journal of Statistical Computation and Simulation, 76(12):1065–1078, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tse:1984:ECS

[Tse84] Y. K. Tse. An empirical comparison of small sample distributions of estimators of the first order autoregression. Journal of Statistical Computation and Simulation, 19(3):227–236, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsionas:2002:LBC

[Tsi02] Efthymios Tsionas. Likelihood-based comparison of stable Paretian and competing models: Evidence from daily exchange rates. Journal of Statistical Computation and Simulation, 72(4):341– 353, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsionas:2003:BQI

[Tsi03] Efthymios G. Tsionas. Bayesian quantile inference. Journal of Statistical Computation and Simulation, 73(9):659–674, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsui:1993:ECF

[Tsu93] Albert K. Tsui. Empirical comparison of finite sample distribution of the least squares estimator in a non-stationary autoregressive model. Journal of Statistical Computation and Simulation, 48(1–2):1–9, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Thomas:1982:CBP

[TT82] Marlin A. Thomas and Audrey E. Taub. Calculating binomial probabilities when the trial probabilities are unequal. Journal of Statistical Computation and Simulation, 14(2):125–131, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comment [Bea85].

Tan:1986:SRP

[TT86] W. Y. Tan and M. A. Tabatabai. Some robust procedures for comparing several straight lines under heteroscedasticity and non-normality. Journal of Statistical Computation and Simulation, 23(4):251–264, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tan:1988:MWR

[TT88] W. Y. Tan and M. A. Tabatabai. A modified Winsorized regression procedure for linear models. Journal of Statistical Computation and Simulation, 30(4):299–313, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tan:2007:EMA

[TTF07] M. Tan, G.-L. Tian, and H.-B. Fang. An efficient MCEM algorithm for fitting generalized linear mixed models for correlated binary data. Journal of Statistical Computation and Simulation, 77(11):929–943, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tian:2009:CSC

[TTNC09] M. Tian, M. L. Tang, H. K. T. Ng, and P. S. Chan. A comparative study of confidence intervals for negative binomial proportion. Journal of Statistical Computation and Simulation, 79 (3):241–249, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tuenter:2001:ADP

[Tue01] Hans J. H. Tuenter. An algorithm to determine the parameters of S_{U}-curves in the Johnson system of probability distributions by moment matching. Journal of Statistical Computation and Simulation, 70(4):325–347, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tutz:1990:SCR

[Tut90] Gerhard Tutz. Smoothed categorical regression based on direct kernel estimates. Journal of Statistical Computation and Simulation, 36(2-3):139–156, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Terza:1991:CBN

[TW91] Joseph V. Terza and Ulrich Welland. A comparison of bivariate normal algorithms. Journal of Statistical Computation and Simulation, 39(1–2):115–127, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Terpstra:2008:CIP

[TW08] Jeff T. Terpstra and Ping Wang. Confidence intervals for a population proportion based on a ranked set sample. Journal of Statistical Computation and Simulation, 78(4):351–366, 2008. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsai:2007:AEP

[TWLC07] Tzong-Ru Tsai, Shuo-Jye Wu, Jyh-Jiuan Lin, and Yi-Ju Chen. Alternative estimation procedure in SPC when the process data are correlated. Journal of Statistical Computation and Simulation, 77(7):575–583, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Taam:1986:CSS

[TY86] Winson Taam and Brian S. Yandell. C263. Small-sample properties of spatial autocorrelation. Journal of Statistical Computation and Simulation, 26(1–2):127–129, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tong:1990:TTT

[TY90] Howell Tong and Iris Yeung. On testes for threshold-type nonlinearity in irregularly spaced time series. Journal of Statistical Computation and Simulation, 34(4):177–194, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tse:2002:SAE

[TYY02] Siu-Keung Tse, Hak-Keung Yuen, and Chunyan Yang. Statistical analysis of exponential lifetimes under an integrated Type-II interval censoring scheme. Journal of Statistical Computation and Simulation, 72(6):461–471, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tse:1997:TCH

[TZ97] Y. K. Tse and X. L. Zuo. Testing for conditional heteroscedasticity: some Monte Carlo results. Journal of Statistical Computation and Simulation, 58(3):237–253, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tse:2004:MCI

[TZ04] Y. K. Tse and Xibin Zhang. A Monte Carlo investigation of some tests for stochastic dominance. Journal of Statistical Computation and Simulation, 74(5):361–378, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tzavelas:2009:MLP

[Tza09] George Tzavelas. Maximum likelihood parameter estimation in the three-parameter gamma distribution with the use of Mathematica. Journal of Statistical Computation and Simulation, 79(12):1457–1466, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Upadhyay:2001:FPA

[UP01] S. K. Upadhyay and M. Peshwani. Full posterior analysis of three parameter lognormal distribution using Gibbs sampler. Journal of Statistical Computation and Simulation, 71(3):215–230, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vanbelle:2008:BMC

[VA08] S. Vanbelle and A. Albert. A bootstrap method for comparing correlated kappa coefficients. Journal of Statistical Computation and Simulation, 78(11):1009–1015, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Valle:2007:THS

[Val07] Alessandra Dalla Valle. A test for the hypothesis of skewnormality in a population. Journal of Statistical Computation and Simulation, 77(1):63–77, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vanhonacker:1984:FSE

[Van84] Wilfried R. Vanhonacker. Finite sample efficiency in a general linear model with serially correlated disturbances: A comparison disturbances: A comparison of estimators given both temporally aggregated and disaggregated data. Journal of Statistical Computation and Simulation, 19(3):195–203, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Vanhonacker:1987:EDD

[Van87] Wilfried R. Vanhonacker. Estimating the duration of dynamic effects with temporally aggregated observations. Journal of Statistical Computation and Simulation, 27(3):185–209, 1987. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

VanDeVen:1993:ESP

[Van93] R. Van De Ven. Estimating the shape parameter for the negative binomial distribution. Journal of Statistical Computation and Simulation, 46(1–2):111–123, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

VanDerGeest:2005:BDD

[Van05] P. A. G. Van Der Geest. The binomial distribution with dependent Bernoulli trials. Journal of Statistical Computation and Simulation, 75(2):141–154, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vardeman:1981:SPB

[Var81] Stepen Vardeman. On the small N performance of bootstrap and Bayes extended and unextended set compound rules for classification between N(−1, 1) and N(1, 1). Journal of Statistical Computation and Simulation, 13(3–4):255–271, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vauchan:1994:EVE

[Vau94] David C. Vauchan. The exact values of the expected values, variances and covariances of the order statistics from the Cauchy distribution. Journal of Statistical Computation and Simulation, 49(1–2):21–32, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vuchkov:1981:EFL

[VB81] I. N. Vuchkov and L. N. Boyadjieva. Errors in the factor levels and estimation of regression model parameters. Journal of Statistical Computation and Simulation, 13(1):1–12, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vuchkov:1983:RED

[VB83] I. N. Vuchkov and L. N. Boyadjieva. The robustness of experimental designs against errors in the factor levels. Journal of Statistical Computation and Simulation, 17(1):31–41, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

vanBuuren:2006:FCS

[vBBGOR06] S. van Buuren, J. P. L. Brand, C. G. M. Groothuis-Oudshoorn, and D. B. Rubin. Fully conditional specification in multivariate imputation. Journal of Statistical Computation and Simulation, 76(12):1049–1064, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Venezuela:2007:DTG

[VBS07] Maria Kelly Venezuela, Denise Aparecida Botter, and Mônica Carneiro Sandoval. Diagnostic techniques in generalized estimating equations. Journal of Statistical Computation and Simulation, 77 (10):879–888, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vassar:1978:MAM

[VC78] Thomas B. Vassar II and Walter H. Carter, Jr. Minimizing average mean squared error: a comparison of preliminary test estimation to other procedures. Journal of Statistical Computation and Simulation, 7(1):55–69, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vasconcellos:2008:NUE

[VD08] Klaus L. P. Vasconcellos and Gilson B. Dourado. Nearly unbiased estimation in a biparametric exponential family. Journal of Statistical Computation and Simulation, 78(5):387–404, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

vanderTouw:1997:LTN

[vdTGL97] J. W. van der Touw, R. F. Galbraith, and G. M. Laslett. A logistic truncated normal mixture model for overdispersed binomial data. Journal of Statistical Computation and Simulation, 59(4):349–373, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Venter:1989:EPC

[Ven89] J. H. Venter. Estimating the probability of correct selection. Journal of Statistical Computation and Simulation, 32(3):165– 176, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vijapurkar:2001:AFF

[VG01] Ujjwala P. Vijapurkar and Carol A. Gotway. Assessment of forecasts and forecast uncertainty using feneralized linear regression models for time series count data. Journal of Statistical Computation and Simulation, 68(4):321–349, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vasconcellos:2005:CES

[VG05] Klaus L. P. Vasconcellos and Sydney Gomes Da Silva. Corrected estimates for Student t regression models with unknown degrees of freedom. Journal of Statistical Computation and Simulation, 75(6):409–423, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vallejos:2006:BAC

[VGD06] Ronny O. Vallejos and Gonzalo García-Donato. Bayesian analysis of contaminated quarter plane moving average models. Journal of Statistical Computation and Simulation, 76(2):131–147, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Viroli:2007:FIF

[Vir07] Cinzia Viroli. Fitting the independent factor analysis model using the MCMC algorithm. Journal of Statistical Computation and Simulation, 77(9):725–737, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vidmar:1996:MCS

[VM96] Thomas J. Vidmar and Joseph W. Mckean. A Monte Carlo study of robust and least squares response surface methods. Journal of Statistical Computation and Simulation, 54(1-3):1– 18, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

VanDorp:2000:SPB

[VM00] René J. Van Dorp and Thomas A. Mazzuchi. Solving for the parameters of a beta a distribution under two quantile constraints. Journal of Statistical Computation and Simulation, 67(2):189– 201, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Velleman:1989:GID

[VP89] Paul F. Velleman and Pratt Paul. A graphical interface for data analysis. Journal of Statistical Computation and Simulation, 32 (4):223–228, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

VanderLaan:2003:NPA

[VPB03] Mark Van der Laan, Katherine Pollard, and Jennifer Bryan. A new partitioning around medoids algorithm. Journal of Statistical Computation and Simulation, 73(8):575–584, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Volt:1992:TSA

[VR92] Eberhard O. Volt and Puilip F. Rust. Tutorial: S-system analysis of continuous univariate probability distributions. Journal of Statistical Computation and Simulation, 42(3–4):187–249, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Venter:1991:EMP

[VS91] J. H. Venter and S. J. Steel. Estimation of the mean of the population selected from K populations. Journal of Statistical Computation and Simulation, 38(1-4):1–14, 1991. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Venter:1993:SSE

[VS93] J. H. Venter and S. J. Steel. Simultaneous selection and estimation for the some zeros family of normal models. Journal of Statistical Computation and Simulation, 45(3–4):129–146, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Venter:1994:PTT

[VS94] J. H. Venter and S. J. Steel. Pre-test type estimators for selection of simple normal models. Journal of Statistical Computation and Simulation, 51(1):31–48, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

VanDerLaan:1978:ECP

[VW78] P. Van Der Laan and J. Weima. Experimental comparison of the powers of the two-sample tests of Wilcoxon and Student under logistic parent distributions. Journal of Statistical Computation and Simulation, 8(2):133–144, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Velu:1985:T

[VW85] Raja P. Velu and Dean W. Wichern. Tutorial. Journal of Statistical Computation and Simulation, 20(4):333–340, 1985. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Winslow:1972:SSS

[WA72] S. S. Winslow and J. C. Arnold. A stochastic simulation study of a rank-like test for dispersion which is distribution-free under location differences. Journal of Statistical Computation and Simulation, 1(4):315–329, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:1990:CUM

[Wan90] C. Ming Wang. Computing uncertainty measures of location estimates for autonomous vehicles. Journal of Statistical Computation and Simulation, 36(2-3):69–89, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:1992:ACI

[Wan92] C. Ming Wang. Approximate confidence intervals on linear combinations of expected mean squares. Journal of Statistical Computation and Simulation, 43(3–4):229–241, 1992. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wan:1999:NAU

[Wan99] Alan T. K. Wan. A note on almost unbiased generalized ridge regression estimator under asymmetric loss. Journal of Statistical Computation and Simulation, 62(4):411–421, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Wang:2008:GFT

[Wan08a] Bingxing Wang. Goodness-of-fit test for the exponential distribution based on progressively Type-II censored sample. Journal of Statistical Computation and Simulation, 78(2):125–132, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2008:CIM

[Wan08b] Hsiuying Wang. Confidence intervals for the mean of a normal distribution with restricted parameter space. Journal of Statistical Computation and Simulation, 78(9):829–841, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2008:LPU

[Wan08c] Lichun Wang. Life prediction under random censorship. Journal of Statistical Computation and Simulation, 78(11):1033–1044, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comments [Lia11].

Wang:2008:CMP

[Wan08d] Lihong Wang. Change-in-mean problem for long memory time series models with applications. Journal of Statistical Computation and Simulation, 78(7):653–668, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wilson:1984:VRQ

[WAP84] R. James Wilson, A. Alan, and B. Pritsker. Variance reduction in queueing simulation using generalized concomitant variables. Journal of Statistical Computation and Simulation, 19(2):129– 153, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Warren:1974:CRF

[War74] W. G. Warren. The comparison of regressions: A formulation suited to conversational mode operation. Journal of Statistical Computation and Simulation, 3(1):71–79, 1974. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wardrop:1984:BR

[War84] Daniel M. Wardrop. Book review. Journal of Statistical Computation and Simulation, 20(1):85–86, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Waterman:1977:CLS

[Wat77] M. S. Waterman. C1. Least squares with nonnegative regression coefficients. Journal of Statistical Computation and Simulation, 6(1):67–70, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Weed:1973:SOS

[WB73] Harrison D. Weed, Jr. and Ralph A. Bradley. Sequential onesample grouped signed rank tests for symmetry: Monte Carlo studies. Journal of Statistical Computation and Simulation, 2 (2):99–137, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:1994:TBE

[WB94] Ping X. Wu and M. Ishao Bhatti. Testing for block effects and misspecification in regression models based on survey data. Journal of Statistical Computation and Simulation, 50(1–2):75– 90, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Willemain:2002:DND

[WB02] Thomas R. Willemain and Mihoko V. Bennett. The distribution of node degree in maximum spanning trees. Journal of Statistical Computation and Simulation, 72(2):101–106, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2009:FSP

[WB09] Zijun Wang and David A. Bessler. Finite sample performance of the model selection approach in co-integration analysis. Journal of Statistical Computation and Simulation, 79(4):349–360, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wyckoff:1980:SCC

[WBE80] James Wyckoff, Lee J. Bain, and Max Engelhardt. Some complete and censored sampling results for the three-parameter Weibull distribution. Journal of Statistical Computation and Simulation, 11(2):139–151, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Webster:1991:SSD

[WC91] Karen M. Webster and Karen M. Campbell. Simulation studies: do data generation techniques play favourites? Journal of Statistical Computation and Simulation, 39(3):163–174, 1991. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2007:SIB

[WCC07] Shuo-Jye Wu, Yi-Ju Chen, and Chun-Tao Chang. Statistical inference based on progressively censored samples with random removals from the Burr Type XII distribution. Journal of Statistical Computation and Simulation, 77(1):19–27, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Whiteside:1994:SLP

[WCEC94] Mary M. Whiteside, Byoung Choi, Mark Eakin, and Henry Crockett. Stability of linear programming solutions using regression coefficients. Journal of Statistical Computation and Simulation, 50(3–4):131–146, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Williford:1979:FLB

[WCF79] W. O. Williford, Michael C. Carter, and Janete E. Field. A further look at the Bayesian approach to calibration. Journal of Statistical Computation and Simulation, 9(1):47–67, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Willemain:1993:MGA

[WD93] R. Thomas Willemain and A. Philip Desautels. A method to generate autocorrelated uniform random numbers. Journal of Statistical Computation and Simulation, 45(1–2):23–31, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Whiteside:1975:CSN

[WDB75] M. Whiteside, B. S. Duran, and T. L. Boullion. A comparison of some nonparametric tests for scale. Journal of Statistical Computation and Simulation, 4(2):121–132, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

West:1986:ARA

[WDR86] P. W. West, A. W. Davis, and D. A. Ratkowsky. Approaches to regression analysis with multiple measurements from individual sampling units. Journal of Statistical Computation and Simulation, 26(3–4):149–175, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wechsler:1994:CBN

[Wec94] Sergio Wechsler. C425. Bayesian note on game theory. Journal of Statistical Computation and Simulation, 49(3–4):251–252, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Whiteside:2000:MCC

[WECC00] Mary Whiteside, Mark Eakin, Byong Choi, and Henry D. Crockett. Modified chance constrained linear programming under uncertainty. Journal of Statistical Computation and Simulation, 67 (3):255–287, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wegman:1972:NPD

[Weg72] Edward J. Wegman. Nonparametric probability density estimation: Density estimation methods. Journal of Statistical Computation and Simulation, 1(3):225–245, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Weintraub:1979:NLL

[Wei79] Sol Weintraub. Non-linear least squares fit of the chi-square and t distributions. Journal of Statistical Computation and Simulation, 10(1):79–84, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Welch:1982:ACT

[Wel82] William J. Welch. Algorithmic complexity: three NP-hard problems in computational statistics. Journal of Statistical Computation and Simulation, 15(1):17–25, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

West:1972:ANN

[Wes72] Eric N. West. Approximations for the non-null distribution of the Pearson χ^{2} statistic. Journal of Statistical Computation and Simulation, 1(4):377–393, 1972. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Westlund:1985:PST

[Wes85] Anders Westlund. On the power of some tests for examining the stability of regression coefficients. Journal of Statistical Computation and Simulation, 21(3–4):275–289, 1985. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wetzel:1996:GDM

[Wet96] Nathan R. Wetzel. Graphical data modeling methods using Ceres plots. Journal of Statistical Computation and Simulation, 54(1-3):37–44, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Whitlark:1992:WHM

[WG92] David Whitlark and Michael Geurts. When and how much do average forecasts improve predictive accuracy. Journal of Statistical Computation and Simulation, 42(1–2):39–45, 1992. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wong:2004:DMA

[WGC04] H. B. Wong, F. F. Gan, and T. C. Chang. Designs of moving average control chart. Journal of Statistical Computation and Simulation, 74(1):47–62, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Whitcher:2000:MDL

[WGP00] Brandon Whitcher, Peter Guttorp, and Donald B. Percival. Multiscale detection and location of multiple variance changes in the presence of long memory. Journal of Statistical Computation and Simulation, 68(1):65–87, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wheeler:1975:ASG

[Whe75] Donald J. Wheeler. An approximation for simulation of gamma distributions. Journal of Statistical Computation and Simulation, 3(3):225–232, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:1980:ESU

[WHF80] George H. K. Wang, Michael Hidiroglou, and Wayne A. Fuller. Estimation of seemingly unrelated regression with lagged dependent variables and autocorrelated errors. Journal of Statistical Computation and Simulation, 10(2):133–146, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

White:1994:CVN

[Whi94] A. P. White. Cross validation of nearest neighbour discriminant analysis — a warning to SAS users. Journal of Statistical Computation and Simulation, 49(3–4):129–140, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Whitaker:1995:NSA

[Whi95] D. Whitaker. A nested simulated annealing algorithm. Journal of Statistical Computation and Simulation, 53(3–4):233–241, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wilson:1975:SDC

[Wil75] Kellogg Wilson. The sampling distribution of conservational, conservative and corrected F -ratios in repeated measures designs with heterogeneity of covariance. Journal of Statistical Computation and Simulation, 3(3):201–215, 1975. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wilson:1979:SEC

[Wil79] G. Tunnicliffe Wilson. Some efficient computational procedures for high order ARMA models. Journal of Statistical Computation and Simulation, 8(3–4):301–309, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wilson:1989:SEA

[Wil89] J. D. Wilson. A smoothed em algorithm for the solution of Wicksell’s corpuscle problem. Journal of Statistical Computation and Simulation, 31(4):195–221, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Williamson:1998:CNR

[Wil98] Patrica Pepple Williamson. C449. A note on the random generation of Dirichlet variates. Journal of Statistical Computation and Simulation, 60(2):168–173, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Williamson:2001:BAS

[Wil01] Patricia Pepple Williamson. A Bayesian alternative to the sign test in the presence of ties. Journal of Statistical Computation and Simulation, 68(2):135–152, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Williamson:2004:CTP

[Wil04] Patricia Pepple Williamson. Comparison of testing procedures utilizing P -values and Bayes factors in some common situations. Journal of Statistical Computation and Simulation, 74(11):833– 850, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Williamson:2007:BET

[Wil07] Patricia Pepple Williamson. Bayesian equivalence testing for binomial random variables. Journal of Statistical Computation and Simulation, 77(9):739–755, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wilcox:2008:SSS

[Wil08] Rand R. Wilcox. Some small-sample properties of some recently proposed multivariate outlier detection techniques. Journal of Statistical Computation and Simulation, 78(8):701–712, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wingo:1975:UIP

[Win75] Dallas R. Wingo. The use of interior penalty functions to overcome lognormal distribution parameter estimation anomalies. Journal of Statistical Computation and Simulation, 4(1):49–61, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Warwick:2005:CRT

[WJ05] J. Warwick and M. C. Jones. Choosing a robustness tuning parameter. Journal of Statistical Computation and Simulation, 75(7):581–588, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

West:1972:CLR

[WK72] Eric N. West and Oscar Kempthorne. A comparison of the χ^{2} and likelihood ratio tests for composite alternatives. Journal of Statistical Computation and Simulation, 1(1):1–33, 1972. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Waller:1975:CBV

[WK75] R. A. Waller and K. E. Kemp. Computations of Bayesian tvalues for multiple comparisons. Journal of Statistical Computation and Simulation, 4(3):169–171, 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wallenius:1990:EMA

[WK90a] K. T. Wallenius and A. S. Korkotsides. Exploratory model analysis using cdf knotting with applications to distinguishability, limiting forms, and moment ratios to the three parameter Weibull family. Journal of Statistical Computation and Simulation, 35(3–4):121–133, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:1990:CAB

[WK90b] Morgan Wang and William J. Kennedy. Comparison of algorithms for bivariate normal probability over a rectangle based on self-validated results from interval analysis. Journal of Statistical Computation and Simulation, 37(1–2):13–25, 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wright:2002:SVC

[WK02] Kevin Wright and William J. Kennedy. Self-validated computations for the probabilities of the central bivariate chi-square distribution and a bivariate F distribution. Journal of Statistical Computation and Simulation, 72(1):63–75, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Watteel:2003:NEC

[WK03] R. N. Watteel and R. J. Kulperger. Nonparametric estimation of the canonical measure for infinitely divisible distributions. Journal of Statistical Computation and Simulation, 73(7):525– 542, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wilcox:2006:DHS

[WK06] Rand R. Wilcox and H. J. Keselman. Detecting heteroscedasticity in a simple regression model via quantile regression slopes. Journal of Statistical Computation and Simulation, 76(8):705– 712, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wozniak:1990:GFT

[WL90] Patricia J. Wozniak and Xiaozhu Li. Goodness-of-fit for the two-parameter Weibull distribution with estimated parameters. Journal of Statistical Computation and Simulation, 34(2-3):133– 143, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Whitaker:2007:EMS

[WL07a] S. Whitaker and T. C. M. Lee. An effective method for selecting the number of components in density mixtures. Journal of Statistical Computation and Simulation, 77(10):907–914, 2007. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2007:PIO

[WL07b] Tsong-Huey Wu and Hai-Lin Lu. Prediction intervals for an ordered observation from the logistic distribution based on censored samples. Journal of Statistical Computation and Simulation, 77(5):389–405, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Weber:2006:MKS

[WLK06] Michael D. Weber, Lawrence M. Leemis, and Rex K. Kincaid. Minimum Kolmogorov–Smirnov test statistic parameter estimates. Journal of Statistical Computation and Simulation, 76 (3):195–206, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Walde:2008:PCB

[WLT08] Janette Walde, Mario Larch, and Gottfried Tappeiner. Performance contest between MLE and GMM for huge spatial autoregressive models. Journal of Statistical Computation and Simulation, 78(2):151–166, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:1990:NTT

[WM90] Dong Qian Wang and J. S. Maritz. Note on testing a three state Markov chain for independence. Journal of Statistical Computation and Simulation, 37(1–2):61–68, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Weston:1991:CPN

[WM91] Susan A. Weston and William Q. Meeker. Coverage probabilities of nonparametric simultaneous confidence bands for a survival function. Journal of Statistical Computation and Simulation, 38 (1-4):83–97, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:1995:RCR

[WM95] Y. Wu and C. A. McGilchrist. Recursive computation for repeated measures models with patterned variance matrices. Journal of Statistical Computation and Simulation, 51(2-4):333–341, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Williams:1999:FN

[WM99] H. P. Williams and A. G. Munford. Formulae for the L_{0}, L_{1} and L_{∞} norms. Journal of Statistical Computation and Simulation, 63(2):121–141, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:1995:FMA

[WmGT95] Yongxiao Wang and Jeremy m. G. Taylor. Flexible methods for analysing longitudinal data using piecewise cubic polynomials. Journal of Statistical Computation and Simulation, 52(2):133– 150, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wright:2007:EPA

[WNB07] S. E. Wright, R. B. Noble, and A. J. Bailer. Equal-precision allocations and other constraints in stratified random sampling. Journal of Statistical Computation and Simulation, 77(12):1081– 1089, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wolfinger:1993:GLM

[WO93] Russ Wolfinger and Michael O’Connell. Generalized linear mixed models a pseudo-likelihood approach. Journal of Statistical Computation and Simulation, 48(3–4):233–243, 1993. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

West:1997:CTE

[WO97] Webster R. West and Todd R. Ogden. Continuous-time estimation of a change-point in a Poisson process. Journal of Statistical Computation and Simulation, 56(4):293–302, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wong:1985:BTP

[Won85] M. Anthony Wong. A bootstrap testing procedure for investigating the number of subpopulations. Journal of Statistical Computation and Simulation, 22(2):99–112, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wong:1995:GAC

[Won95] Weng Kee Wong. A graphical approach for the construction of constrained a and l-optimal designs using efficiency plots. Journal of Statistical Computation and Simulation, 53(3–4):143– 152, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Woodfield:1985:SWW

[Woo85] Terry J. Woodfield. The selection of window widths in kernel nonparametric regression. Journal of Statistical Computation and Simulation, 23(1–2):113–122, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Worton:1989:OSP

[Wor89] Bruce J. Worton. Optimal smoothing parameters for multivariate fixed and adaptive kernel methods. Journal of Statistical Computation and Simulation, 32(1–2):45–57, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wozniak:1994:PGF

[Woz94] Patricia J. Wozniak. Power of goodness of fit tests for the two-parameter Weibull distribution with estimated parameters.

Journal of Statistical Computation and Simulation, 50(3–4):153– 161, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Weinberg:1981:SIC

[WP81] Roger Weinberg and Yogesh C. Patel. Simulated intraclass correlation coefficients and their z transforms. Journal of Statistical Computation and Simulation, 13(1):13–26, 1981. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Williams:1994:SPC

[WR94] Thomas R. Williams and George C. Runger. Statistical process control using level crossings. Journal of Statistical Computation and Simulation, 51(1):7–20, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Willemain:1998:SPC

[WR98] Thomas R. Willemain and George C. Runger. Statistical process control using run sums. Journal of Statistical Computation and Simulation, 61(4):361–378, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wright:1995:PCI

[Wri95] Peter A. Wright. A process capability index sensitive to skewness. Journal of Statistical Computation and Simulation, 52(3): 195–203, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Westlund:1982:REC

[WS82] Anders H. Westlund and Hans Stenlund. On robust estimation in certainty equivalence control. Journal of Statistical Computation and Simulation, 14(3–4):241–269, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wypij:1990:IEM

[WS90] David Wypij and Thomas J. Santner. Interval estimation of the marginal probability of success for the beta-binomial distribution. Journal of Statistical Computation and Simulation, 35 (3–4):169–185, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:1992:IPB

[WS92] Berlin Wu and Neng-Hui Shih. On the identification problem for bilinear time series models. Journal of Statistical Computation and Simulation, 43(3–4):129–161, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wong:2000:GFT

[WS00] W. K. Wong and C. H. Sim. Goodness-of-fit test based on empirical characterisitc function. Journal of Statistical Computation and Simulation, 65(1-4):243–269, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wludyka:2004:RSA

[WS04] Peter Wludyka and Ping Sa. A robust I-sample analysis of means type randomization test for variances for unbalanced designs. Journal of Statistical Computation and Simulation, 74 (10):701–726, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2009:CMF

[WSPC09] Chien-Wei Wu, Ming-Hung Shu, W. L. Pearn, and Feng-Tsung Cheng. A comparison of methods forcing loss-based capability index. Journal of Statistical Computation and Simulation, 79 (9):1129–1141, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wojciechowski:2006:MTT

[WT06] William C. Wojciechowski and James R. Thompson. Market truths: theory versus empirical simulations. Journal of Statistical Computation and Simulation, 76(5):385–395, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2001:EBM

[Wu01] Y. Wu. An m-estimation-based model selection criterion with a data-oriented penalty. Journal of Statistical Computation and Simulation, 70(1):71–87, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2002:NTM

[Wu02] Dongfeng Wu. Norm thresholding method in wavelet regression. Journal of Statistical Computation and Simulation, 72(3):233– 245, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2003:ETP

[Wu03] Shuo-Jye Wu. Estimation for the two-parameter Pareto distribution under progressive censoring with uniform removals. Journal of Statistical Computation and Simulation, 73(2):125–134, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:1979:DDS

[WV79] Min Chiang Wang and John Van Ryzin. Discrete density smoothing applied to the empirical Bayes estimation of a Poisson mean. Journal of Statistical Computation and Simulation, 8 (3–4):207–226, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:1995:BCI

[WW95] Yuedong Wang and Grace Wahba. Bootstrap confidence intervals for smoothing splines and their comparison to Bayesian confidence intervals. Journal of Statistical Computation and Simulation, 51(2-4):263–279, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wimmer:2003:BGV

[WW03] Gejza Wimmer and Viktor Witkovský. Between group variance component interval estimation for the unbalanced heteroscedastic one-way random effects model. Journal of Statistical Computation and Simulation, 73(5):333–345, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2006:TEC

[WW06] Lihong Wang and Jinde Wang. Testing and estimating for change in long memory parameter. Journal of Statistical Computation and Simulation, 76(4):317–329, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wimmer:2007:ULC

[WW07] Gejza Wimmer and Viktor Witkovský. Univariate linear calibration via replicated errors-in-variables model. Journal of Statistical Computation and Simulation, 77(3):213–227, 2007. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2009:EHT

[WWL09] Shu-Fei Wu, Chin-Chuan Wu, and Huan-Min Lin. The exact hypothesis test for the shape parameter of a new two-parameter distribution with the bathtub shape or increasing failure rate function under progressive censoring with random removals. Journal of Statistical Computation and Simulation, 79(8):1015– 1042, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wong:2004:LBI

[WWS04] Augustine C. M. Wong, Jianrong Wu, and Xiang Sun. Likelihood based inference for the ratio of gamma means. Journal of Statistical Computation and Simulation, 74(3):171–181, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:1995:LSM

[WZ95] Cheng-Sheng Peter Wu and Yvonne Zubovic. A large-scale Monte Carlo study of the Buckley–James estimator with censored data. Journal of Statistical Computation and Simulation, 51(2-4):97–119, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2009:RAW

[XDZ09] Xingzhong Xu, Xiaobo Ding, and Shuran Zhao. The reduction of the average width of confidence bands for an unknown continuous distribution function. Journal of Statistical Computation and Simulation, 79(4):335–347, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xiong:1999:SSM

[Xio99] C. Xiong. Step stress model with threshold parameter. Journal of Statistical Computation and Simulation, 63(4):349–360, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xiong:2009:SCI

[XM09] Shifeng Xiong and Weiyan Mu. Simultaneous confidence intervals for one-way layout based on generalized pivotal quantities. Journal of Statistical Computation and Simulation, 79(10): 1235–1244, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xie:2003:EAS

[XPC03] Min Xie, Wolfgang Preuss, and Li-Rong Cui. Error analysis of some integration procedures for renewal equation and convolution integrals. Journal of Statistical Computation and Simulation, 73(1):59–70, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xie:2007:DAC

[XW07] Feng-Chang Xie and Bo-Cheng Wei. Diagnostics analysis in censored generalized Poisson regression model. Journal of Statistical Computation and Simulation, 77(8):695–708, 2007. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xie:2009:DGP

[XW09] Feng-Chang Xie and Bo-Cheng Wei. Diagnostics for generalized Poisson regression models with errors in variables. Journal of Statistical Computation and Simulation, 79(7):909–922, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xie:2009:FUL

[XYR09] Yingfu Xie, Jun Yu, and Bo Ranneby. Forecasting using locally stationary wavelet processes. Journal of Statistical Computation and Simulation, 79(9):1067–1082, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xiang:2008:IDS

[XYT08] Liming Xiang, Kelvin K. W. Yau, and S. K. Tse. Influence diagnostics for stratified ordinal contingency tables. Journal of Statistical Computation and Simulation, 78(5):405–415, 2008. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Young:1987:BCP

[YAAB87] David H. Young, Akram M. Al-Abood, and Saad T. Bakir. Bias corrected Pearson estimators for the shape parameter in gamma regression. Journal of Statistical Computation and Simulation, 27(3):233–241, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:1998:RUC

[Yan98] Zhenlin Yang. On robustness of usual confidence region under transformation misspecification. Journal of Statistical Computation and Simulation, 61(1–2):175–190, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yanagisawa:1999:GDH

[Yan99] Yukio Yanagisawa. Generalization for the distribution of Hartley’s and Nair’s Fmax statistics. Journal of Statistical Computation and Simulation, 62(4):305–318, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Young:1985:PTR

[YH85] David Young and Mohamed Abdel Hameed. A power transformation regression model for censored failure time data. Journal of Statistical Computation and Simulation, 22(3–4):203–215, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Younan:2001:SCP

[YI01] Wafik Youssef Younan and Maged George Iskander. A suggested computational procedure to Peizer and Pratt’s approximation to the incomplete beta and gamma ratios. Journal of Statistical Computation and Simulation, 68(2):127–134, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yeasmin:2004:TSI

[YK04] Mahbuba Yeasmin and Maxwell L. King. Test sizes and the importance of finding global maxima of the likelihood function. Journal of Statistical Computation and Simulation, 74(7):495– 500, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yamadu:1986:NPC

[YKB86] Sumasu Yamadu, Kuniharu Kishida, and Keisuke Bekki. Notes on poles and convergence rate of autoregressive model. Journal of Statistical Computation and Simulation, 25(3–4):273–293, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:1985:CAC

[YL85] Mark C. K. Yang and Pan-Yu Lai. A computational algorithm for the coverage probability of a first order autogressive process. Journal of Statistical Computation and Simulation, 21(1):59– 74, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Loh:1987:SML

[yL87] Wei yin Loh. Some modifications of Levene’s test of variance homogeneity. Journal of Statistical Computation and Simulation, 28(3):213–226, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2001:FRE

[YL01] Zhenlin Yang and Ronnie T. C. Lee. On the failure rate estimation of the inverse Gaussian distribution. Journal of Statistical Computation and Simulation, 71(3):201–213, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:1990:CPO

[YM90] Mark C. K. Yang and Charles T. C. Mo. Coverage probability to the original data by prediction intervals in simple linear regression. Journal of Statistical Computation and Simulation, 36 (2-3):91–104, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yau:1996:SSG

[YM96] K. K. W. Yau and C. A. Mcgilchrist. Simulation study of the GLMM method applied to the analysis of clustered survival data. Journal of Statistical Computation and Simulation, 55 (3):189–200, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yau:1999:SSB

[YM99] K. K. W. Yau and P. S. F. Ma. A simulation study for the binomial-logit model with correlated random effects. Journal of Statistical Computation and Simulation, 63(2):169–186, 1999. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Young:1993:BDC

[You93] K. D. S. Young. Bayesian diagnostics for checking assumptions of normality. Journal of Statistical Computation and Simulation, 47(3–4):167–180, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Youssef:2008:NEU

[You08] Ahmed H. Youssef. A new estimator for the unit root. Journal of Statistical Computation and Simulation, 78(6):515–526, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2006:MEE

[YR06] S. F. Yang and M. A. Rahim. Multivariate extension to the economical design of control chart under Weibull shock model. Journal of Statistical Computation and Simulation, 76(12):1035– 1047, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Young:1990:RUM

[YST90] Dean M. Young, John W. Seaman, Jr., and Danny W. Turner. A ratio-of-uniforms method for generating exponential power variates. Journal of Statistical Computation and Simulation, 35 (1–2):13–18, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yuen:1996:PEW

[YT96] Hak-Keung Yuen and Siu-Keung Tse. Parameters estimation for Weibull distributed lifetimes under progressive censoring with random removals. Journal of Statistical Computation and Simulation, 55(1–2):57–71, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yu:2002:HFA

[YW02] Qiqing Yu and George Wong. How to find all Buckley–James estimates instead of just one? Journal of Statistical Computation and Simulation, 72(6):451–460, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yu:2009:ABJ

[YW09] Qiqing Yu and George Y. C. Wong. An algorithm for the Buckley–James estimator with interval-censored data. Journal of Statistical Computation and Simulation, 79(11):1341–1353, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2003:EEW

[YX03] Zhenlin Yang and Min Xie. Efficient estimation of the Weibull shape parameter based on a modified profile likelihood. Journal of Statistical Computation and Simulation, 73(2):115–123, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2007:UCI

[YXW07] Zhenlin Yang, Min Xie, and Augustine C. M. Wong. A unified confidence interval for reliability-related quantities of twoparameter Weibull distribution. Journal of Statistical Computation and Simulation, 77(5):365–378, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yazici:2007:CVT

[YY07] Berna Yazici and Senay Yolacan. A comparison of various tests of normality. Journal of Statistical Computation and Simulation, 77(2):175–183, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zoubeidi:2009:ENP

[ZA09] Taoufik Zoubeidi and Mustapha Abane. Efficient nonparametric sequential procedures for comparing the rates of change of two treatments. Journal of Statistical Computation and Simulation, 79(12):1441–1456, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zacks:1980:CSI

[Zac80] S. Zacks. C63. On some inverse moments of negative-binomial distributions and their application in estimation. Journal of Statistical Computation and Simulation, 10(2):163–165, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zanakis:1979:SSS

[Zan79] Stelios H. Zanakis. A simulation study of some simple estimators for the three-parameter Weibull distribution. Journal of Statistical Computation and Simulation, 9(2):101–116, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zantek:2008:MCM

[Zan08] Paul F. Zantek. A Markov-chain method for computing the runlength distribution of the self-starting cumulative sum scheme. Journal of Statistical Computation and Simulation, 78(5):463– 473, 2008. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2003:BMN

[ZC03] Chunming Zhang and Bin Cheng. Binning methodology for nonparametric goodness-of-fit test. Journal of Statistical Computation and Simulation, 73(1):71–82, 2003. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2000:SCC

[Zha00] Nien Fan Zhang. Statistical control charts for monitoring the mean of a stationary process. Journal of Statistical Computation and Simulation, 66(3):249–258, 2000. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2002:EAS

[Zha02] Biao Zhang. An em algorithm for a semiparametric finite mixture model. Journal of Statistical Computation and Simulation, 72(10):791–802, 2002. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2006:STU

[Zha06] Biao Zhang. A score test under a semiparametric finite mixture model. Journal of Statistical Computation and Simulation, 76 (8):691–703, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zheng:1988:ESR

[Zhe88] Youlu Zheng. Evaluation and simulation of random sequential packing density for a small square. Journal of Statistical Computation and Simulation, 29(2):105–115, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zimmerman:1989:CES

[Zim89] Dale L. Zimmerman. Computationally exploitable structure of covariance matrices and generalized covariance matrices in spatial models. Journal of Statistical Computation and Simulation, 32(1–2):1–15, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zimmerman:2004:ISS

[Zim04] Donald W. Zimmerman. Inflated statistical significance of Student’s t test associated with small intersubject correlation. Journal of Statistical Computation and Simulation, 74(9):691–696, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zanakis:1986:RML

[ZK86] Stelios H. Zanakis and Jerzy Kyparisis. A review of maximum likelihood estimation methods for the three-parameter Weibull distribution. Journal of Statistical Computation and Simulation, 25(1–2):53–73, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhao:1996:SSE

[ZL96] Yuejen Zhao and Andy H. Lee. A simulation study of estimators for generalized linear measurement error models. Journal of Statistical Computation and Simulation, 54(1-3):55–74, 1996. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2007:AIL

[ZL07] Y. Zhang and W. E. Leithead. Approximate implementation of the logarithm of the matrix determinant in Gaussian process regression. Journal of Statistical Computation and Simulation, 77(4):329–348, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhou:2009:MCE

[ZL09] X. C. Zhou and X. S. Liu. The Monte Carlo EM method for estimating multivariate Tobit latent variable models. Journal of Statistical Computation and Simulation, 79(9):1095–1107, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zimmerman:1998:CAL

[ZNAEB98] Dale L. Zimmerman, Vicente Nuúñez-Antón, and Hammou El-Barmi. Computational aspects of likelihood-based estimation of first-order antedependence models. Journal of Statistical Computation and Simulation, 60(1):67–84, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zayed:1997:RRC

[ZQ97] Hany Zayed and Dana Quade. On the resistance of rank correlation. Journal of Statistical Computation and Simulation, 58 (1):59–81, 1997. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zaid:1993:CNG

[ZR93] Khaled Zaïd and André Rogatko. C391. Note on a general algorithm for constructing Bayesian credible regions. Journal of Statistical Computation and Simulation, 44(3–4):251–255, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2007:EAR

[ZR07] Zhiwei Zhang and Howard E. Rockette. An EM algorithm for regression analysis with incomplete covariate information. Journal of Statistical Computation and Simulation, 77(2):163–173, 2007. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2005:IGM

[ZS05] Hongmei Zhang and Hal Stern. Investigation of a generalized multinomial model for species data. Journal of Statistical Computation and Simulation, 75(5):347–362, 2005. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2001:LIB

[ZT01] Xibin Zhang and Yiu Kuen Tse. Local influence on bandwidth estimation for kernel smoothing. Journal of Statistical Computation and Simulation, 70(4):349–370, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zuehlke:1996:INL

[Zue96] Thomas W. Zuehlke. Identification by non-linearity in censored regression models. Journal of Statistical Computation and Simulation, 54(4):289–304, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zurbenko:1993:WCR

[Zur93] I. G. Zurbenko. On weakly correlated random numbers generator. Journal of Statistical Computation and Simulation, 47 (1–2):79–88, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2001:FSD

[ZW01] Jin Zhang and Yuehua Wu. A family of simple distribution functions to approximate complicated distributions. Journal of Statistical Computation and Simulation, 70(3):257–266, 2001. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2004:NDL

[ZY04] Jin Zhang and Keming Yu. The null distribution of the likelihood-ratio test for two upper outliers in a gamma sample. Journal of Statistical Computation and Simulation, 74(6):461– 467, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2009:IMD

[ZZCS09] Wei Zhang, Ying Zhang, Kathryn Chaloner, and Jack T. Stapleton. Imputation methods for doubly censored HIV data. Journal of Statistical Computation and Simulation, 79(10):1245–1257, 2009. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.